A Complete Bibliography of Publications in The Journal of Typographic Research and an

Extensive Bibliography of Publications in Visible Language

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 1.49

Title word cross-reference

$1.95 [152, 71]. $10.00 [83, 71]. $11.95 [1299, 1301]. $12.50 [152]. $12.95 [1141, 83]. $125.00 [1140]. $15.00 [152, 152]. $16.50 [152]. $17.00 [1331]. $17.95 [1512]. $18.00 [242]. $18.95 [1379, 1300]. $19.50 [1142, 1122]. $19.95 [1245, 1296, 1080]. $2.00 [83]. $2.45 [71]. $2.85 [56]. $20.00 [1510]. $21.95 [1647]. $22.95 [1277]. $24.00 [275]. $24.95 [1142, 1552, 1555, 1621, 1646]. $25.00 [1389, 1601]. $26.00 [1624]. $26.95 [1246, 1376, 1622]. $27.99 [1375]. $29.95 [1461, 1557, 1318]. $3.00 [83]. $3.75 [152]. $30.00 [1123, 1571, 1462]. $34.95 [1461]. $35 [1390, 1256]. $35.00 [1375, 1377, 1141, 1551, 1554, 1511]. $39.00 [1626]. $39.95 [1375, 1572, 1344, 1625]. $4.95 [263, 71]. $40.00 [1645, 1298, 1649, 1623].

$44.95 [1276]. $45 [1391, 1258]. $45.00 [1573, 1218, 1329]. $48.00 [1553]. $482.00 [1275]. $49.95 [1648]. 4 × 7 [416]. $5.50 [116]. $50.00 [1375, 1600]. $532.00 [1275]. $55.00 [1319]. $56.50 [287]. $60 [1429]. $7.95 [153]. $75 [1257]. $9.00 [152]. $90.00 [1311]. • [1050]. $I8.00 [1599]. × [1257, 1344, 1391, 489, 275, 521, 242].

*A [1277].

0 [1556, 457, 1142, 1389, 1122, 1123, 1141, 1246, 1379, 1429, 1376]. 0-201-51048-0 [1122]. 0-226-06715-7 [1389]. 0-226-46867-4 [1599]. 0-253-33776-3 [1375]. 0-262-02545-0 [1512]. 0-262-07248-3 [1462]. 0-262-12271-5 [1511]. 0-262-13467-5 [1553]. 0-262-13472-1 [1510]. 0-262-18186-X [1319]. 0-262-20168-2 [1552]. 0-292-71334-7 [1573]. 0-300-07461-1 [1297]. 0-300-11775-2 [1554]. 0-393-73034-4 [1376]. 0-415-17882-7 [1375]. 0-415-18685-4 [1379]. 0-471-97858-2 [1318]. 0-500-01925-8. [1375]. 0-584-10354-9 [420]. 0-670-84302-4 [1246]. 0-672-48543-5 [1142]. 0-674-90215-7 [1277]. 0-684-84001-4. [1300]. 0-688-141188-8 [1331]. 0-691-01544-9 [1141]. 0-7007-1135-X [1328]. 0-7123-4588-4 [1310]. 0-7484-0672-7 [1276]. 0-8014-3559-5 [1329]. 0-85331-360-1 [489]. 0-86142-110-8 [1378]. 0-902703-12-9 [457]. 0-907259-05-7 [1123]. 0-9072759-08-1 [1298]. 0-9729696-2-4 [1509]. 00-950 [1375]. 0396-3004 [1330].

1 [1556, 1391, 1390, 1245, 1429]. 1-52730-637-8 [1390]. 1-55860-936-9 [1461]. 1-56898-047-7 [1258]. 1-56898-048-5 [1245]. 1-56898-089-2 [1299]. 1-56898-273-9 [1391]. 1-57230-637-8 [1375]. 1-58093-003-4 [1311]. 1-80559-71-4 [1296]. 1-85414-286-0 [1556]. 1-85973-424-3 [1461]. 1-884718-15-9 [1257]. 1-884718-93-0 [1344]. 1-887123-01-6 [1256]. 1-932026-01-0 [1429]. 1-949484-09-2 [1846]. 1/2 [489, 275, 521, 242]. 1/4 [275]. 10.95 [1217]. 107pp [153]. 1260-8750 [1330]. 128pp [116]. 15 [334]. 15th [1048]. 16.00 [702]. 160/ [164]. 160/- [164]. 16th [287, 1810]. 1700s [1219]. 1702 [984]. 17th [287]. 180mms [1257]. 1930s [1102, 1314]. 1945 [1344]. 1960s [1557, 1816]. 1968 [60]. 1969 [126, 140]. 1970s [1816]. 1971 [1697, 1696]. 1988 [985]. 1990s [1356]. 19th [1793]. 1Os [31].

2-00-993486-5 [1624]. 2000 [1320, 1378]. 200083-88277-36-7 [1375]. 2013 [1688, 1697, 1695]. 20th [1375]. 210mm [1391]. 25th [269].

3-11-011129-2 [1275]. 3-11-014744-0 [1275]. 3-7643-0711-0 [407]. 3-9520297-1-8 [1375]. 3/4 [242]. 300 [1377]. 300-07570-7 [1377]. 30s [83]. 33.2 [1367]. 35 [1381]. 36 [1414, 35]. 39 [1480].

4 [1246, 1379, 1376]. 4-1 [489]. 4-1/2 [489]. 4.50 [312]. 40 [1502]. 404 [564]. 42 [1565]. 44 [1608]. 45 [1627, 1650]. 4th [242].

5 [1142, 1245]. 5.5.0 [153]. 5.80 [218]. 505 [41]. 50s [153].

6 [1065]. 62 [126]. 6o [1299].

7 [1377, 1389, 1123].

8 [1390]. 80p. [263]. 84/ [164]. 84/- [164]. 88-87469-00-8 [1375].

9 [457, 1391, 1141]. 9.50 [420]. 9.95 [1310]. 90-6617-212-6 [1366]. 91 [1025, 1140]. 978-0-12-370896-0 [1572]. 978-0-230-52716-4 [1555]. 978-0-262-01548-6 [1625]. 978-0-262-01742-8 [1647]. 978-0-262-11298-7 [1551]. 978-0-262-51240-4 [1601]. 978-0-393-07222-8 [1622]. 978-0-393-73293-1 [1621]. 978-0-935640-98-4 [1645]. 978-1-58729-906-3 [1646]. 978-1-59253-595-8 [1600]. 978-1-84150-195-6 [1571]. 978-1-84150-240-3 [1623]. 978-1-84788-576-0 [1648]. 978-1-949484-09-0 [1846]. 978-90-6369-259-9 [1626]. 978-90-6369-271-1 [1649].

A. [702, 287]. A.TYP.I [69]. A.Typ.I. [67]. Abbot [559]. Abbott [1258, 1299]. ABC [366, 152, 988]. Abel [1626]. Abeyance [634]. Ability [1801, 819, 1243, 716]. Abrams [1509]. Abstract [475, 596]. Abstract/Biography [475]. Abstraction [1179, 1188, 1525]. Abstractions [775]. Abstracts [22, 32, 137, 513, 579, 858, 957, 963, 46, 58, 74, 87, 95, 106, 117, 127, 141, 155, 165, 179, 191, 198, 208, 220, 231, 245, 255, 266, 276, 289, 302, 314, 323, 335, 354, 370, 380, 390, 399, 408, 422, 435, 445, 458, 476, 491, 501, 522, 541, 552, 599, 619, 626, 640, 653, 670, 847, 346, 562, 577]. academia [1778]. Academic [1831, 1738, 1727]. Academy [1318]. Acadian [1113, 938]. Access [821, 1342, 1779]. Accessible [1662]. Accountability [1633]. Accuracy [65]. Accurate [1827]. Achievement [949]. Acknowledgement [1823]. Acknowledgements [512, 1395]. Acquisition [282, 730, 384]. acting [1749, 1551]. Activation [820]. Active [1317, 830]. Activity [1551, 651, 1462]. Acts [1478]. Adaptation [936]. Adapted [550]. added [1418]. Addendum [648]. Addison [1122]. Addison-Wesley [1122]. Address [249]. Adjustments [44]. Adobe [1142]. Adoption [810]. Adornment [1766]. Ads [1255]. Adult [519, 865, 710]. Adults [431, 1386]. Advancing [1494]. Adventure [387]. Advertisements [53, 1518]. Advertising [216, 439, 536]. Advisors [976]. Advisory [535, 952, 942, 968, 995, 1012, 1037, 1026, 1054, 1077, 1125, 1115, 1135, 1145, 1151, 1160, 1169, 1176, 1195, 1210, 1251, 1225, 1238, 1261, 1271, 1279, 1313, 1291, 1312, 1339, 1347, 1361, 1370, 1385, 1394, 1407, 1431, 1422, 1441, 1450, 1456, 1465, 1473, 1482, 1491, 1497, 1504, 1515, 1520, 1546, 1530, 1567, 1585, 1592, 1603, 1617, 1639, 542, 604, 605, 631, 642, 657, 671, 684, 685, 694, 705, 721, 734, 747, 757, 769, 784, 815, 786, 795, 806, 828, 860, 863, 871, 883, 896, 909, 921, 1659]. Advisory [1653, 1699, 1674, 1680, 1690, 1712, 1723, 1746, 1754]. Aesthetic [1433]. Aesthetics [108, 1487, 962]. Affecting [1818]. Affix [341].

Affordance [1410]. Africa [1703]. African [1242]. After [1556, 1531, 248, 1534, 1846]. AG [1624]. Age [1599, 1264, 1801, 1633, 960, 133, 1306, 1722, 1245, 1429, 1362, 820, 1534, 811, 1555]. Agents [1685]. Ago [1303]. AI [1841]. Aid [80, 982, 853, 183]. Aided [568]. Aids [52]. AIGA [1296]. AIME [1715]. Akademie [1366]. AKENATON [1330]. akenaton docks [1330]. Akros [1378]. al [1375]. Alan [1344]. Alarm [1109]. Albers [1726]. Albert [236, 242]. Alberto [1246, 229]. Albinson [1645]. Aldus [846]. Alex [1376]. Alexander [1375]. Algorithm [173]. Alice [1095]. ALIRE [1330, 1330]. Alison [1090]. Alix [1388]. All-display [53]. Allen [312]. Allmusic [1540]. Allworth [1296]. Along [1002]. Aloud [1800, 1813]. Alphabet [214, 387, 1508, 3, 202, 482, 135, 62, 405, 937, 1040, 1173, 393, 1049, 1038, 104, 1379]. Alphabetese [500]. Alphabetic [1218, 160, 63, 912]. Alphabetical [1404, 90]. Alphabets [19, 318, 152, 24]. Alphameric [163]. Alphanumeric [506]. also [116]. Alston [31]. Alternative [375, 962]. Alto [1733]. am [1001]. Ambrose [275]. America [762]. American [444, 163, 1190, 585, 1341, 1231, 947, 1242, 71]. Amorous [499]. amply [1080]. Amsterdam [1626, 1649]. AnAesthetic [1433]. analyses [1810]. Analysis [149, 1282, 1578, 263, 510, 1440, 1416, 1419, 1495, 568, 203, 1781]. Analytical [1254, 1495, 1640]. Analyzing [1776, 899, 1499]. Anamorphosis [64]. Anatomie [242]. Anatomy [1550, 1043]. Ancestor [1389]. Ancient [211, 1108, 610, 285, 621, 320]. Ancillary [25]. Andersen [1621]. Andrew [1645]. Anfsätze [407]. Angeles [1436]. Anne [1184, 1492]. Anno [387]. Annotated [1401, 1399, 1403, 1397]. Announcement [291, 1250]. Announcements [142]. Annual [60, 126]. Anomalies [1715]. Ansgewählte [407]. ante [1508]. Anthenum [985]. Anthology [83, 1296, 83, 1319]. Anthony [1277]. Anthropology [1820, 1821, 1823]. Anti [1064]. Anti-Book [1064]. Antibiotic [1701, 1815]. Anticipating [1806]. Antinomies [1586]. Antoine [842]. Antwerp [844]. Any [65, 1795]. Aphasia [1799, 1785]. Apollinaire [998, 481]. Apostrophe [497]. appearances [1252]. Applebroog [1074]. Applications [856, 1806, 789]. Applying [1829]. Appraisal [891]. Apprenticeship [1215]. Approach [1693, 124, 112, 1366, 1462, 547, 76, 902, 1215, 557, 406, 833, 1704, 1728, 1640, 1300]. Approaches [899, 1683]. Approaching [763]. appropriate [1730]. Appropriateness [1605]. approx [242]. approximately [1275]. Apps [1837]. April [140]. Arabetic [1454, 1500, 1550]. Arabic [1791, 1453]. archaeological [1828]. Archbishop [517]. Architectural [1391, 1245, 1299, 1757]. Architecture [1621, 1682, 1311]. Archive [1444]. Archives [285, 1744]. area [328]. Argument [406]. arithmetic [1728]. Arithmetical [729]. Arnold [31, 485]. Art [1375, 1573, 1645, 793, 1004, 153, 1177, 1119, 1031, 1010, 845, 1306, 1065, 1722, 83, 1097, 1487, 1066, 1512, 1060, 878, 1193, 1076, 1240, 454, 1185, 1293, 1798, 1788, 368, 1557]. Arthur [164]. Article [737, 991, 1367, 1143]. Articles [22, 32, 513, 46, 58, 74, 87, 95, 106, 117, 127, 141, 155, 165, 179, 191, 198, 208, 220, 231, 245, 255, 266, 276, 289, 302, 314, 323, 335, 354, 370, 380, 390, 399, 408, 422, 435, 445, 458, 476, 491, 501, 541, 552, 599, 619, 626, 640, 653, 670, 346, 562, 577]. Artisan [1598]. Artist [1375, 1486, 1535, 1063, 1256, 1073, 1076, 1019]. Artistic [213]. Artists [1476, 1086]. Arts [113, 1068, 174, 645]. ASCII [1293]. Ashall [1257]. Ashley [788]. Asian [1620]. Askim [152]. Aspect [388]. Aspects [516, 809, 353, 808]. Aspen [778]. Aspires [1540]. Assent [538]. assessed [1781]. Assessment [1661, 162]. Assessments [1056]. Assimilation [1428]. Assisted [76]. Association [1429, 153]. Associativity [1105]. Assumptions [1587, 1715]. Asymmetric [104]. Atchley [152]. Attempting [1232]. Attention [1599, 1795, 1808, 1066]. Attributes [294]. ATypI [1429, 1308]. Auckland [521]. Audible [1353]. Audience [1604, 966]. Audience/online [1604]. Audio [974]. Audio-Visual [974]. Augmented [1842]. Aura [1352]. Austin [1573]. Authenticity [751]. Author [45, 460, 1061, 1404, 1448, 1544, 1672, 1673, 31]. Authors [11, 23, 33, 48, 59, 75, 89, 96, 107, 118, 128, 143, 156, 166, 180, 190, 199, 209, 221, 232, 246, 256, 267, 277, 290, 303, 315, 324, 336, 347, 355, 371, 381, 391, 400, 409, 423, 436, 446, 459, 492, 503, 523, 553, 563, 578, 627, 641, 654, 668, 678, 682, 692, 703, 720, 733, 743, 755, 768, 783, 794, 804, 814, 826, 848, 869, 880, 894, 906, 918, 929, 940, 951, 993, 1330]. autism [1805]. Automated [112, 195]. Automatic [1835]. Automobile [1606]. Autumn [564]. Available [1375]. Avant [965, 1375, 960, 1478, 1070, 1226, 959, 1306, 1320]. Avant-Garde [965, 960, 1070, 1226, 959, 1375, 1478, 1306]. Avant-Gardes [1320]. Avatars [1244]. Avenue [1378]. Awareness [789]. Axis [1294]. Azazian [365].

B [1601, 498, 360, 1275, 702, 14, 152]. b/w [702]. Babel [187, 1108]. Back [97, 192, 862, 870, 895, 907, 919, 931, 944, 295, 479, 536, 1444, 544, 1286]. Background [584, 587]. Backgrounds [740]. Backward [1669]. Backwards [905, 1611]. Bahia [915]. Balance [1606]. Baldwin [241]. Ballot [1138]. Ballpoint [550]. Baltimore [71]. Bandwidth [1051]. Banham [1436]. Bank [485]. Banks [1408]. Bann [83]. Bar [638]. Barcelona [1356]. Barnhurst [1390]. Baron [1379]. Baroque [888]. Barriers [1705]. Barthes [496, 497, 500, 630, 639]. Bartram [1344]. Bas [1626]. Based [294, 1355, 1806, 1832, 1811, 753, 1419, 1805]. Basel [1624, 407]. Basic [473, 375, 1799]. Baskerville [150, 420]. Bastard [1060]. Battuta [1589]. Bauhaus [1846]. Baurmann [1275]. Bayer [1052]. Bazaar [1171]. Be [1015, 1228, 1594, 1829, 1776, 1557, 1226, 66, 1630]. Beck [1217, 1556]. Beckett [877]. Becoming [637, 1305]. Becoming-zoa [1305]. Been [1308, 1171]. Bees [634]. Before [1846, 790, 383, 1038]. Beginning [1015, 1831, 1048, 238, 15]. Behavior [101, 1138]. Behaviors [483]. Behavioural [566]. Behind [1500, 1688, 1755, 776, 1695]. Beier [1649]. Beirut [1656]. Beiträge [346, 562, 577]. Belle [1017, 1011]. Ben [1478]. benefits [1707]. Bengali [1328]. Benign [1587]. Bennett [484]. Berg [1461, 1648]. Berlin [1275, 961]. Berry [1429, 242]. Bespoke [1767].

Besserat [1573, 1331]. Best [1688, 1695, 1783]. Better [152, 1213, 1780, 1830]. Between [1821, 1170, 983, 1402, 296, 1198, 966, 279, 1188, 1357, 1493, 1760, 1718, 1306, 1426, 1741, 927, 1814, 1772, 710]. Beverly [1600, 1688]. Beware [625]. Beyond [1717, 1820, 1596, 1582, 1206, 1108, 296, 1667, 494, 1569]. bibl [287]. Bible [698]. Biblica [698]. Bibliographical [39, 111]. Bibliographies [1383]. Bibliography [31, 857, 1397, 766, 1489, 228, 330, 376, 16, 667]. Bibliophiles [1507]. Biculturalism [1107]. Bilingual [1656, 1108, 935, 1110, 1106, 1214]. Bilingualism [1112, 1111, 609]. Binding [1507]. Bing [1533]. Biographical [1300]. Biographies [859]. Biography [475, 511, 1405, 1672, 1673]. Birken [892]. Birkhäuser [1624, 407]. birthday [280]. BIS [1626, 1649]. Bitten [269]. Black [1811, 1551, 1552, 1553, 1557, 1573, 1599, 1601, 1622, 206, 1511, 1512, 1646, 1647, 1080, 1623, 1275, 1141]. Blackbird [1358]. Blackboard [1013]. Blake [1006, 1469]. Blandford [242]. Blauvelt [1645]. BLEND [801]. Blind [1139, 49]. Bloomington [1375]. Blumenthal [56]. Blunt [1666]. Blush [1105]. Board [535, 952, 942, 968, 995, 1012, 1037, 1026, 1054, 1077, 1125, 1115, 1135, 1145, 1151, 1160, 1169, 1176, 1195, 1210, 1251, 1225, 1238, 1261, 1271, 1279, 1313, 1291, 1312, 1339, 1347, 1361, 1370, 1385, 1394, 1407, 1431, 1422, 1441, 1450, 1456, 1465, 1473, 1482, 1491, 1497, 1504, 1515, 1520, 1546, 1530, 1567, 1585, 1592, 1603, 1617, 1639, 542, 604, 605, 631, 642, 657, 671, 684, 685, 694, 705, 721, 734, 747, 757, 769, 784, 815, 786, 795, 806, 828, 860, 863, 871, 883, 896, 909, 921, 1659]. Board [1653, 1699, 1674, 1680, 1690, 1712, 1723, 1746, 1754]. Bodies [1152]. Bodoni [150]. Body [1041, 1294]. Boggs [1192, 1191]. Bohême [876]. Bolter [1512]. Bone [1501]. Boniface [517]. Bonnie [1551, 241]. Bonsiepe [1366]. Book [21, 31, 239, 1320, 1551, 1552, 1553, 1554, 1555, 1556, 1557, 1571, 1572, 1573, 1599, 1600, 1601, 1621, 1622, 1645, 1696, 116, 218, 206, 312, 334, 420, 457, 1142, 1217, 1257, 1310, 1391, 1461, 104, 153, 702, 984, 489, 1218, 164, 108, 876, 345, 56, 263, 1366, 1688, 1695, 636, 985, 1774, 1462, 1377, 1072, 1231, 1336, 275, 1068, 1389, 152, 1256, 1065, 1722, 1457, 83, 1378, 1697, 778, 1122, 1744, 1326, 1328, 1066, 1390, 1123, 1141, 1150, 1246, 1245, 1258, 1277, 1276, 1296, 1298, 1299, 1318, 1297, 1300, 1319, 1311, 1301, 1329]. Book [1331, 1379, 1429, 1509, 1510, 1511, 1512, 1625, 1624, 1626, 1646, 1647, 1648, 1649, 1846, 1064, 1060, 872, 1073, 1140, 879, 1330, 1507, 1075, 1074, 1076, 1080, 287, 248, 1376, 1623, 1018, 521, 71, 1275, 1698, 1731, 242, 407, 717, 1569, 142, 1601, 1257, 1722, 444]. Bookbinding [1071]. Bookcover [1325]. Bookmaking [116]. Books [72, 178, 1081, 1144, 1220, 1247, 1375, 1463, 1555, 1571, 1344, 1820, 924, 56, 340, 1008, 1063, 152, 1256, 1273, 778, 1498, 287, 1214, 1640, 521, 71, 1069, 1698, 916, 1743, 1298, 1140, 1375, 1256]. Bookshelves [1053]. Border [1335]. Borderland [1824]. Borders [1596]. Borgia [1765]. Bornhurst [1375]. Boston [328, 946]. Bottéro [1389]. Bound [1268, 1552]. Boundaries [1821, 1616, 1162]. bow [1494]. Bowel [1807]. Box [1375, 241]. Boyer [1245]. Braille [1803, 812]. Brain [1641, 71, 912, 1711]. Brains [1622]. Brainstorm [1829]. Brainstorming [1829]. Brainy [1711]. Branczyk [1375]. branding [1781]. Brandmarks [1829]. Brazilian [1127]. Brenda [1461, 1122]. Bridge [983]. Bridging [850]. Brief [358, 216, 1714, 345, 344, 364, 284, 321, 217, 1706, 310]. Brightness [608]. Bristol [1571, 1623]. Britain [83]. British [1257, 1310, 1344, 887, 1173]. Broadening [1051]. broadsides [1791]. Brogue [719]. Broken [159]. Bronze [252]. Bror [784]. Brown [1434]. Browser [1570]. Browsing [1505]. Bruce [227]. Buches [407]. Buchstaben [242]. Building [1402, 1748, 1400, 1398]. Burke [1697, 1367]. Burma [410]. Burnt [1435]. Bush [1492]. Business [1510, 985, 1199, 1237, 1287]. Butor [532]. Byzantine [425].

C [31, 1341, 517, 152, 360]. Cadmonian [251]. Cage [1087]. Cain [1388]. Cal [1080]. Calculating [1728]. calendar [1794]. California [1142, 817]. Call [1025, 1368, 1384, 1392, 1559, 1663, 1687, 1825, 1751, 1752]. called [687]. Calligrammes [481]. Calligraphy [1500, 485, 211, 204, 183, 288, 38]. calls [1761]. Cambridge [1551, 1552, 1553, 1557, 1601, 1462, 83, 1277, 1319, 1510, 1512, 1625, 1647]. Came [1594]. Camus [236]. Can [269, 1729, 1228, 1147, 1788]. Canadian [1142]. Cancer [1662, 1807]. Cannons [551]. Cannot [1626]. Canon [1748, 499]. Can’t [636]. Capacity [1834]. Capelli [1375]. Capital [214, 1556, 1217]. Capitalization [271]. caps [417]. Captive [1017, 1011]. Captivity [1011]. Card [949]. Carer [1661]. Caribbean [1538]. Carne [377]. Carole [1184]. Carr [1622]. Carter [457]. Cartographic [124, 99]. Cartographies [1509]. Cartography [1717, 149, 183]. Case [529, 1580, 148, 1684, 647, 1072, 901, 1299, 1444, 925, 406, 1113, 1698, 855, 1794, 1815, 1705, 1311, 1779, 1726]. Cases [1656]. Caslon [150]. Castelló [1461]. Casting [840]. Cat [1006]. Cataloging [1076]. Catalogue [967, 457]. Catalogues [799]. Categorical [1619]. Catfish [275]. Cathode [7]. Catich [275]. Cause [1055]. CBS [35]. CD [1330]. CD-ROM [1330]. Celebrating [1577]. Cell [1461]. Center [1375, 1645]. Centered [1462, 1704, 1788, 1276]. Centuries [287]. Century [924, 1070, 1048, 1256, 1810, 946, 1635, 1172, 1571, 1791, 1306, 1793, 1375, 116, 702, 997, 104, 197]. Cesare [817]. châteaux [876]. Chado [1374]. Chaldean [690]. challenge [1788]. Challenges [1809, 1637]. Ch’an [204]. Chancery [702]. Change [798, 1633, 1685, 1588, 712, 1571, 1276]. Changed [797]. Changes [122, 79, 661, 1677]. Changing [1596, 259, 1104, 1724]. Character [294, 1056, 14, 6, 1328, 102, 330]. Characteristics [345, 101, 1810]. Characterization [393]. Characters [163, 740, 539, 1079, 5, 1365, 100, 414, 506, 330, 773, 573, 417, 415]. Charles [876]. Charts [1755, 1777]. Chasing [1646]. Chasm [710]. ChatGPT [1841]. Chaucer [538]. Check [1121, 569]. Checklist [767, 967, 287]. Chew [269]. Chiasmus [636]. Chicago [1599, 1389, 83]. Chicano [939]. Child [482]. Children [823, 344, 679, 822, 442, 546, 950, 1242, 1498, 575, 1837, 1214, 1640, 454, 1782].

China [456, 1647]. Chinese [1056, 1196, 890, 1365, 414, 415, 773, 1518]. Chips [1194]. Chizlett [1139, 1143]. Choice [1197]. Chopin [764]. Chris [1648]. Christian [116]. Christine [1245]. Christopher [1697]. Chronological [25]. Chronology [760]. Chunking [791]. Cinema [763, 1325]. Cinematic [527, 1716]. Circle [1095]. Circulation [434, 488, 525, 555, 629, 742, 805, 882]. Citrakāvyas [891]. City [1326, 1625, 1646, 1686, 484, 311]. Civil [119, 146]. cixous [529]. Claims [1124]. Clarisse [1389]. Clark [457]. Class [597, 1668, 341]. Classic [1766]. Classification [570, 159, 186]. Classroom [1293, 1254, 1667, 1499, 789]. Claudel [481]. Claudian [252]. Clay [521]. clinicians [1709]. Clive [1139, 1143]. Clocks [1109]. Cloth [164, 1268, 405, 1257, 1462, 83, 1122, 1141, 1246, 1245, 1329, 1331]. Cloth-Bound [1268]. Clothbound [56, 287]. Clues [80, 100, 1715]. Cluster [519]. Co [1708, 1815, 1837, 71]. Co-design [1837]. Co-designing [1708, 1815]. Code [939, 102]. Code-switching [939]. Codes [820]. Codex [1032, 1765, 297]. Codexspace [1206, 1252]. Coding [823]. Coffee [779]. Cognition [1691, 1295]. Cognitive [651, 1419, 1828]. Cohen [1326]. Coin [1092]. Colin [1572, 1624, 1408]. Collaboratel [926]. Collaborates [1598]. Collaboration [1068, 1569, 1443, 1444, 1833, 1275, 1623]. Collaborative [1840]. Collage [1098]. Collages [362]. Collection [1034, 1349]. Collections [1773]. Collective [1447]. College [457, 153, 275]. Collier [241]. Collisions [934]. Colloquial [925, 717]. colonial [1757]. Colophon [526, 601, 849, 955, 1047, 1529, 1652]. Color [877, 42, 1191, 1255, 1375, 1554, 1556, 1572, 1600, 1621, 1645, 1461, 984, 1377, 1122, 1390, 1258, 1276, 1297, 1311, 1329, 1331, 1429, 1509, 1625, 1624, 1626, 1648, 1649, 1330, 1275, 1366, 1328]. Colored [740]. colors [1298, 1299, 1140]. Colour [334, 43]. Columbian [1283]. Comenius [1100]. Comics [1611]. Coming [960]. Comma [1058]. Commemoration [1408]. Comment [187, 188, 286, 388, 367, 138, 151, 114]. Commentary [1344, 56, 19, 70, 82, 681, 1751, 1822]. Comments [738, 611, 737, 1022, 1082]. Commerce [457]. Commercial [1737]. Commitment [646]. Committee [67]. communicate [1786]. Communicating [1461, 731, 1412, 1796]. Communication [1368, 1384, 1383, 1559, 1663, 1752, 798, 1612, 99, 1736, 1633, 1821, 1644, 1052, 1427, 1685, 1790, 1701, 12, 1598, 1789, 1051, 1704, 1705, 1276, 1300, 1423, 1416, 1768, 1579, 1417, 1840, 1032, 1568, 796, 1583, 1494, 1587, 1630, 1778, 1590, 1245]. Communications [358, 506, 1241]. Communicative [1620]. Communities [1598, 1560]. Community [1093, 1086]. companion [1786]. Company [1621, 1622, 1331]. Comparative [1301, 1810]. Compared [843]. Comparison [1501, 972, 3, 664, 1426]. Competence [571]. Competency [707]. Competing [1433]. Compiled [1275]. Complaint [1185]. Compleat [312]. Complex [1682, 1780]. Complexity [1835]. Components [676, 1440, 168, 506, 1814]. Composer [27]. Composing [750]. Composition [112, 44, 902, 1643]. Comprehensibility [1812].

Comprehension [825, 1838, 452, 1678]. Computational [677]. Computer [798, 65, 133, 466, 333, 34, 76, 1523, 91, 851, 850, 331, 114, 568, 811, 17, 1605, 569, 1069, 164, 801]. Computer-Aided [568]. Computer-drawn [34]. Computer-hyphenation [65]. Computer-Produced [466]. Computerized [201, 78]. Computers [855, 1078, 902, 1485, 1809, 1069, 1122]. Concept [737, 695, 621]. conceptions [1778]. Concepts [1506]. Conceptual [1242]. Concerning [463, 1048]. Concrete [338, 351, 1843, 83, 234, 28, 83]. Concretism [83]. Concurrency [741]. Condensed [991]. Condition [1540]. conference [1308, 707, 1666]. Conferencing [1597]. Conflicting [647]. Conflicts [939]. Conformity [646]. Confrontation [1595]. Congeniality [29, 40]. Congress [69, 69]. Connected [483]. Connection [394, 1461]. Connotative [1702]. Consciousness [1392]. Consensus [1763]. Consequently [1811]. Conservatism [622]. Considerations [548]. Consistency [147]. Constant [1588]. Constellation [1771]. Construction [688, 1179, 1326, 1171, 1657]. Constructional [575]. Constructivist [825]. consultancy [1311]. Consumer [216, 1644]. consumers [1786]. Contact [934]. Contemporary [1375, 1476, 873, 1314, 1073, 132, 376, 1354, 1683, 1349]. Content [1816, 1580]. Contents [461]. Context [980, 559, 539, 679, 80, 853, 540, 1725, 1568, 1607]. contexts [1696, 1697, 1678]. Contextualizing [1294]. Continuous [1811]. Contour [206]. Contra [1006]. Contra-Diction [1006]. Contract [639]. Contractions [925]. Contrast [608]. contribution [1727]. Contributors [598, 1742]. Control [1272, 131]. controlled [1710]. Conventions [1827]. Converging [939]. Conversation [1095, 1090, 1655, 648]. Conversational [1789]. Cooking [1308]. Cooper [1301]. Copeland [152]. Copenhagen [116]. Copies [536, 1286]. Coppersmith [71]. Copying [1057, 575]. Copyright [1286]. Core [1634]. Cornell [1329]. Cornoldi [817]. Corporation [152]. Corpus [887]. Correction [8, 1380]. Correspondence [20, 669, 1199, 1237, 1287, 1694, 398, 84, 177, 613, 207, 301, 612, 254, 85, 105, 628, 313, 490, 433, 73, 389, 782, 369, 379, 244, 189, 322, 1449, 238, 86, 230, 797, 154, 219]. Correspondences [508, 432]. Cosmopolitanism [1107]. Cosmos [1029]. Costume [697]. Could [1630]. Coulmas [1275]. Council [457]. Count [623, 1777]. Counting [1228, 1777, 1331]. Countries [104]. Country [1787]. Coupons [1194]. Course [419, 272, 761]. Cover [252, 262, 311, 636, 771]. Covering [1071]. Covers [1820, 630]. Crâne [377]. Craft [307, 545]. Crafts [645]. Cramp [1044]. Created [1484]. Creating [185]. creation [1273, 1744]. Creative [1119, 1789]. Creativity [1049]. Credits [544]. Cresci [214]. Crisis [1091]. Critic [1152]. Critical [1717, 1719, 1687, 1344, 1713, 1644, 1668, 1254, 891, 1722, 1373, 1648, 1492, 1123]. Criticism [978, 866, 754, 1162]. Critique [1620]. Critiques [1668]. Critiquing [1748]. Crombie [1457]. Cross [1375, 1595, 972, 1426]. Cross-Cultural [1426]. Cross-cultures [1595]. Cross-Media [1375]. Cross-Model [972]. Crossing [638]. Crossroads [1030]. Crowbar [1269].

CRT [868, 800, 17]. Crutcher [1699]. Cruz [817]. CT [1554]. Cuban [1128]. Cubism [258]. Cubo [1060]. Cubo-Futurist [1060]. Cue [298]. Cueing [1413, 1120]. Cues [344, 982, 364, 611, 1020, 674, 663]. Cultivating [1409, 1396]. Cultural [1384, 1383, 1645, 1820, 1426, 1423, 1804, 1678]. Culture [873, 1440, 1594, 1720, 138, 1809, 1296, 1429, 1162, 1568, 1766, 1624]. Cultures [962, 1595]. Cumulative [460]. Cuneiform [1108, 688, 689, 691, 1412]. CurioCity [1317]. Curious [359, 1277]. Currency [1190, 1285]. Curricula [94]. Curriculum [1595, 946]. cursive [613]. Curzon [1328]. Customers [1213]. Customized [1273]. Cut [1152, 1810]. Cutter [984]. Cybercities [1245]. Cyberpl [1461]. Cybertalk [1461]. Cybertext [1206]. Cyrillic [451].

D [104, 1429, 360, 71]. D. [1341]. Dada [961]. Dadaist [962]. Dal [116]. Damage [111]. Damned [1102, 1281, 1143]. Dana [152]. Danet [1461]. Danger [1485]. Danish [1794]. Darwin [1771]. Data [1693, 1662, 1731]. Database [1412]. Dataspark [1662]. Davenport [275]. David [1621, 104, 1512, 1646, 1792, 489, 1180]. Davies [1624]. Davis [1655]. Day [1265, 710]. Daydream [637]. Dé-Collage [1098]. Dead [1555]. Deadpan [1434]. Deaf [823, 442, 1813]. Death [1006, 1534]. Deborah [309]. Decades [56]. Decide [1103]. Deciphering [1031]. Decipherment [349]. Decision [1818]. Declaration [430]. Decoder [989]. Deconstruction [1165]. decoration [1757]. Decorative [484, 1782]. Deep [731]. Defense [622]. Defining [1797, 1799]. Definition [886, 98]. Defoe [637]. Degradation [587]. Deguy [1067]. Deguy/Dorny [1067]. deletions [901]. delightful [241]. Demand [1273]. Dementia [1661]. Democratization [1084]. Demonstration [839]. Demotic [1174]. Denial [538]. Denise [1573, 1331]. Denotation [518]. dependent [1293]. Depiction [1526]. Depression [1341]. Depth [1106]. Derrida [530]. Descartes [1204]. Describing [188, 330, 1640]. Description [664, 39, 111]. Desiderata [393]. Design [1550, 1776, 855, 54, 261, 253, 243, 477, 600, 596, 954, 1383, 1401, 1399, 1403, 1559, 1576, 1600, 1645, 1696, 1687, 1694, 1156, 1656, 334, 1580, 1733, 1670, 1593, 1633, 1821, 1831, 1596, 1161, 1153, 61, 1050, 1718, 1714, 1052, 1397, 1402, 1682, 1446, 133, 235, 1578, 832, 1684, 1085, 1631, 473, 1638, 212, 1289, 1823, 845, 1667, 1748, 1790, 1774, 41, 451, 843, 1462, 1845, 1400, 856, 1818, 1314, 361, 1139, 1642, 100, 1838, 1836, 131, 1165, 1006, 1789, 1440, 1155, 1418, 1445, 76, 1523, 1597, 1697, 1051, 1420, 1548, 581, 1634, 121]. Design [851, 1709, 1755, 1832, 580, 1276, 1297, 1409, 1396, 1429, 1416, 1443, 1510, 1511, 1512, 1560, 1624, 1626, 1647, 1629, 1725, 1846, 801, 151, 1468, 321, 595, 1140, 1444, 548, 1581, 1138, 1507, 1505, 1637, 1710, 1447, 1419, 1833, 1686, 132, 1166, 1425, 1632, 1162, 1376, 1586, 1386, 1495, 1780, 1829, 590, 1706, 1398, 1840, 1837, 186, 1243, 1018, 558, 1660, 17, 1640, 1643, 557, 701, 586, 1172, 1584, 1620, 37, 624, 1583, 1137, 833, 1049, 1494, 1657, 1698, 1726, 1747, 1763, 1830, 1492, 1807, 407, 1163, 1587, 1630, 1368, 1724, 1778, 1730]. design [1741, 201, 1804, 1786, 1311, 1788, 1459, 1590, 1764, 1159, 1375, 153, 152, 1551, 1554, 1572, 1600, 1645, 1391, 1366, 1377, 1296, 1731, 1344, 1688, 1695, 1258, 1296, 1648, 1623]. Design-Dragon [1584]. Design-led [1670]. Designalltag [1375, 1375]. Designed [1429]. Designer [1571, 520, 367, 487, 259, 1235, 1053, 1749]. Designers [738, 1685, 1619, 1426, 1837, 1698, 1708]. Designing [1661, 1684, 1462, 1787, 135, 1569, 1704, 1648, 1649, 1516, 1514, 1417, 1214, 698, 1703, 1763, 1708, 1815, 1750, 1298]. Designs [1378]. Detecting [1172, 1715]. Deutsch [522]. Developing [260, 1317, 1316, 1416, 1706, 1827]. Development [378, 35, 184, 1015, 1057, 1562, 1684, 226, 1707, 1606, 225, 574]. Developmental [824]. Developments [843, 79]. Devices [131]. Diacritical [3]. Diacritics [170]. Diagnostic [1045, 1818]. Diagram [461, 1771, 1201, 55]. Diagrammatic [235, 868]. Diagramming [1562, 1364]. Diagrams [1562, 1099, 1551]. Diagraphic [1120]. Dialect [718, 719]. Dialectics [1483]. Dialogue [1091]. Dials [1606]. Diane [1512]. Dick [1646, 1290]. Diction [1006]. dictionaries [1616]. Dictionary [1320]. did [521]. Didactic [1340]. Dietary [1522]. Difference [1117, 1760]. Differences [927]. Different [1580, 1778, 561, 820, 1495, 567, 1678]. Differential [583]. Differentials [40]. Differentiation [1671]. differently [1749]. Difficulties [724]. Digital [1717, 1719, 1555, 1687, 1718, 1714, 1734, 1254, 1306, 856, 1272, 1740, 1273, 1720, 1188, 1512, 837, 974, 1534, 1025, 1715]. Digits [416]. Diglossia [1107]. Digraph [431]. Digraphia [439]. Dilemma [559]. Dimension [607]. Dimensional [582, 1783, 352, 1299]. Dimensions [1384, 1383, 1392, 1295, 1093, 1120, 1440, 1423, 1073]. direct [1378]. Direction [691, 98]. Directional [147, 574]. Directories [1255]. Directors [738]. Directory [1600]. Diringer [104]. Dirty [1843]. Dis [1252]. Disabilities [1838, 608, 1704, 1836]. Disadvantaged [238]. Disambiguating [1367, 1304]. Disciplinary [1152, 1707]. Discourse [1461, 1402, 1229, 751, 1560, 499]. Discriminability [1079]. Discrimination [274, 15]. Discs [779]. Discursive [1153]. Discusses [1087]. Discussion [711, 713, 715, 140]. Disfluent [1015]. Disharmonies [645]. Disorders [825, 680, 824]. Display [1759, 741, 1845, 1810, 509, 854, 699, 1760, 53, 1652]. Displayed [416]. Displays [740, 1522, 588, 310]. Distributed [489]. Distribution [1566, 607, 773]. Disuse [170]. Divided [194]. Do [564, 1619, 537, 851, 1039, 1829, 1069]. Do/Be/Feel [1829]. DOC [1330]. Docs [1841]. Document [1458, 1273]. Documentary [1775]. Documentation [1776, 1013, 904, 1788]. Documentracings [465]. Documents [1644, 1411, 775, 1041, 1677]. does [1494]. Doing [1622, 1434]. Dole [362]. Domain [1329]. Dominic [817]. Donald [737, 1182]. Dooley [719]. Dorny [1067]. Dorny/Deguy [1067]. Dot [1233, 416]. dotlinepixel [1375]. Double [1576, 1112]. Doubly [1281]. Douglas [1843, 780]. Dowling [1646]. Down [1335, 1776, 453]. Dr. [1367]. Drachms [270]. Dragon [1584]. Drawing [206, 1079, 739, 1826]. DrawIt [1764]. Drawn [1541, 34, 1764]. Dreams [1597]. Dresden [242]. Drucker [1218, 1256]. Drug [26, 1386]. DTL [1563]. Duchamps [474]. Due [79]. Duncan [1378]. Dundan [1378].

Dunne [719]. Dupin [926]. Durability [197]. Duras [533]. Durch [474]. During [510, 1705, 1356, 1793]. Dutch [936]. Dynamic [547, 509, 1798, 1466, 854]. Dynamics [1425]. Dynamism [1770]. Dyslexia [824, 1842]. Dyson [1655].

E-Inclusion [1799]. E. [206]. each [206]. Earl [1367]. Early [1190, 1614, 677, 1070, 739, 1635, 1828, 1498, 973, 1253, 1219, 1306, 1757]. Earth [1040]. Easier [282]. Easy [136, 1262, 1812, 985, 1836]. Easy-to-Read [1812]. Ebsen [616, 614, 615, 617]. Eccentric [1314]. Echoing [1825]. Eckhardt [985]. Eco [1745]. Eco-System [1745]. Ecological [1462, 1818]. Ecologically [1364]. Economic [1341]. Economics [1599]. Economy [283]. Ed [83, 71, 242, 262, 1533]. Edge [472, 462, 1055]. Edges [1713]. Edible [1334]. Edinburgh [457]. Edited [817, 1846]. Editing [792]. Edition [1687, 1110, 1217, 56]. Editions [83, 1318]. Editor [54, 758, 1318, 536, 1552, 457, 1429, 1140, 47, 1367, 1236, 969, 1470, 1681, 1224, 686, 673, 1027, 57, 98, 182, 233, 421, 736, 1666]. Editore [1375]. Editorial [1199, 1237, 1287, 659, 1116, 1124, 1449]. Editors [1319, 1509, 1624, 1623, 1375, 1296, 1648, 1275, 1405, 1839]. edns [287]. Eds [1697]. Education [855, 175, 343, 1636, 1631, 281, 1638, 94, 1445, 1597, 772, 581, 1709, 1629, 151, 1637, 1089, 1745, 1632, 1241, 376, 1620, 1630, 1724, 1805, 707]. Educational [1694, 521]. Edutaining [1589]. Edward [457, 218, 275]. Effect [1801, 647, 1835]. Effecting [1315]. Effective [1427, 1688, 1420, 1416, 1763, 1695, 1376]. Effectiveness [1255]. Effects [483, 1055, 1079, 972, 1274, 518, 7, 1365, 791, 1014, 914, 1138, 240, 701, 2, 796, 661, 1616, 1784]. Efficiency [26, 791]. Efficient [1196]. Egypt [610]. Egyptian [41]. Ehlich [1275]. Eighteenth [997, 946]. eInk [1784]. Eisenberg [1275]. Ejlers [116]. Elam [1389]. Elderly [1045, 1780]. Electric [1722]. Electronic [1411, 1179, 778, 1150, 1245, 1507, 1425, 1181, 802, 1376]. Electronically [1612]. Element [258]. elements [1836]. Elias [237]. Elkins [1329]. Ellen [1645, 1258]. Else [1092, 83, 1509]. Else/Where [1509]. Email [1264, 1379, 1330]. Embellished [1759]. Embodied [264]. Embodiments [1172]. Emergence [169, 739, 279]. Emergencies [1705]. Emergency [1779]. Emergent [950]. Emerging [1645, 327]. eminent [1742]. Emmett [83]. Emotion [1295]. Emotional [725]. Emotional Digital [1375]. Emotions [1642]. Empathic [1823]. Empire [494, 1239, 496]. Empire/Gaining [1239]. empirical [1642, 1812, 1727]. Empowerment [712]. Enacted [913]. Encyclopaedia [242]. Encyclopedia [1257]. End [1334]. Ended [780]. Engagement [1457, 1793]. England [31, 94]. English [31, 251, 1379, 1508, 1136, 887, 901, 903, 772, 414, 396, 1424, 440, 696, 622, 898, 443, 586, 900]. English-Language [887]. English-Speaking [414]. Enhancing [1243, 1807]. Enough [176]. Enquiry [1190]. Environment [426, 385, 1581, 1683]. Environmental [1663, 1685]. Environments [1244, 1299]. Envisioning [1633, 1638, 1629, 1632]. Ephemera [56].

Epistemology [914]. Equipment [25]. Era [1831, 778, 837]. Ergonomics [576, 1606]. Eric [152, 1697]. Erik [1142, 1511, 116]. Errata [564]. Español [522]. essay [457, 1123]. Essays [407, 1349]. essence [273]. Essential [429]. Establishment [1086]. Esthetic [78]. Ethnographic [1363]. Eugene [83]. EUR [1461]. European [1341, 1453]. Evaluating [1703]. Evaluation [1838, 308, 665, 687]. Evans [1648]. Event [1620]. Events [1090]. Everett [1300]. Evers [1626]. Evidence [1761, 1832, 111]. Evidence-Based [1832]. Evocative [1552]. Evolution [36, 1412, 1328]. Evolutionary [1033]. evolved [1379]. evolving [1803]. exactly [1760]. examination [1642]. examples [1219]. Excellence [1745]. Excerpt [387, 495, 366, 333, 285, 103, 320, 299, 444, 352]. Excerpts [1541]. Excès [497]. Exchange [1194, 1193]. Exclusive [1626]. Exercises [37]. Exhibit [1075]. Exhibition [115, 767, 56, 1845, 30, 1097]. Exhibitions [1684]. Exhibits [1242]. Exile [1114]. Existential [1457]. Expanded [1166]. Expected [1320, 1715]. Experience [1295, 1374, 1709, 1788]. experience-design [1788]. Experiences [1684, 1241, 1803]. Experiential [1686]. Experimental [54, 253, 149, 1356, 1845, 42, 1387, 1314, 1355, 272, 1351, 557]. experimentalism [1349]. Experiments [564, 205, 1336, 537, 71]. Explain [1100, 1232]. Explanation [1170]. Explanations [753]. Explicit [279, 1104]. Exploits [1114]. exploration [1706]. Explorations [1809]. Exploring [1821, 1741, 1789, 1785, 1241, 1784, 1814]. Expression [516, 213, 1424, 182, 1506]. Expressively [1800, 1813]. Extending [1272]. extensively [1429]. Eyck [1366]. Eye [718, 675, 868, 1806, 1186, 71]. Eyes [1139].

F [1275, 242]. F. [420]. Faber [334, 702]. Face [1112, 1607, 1137, 1171]. Facial [973]. Facilitate [663]. facilitators [1708]. Facilities [378]. Facility [567]. Fact [1755]. Factors [1315, 1818, 675, 1419, 801]. Facts [911]. Faculty [1832, 1620]. Fagan [1389]. Failure [1580, 1588, 1832, 1579, 1577, 1584, 1590, 1583]. Failures [1559]. Fair [778]. Fallout [1096]. Family [1822, 1548, 1060]. Fanzines [1173]. Fashion [121, 1786]. Favourite [1006]. Faxcett [1554]. Faxcett-Tang [1554]. Features [1692, 193, 1507, 1837, 1798, 1804, 1727]. feedback [1624]. Feel [1829]. Ferguson [249]. Fernand [1008]. Festschrift [484]. few [1646]. Fiction [1618]. Field [1598]. Fife [1378]. Fifteenth [197]. Fifteenth-century [197]. Fifty [1311]. figs [489]. Figure [610, 608]. Figure/ground [608]. Figured [886]. figures [1275, 984]. Film [1179, 1534, 1607, 1212]. Finamore [1296]. Find [1142, 417]. Finding [1445, 586]. Findings [1014, 789]. Finley [719]. Fiona [1328]. Fire [1774]. First [90, 1677, 271, 1, 1742, 287]. Fit [549, 851]. Fitting [840]. Five [135, 1311, 238, 1140]. Five-Year-Olds [238]. Flaming [1537]. Flanders [637]. flatlands [1794]. Flemish [1810]. Fleurs [1072]. flexible [1273]. Flood [1119]. Floor [903]. Floppy [779]. Flow [1770, 1041]. Fluency [1057]. Fluxacademy [1089]. FluxBase [1097]. Fluxkids [1477]. Fluxus [1318, 1088, 1478, 1476, 1085, 1096, 1093, 1489, 1488, 1092, 1484, 1090, 1086, 1475, 1087]. Flying [1189]. Focus [1657]. Focused [1836, 1750]. Focusing [1838]. Folio [271, 287]. Font [1733, 1801, 1563, 737, 1451, 737, 695]. Fonts [1563, 34, 93]. Fontstyle [110]. Footnote* [1277]. Force [798]. Foreign [936]. Foreword [1152, 1161, 1170, 758, 885, 933, 1389, 1349, 686, 1027]. Forlag [116]. Form [1375, 1580, 1000, 983, 147, 843, 1795, 1457, 341, 1390, 418, 1579, 582, 590]. Formal [283, 451, 258, 218]. Format [51, 1554, 1556, 1273, 1140, 1080]. Formation [1466]. Formative [1029, 1030]. Forming [1447]. Forms [321, 624]. Formulation [279]. Forsenic [570]. Forth [295]. Forthcoming [1383]. Forward [1669]. Found [1533, 1326]. Foundation [1400, 1447, 1830]. Four [249, 56, 350]. Fournier [312]. Fourth [949]. Fragen [407]. Fragments [52, 499]. Frame [1099, 1104]. Framework [388, 1419, 1778]. Francais [522]. France [225]. Francis [1276]. Francisco [1461, 71]. Francois [1389]. Francs [407]. ‘Franglais’ [938]. Frank [457, 263]. Franked [638]. Franklin [1190]. Frascara [1276, 1672]. Frase [611]. Frederick [312]. Free [1000, 1300, 1454, 1375]. Freedom [999, 916]. Freeman [71]. French [22, 32, 513, 858, 46, 58, 74, 87, 95, 106, 117, 127, 141, 155, 165, 179, 191, 198, 208, 220, 231, 245, 255, 266, 276, 289, 302, 314, 323, 335, 354, 370, 380, 390, 399, 408, 422, 435, 445, 458, 476, 491, 501, 541, 552, 599, 619, 626, 640, 653, 670, 847, 346, 562, 577]. frenetic [365]. Frequency [623]. Freud [636, 632, 633]. Friedman [1318]. Friedrich [830]. friendly [1454, 1451]. Friends [1095]. Front [51, 1475]. Front-page [51]. Frontiers [71]. Früchte [893]. ful1 [1331]. full [1375, 1554, 1556, 1600, 1621, 1645, 1461, 1122, 1297, 1311, 1509, 1625, 1624, 1626, 1649, 287]. Function [1559, 9, 822, 1101, 341, 1029, 1424]. Functioning [867]. Fundamental [843]. Funny [1194, 1171]. Further [941]. Future [1636, 1638, 966, 1136, 286, 778, 1629, 1561, 1632, 1840, 833]. Futures [1478]. Futurist [1060, 406].

G [1375, 1192, 1328, 1390, 1191]. G. [784]. Gabberjabbs [1065]. Gabriele [1375]. Gaining [1239]. Galliard [838]. Game [1317, 1548, 593]. Games [1487]. Gamification [1837]. Ganis [1698]. Gap [850]. Garcia [1180]. Garde [965, 960, 1070, 1226, 959, 1375, 1478, 1306]. Garden [1181]. Gardes [1320]. Garland [1217]. Garment [1786]. Garner [1648]. Garofalo [1021, 1020]. Gather [1661]. Gay [1462]. Gaze [1154, 1806]. Gaze-Based [1806]. Gem [1307]. Gender [1547]. Gendered [1822]. General [130, 144, 157, 167, 200, 210, 223, 247, 265, 278, 292, 304, 317, 325, 337, 348, 357, 372, 382, 392, 402, 411, 424, 437, 448, 480, 502, 524, 554, 556, 565, 602, 606, 620, 643, 658, 672, 683, 704, 706, 722, 735, 748, 756, 770, 816, 787, 803, 807, 829, 861, 864, 881, 884, 897, 908, 922, 943, 953, 1024, 514, 536, 457, 1796, 517]. Generator [112]. Genesis [119, 146, 251]. Genome [1426]. Genre [886]. Geoff [1846]. Geoffrey [1257]. Geometric [418]. Georg [893]. George [152, 1301, 249]. Geri [1462]. German [22, 32, 513, 858, 46, 58, 74, 87, 95, 106, 117, 127, 141, 155, 165, 179, 191, 198, 208, 220, 231, 245, 255, 266, 276, 289, 302, 314, 323, 335, 354, 370, 380, 390, 399, 408, 422, 435, 445, 458, 476, 491, 501, 541, 552, 599, 619, 626, 640, 653, 670, 847, 346, 562, 577]. Germany [152, 960]. Geschichte [242]. Gestalt [407]. Gestures [591]. Get [1420]. Ghana [1702]. Giampietro [1645]. Giese [1275]. Gifts [645]. Gilliland [263]. Ginger [1142]. Giorgio [1698]. give [746]. given [484]. Glaister [1257]. Glass [1154]. Glen [1378]. Glide [1576]. Global [1576, 1600, 1593, 1093, 594, 1597, 1181, 1840, 1494]. Globalization [1568]. Gloss [1340]. Gluck [1275]. glyphic [1756]. Goals [647]. Godard [1325]. Going [1171]. Gold [1193]. Gomez [1555]. Good [1633, 1372, 927]. Goodhumor [405]. Google [1841]. Gothic [169, 451, 297]. Goudy [150]. Gourdie [152]. Grade [949]. Graduate [1636, 707]. Graffiti [311, 353]. Grafton [1277]. Grammarly [1841]. Granary [1256]. Granjon [838]. Grapheme [483, 238]. Graphemes [63]. Graphic [1645, 1156, 1580, 216, 1161, 1153, 972, 1045, 520, 1120, 1748, 1774, 1845, 934, 856, 1426, 1838, 1165, 1155, 1428, 1418, 581, 851, 580, 1276, 1296, 1846, 595, 1444, 584, 1833, 1686, 1166, 852, 1162, 1586, 274, 1837, 110, 1525, 1749, 37, 1836, 938, 1600, 1645, 1297]. Graphicacy [306]. Graphical [539, 775]. Graphics [1050, 718, 1100, 851, 850, 852, 217, 1796, 1311]. Graphological [516, 809]. Graphs [1104]. Gratisk [152]. Grave [1531]. gravestone [328]. Gray [1034, 206, 1006]. Grazhdanskii [119, 146]. Greater [1686]. Greece [621, 1389]. Greek [270, 320, 1253]. Greenwood [984]. Gregory [71, 1655]. Grey [466]. Grid [1601]. Grids [1671]. Grillet [635, 1017, 1011]. Gromala [1512]. Grotesk [1652]. ground [608]. groundwork [1749]. Group [37]. Grown [1532]. Gruyter [1275]. Gu [1533]. Guemil [1779]. Guest [1405, 969, 1681, 673, 1839]. Gui [1366]. Guide [45, 1310, 520, 152, 1783]. Guilford [1375, 1390]. Guillaume [481]. Guinea [439]. Guinta [844]. Gunter [1275]. Gunther [1275]. Gutenberg [1378]. György [984].

’h [1670, 164, 1275]. H. [104]. Hackiana [841]. Haiman [984]. half [1571]. Hall [1509]. Hamady [1065]. Hand [1375, 702, 591, 91, 840, 1043, 9, 985, 152, 31]. Hand-printed [91]. hand-writing [985]. hand-written [152]. Handbook [702, 987, 1275]. handedness [1452]. Handicap [1452]. Handicapped [4]. Handsomely [1635]. Handwriting [570, 169, 549, 343, 550, 571, 576, 308, 545, 547, 1014, 772, 551, 114, 1039, 376, 568, 1042, 566, 567, 203, 1781, 50, 1375, 152]. Handwritten [573]. Hangeul [1770]. Hannah [1601]. Hans [1391]. Happens [1230]. Hard [1552, 1776, 1375]. Hardbound [1551, 1553, 1555, 1556, 1557, 1573, 1600, 1622, 1389, 1429, 1625, 1649, 1375, 1391, 1377, 1256, 1328, 1258, 1277, 1298, 1319, 1311, 1510, 1511, 1275]. Hardcover [1390, 1647]. Hardware [8]. Harl [152]. Harpers [1171]. Harris [1375]. Harrow [1556]. Harry [457, 1683]. Harsy [842]. Hartley [1021]. Hartmut [1275]. Harvard [1277, 707]. Hastings [153]. Haven [1554, 1377, 1297]. Hays [1331]. HCI [1806]. Head [1683]. heading [1379]. Headings [175]. Headline [1518]. Health [1707, 1709, 1833, 1840, 1703, 1782]. Healthcare [1708]. Heard [972]. heavily [1554]. Hebrew [1792, 793, 774, 1111, 607]. Heft [1683]. Heinemann [521]. Helene [1462, 529]. Heller [1296]. Help [546, 1729, 1788]. Helping [902]. Helvetica [1607, 1625]. Helwig [892]. Hembrooke [1462]. Hendel [1377]. Hendrik [1810]. Henri [764]. Her [164]. Herbert [153, 164, 1052]. Here [1303]. Heritage [1351]. Herman [1018]. Hermeneute [749]. Hermeneutic [751]. Hermeneutics [1752, 754, 914]. Herrenschmidt [1389]. Herrick [1367]. Heuristics [1818]. Hidden [1172]. Hiebert [1297]. Hieroglyph [610]. Hieroglyphic [1031, 1033]. Hieroglyphics [774]. Higgins [1601, 1290]. High [1581]. High-stakes [1581]. Highlighting [1255]. Hill [71]. Hindu [1453]. Hire [1157]. Hispanic [1032]. Historical [1761, 1718, 1085, 1084, 286, 1253]. Histories [1688, 1790, 1164, 1157, 1695, 1311]. Historiographic [1154]. Historiography [1478, 1479]. History [1310, 1218, 1479, 1667, 1748, 1165, 1155, 1246, 1300, 1331, 1846, 691, 595, 1166, 443, 1172, 1747, 1538, 1378, 1744, 1217, 1390, 1123, 1277]. Hittite [689]. HIV [1702]. Hobby [987]. Hobby-Horse [987]. Hochuli [1298]. Hofstadter [780]. Holes [1498]. Holopoetry [1207]. Holt [263]. Homage [229]. homes [1780]. Homogenized [1517]. Homonymy [724]. Hong [1758, 1741]. honor [484]. Horizon [1263]. Horse [987, 1065]. Hospice [1706]. Hospital [1693]. Hospitalized [1661]. Hospitals [1767, 1709]. hosted [707]. House [153, 123]. House-style [123]. Hudson [1375, 1218]. Human [1050, 1841, 1295, 1093, 506, 801, 1041, 1172]. Humanist [331]. Humanistic [778]. Humanities [1717, 1687, 1718, 1714, 1720, 1716, 1715]. Humans [1750]. humans* [1841]. Humphries [218, 153, 489, 164]. Hundred [429]. Hungarian [984, 1354]. Hutchins [1095]. Hutt [312]. Hygiene [1703]. Hypertext [1201, 1240]. Hyphen [1697, 194, 1123, 1298]. Hyphenation [173, 65].

I. [83, 71]. IA [1646]. Ian [1645]. Iannuzzi [1375]. IBM [27]. Ibn [1589]. ICOGRADA [69]. ICOGRAPHIC [243]. Icon [633, 1605, 1494, 1657, 1830, 1678]. Iconic [1771, 1768, 1769]. Iconicity [867]. Iconography [1757]. Icons [727, 1705, 1779]. Id [139]. Ida [1074]. Idea [1718]. Identification [147, 102, 39]. Identifying [1776]. Identities [1798, 1375]. Identity [1428, 1172]. Idéogrammes [481]. Ideograms [24]. Ideology [1264, 1461]. II [88, 146, 18, 487]. III [129]. Ikon [425]. Illinois [116]. Illuminography [1758]. Illusion [1643]. Illustrated [1553, 1389, 1390, 1799, 879, 1375, 1552, 1554, 1557, 1645, 1217, 1310, 1344, 1391, 1461, 1218, 1366, 1462, 1377, 152, 1256, 1328, 1246, 1258, 1276, 1298, 1299, 1297, 1319, 1311, 1329, 1429, 1647, 1648, 1649, 1330, 1080, 1275]. Illustration [1008, 1572, 1573, 152, 1785, 1331, 1625]. Illustrations [457, 1006, 1556, 1571, 1599, 1600, 1601, 1621, 206, 702, 152, 1123, 1318, 1511, 1512, 1624, 1626, 1646, 1140, 287, 1623, 242, 1509]. illustrator [1331]. Image [1816, 1702, 1718, 1179, 1017, 1775, 1771, 1540, 1011, 1676, 1351, 1799, 587, 1571, 1766, 1349, 1536]. Imagery [1517, 1698]. Images [350, 1818, 729, 1772, 1007, 1716, 1358, 1329]. Imaginary [1824]. Imagination [1218, 387]. imaging [1025]. Imagining [1319]. Immersive [1598]. Immigrants [1428]. Impact [1264, 1461, 1808, 1676, 1060, 1710, 1796]. Impairment [1548]. Imperative [866]. Imperialism [990]. Implementation [688]. Implications [1791, 1759, 680, 100, 258]. Implicit [1104]. Important [215, 1698]. Impoverishment [1341]. Impressions [1671, 1677]. Imprint [632]. Impromptus [1635]. Improve [1668, 1815, 1709, 1342, 1834, 985]. Improvement [557]. Improves [1059]. Improving [904, 701, 1657]. Improvisation [1691]. In/visible [1615]. inappropriate [1815]. Inarticulable [1353]. inches [1344, 489, 275, 521, 242]. Incident [1644]. included [1330]. Inclusion [1799]. Inclusive [1836, 1842, 1243]. Indentation [240]. Independence [430]. Index [88, 129, 181, 222, 268, 316, 356, 401, 447, 460, 504, 543, 603, 656, 693, 744, 785, 827, 930, 964, 994, 1023, 1083, 1134, 1167, 1222, 1259, 1288, 1321, 1345, 1381, 1414, 1480, 1502, 1528, 1565, 1608, 1627, 1650, 173, 287]. Indexes [799]. India [1720]. Indian [1424]. Indiana [1375]. Indicate [194]. Indices [1174]. Indigenous [1594, 1730]. Individual [1657]. individuals [1805]. industrial [1311]. Inescapable [1295]. Infancy [973]. Infections [1815]. Inference [305]. Influence [1670, 1812, 1677]. infographics [1796]. Information [130, 144, 157, 167, 163, 200, 210, 223, 247, 265, 278, 292, 304, 317, 325, 337, 348, 357, 372, 382, 392, 402, 411, 424, 437, 448, 480, 502, 524, 554, 556, 565, 602, 606, 620, 643, 658, 672, 683, 704, 706, 722, 735, 748, 756, 770, 816, 787, 803, 807, 829, 861, 864, 881, 884, 897, 908, 922, 943, 953, 996, 1024, 1158, 1168, 1175, 1187, 1211, 1223, 1249, 1260, 1270, 1278, 1286, 1322, 1302, 1332, 1338, 1346, 1360, 1369, 1382, 1393, 1406, 1415, 1421, 1430, 1438, 1455]. Information [1464, 1472, 1481, 1490, 1496, 1503, 1513, 1519, 1527, 1545, 1558, 1564, 1574, 1591, 1602, 1609, 1628, 1651, 1658, 1665, 1679, 1689, 1700, 1732, 1753, 1762, 1580, 298, 1604, 1688, 1411, 741, 284, 518, 1343, 664, 1100, 1372, 1722, 854, 1511, 584, 1419, 607, 1213, 1750, 773, 514, 536, 1807, 1695, 1796, 1786, 1711, 1788, 1599]. Informative [1782]. Informing [1781]. Initial [3, 135, 168]. Initiative [1833]. Initiatives [1840]. Inner [867, 1066]. Innis [917]. Innovation [1190]. Innovations [843]. Innuendo [1139]. Input [1661, 1806]. Inscription [374, 1008, 160]. Inscriptions [1004, 329, 1003]. Installation [385]. Instant [986]. Institute [1159, 778]. Instruction [949, 1432, 947, 664, 1215, 359, 710]. Instructional [923, 661]. Instructions [1521, 1729]. Integration [1676, 936]. Integrations [1623]. Intellect [1571, 1623]. Intellectual [1838, 1836]. Interact [1830]. Interaction [1551, 1576, 1593, 1499, 1548, 1511, 1512, 224, 1772, 1243, 1830]. Interactions [1718, 1604, 1386]. Interactive [1618, 1684, 1410, 1097, 1242, 1201, 1089, 1425, 1386, 1705]. Interactivity [1181]. Interdependencies [594]. Interdisciplinary [1832, 1833]. Interest [1676, 1409, 1396]. Interests [976]. Interface [1050, 1440, 1523, 868, 1505, 1366]. Interfaces [1719]. Intermedia [1089]. Internalization [440]. International [243, 239, 1598, 778, 1429, 1275, 1025, 83, 1349, 1311]. Internationale [1429]. Internet [1622, 1846, 1461, 1517]. Interpretation [1559, 1818, 1795, 750, 1117]. Interpretations [1050, 911]. Interpreting [914, 1172]. Interrelationships [739]. Interrogating [1820]. Interrogation [635]. Interscriptual [1055]. Intersection [1809]. Intertextual [1185]. Intertitle [1325]. Interval [1806]. Interview [1192, 1704, 1184, 1180, 1182, 764, 1178]. Intonation [1058, 344]. Intrigue [1826]. Introduction [831, 1177, 969, 1688, 1474, 1003, 1063, 1200, 749, 351, 945, 910, 1280, 1323, 1333, 1423, 1442, 1629, 1126, 872, 674, 287, 898, 1839, 438, 644, 566, 796, 1819, 1695, 1648]. Introductions [709, 1485, 708]. Introductory [281]. Invades [1508]. Invented [1040]. Invention [31, 559, 309]. Investigating [1671, 802]. Investigation [591, 131, 590, 1293, 15]. Investigation/Hand [591]. Investments [1580]. Invisibility [1611, 1362]. Invisible [636]. Iowa [275, 1646]. Irish [1822]. Isabel [1688, 1695]. ISBN [1375, 1551, 1552, 1553, 1554, 1555, 1556, 1571, 1572, 1573, 1599, 1600, 1601, 1621, 1622, 1645, 420, 457, 1142, 1257, 1310, 1344, 1391, 1461, 489, 1366, 1462, 1377, 1389, 1256, 1378, 1122, 1328, 1390, 1123, 1141, 1246, 1245, 1258, 1277, 1276, 1296, 1298, 1299, 1318, 1297, 1300, 1319, 1311, 1329, 1331, 1379, 1429, 1509, 1510, 1511, 1512, 1625, 1624, 1626, 1646, 1648, 1649, 1846, 1376, 1623, 1275, 407]. ISBN-13 [1551, 1846]. Island [1140, 1581]. Ismar [1792]. Isn’t [1584]. Isolate [1059]. Isolated [823]. Isometric [582]. Isotype [1549, 1755, 1777, 1696, 1697]. ISSN [1330]. Issue [269, 1559, 1576, 1663, 1761, 817, 872, 808, 438, 566, 796, 1368, 140, 910]. Issues [862, 944, 1036, 1383, 1575, 1273, 976, 870, 895, 907, 919, 931, 1613, 1786]. Italian [287, 924, 1129, 406]. Italic [1014, 114]. Italicization [1524]. Italy [817]. items [287]. Ithaca [1329]. IV [181]. ix [152, 401]. Izenour [1434].

J [31, 1556, 1600, 457, 1258, 1299, 71, 1275]. J. [1192, 1191]. Jackhammers [1109]. James [1329, 1021]. Jameson [1185]. Jan [104, 1366, 407, 152, 1216, 407]. Janet [1509]. Janson [837]. Japan [1056, 1804, 1493]. Japanese [1508, 809, 847, 212, 1374, 1426, 937, 813, 808, 1197, 590, 38, 810, 812, 811]. Jasper [1009]. Jaspert [242]. Jaume [1461]. Jay [1512]. Jean [1389, 1325]. Jean-Luc [1325]. Jean-Pierre [1389]. Jed [1319]. Jeff [1555]. Jem [1326]. Jerome [1141]. Jersey [1141]. Jerusalem [778]. Johann [559, 892]. Johanna [1256, 1218]. John [1375, 263, 83, 1390, 1429, 420, 1457, 1510, 1087]. Johns [1009]. Johnson [242]. Johnstone [218]. jointly [778]. Jonas [1511]. Jorge [1276, 1672]. Josef [1726]. Joseph [56]. Jost [1298]. Journal [22, 32, 45, 54, 122, 192, 513, 996, 1158, 1168, 1175, 1187, 1211, 1223, 1249, 1260, 1270, 1278, 1322, 1302, 1332, 1338, 1346, 1360, 1369, 1382, 1393, 1406, 1415, 1421, 1430, 1438, 1455, 1464, 1472, 1481, 1490, 1496, 1503, 1513, 1519, 1527, 1545, 1558, 1564, 1574, 1591, 1602, 1609, 1628, 1651, 1658, 1665, 1679, 1689, 1700, 1732, 1753, 1762, 46, 58, 74, 87, 95, 106, 117, 127, 141, 155, 165, 179, 191, 198, 208, 220, 231, 245, 255, 266, 276, 289, 302, 314, 323, 335, 354, 370, 380, 390, 399]. Journal [408, 422, 435, 445, 458, 476, 491, 501, 541, 552, 599, 619, 626, 640, 653, 670, 140, 801, 346, 562, 577, 182, 1296, 97]. Journalism [94]. Journals [1725, 1660]. judge [636]. Judging [1820]. Judgment [414]. Judgments [753, 1677]. Judith [1655]. Jules [1010]. Jumper [741]. June [1178]. Just [1667]. Justice [1647]. Jutta [1375].

Kaplan [1846]. Kapr [242]. Kaptelinin [1551]. Katakana [590]. Kaufmann [1572, 1461]. Kearney [1843]. Keere [1810]. Keiichi [1623]. Keith [1699]. Ken [1318, 1217]. Kenneth [1297]. Kevin [1375, 1390]. Killer [634]. Kilpatrick [457]. Kinbote [1114]. Kindel [1697]. Kindersley [489, 836]. Kinetic [1808]. Kingdom [1378, 1328, 1318]. Kinross [1123, 1298]. Kiosk [1258, 1299]. Kirkcaldy [1378]. Kirschenbaum [1010]. Kis [984, 837]. Kit [430]. Kitchen [1308]. Klaassen [1626]. Klee [24]. Klonarides [1184]. Knacks [645]. Knoll [1257, 1344]. Knoop [1275]. Know [564, 537, 1039]. Knowing [507]. Knowledge [1670, 822, 1845, 819, 1749, 1730, 1744]. Knowles [1090]. Knuth [737]. Kombinatorik [474]. Kong [1758, 1741]. Korea [1056]. Korean [1426, 1324]. Kosslyn [1698]. Kostelanetz [1320]. Kotz [1557]. Kunst [242]. Kurzfassung [346, 562, 577]. Kuspit [1182].

L [104, 71, 1698]. Lab [331]. Label [1833, 1786]. Labeling [26]. Laboratory [707]. Labyrinth [1218, 889]. Labyrinths [888]. Lady [1378]. Lambert [1388]. Landscape [1343]. Language [31, 1375, 1557, 1751, 1752, 1531, 395, 798, 1461, 1050, 899, 108, 496, 235, 1822, 873, 726, 250, 264, 739, 279, 887, 339, 228, 730, 532, 632, 10, 18, 1440, 509, 970, 388, 367, 487, 1548, 251, 1141, 1319, 1624, 1654, 224, 1710, 1812, 1216, 582, 271, 120, 867, 586, 1371, 609, 796, 182, 403, 1506, 1353, 261, 453, 365, 1758, 1741, 1814, 272, 434, 488, 525, 555, 629, 742, 805, 882, 261, 291, 269, 870, 895, 907, 919, 931, 920, 1036, 1368, 1528, 1575, 1663, 1742, 1751, 1752, 1761, 1367, 1748, 976]. Language [1429, 1080, 182, 1747]. Languages [1108, 282, 939, 463, 934, 188, 1424]. Lanham [1599]. Large [1554, 1140, 1556, 1080]. last [1743]. late [613, 1030]. lateinischen [242]. Latencies [341]. Latin [202, 1805]. Latin-script [1805]. laude [559]. Laurel [1122]. Lavender [1389]. Law [621]. Laws [899, 1510, 1553]. Layout [1656, 853, 1677, 1342, 17, 661]. Layouts [1343]. Leading [134, 1055]. Leading-edge [1055]. Leaflets [1343]. Learn [508, 1634]. Learned [834, 1590]. Learning [1227, 1317, 1604, 1684, 263, 982, 442, 545, 509, 663, 1475, 1837, 299, 1436, 1707, 1434]. learnings [1710]. Learns [482]. Lebens [893]. Lebens-Früchte [893]. led [1670, 1771, 1768, 1769]. Leeds [31]. Left [1452]. Left-handedness [1452]. Legacy [1479, 1483, 1777]. Legate [517]. Léger [1008]. Legibility [137, 124, 99, 149, 1759, 70, 520, 1315, 308, 1014, 416, 1797, 588, 93, 587, 701, 67, 1802, 1727, 1649]. Legible [1365]. L’Empire [495]. Length [1274, 777, 518, 791, 1467, 341, 1728]. Lengths [928]. Leroy [93]. Lessons [834, 1581, 985]. Let [546]. Letter [529, 493, 489, 1367, 877, 202, 100, 425, 76, 1811, 500, 359, 648]. Letterform [1738, 98, 1802]. Letterforms [113, 213, 343, 171, 404, 363, 174, 270, 297, 79, 833]. Lettering [342, 152, 206, 152]. Letterisme [766, 759, 761, 764]. Letterist [763]. Letterists [765]. Letterpress [1669, 1457]. Letters [533, 1218, 1470, 1224, 375, 532, 10, 18, 34, 1299, 1649, 607, 638, 1143, 454, 1710, 1735, 152, 536]. level [662, 1458]. Levels [649]. Leveraging [1636]. Lex [633]. Lexical [821, 822]. Leyden [841]. L’Histoire [876]. Librarian [1076]. Library [457, 1257, 1310, 1344, 799, 56, 152, 1181]. Lies [1102, 1143]. Life [1391, 1191, 1510]. Ligature [132]. Light [1180]. Like [1532, 1541, 1086]. Like-minded [1086]. Liminal [1186]. Limit [1162]. Limitations [764]. Limited [31, 1217, 1065, 1624]. Limits [1371]. Lincoln [93]. Lincoln/MITRE [93]. Lindegren [152]. Line [877, 8, 5, 1274, 518, 791, 385, 928, 1467, 1728, 1498, 1294, 206]. Lines [1770, 296, 1002, 1475]. Ling [1461]. Linguistic [193, 388, 1804]. Linguistics [82, 901, 1060]. Linking [1694]. Lip [971]. Lipread [972]. Lipreading [970]. Lips [974]. Lissitsky [979]. Lissitzky [77]. List [68, 1404, 1401, 1399, 1403, 25]. Listening [510, 505]. Lists [1342]. Literacy [1392, 327, 515, 1056, 282, 700, 977, 950, 168, 388, 865, 914, 1560, 707, 691, 712, 710, 912, 1828, 1805]. Literary [1056, 1206, 866, 1149, 1183, 1107, 719]. Literate [1228]. Literature [1088, 1488, 1231, 1107, 1467]. Literatures [449]. live [1788]. Living [1181]. Liz [1557]. Loanwords [936]. Localization [1702]. Locating [520, 1720]. Logical [875]. logos [1375]. Logotypes [1284]. London [1375, 218, 206, 312, 334, 420, 1257, 1344, 153, 702, 489, 164, 263, 152, 83, 1697, 1123, 1276, 1298, 1379, 1648, 242]. Long [1171]. Longest [1811]. Look [1669, 610, 1188, 1711]. Looked [1557]. Looking [1154, 1669, 561, 1444]. Lorraine [1647]. Losing [1777]. Loss [1239]. Lost [1055, 1533, 1326]. Lothian [457]. Loud [980, 979]. Love [532]. Loved [474]. Low [104, 1740]. Lower [529, 1794]. Lower-Case [529]. Löwgren [1511]. Ltd [312, 702, 489]. Luc [1325]. Lucus [1626]. Ludwig [1275]. LUMINARIO. [287]. Luminous [1177]. Lund [218, 153, 489, 164]. Lupton [1645, 1258]. Lyon [842].

M [1142, 1367, 275, 1300, 521, 71, 1698]. M. [1245]. M.I.T [83]. MA [1551, 1552, 1553, 1557, 1600, 1601, 1688, 1510, 1512, 1625, 1647]. Maastricht [1366]. Machinations [1202]. Machine [1841]. Machines [25, 878, 49]. Macmillan [1320, 1555]. Macrotypography [1812]. Made [136, 1069, 1819]. Maeda [1510]. Magazine [216, 83, 153]. Magritte [1011, 1007]. Maier [152]. Maine [319]. Majesty [164]. Major [1190, 1059].

Make [8]. Make-up [8]. makers [1819]. Makes [103]. Making [1719, 1687, 1344, 1713, 971, 1662, 1563, 1667, 279, 1773, 1720, 840, 986]. Mal [1072]. Malignant [1587]. Mall [1589]. Mallarmé [998, 533, 108]. Mana [1793]. Manage [1588]. Management [434, 488, 525, 555, 629, 742, 805, 882]. Mandala [1570]. Manguel [1246]. Manifesto [1359, 403]. Manipulating [1411]. Manipulation [1070, 589]. Manual [1539, 1804]. Manualium [559]. Manuscript [920, 1221, 1248]. Manuscripts [1035, 1062, 386, 320, 1757, 536]. Manutius [846]. many [206, 1766, 1140, 1080]. Maori [1793]. Map [1717, 1217, 1682]. Mapping [664, 1426, 1509]. Maps [1556, 1682, 1783]. Marathon [1646]. Marcel [564]. Margin [1019]. Marginalia [1469]. Margins [9]. Marie [1296, 521]. Mark [1388, 842, 846, 830, 844, 841, 835]. Markers [981]. Marketing [3]. Marks [1614, 92, 1613, 1616]. Marot [373]. Marschall [1275]. Marshall [114]. Marx [917]. Mary [1655]. Masks [1157]. Mass [83, 1276, 1811]. Massachusetts [1462]. Master [250, 1563]. matatuhi [1793]. Material [13, 820, 1791]. Materiality [1531, 1691]. Materials [284, 1785, 1727]. Mathematical [724, 726, 727, 730, 732]. Mathematics [1736, 731, 723]. Matrix [840, 416]. matter [479]. Matters [1703]. Maurice [527, 377]. Maxim [1301, 1429]. Maxwell [1556]. May [1226]. Maya [1501, 211, 1031, 1756, 1766]. Mayakovsky [979]. McCaffery [1319, 1843, 1066]. McGann [1141]. McGraw [71]. McGraw-Hill [71]. McLean [334, 153]. McLuhan [917, 114]. McMurtrie [1341]. MDN [242]. Me [1705, 1842, 1069]. Mean [851, 1842, 1494]. Meaning [1091, 472, 983, 473, 507, 1314, 462, 1657, 1830, 1779]. Meaningful [298]. Meaningfulness [414]. Means [426, 1646]. Meant [714]. Measurement [124]. Measuring [65]. mechanisms [1798]. Media [1375, 1791, 1374, 1410, 1194, 1676, 1379, 1419, 1208, 182, 1368, 1306]. Medial [388]. Mediated [798, 1612, 560]. Medical [1710]. medication [1729]. Medicines [1587]. Medieval [172]. Meditation [983]. Medium [99, 213, 1201, 802, 796, 1784]. Meets [1165, 1418]. Meirelles [1688, 1695]. Meisterbuch [152]. Melville [1646]. Members [68, 784]. Memoir [1292]. memoriam [784]. Memory [818, 927, 1240]. Men [236]. Mental [729, 1698]. Merald [932]. Meredith [1655]. Merror [618]. Mesoamerica [697, 1029, 1028]. Mesoamerican [1033, 1030]. Mesopotamia [1389, 1108]. Message [55]. Messy [1166]. Meta [737, 1714, 737, 695]. Meta-Font [737, 737, 695]. Metadiscourse [981]. Metafont [834]. Metalinguistic [789]. Metamathematics [737]. Metaphor [338, 1525]. Metaphoric [1524, 1765]. Metaphors [1522, 1521]. Metaphysics [737]. Metastructural [1425]. Method [61, 1079, 1274, 308, 974, 1525, 40, 1763, 1764]. Methodological [19, 1683]. Methodologists [1446]. Methodology [1410, 1768, 1643]. Methods [1401, 442, 1254, 1411, 843, 386, 1014, 1706, 1838, 1709]. Metonymic [1765]. Metzger [1553]. Mexican [1282, 1130]. Mexico [1335, 1032]. Michael [1310, 1331]. Michel [532]. Michele [1375]. Micrography [793]. Middlesex [1217]. Milan [1253]. Miland [287]. Mile [1581]. Miller [1258, 1299, 1652]. Mind [868]. minded [1086]. Minimal [1365]. Minneapolis [1645, 1509]. Minnesota [1509]. Mirage [1387]. Mirrors [1512]. Mis [1678]. Missing [1631]. Mistress [250]. MITRE [93]. Mixing [649]. Mixtec [1439, 1032]. MN [1645]. Mobile [1461, 1837]. Moby [1646]. Moby-Dick [1646]. Modality [981]. Model [972, 571, 505, 902, 540, 1242, 1181, 120, 1798]. Modelling [1829]. Models [905, 680, 1638, 431]. Modern [1547, 1614, 164, 838, 1010, 1123]. Modernism [1085, 1216, 1141]. Modernist [1341, 1171]. Modest [1190]. Modifications [451]. Molee [237]. Moll [637]. Mollerup [1731]. Moment [1303]. Monacelli [1311]. Monastery [172]. Monastic [559]. Money [1189, 1541, 1194, 1188]. Monika [1624]. Monks [204]. Mono [1112]. Monospaced [1121]. Morgan [1572, 1461, 56]. Morpho [938]. Morpho-graphic [938]. Morphological [821]. Morrow [1331]. Most [1698, 746]. Motion [1410, 1796]. Motivational [1829]. Motives [974]. motor [101]. MOTS [1330]. MOTS-VOIR [1330]. Mould [840]. Mountain [1142]. Mouth [975]. Movement [1274, 1524]. Movements [971, 675, 574, 568, 973, 567, 1306]. Moving [206, 1667, 800]. Mr. [1217, 719]. Much [1039]. Muller [312, 420]. Multi [1693, 662, 540, 1458]. Multi-Context [540]. Multi-Hospital [1693]. Multi-level [662, 1458]. Multilingual [1114, 1301]. Multime [1357]. Multimedia [1072, 1243, 1412]. Multimodal [1499, 1825]. Multimodality [1756]. Multinlingualism [609]. Multiple [1580, 1830]. Multiplicities [1362]. Multisensory [1694]. MultiType98 [1308]. Multiyear [1833]. Mundurukú [185]. Murmure [926]. Musée [1075]. Museum [1684, 152, 1783]. Music [250, 339, 1485, 666]. Must [1015]. Mutamathil [1470, 1454]. My [1641]. Myth [959, 1084, 1512]. Mythical [1282].

Nabokov [1114]. Nachtwey [1375]. NAEP [949]. Nairn [1378]. Nam [1178]. Name [697]. Names [202, 359]. Naomi [1379]. Napier [457]. Nardi [1551]. Narrative [1155]. Narratives [913, 1172]. Narrativity [1017]. National [163, 1444, 1285]. Native [585]. Natural [1534, 1538]. Nature [914]. Navigating [1595]. Navigation [1683]. Neat [1166]. Necessary [1612]. Need [1420]. Needs [1252, 710, 98]. Neighborhoods [1686]. neon [1758]. Nerone [1375, 1390]. Nested [1105]. Netherlands [123, 287]. Netinguistics [1461]. Network [397]. Networking [1181]. Networks [1509]. Neurath [1102, 1549, 1777]. Neutrality [1549]. Never [1698]. News [35, 1149, 1390, 1375]. Newspaper [51]. newspapers [1741]. Next [1838]. Nicholas [1622, 984]. Nicolette [206]. Nieuwkoop [287]. Nieuwkoop/Netherlands [287]. Nieuwsblad [44]. Nine [1613, 1630]. Nine-year-old [1613]. nineteenth [1791]. nineteenth-century [1791]. NL [1626, 1649]. No [613, 83, 1318, 1846, 875]. Nodier [876]. Noise [587]. Noise/Legibility [587]. Nomenclature [404, 1042]. Non [1266, 519, 1517, 1073]. Non-Book [1073]. Non-Proficient [519]. Non-Profit [1517]. Non-Standard [1266]. Nonverbal [1494]. Normal [680, 1724]. Norton [1621, 1622, 1376, 1435]. Nostrand [1080]. Not-So-Easy [1262]. Notation [666]. Notations [1046]. Note [1645, 1085, 1289, 665, 673, 1235, 1253, 1, 421, 736, 1666, 544]. Note-Taking [665]. Notes [477, 478, 600, 920, 954, 1035, 1061, 1062, 1221, 1248, 1544, 1681, 583, 1094, 1076, 443, 1448, 457]. Nothing [496]. Nouveau [83]. Novelist [648]. Ntaria [1787]. Number [1755, 1]. Number-Fact [1755]. Numbers [97, 192, 635, 467, 544]. numeracy [1828]. Numerals [416, 790, 1453]. Numerical [687]. Nuremberg [830]. NY [1621, 1622, 1846].

O [109]. Oak [1257, 1344]. Oaxaca [1032]. Object [1071, 1177]. Objective [308]. Objects [1552, 1773]. Observations [463, 1475]. Observers [414]. Obtainable [1378]. Occasional [1832]. occidentaux [481]. Ockerse [1140]. OCR [14]. OCR-B [14]. October [1330]. Ode [300, 1006]. Off [269]. Office [164]. Official [1194]. Oh [110]. old [1613, 251]. Older [1386]. Olds [238]. Olmec [1029]. Olmec-style [1029]. On-Line [8]. One [793, 1822, 257, 76, 429, 1375, 1366, 1377, 1328, 1390, 1276, 1297, 1329, 1330, 99]. Online [1241, 1781, 1604, 1461]. Only [1537]. Ontology [634]. Open [1091, 345, 1841, 1779, 780, 1626]. Open-access [1779]. Open-Ended [780]. Opening [1153, 745]. Operational [1252]. opinions [1805]. Opportunities [1789]. Optical [14, 6, 1643, 489]. Optimal [1467, 586]. Optophonetic [1357]. Oracle [1501]. Oral [1015, 999]. Orality [1146]. Order [61, 1005]. Ordinary [1435]. Organization [1644, 52, 1598]. Organizations [1517]. Oriental [1004]. Orientation [345]. Oriented [378]. Origin [1282, 275, 224]. Original [1792, 273]. Originally [152]. Originals [484]. Origins [1816, 250, 160, 691, 710]. orphan [1738]. Ortho [938]. Orthographic [326, 1015, 1057, 1739, 819, 396, 440, 237, 1794]. Orthography [185, 622, 811]. Osley [702, 287]. Other [738, 1295, 887, 1194, 1616]. Otomi [1757]. Otto [152, 1275, 1102, 1549, 1777]. Our [1555, 1622, 329]. Outcomes [1692]. Overcome [1705]. Overlapping [1147]. Overview [1397, 1477, 1028]. Ovink [784]. Own [761]. Ownership [434, 488, 525, 555, 629, 742, 805, 882]. Oxford [206, 1461, 153].

P [1375, 457, 1275, 242]. Padova [817]. Page [112, 923, 201, 853, 195, 484, 51]. Page-design [201]. pager [1298]. Pages [253, 1375, 1551, 1552, 1553, 1554, 1555, 1556, 1557, 1566, 1571, 1572, 1573, 1599, 1600, 1601, 1621, 1622, 1645, 218, 206, 457, 1257, 1344, 1391, 1461, 984, 1218, 56, 263, 1366, 1462, 1377, 275, 1389, 152, 1256, 83, 1122, 1328, 1390, 1123, 1246, 1245, 1258, 1277, 1276, 1296, 1299, 1318, 1297, 1300, 1319, 1311, 1301, 1329, 1331, 1379, 1429, 1509, 1510, 1511, 1512, 1625, 1624, 1626, 1646, 1647, 1648, 1649, 1330, 1080, 1376, 1623, 71, 1275, 242]. Paik [1178]. Pain [1661]. paint [1756]. Painted [1009]. Painter [1010]. Painting [875, 1005]. Paintings [1198]. Palimpsest [1561]. Palindromes [418]. Paper [1266, 1582, 92, 800, 1193, 1453, 263, 611, 83, 1784, 1123, 1141, 287, 71]. Paperback [1601, 1645, 1310, 1646, 1648, 1375, 1257]. Paperbacks [206].

Paperbound [1599, 56]. Papers [1025, 1368, 1384, 1392, 1559, 1663, 1687, 681, 1751, 1752, 1761]. Paradigm [1476, 1797]. Paragraph [518]. Paralinguistics [1418]. Parallel [506]. Parameters [575]. Paranoia [1485]. Pardoe [420]. Parrinder [1624]. Part [750, 1635, 99, 119, 146, 10, 18]. Parthian [270]. Partially [560]. Participatory [1596, 1840, 1709]. Particular [1637]. Partnerships [1707]. parts [746]. Party [94]. Passenger [378]. Passenger/Pedestrian [378]. Past [329, 1561, 1721]. Pastoral [892]. Pat [1391]. Path [966, 1357]. Paths [888]. Patient [1343]. patients [1729, 1709]. Pattern [887, 891]. Patterned [890, 665]. Patterns [1274, 440, 1621]. Paul [1621, 1625, 484, 481, 24]. Pedagogy [1846]. Pedestrian [378]. Pen [572, 550, 548]. pence [206]. Penguin [71]. penmanship [1219, 218]. Penrose [60, 126]. Pens [549]. Pentagram [1311, 1311]. People [564, 1661, 1838, 537, 1090, 1704, 1836, 1785]. Perception [1619, 226, 970, 1844, 1245, 215, 973, 71, 1802]. Perceptions [1109, 1686]. Perennial [1637]. Performance [326, 131, 1074, 575, 1779, 1352]. Period [1029, 1810, 1757]. Periods [1612]. Peripherals [164]. Perishable [1065]. Persian [1809]. Persistent [1832]. Personal [1161, 1781]. Perspective [825, 1771, 1372, 1511, 1828]. Perspectives [1402, 1400, 6, 1447, 1495, 1398]. Peso [1285]. Peter [1600, 1509, 1626, 719]. Peypus [830]. Pharmaceutical [1797]. PhD [1159]. Phenomenological [294]. Phenomenon [751]. Phillips [164]. Philosophical [152]. Philosophy [1399, 1258, 1516, 1514, 1398, 1018]. Phoenix [837]. Phone [1461]. Phoneme [238]. Phonemes [506]. Phonemic [696]. Phonological [326, 441]. Photo [171, 1294]. photo-text [1294]. Photo-typography [171]. Photocomposition [78]. Photograph [788, 332]. Photographic [1776, 363]. Photography [1776]. photos [1080]. Phrases [1059]. pica [417]. Pictogram [1838]. Pictograms [1816, 1729, 1838, 1797, 1763, 1836]. Pictographs [24]. Pictoral [1032]. Pictorial [1833, 1758]. Picture [892, 1532, 1198, 1103]. Picture-Rhymes [892]. Pictures [1119, 1372, 1755, 1493, 66, 1782]. Pieces [1262]. Pien [1533]. Pierpont [56]. Pierre [1389]. Pig [439]. Pintaderas [1283]. Pioneer [1726]. Pioneers [164]. Piotr [1375]. Place [158, 1445, 1327, 1846]. Places [1776]. Plain [899, 901, 898, 903, 900]. Plana [1461]. Plane [1532]. planning [853]. Plant [903]. Plates [1765, 218, 984]. platform [1781]. Play [1461, 1487]. Plays [1570]. Please [985]. pleasure [985]. Plot [1618]. plus [287]. Pnin [1114]. Poem [997, 886, 891, 1001, 385, 889, 1357, 234, 415]. Poems [418, 1380]. Poetic [1202, 1070, 1336, 528, 867]. poetical [1349]. Poetics [1156, 1355, 1340]. Poetrait [874]. Poetry [1131, 338, 1691, 978, 373, 1356, 1350, 1130, 939, 426, 1128, 1205, 887, 1314, 1212, 351, 419, 1843, 1127, 1355, 83, 1230, 1234, 1129, 1132, 1126, 1133, 1351, 1357, 225, 1354, 1208, 28, 874, 66, 83]. Poetry-film [1212]. Poets [481, 878]. Poggenpohl [1623]. Point [1532, 832, 759]. Pointer [1562]. points [140]. Poland [1375]. Policy [122, 327, 609]. Polish [1375]. Political [1161].

Politics [1139, 962, 1113]. Poly [1352]. Poly-performance [1352]. Polyglotta [1301]. Polyphonic [250]. Polypoetry [1359, 1352, 1358]. Polyscriptal [1056]. Pompidou [1075]. Pompino [1275]. Pompino-Marschall [1275]. Pond [319]. Ponge [528]. Poor [927]. Popular [986]. Population [1686]. portas} [1508]. Portrait [1343, 874]. Portuguese [1131, 888]. Posed [559]. Positioning [1811]. Possible [506, 753]. Post [1149]. Post-Literary [1149]. Postage [609]. Posteguillo [1461]. posters [1782]. Postlinguistic [1691]. Postmaster [517]. Postmodernism [1229]. Potential [1101]. Potentialities [1206]. pouitry [352]. Power [1606]. pp [31, 564, 153, 702, 489, 164, 83, 287]. pp. [521]. Practical [5, 463, 6]. Practice [1403, 1692, 1402, 1106, 662, 1771, 1445, 1815, 580, 1258, 1560, 1768, 1769, 1162, 1492, 1778, 1730, 1298]. Practice-led [1771, 1768, 1769]. Practices [1702, 1161, 1688, 237, 1724, 1695, 1715, 1788]. Practicing [1443]. Praxis [948]. Pre [1283, 1032]. Pre-Columbian [1283]. Pre-Hispanic [1032]. Precedents [1084]. Precursors [530]. Predicting [431]. Preface [956, 1348, 1396]. Prefatory [1]. Preference [284, 574]. Preferences [134, 519, 1604, 661]. Preliminary [3, 1042, 1814, 1706]. Premises [1772]. Preparation [920, 1035, 1062, 1221, 1248]. Prereading [101]. Prerequisite [958]. Prescription [1701]. Prescriptive [901]. Present [710]. Present-day [710]. Presentation [801, 1213, 163]. Presentations [1675]. Preserving [1324]. Press [291, 844, 841, 1375, 1551, 1552, 1553, 1554, 1557, 1573, 1599, 1601, 116, 206, 1142, 1257, 1344, 1391, 153, 984, 56, 263, 1462, 1377, 275, 1389, 152, 1065, 83, 1697, 1328, 1390, 1123, 1141, 1245, 1277, 1296, 1299, 1297, 1300, 1319, 1311, 1329, 1509, 1510, 1511, 1512, 1625, 1646, 1647, 878, 242, 1846, 1092]. Preventing [1386]. Prevention [1702]. Price [287, 242]. primer [1750]. Priming [972]. Princeton [1391, 1141, 1245, 1299]. Principle [1269]. Principles [1401, 1694, 1788]. Print [1532, 507, 1594, 138, 215, 17, 29, 1518, 870, 895, 907, 919, 931, 56, 1555]. Printed [31, 1059, 539, 344, 1101, 1328, 195, 457, 91, 1140, 1793]. Printer [984, 333, 842, 846, 830, 844, 841, 835]. Printing [31, 457, 1310, 104, 489, 559, 1048, 152, 419, 286, 309, 878, 4, 1794, 334, 1378]. privately [457]. Problem [99, 718, 226, 726, 650]. Problems [19, 741, 1148, 1155, 691, 1780, 900, 789, 1137, 1613]. proceedings [707]. Process [1397, 145, 676, 1818, 750, 560, 989, 1829, 1750]. Processes [1776, 817, 1835, 912, 1711]. Processing [163, 677, 1739, 510, 1372, 681]. produce [1788]. Produced [152, 466]. Product [1579]. Production [478, 1645, 843]. Proficient [519]. Profit [1517]. Program [260, 65, 904, 1703]. Programmed [664]. Programs [1598]. Progress [81, 29]. Project [597, 1715, 1387]. Projections [1823, 1524]. Projects [1535, 1583]. Promote [1702]. Promoting [898]. Prompt [1837]. Pronounce [431]. Pronunciation [519, 441]. Proof [1121]. Proof-reading [1121]. Propaganda [1102]. Proper [393]. Properties [1799]. Prophetess [309]. Proportional [13]. Proportionally [1121]. Proposal [235, 1781].

Proposed [163, 110]. Proposition [343, 1198]. Prose [679, 518, 217]. prosodic [1814]. Prosody [1800, 396, 1814, 1813]. Prospect [947, 1505]. Protection [845]. Proto [1086]. Proto-Fluxus [1086]. Prototype [201]. Prototypes [1582]. Prototyping [1721]. Providence [1140]. Provocation [1432]. Provoke [1830]. Proximité [926]. Pseudo [1004]. Pseudo-Inscriptions [1004]. Psycholinguistic [263, 145, 168]. Psychological [70, 817, 71]. Psychology [1078, 1818, 71]. PTT [123]. Public [1776, 1508, 1460, 1662, 1229, 439, 1833, 471, 1459, 1796]. Publicaciones [1461]. publication [1320]. Publications [1378, 376]. Publish [291]. published [116, 152]. Publishers [1600, 489, 1688, 1695, 1122, 1626, 1649, 287]. Publishing [1556, 1217, 1344, 152, 114, 249, 1376]. Punch [984]. Punch-Cutter [984]. Punchcutting [839]. Punctuation [1612, 1611, 1615, 1616, 1610, 810]. Puns [1229]. Purpose [1085]. Pushing [1162].

Qualitative [1387, 1620]. Qualities [1844]. Quality [1693, 173, 753, 1784]. Quantifying [61]. Quarterly [243]. Quasi [1194]. Quasi-Official [1194]. Quest [1792]. Questioned [1041]. Questionnaire [1661]. quick [985]. Quillbot [1841]. Quipu [395]. Quotations [652]. Qwerty [280].

R [71, 1275]. R. [31]. Radial [361]. Radix [752]. Railway [1816]. Randolph [493]. Range [548]. raster [1025]. Rasula [1319]. Rates [536, 1286]. Rationale [664]. Rationalizing [1468]. Ravensburg [152]. Ray [7]. Re [1824, 162, 687, 1459, 564]. Re-assessment [162]. Re-design [1459]. Re-evaluation [687]. Reaction‘ [1314]. Read [507, 509, 1812, 1698, 66, 716, 263]. Read/Poetry [66]. Readability [1059, 327, 9, 13, 588, 29, 263]. Readable [879, 534]. Readable/Writable/Visible [534]. Reader [134, 51, 80, 464, 1808, 1671, 1803, 1022, 1676, 1451, 103, 753, 1082, 1677, 1318]. Reader-friendly [1451]. Readers [519, 946, 913, 1813, 15]. Readership [53]. Readiness [412]. Reading [483, 564, 941, 1694, 818, 326, 1015, 1641, 1801, 1800, 825, 980, 677, 799, 823, 196, 680, 296, 1274, 344, 50, 822, 263, 145, 257, 1739, 676, 1072, 486, 888, 43, 791, 752, 979, 1740, 510, 537, 1001, 560, 505, 819, 817, 970, 824, 608, 675, 1467, 140, 927, 1326, 674, 1498, 238, 1342, 1834, 49, 383, 974, 1835, 1173, 195, 359, 412, 681, 2, 172, 789, 1791, 1121, 136, 1824, 1784, 1813, 1727, 1034, 1246, 707, 273, 1649]. Readings [71, 71]. Real [1420, 992]. Realist [1578]. Realities [1597, 1832]. Reality [924]. Realization [1267]. Reappraisal [269]. Reappropriation [1822]. Reasoning [1389]. Rebuilding [1686]. Recall [52, 981, 1729, 1784]. Received [72, 178, 1081, 1144, 1220, 1247, 1375, 1463]. Reception [1341, 1387]. Receptivity [1806]. Recognition [294, 679, 822, 14, 364, 1365, 100, 540, 91, 6, 341, 674, 330, 203]. Recognizing [92, 1730, 1583]. Recollect [1146]. Recommendation [1833]. Reconnect [1686]. Reconsidered [540]. Recording [728]. Records [349, 455]. Recovery [1807]. recruitment [1710]. Redefining [1612]. Redesign [1460]. Reduce [1701]. reducing [1815]. Referee [802]. Reference [1476, 284, 1099, 240]. Reflecting [1747]. Reflection [472]. Reflections [1820, 1316, 1363, 1337, 1655, 879, 973, 1630]. Reflective [1660]. Reflexivity [1163]. reform [1794]. Regarding [1143]. Regional [457]. Regularized [19]. Regulation [1606]. Reinhold [152]. Reinvention [486]. Related [378, 820, 575]. Relating [1518]. Relation [1411, 610, 750, 973, 1814]. Relations [450, 1357, 1766]. Relationship [1085, 279, 1741]. Relationships [326, 1306]. Relative [93]. Relearning [1432]. Relevance [441]. Relevancy [1645]. Relevant [1636, 1664]. Reliable [1664]. Religion [915, 1389]. Remain [1626]. Remaining [900]. remarks [745]. Remembering [825, 1290]. Remembrances [1742]. Reminiscence [1094]. Remix [1722]. Renaissance [1004, 340, 386, 90, 551]. Renascence [749]. Repertory [1671]. Repetition [972]. Replies [249]. Reply [616, 614, 781, 1139, 615, 617, 1470]. Report [112, 949, 43, 1299, 98, 182, 233, 1308, 778]. Representation [1105, 1773, 1825, 1193, 696, 110, 1730]. Representational [1252]. Representations [1266, 1523, 1605]. Reprints [536]. Reproduction [1306, 1124]. request [1375]. Requirements [802]. Research [137, 394, 1034, 1383, 1401, 1694, 1692, 216, 1670, 1738, 1055, 1821, 1831, 1596, 1397, 345, 19, 344, 50, 81, 1604, 70, 520, 364, 843, 12, 284, 1078, 1400, 43, 386, 1771, 950, 1420, 1832, 976, 1409, 1396, 1416, 1655, 1768, 584, 1710, 1447, 1833, 1769, 217, 1623, 1780, 16, 71, 29, 40, 1275, 1620, 310, 98, 182, 789, 1137, 67, 1664, 1726, 1368, 1778, 1764, 1258, 45, 192]. researchers [1708]. Researching [766]. Resistance [1152, 1757]. Resonance [1117]. resources [1785]. Responding [1624]. Responds [1021]. Response [1471, 125, 51, 101, 1020, 780, 1367]. Responses [725, 1498]. Responsibilities [259]. Restall [457]. Restaurant [1334]. Restoration [363]. Résumé [513, 46, 58, 74, 87, 95, 106, 117, 127, 141, 155, 165, 179, 191, 198, 208, 220, 231, 245, 255, 266, 276, 289, 302, 314, 323, 335, 354, 370, 380, 390, 399, 408, 422, 435, 445, 458, 476, 491, 501, 541, 552, 599, 619, 626, 640, 653, 670]. Rethinking [1771, 1375]. Retrieval [1411, 284, 518]. Retrospect [947]. Revelation [1543]. Revelations [1660]. Reverie [1268]. Reversible [1314]. Review [31, 115, 1320, 1551, 1552, 1553, 1554, 1555, 1556, 1557, 1571, 1572, 1573, 1599, 1600, 1601, 1621, 1622, 1645, 1696, 116, 218, 206, 312, 334, 420, 457, 1142, 1257, 1310, 1344, 1391, 1461, 702, 984, 489, 1218, 56, 1366, 1688, 1695, 985, 1774, 1462, 1078, 1377, 275, 1389, 1256, 83, 30, 1378, 1697, 1467, 1122, 1328, 1390, 1123, 1141, 1150, 1246, 1245, 1258, 1277, 1276, 1296, 1298, 1299, 1318, 1297, 1300, 1319, 1311, 1301, 1329, 1331, 1379, 1429, 1509, 1510, 1511, 1512, 1625, 1624, 1626, 1646, 1647, 1648, 1649, 1846]. Review [1140, 1330, 1080, 287, 1376, 1623, 1018, 521, 1275, 1731, 407]. Reviewing [1779]. Reviews [21, 1217, 104, 153, 164, 263, 152, 83, 71, 1698, 242]. Revision [649]. Revival [838]. Revolt [915]. Revolution [36, 1647].

Rewriting [1624]. Reyner [1436]. Rhetoric [1284, 1549, 1280, 1285, 645]. Rhetorical [1282, 1281]. Rhode [1140]. Rhymes [892]. Rhythm [1811]. Rhythms [509]. Rich [1461, 1505]. Rich-Prospect [1505]. Richard [1320, 1375, 1599, 1377]. Riders [1141]. RIDT [1025]. RIDT-91 [1025]. Right [9]. Right-hand [9]. Righting [618]. Rimbaud [633]. Rinehart [263]. Ripples [1263]. Risk [1583]. Risk-of-failure [1583]. Road [1090]. Rob [1645]. Robbe [635, 1017, 1011]. Robbe-Grillet [635, 1017, 1011]. Robert [838]. Roberts [1556]. Robin [1123, 1298]. Roche [527, 377]. Rockport [1600, 1688, 1695]. Roel [1626]. Roger [1554]. Rogers [1300, 227]. Roi [876]. Roland [630, 639]. Role [1596, 1748, 188, 286, 1342, 359, 1657]. ROM [1330]. Roman [1367, 404, 152, 780, 613, 1304, 162, 1810, 937, 607]. Rosemary [1375, 1571]. Ross [1328]. Rotterdamsch [44]. Routledge [1375, 1379]. Roxane [1315]. Roy [1375]. Royal [153]. RTA [397]. Ruari [153, 334]. Rubbing [252]. Ruecker [1673]. Ruedi [1375]. Ruegg [1375]. Rulership [1029]. Rules [1816]. Rune [319]. Runs [761]. Runyon [493]. Ruscha [262]. Russell [83]. Russian [451, 119, 146]. Rypson [1375].

S [1034, 702, 498, 1192, 1379, 1330, 1191, 287, 1309, 497, 1330]. S/Z [497]. §45.00 [1297]. Saccadic [699]. Sadek [1301]. Safer [1702]. Safety [1579, 1729, 1782]. Said [714]. Salomon [1621]. Same [1778]. Samuel [877]. San [1461, 71]. Sang [1314, 1380]. Sans [451]. Sans-serif [451]. Sanskrit [891]. Santa [817]. Santiago [1461]. Sassoon [1375, 1571]. Satisfying [1213]. Sato [1623]. Saving [1119]. Sawhorse [987]. says [875]. Scales [466]. Scan [5]. Scandinavian [116]. Scanty [1547]. Scheerer [1275]. Schema [1526]. Schemes [1105]. Schirmer [1320]. Schirmer/Macmillan [1320]. Schleger [1391]. Schmandt [1573, 1331]. Schmandt-Besserat [1573, 1331]. Scholarship [1256, 1832, 1744]. Schønheit [242]. School [94, 707, 1140, 710, 1598]. Schools [946]. Schriff [152]. Schriftkunst. [242]. Schwartz [611]. Science [570, 1724]. Scientific [1670, 449, 71, 833, 1506]. Scores [928]. Scott [1434]. Screen [1274, 1413, 923, 800, 1307]. Screenfolds [1439]. Screenspace [1252]. Scribal [924, 193]. Scribe [625, 690]. Scribes [161, 702]. Scrip [1194]. Script [793, 188, 228, 691, 813, 1493, 1197, 810, 1805]. Scripting [852]. Scriptorum [559]. Scripts [1501, 318, 159]. Scrolling [699]. Scrutinized [1559]. Search [1383, 124, 149, 1055, 1174, 1157, 580]. Second [257, 509, 1025]. Secondary [10, 18]. Secret [288]. security [1796]. Seduction [1492]. See [1147]. Seeing [1439, 1106, 1610, 71, 1553]. Seen [66]. Segment [416]. Select [228]. Selected [1209, 765, 71, 407, 667]. Selectric [27]. Self [1267, 1781]. self-assessed [1781]. Self-Realization [1267]. Semantic [823, 258, 40]. semasiography [1757]. Seminar [249, 778]. Semiotic [1774, 406]. Semiotics [581, 1586]. Seniors [1505]. Sense [1749]. Sensibility [1096]. Sensitivity [1468]. Sentence [1136, 510, 1201, 469, 332]. Sentences [689]. sept [876]. Sequences [1775]. Serendipity [1725]. Series [1543, 1541, 1765]. serif [451, 275]. Serifs [1670, 215]. Seriously [1675]. Serres [531]. ses [876]. Setting [1499, 455]. Seven [1613, 416, 1810]. Seven- [1613]. Seven-Segment [416]. Seventh [1265]. Sex [1702]. Sexagesimal [224]. Sfumato [1532]. Shakespeare [1570, 271]. Shallows [1622]. Shape [373, 1299, 406]. Shapes [364, 375]. Sharon [1623]. Shave [410]. Shaw [1625, 176]. Sheep [1142]. Sheet [125]. Sherry [1552]. Short [927, 31]. Short-Hand [31]. Short-Term [927]. Should [1121]. Show [1619, 546, 1100, 1705, 1842]. Shows [1090]. Shrift} [119, 146]. Sides [975, 1092]. Sight [429]. Sighted [560]. Sigmund [892]. Sign [1822, 305, 592, 1101, 593, 1814]. Signage [1547, 1427]. Signature [1478]. Signatures [1485, 569, 1429]. Signes [495]. Significance [777]. signified [875]. Signifier [1353]. Signs [1792, 283, 1774, 610, 468, 494, 1437, 1757, 1758, 496]. Silent [196, 497]. Silhouette [206]. Silhouettes [1547]. Siloam [160]. Simplicity [1510]. Simplified [19]. Simplifying [366]. Simply [1233]. Simulated [1164]. Since [1575, 1344, 187, 813]. Sincere [464]. Singing [1358]. sitec.fr [1330]. Six [1262]. Sixteenth [702, 924, 104]. Sixteenth-Century [924, 104]. Size [1365, 1191]. Sizes [1810]. Skepticism [1435]. sketching [853]. Skill [566]. Skills [101, 1045, 1684, 547, 384]. Skirts [1547]. Skittish [1547]. Sky [1485]. Slash [776]. Slay [1584]. Slide [1675]. slipcase [1140]. Slips [572]. Sloppy [1256]. sm [287]. Small [215]. Smart [1462]. Smells [1485]. Smileys [1612]. Smith [1375, 263, 788]. So-called [687]. Social [1633, 1161, 1085, 426, 1428, 1276, 712, 1751]. Societies [1056]. Society [489, 342, 286, 1461]. Sociolinguistics [1197]. Sofie [1649]. Softbound [1571, 1572, 1621, 1624, 1626, 1623, 1375, 1461, 1366, 1276, 1296, 1299, 1297, 1300, 1301, 1509, 1512, 1140, 1330]. Software [1562, 1707, 1386]. SoHo [1094]. Solving [226, 726, 650]. Some [1004, 282, 1613, 611, 1372, 170, 168, 575, 973, 753, 796, 645, 1375, 1571, 1461, 152, 1258, 1429, 1648, 1275]. Something [83, 1092]. Son [31]. Soul‘ [425]. Sound [1694, 508, 1052, 507, 1355, 1351, 1357, 1349]. Soundings [1002]. Sounds [1350]. Soundscape [1354]. Soup [1173]. Sourcebook [1375]. Sources [702, 1297]. South [1335]. Soviet [1102]. Space [1508, 1013, 148, 1460, 1336, 307, 1843, 1326, 1337, 1323, 1333, 666, 1459]. Space/Book [1336]. Space/Poetic [1336]. spaced [1121]. Spaces [1153, 1616, 1789]. Spacing [1059, 489, 193, 13, 1498]. Spanish [513, 231, 245, 255, 266, 276, 354, 370, 435, 445, 458, 476, 491, 501, 541, 552, 599, 626, 640, 653, 670, 888, 562, 577]. Spatial [516, 345, 611, 1020, 853, 663]. Speak [1789, 1837]. Speaking [975, 414, 1117, 383]. Special [1036, 1559, 1575, 1576, 1663, 1687, 1761, 817, 872, 808, 1241, 438, 925, 566, 796, 1368, 910, 1805]. specimens [242]. Spectacle [1149]. speculation [1802]. Speculative [1714]. Speech [1614, 978, 971, 1613, 510, 970, 1147, 714, 974, 973, 452, 453]. Speed [136, 2, 1784]. Speed-reading [136]. Speedometer [1606]. Spell [507, 442]. Spellers [927]. Spelling [508, 947, 441, 696, 443, 925, 444, 1137]. Spellings [431]. Spencer [153, 164]. Spiekermann [1142]. Spiel [474]. Spiral [56]. SPIRALS [1140]. Spirit [319]. sponsored [778]. Spontaneous [452]. St [275]. Stage [505, 902]. Stains [779]. Stakeholder [1578]. stakes [1581]. Stamps [1283, 609]. Stan [1673]. Standard [1266, 65, 148, 168, 13, 102, 163]. Standardization [1838, 63]. Standardized [14]. Standards [5]. Stanley [71]. Starch [928]. starting [140]. State [1031, 845, 1798]. state-of-the-art [1798]. Statement [434, 488, 525, 555, 629, 742, 805, 882]. Statements [1486, 428]. States [1257, 1133, 1086]. Stationery [164]. Statistical [1693]. Statistics [1102, 1143]. Statu [875]. Stealing [1142]. Steiner [249]. Stele [630, 639]. Stem [341]. Step [1838]. Stephen [83, 1698]. Steps [203]. Stereo [1112, 1109]. Stereogrammic [450]. Steve [1319, 1648, 1843, 1066]. Steven [1296]. Stevens [361]. Stevin [1100]. Still [1683, 1049, 895, 907, 919, 931]. Stimuli [972, 1799, 274]. Stimulus [820]. Stolterman [1511]. Stone [236, 916]. Stonecuttings [836]. Stones [319]. Stop [1142]. Stopped [1641]. stories [1787]. Story [43, 1573]. Storyboard [1212]. Straightness [9]. Strassburg [1048]. Strategic [1684]. Strategies [1776, 1668, 226, 1773, 1084, 1252, 1208]. Strategy [624]. Stream [761]. Street [1090]. Strip [1433]. Strokes [1770]. Strøm [374]. Stroop [1055]. Structural [823, 1060]. Structure [809, 469, 819, 753]. Structures [1091, 821, 469, 731, 332]. Structuring [426]. Student [1761, 1424, 1660, 753]. Students [1842, 1620]. Studies [442, 26, 1110, 42, 1445, 1299, 566, 778]. Studio [1667, 152, 1293]. Study [338, 1682, 5, 1684, 473, 1570, 647, 1315, 1387, 1014, 405, 1300, 1301, 449, 1444, 1710, 1660, 855, 1741, 1815, 1705, 1779, 1814, 1812, 1780, 1727]. Studying [1834]. Style [1454, 1599, 1470, 1606, 927, 1113, 123, 1029]. Styles [1660, 1785]. Stylistics [538]. Subject [449]. Subjective [284]. Subjectivity [1244]. Subjugation [1011]. Subscription [122, 1286, 261, 536]. Subscriptions [1124]. Substance [1599]. Substitutions [483]. Subversion [1003]. Subversive [1076]. Subverting [1538]. Subway [311, 1625]. Success [1832]. Sudden [1537]. Sue [1697]. Sufficient [1630]. Suggestions [941, 1475]. Summation [298]. Sunday [413]. Sundial [470]. Super [762, 1652]. Super-Writing [762]. Supplementary [1353]. supportive [1781]. Supports [1800, 1813]. Surface [1170, 824, 731]. Surgery [1807]. Surrealist [350]. Surrey [1328]. Surrogate [1362]. Surveillance [1662]. Survey [597, 1667, 1758, 1798, 1793]. Suspension [1492]. Sussux [1318]. sustainability [1786]. Sustained [1833]. SW [1624]. Swallow [83]. Swann [1080]. Sweden [152]. Swiss [407]. switching [939]. Switzerland [1375]. syllabary [1804]. Syllable [623]. Symbol [1694, 592, 261, 1573, 1764]. Symbolism [724, 725, 730, 732]. Symbols [378, 1375, 260, 1433, 5, 727, 729, 1029, 1657, 1830]. Synonymy [724]. Syntactic [1117]. Syntax [990, 351]. Synthesis [974]. System [1733, 1196, 3, 688, 201, 224, 801, 937, 780, 808, 1745, 582, 1032, 17, 586, 1492, 853, 1625]. Systems [1656, 8, 19, 7, 506, 489, 1462].

T [242]. Tables [800]. Tablet [252]. Tablets [687]. Tabular [310]. Tacit [1845]. Tactics [624]. Taiwanese [1605]. Taken [1437]. Taking [665]. Tale [1282]. Tales [1016]. Tallone [229]. Tang [1554]. Target [635]. Targeted [1701]. Task [149]. Taxonomy [318]. Taylor [83, 1276]. Te [1118]. Tea [1374]. Teach [1694]. teacher [1805]. Teachers [1042]. Teaching [3, 1254, 545, 547, 135, 1655, 1475, 1495, 94]. Tech [1806]. Technical [775, 79]. Technique [1310]. Techniques [330, 1571]. Technological [843, 1803]. technologically [1454]. technologically-friendly [1454]. Technology [1551, 457, 1264, 327, 1350, 865, 1809, 1510, 132, 589, 1352, 1358, 1511]. Telephone [1255]. Television [1098, 1373, 1182, 93]. Tell [1826, 1498]. telling [1787]. Tells [1016, 1053]. Temple [690]. ten [261, 985]. ten-year [261]. Tendencies [1795, 833]. Tension [1066]. Tension/In [1066]. Tensions [1170]. Teotihuacan [1817]. Teresa [1389]. Term [1367, 1304, 927]. Terminal [868]. Terms [1600]. Territories [1154, 1509]. terror [1796]. Test [1652]. Tested [298, 442]. Testing [1677, 1779]. Texas [1573]. Text [483, 1692, 1071, 1508, 1820, 960, 1059, 175, 1070, 958, 1120, 966, 205, 518, 611, 1020, 923, 1810, 1111, 509, 1418, 959, 91, 1676, 962, 1117, 699, 1342, 1536, 753, 998, 717, 218, 1293, 1760, 661, 152, 1766, 1294, 1677, 1727]. Text-Based [753]. Textbooks [982, 663, 557, 701]. Texts [792, 585, 1108, 890, 1524, 1843, 914, 765, 1812, 913, 1756]. Textuality [1457]. Thames [1375, 1218]. Theater [1122]. Their [158, 1820, 680, 878, 1727]. Them [508, 546]. Theme [472]. Theoretical [765]. Theories [1694, 1688, 442, 1695]. Theory [1399, 1403, 1551, 1152, 677, 1177, 1402, 298, 1774, 12, 662, 750, 1165, 583, 580, 1416, 988, 986, 1417, 987, 1208, 1752, 1298]. there [1795]. Thesis [596, 241]. Things [1552, 1773, 1372, 1005, 1667]. Think [1552, 1846]. Thinkeringspace [1569]. Thinkers [1148]. Thinking [1572, 1227, 1684, 1364, 986, 1453, 1660, 1749]. Thirty [1292]. Thomas [1140]. Thompson [1698]. Thought [1201]. Thoughtful [1511]. Thoughts [1375, 170, 1419, 1042, 1475]. Thousand [793, 1038]. threat [1796]. Three [34, 1014, 691, 1495, 274, 2, 1783, 352, 1581]. three-dimensional [1783, 352]. Three-Year [1014]. Time [1013, 1718, 1523, 1180, 1584, 1185]. Time-Wall [1185]. Times [162]. Tiriti [1118]. Today [1631, 1355, 1646]. Together [1068]. Toilet [1427]. Tokens [1188, 687, 911]. Tom [152]. Tony [564]. Tool [1045, 279, 1704]. Tool-Making [279]. Tool-Using [279]. Tools [1462, 1099, 1706, 1787, 1705]. Top [1743]. Topical [753]. Torturing [1198]. Trace [1041]. Tracing [971]. Tracking [1806]. Tract [1815]. Tractatus [1198]. Tradition [965, 999]. Traffic [131]. Training [264, 1571, 15]. Trans [1707, 1084]. Trans-disciplinary [1707]. Trans-historical [1084]. Transcendence [108]. Transcending [1340]. Transcoding [938]. Transformation [1336]. Transformations [1212, 1114]. Transit [397]. Transition [983, 923, 1654]. Translated [1600]. Translation [858, 1110, 979, 1533, 586]. translator [1389]. transmission [1744]. Transmitter [385]. Transparency [1512]. Transport [1556, 1217].

Transportation [378]. Transportation-Related [378]. Transylvanian [837]. Treasury [152]. Treating [1815]. Treatises [31, 90]. Treatment [1799, 1815]. Treaty [1118]. Trends [216, 1033]. Tribulations [1547]. Tribute [932]. Tripitaka [1324]. Trithemius [559]. Trojan [1065]. trouble [746]. Troxler [1626]. True [1351]. Trying [508]. Tschichold [104, 152, 407, 1216]. Tube [7]. Turkle [1552]. Turned [1542]. Turner [1542]. Turning [1542, 832]. Twentieth [1375, 116, 1070, 1172, 1306]. Twentieth-Century [1070, 1172]. Twenty [1262]. Twenty-Six [1262]. Two [457, 1105, 1108, 975, 1092, 42, 1783, 1735, 582, 1298, 1140, 1780]. Two-Dimensional [582, 1783]. Twyman [1310]. TX [1573]. Type [1454, 1550, 1694, 1142, 1308, 124, 99, 149, 738, 1052, 1470, 133, 832, 404, 843, 119, 146, 1810, 1378, 121, 1429, 197, 1811, 780, 589, 49, 111, 1711, 186, 833, 242, 1376]. Typeface [134, 212, 1316, 845, 451, 1844, 1835, 558, 1253, 698, 1760, 1805]. Typefaces [1652, 1759, 1121, 1619, 135, 159, 39, 40, 242]. Typefoundry [1734]. Types [104, 838, 837, 274, 1454, 1810, 152]. Typeset [1121]. Typesetting [7, 25, 164, 1301]. Typewriters [797]. Typewritten [1121, 13]. Typographer [312]. Typographic [45, 192, 1554, 1692, 1792, 1831, 61, 175, 1070, 1413, 982, 1120, 12, 1020, 1642, 94, 258, 1157, 1671, 1677, 259, 102, 1424, 64, 16, 557, 701, 1727, 1784, 1711]. Typographical [1105, 741, 7, 663, 2, 457, 1343]. Typographics [1375]. Typographie [407]. Typographique [1429]. Typographis [1301]. Typographs [293]. Typography [1500, 892, 99, 36, 1174, 935, 281, 27, 1254, 1341, 82, 1272, 1794, 1808, 561, 94, 77, 300, 1123, 1299, 1016, 103, 1362, 1745, 1834, 1334, 234, 1376, 1842, 1453, 53, 660, 1253, 406, 407, 1025, 1743, 171, 1616, 368, 104, 164, 1080]. Typophiles [484]. Typos [212]. Tz’ihb’ [1756].

U [1275]. Ubac [926]. Ubis [499]. UK [1556, 1571, 1648, 1623]. Unconscious [564, 537, 1182]. Unconsidered [1138]. Uncovering [1410]. Underground [1217, 1556]. Underlining [240]. underlying [1715]. understand [1836]. Understanding [1170, 1562, 727, 1524, 1780, 310, 1778, 1678, 263, 1715]. Understandings [1613]. Understood [1204]. Unfamiliar [431]. Unicode [1429]. Union [1301]. unique [1788]. Unit [584]. United [1257, 1378, 1328, 1318, 1133, 1086]. Universal [716]. Universals [145]. Universe [1334]. Universitat [1461]. Universities [1832]. University [1375, 1554, 1573, 1599, 116, 206, 153, 263, 1377, 1389, 817, 1141, 1277, 1297, 1329, 1509, 1646]. Unjustified [44, 205]. Unpacking [1717]. Unpunctured [998]. UnReading [1469]. Unstable [1771]. Untitled [427]. Upcoming [1036, 1576, 976]. Upon [483, 1375]. Urban [1326]. Urbana [116]. Urinary [1815]. Uruguayan [1132]. Uruk [690]. USA [1846, 94]. Usable [1462]. Use [378, 905, 149, 170, 1522, 194, 1376, 111, 1837, 1805, 1275]. used [1785]. Useful [1670, 1788]. Useless [1612]. User [1050, 1682, 1570, 1440, 1445, 1419, 1053, 1521, 1749, 1704, 1788, 1764, 1276]. user-centered [1704, 1788, 1276]. user-drawn [1764]. User-Interface [1440]. Users [1605]. Uses [10, 18, 666]. Using [508, 1119, 279, 1671, 1705, 1686, 1342, 1780, 1706, 567, 1796, 1709]. Utopia [1231].

V [1297, 222]. Value [515, 1418, 1725]. Value-added [1418]. Values [78]. Vanishing [1532]. Variability [1197, 573]. Variables [2]. Variation [99]. Variations [925]. Various [899, 1299]. Vautier [1478]. VEB [242]. Vegas [1433, 1432, 299, 1435, 1434]. Veil [1500]. Venice [846]. Venturi [1434]. Verbal [373, 298, 1010, 979, 120]. Verlag [152, 1624, 242, 407]. Vernant [1389]. Vernon [71]. Verona [924]. Verse [1000]. Versions [1570]. Versus [13, 1826, 1264, 1112, 1784]. Vertical [1770, 607, 37]. Vervliet [104]. VI [268, 316]. Victor [1551]. Victorian [334]. Video [1177, 1183, 589, 1185]. Video-Art [1185]. Videopoetry [1203]. Vidifont [184]. View [1142, 99, 759, 281, 1166]. Viewdata [740, 588]. Viewing [1373]. Views [1447]. VIII [31, 356]. Viking [152, 1246]. Villon [373]. Vindobonensis [297]. Virtual [1244, 1205, 1597, 1299, 1706]. Virtuoso [1219]. Visibility [740, 1611]. Visible [453, 1751, 1752, 395, 153, 932, 971, 873, 250, 739, 1773, 1292, 532, 632, 1440, 970, 1141, 1654, 879, 534, 609, 796, 403, 1353, 261, 1615, 272, 434, 488, 525, 555, 629, 742, 805, 882, 261, 291, 269, 870, 895, 907, 919, 931, 920, 1036, 1368, 1528, 1575, 1663, 1742, 1751, 1752, 1761, 1367, 1748, 976, 182, 1747]. visiblelanguage [1250]. Vision [305, 1740]. Visionary [876]. visitors [1783]. Vista [152]. Visual [905, 1131, 1384, 1383, 1392, 1600, 1046, 792, 260, 1000, 1800, 1821, 1691, 997, 677, 1282, 1052, 983, 538, 8, 1119, 101, 1130, 463, 1315, 1823, 1790, 1148, 1128, 676, 481, 528, 1773, 689, 1818, 1795, 979, 583, 351, 509, 1798, 1843, 1466, 1522, 1127, 1129, 675, 1704, 251, 1132, 1188, 1245, 1281, 1280, 1423, 1126, 1133, 988, 674, 1813, 1768, 1581, 1357, 225, 607, 234, 974, 1386, 1750, 120, 1525, 71, 1521, 1568, 182, 1518, 1506, 1802, 1587, 1827, 261, 1766, 1785, 1708, 1814, 15, 1572]. Visual-motor [101]. Visualization [1458]. Visualizations [1618, 1695, 1688]. Visualizing [1693, 1570, 1796, 594, 1327]. Visually [4]. VL [1687]. Vocalism [196]. Voice [1266, 1350, 138, 1358]. Voicimage [1349]. VOIR [1330]. Volume [31, 126, 1134, 1167, 1222, 1259, 1288, 1321, 1345, 1528, 515, 1275, 1083, 206, 88, 129, 181, 222, 268, 316, 356, 401, 447, 504, 543, 603, 656, 693, 744, 785, 827, 930, 964, 994, 1023, 1381, 1414, 1480, 1502, 1565, 1608, 1627, 1650, 152]. Volumes [1275, 457, 460]. Volunteerism [1710]. Vostell [1098]. Voting [1138]. Vowel [519, 431]. vs [823, 1783, 779, 1166].

W [1621, 1622, 457, 1376, 1275, 242, 702]. Wake [1096]. Walker [1645, 1697]. Wall [1685, 1185]. Wallace [361]. Walter [1275, 1065]. Wanted [549, 1049]. War [456]. Ware [1572]. Warning [1833]. Warsaw [1375]. Was [1303]. wasn’t [613]. Way [1563, 1364, 1003, 1374, 561, 396, 1335]. Wayfinding [1816, 1682, 1683]. Ways [1228]. Wayshowing [1767]. Web [1419]. Web-based [1419]. Weber [893]. Webliography [1209]. Website [1250]. Wechel [835]. Weingarten [1275]. Welcome [1699]. Wesley [1122]. West [1389, 1318]. Western [891, 1620]. Whale [1646]. Where [1446, 1817, 1418, 1379, 1509]. Whetstone [374]. Which [333, 1001, 1783, 1294, 1498]. White [1843, 1646, 1376, 1551, 1552, 1553, 1557, 1573, 1599, 1601, 1622, 206, 1511, 1512, 1647, 1080, 1623, 1275]. Whiz [901]. Whiz-deletions [901]. Who [564, 537, 1040]. Whole [750, 1835]. Whole-Word [1835]. Whose [1147]. Wildman [83]. Willem [784]. William [83, 1698, 1034, 362, 1469]. Win [261]. Windows [1512]. Winston [263]. within [1499]. Without [508, 507, 660]. Wittenborn [152]. Wittgenstein [1198]. Wizard [139]. Wm [1331]. Wolf [1600, 985, 1098]. Wolfgang [1553]. Woman [1171]. Women [1166]. Wonders [1437]. Wooly [741]. Word [1227, 1015, 1537, 148, 1350, 679, 822, 226, 364, 777, 1336, 52, 1365, 1101, 540, 1011, 341, 455, 674, 321, 471, 590, 1835, 681, 1793, 1494, 153, 1150]. Word-By-Word [1015]. Word/Forms [321]. Worde [489]. Wording [701, 661]. Words [158, 456, 262, 822, 42, 52, 875, 431, 474, 1009, 1005, 194, 429, 1324, 1337, 1323, 1333, 1007, 660, 867, 999, 1557]. Words-in-Freedom [999]. Work [1461, 1460, 1010, 1306, 24, 103, 1424, 227, 240, 746, 1783, 1294]. Working [818, 1045, 1068, 94]. Works [823, 1142]. Workshop [836, 1025]. Workshops [1119]. World [939, 1589, 188, 1522, 1051, 1534, 1371, 1803, 285, 320]. Worldmaking [1099]. worm [613]. Wortgebilde [474]. wrapper [287]. Writable [534]. Write [546, 716, 985, 1756, 521]. write/paint [1756]. Writer [1539, 1044]. Writing [905, 1770, 529, 1375, 1573, 1264, 651, 978, 949, 924, 211, 1668, 762, 1119, 1196, 19, 655, 1031, 647, 349, 915, 1148, 739, 1078, 650, 527, 1817, 82, 1106, 528, 662, 1033, 1389, 752, 1292, 946, 384, 1215, 714, 1828, 1326, 1809, 1452, 1258, 1296, 224, 691, 992, 396, 937, 808, 648, 287, 621, 1030, 1534, 574, 500, 452, 646, 406, 586, 1716, 753, 948, 985, 62, 1375, 1275]. Writing-Books [287]. Writing-Systems [19]. Written [283, 1227, 264, 1766, 1197, 586, 453, 152, 1379, 1793]. Wrolstad [932]. www.id.iit.edu [1250]. www.id.iit.edu/visiblelanguage [1250]. Wynkyn [489].

X [83, 1328, 1319, 447, 460]. Xerox [1733]. XI [504]. XII [564, 275, 543]. xiii [31, 287, 603]. Xiong [1533]. XIth [69]. xiv [263, 656]. XML [1570]. Xu [1533]. XV [693]. XVI [744]. XVII [785]. XVIII [827]. XX [930]. XXI [964]. XXII [994]. XXIII [1023].

Yale [1554, 1377, 1297]. Year [31, 239, 1014, 238, 261, 1613]. Years [793, 1292, 1742, 1743, 1038, 1747]. Yi [1314, 1380]. York [484, 1320, 1375, 1555, 1572, 1621, 1622, 1391, 1461, 153, 1218, 56, 311, 263, 985, 152, 1256, 83, 1122, 1390, 1246, 1245, 1258, 1296, 1300, 1311, 1301, 1329, 1331, 1429, 1625, 1846, 1080, 1376, 71, 1096, 1098]. Young [788, 1476, 950, 1837, 1786]. you’re [1776, 1498].

Z [497]. Zabbal [1389]. Zachrisson [784]. Zapf [1018]. Zealand [1316, 1334]. Zebrackets [1105]. Zen [204]. Zero [110, 1391]. Zhukov [1301, 1429]. zoa [1305]. Zoo [592]. Zurich [1375].

References

Wrolstad:1967:PNF

[1] Merald E. Wrolstad. A prefatory note to the first number. The Journal of Typographic Research, I(1):3–4, January 1967. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V1N1_1967_E.pdf.

Wiggins:1967:ETT

[2] Richard H. Wiggins. Effects of three typographical variables on speed of reading. The Journal of Typographic Research, I(1):5–18, January

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N1_1967_E.pdf.

Fry:1967:DMS

[3] Edward Fry. The diacritical marketing system and a preliminary comparison with the initial teaching alphabet. The Journal of Typographic Research, I(1):19–30, January 1967. ISSN 0449-329X. URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V1N1_1967_E.pdf.

Prince:1967:PVH

[4] J. H. Prince. Printing for the visually handicapped. The Journal of Typographic Research, I(1):31–47, January 1967. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V1N1_1967_E.pdf.

Duncan:1967:LSS

[5] C. J. Duncan. Line scan standards for characters and symbols: a practical study. The Journal of Typographic Research, I(1):49–62, January

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N1_1967_E.pdf.

Nadler:1967:PPO

[6] M. Nadler. The perspectives for practical optical character recognition. The Journal of Typographic Research, I(1):63–68, January

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N1_1967_E.pdf.

Holland:1967:TEC

[7] F. C. Holland. Typographical effects by cathode ray tube typesetting systems. The Journal of Typographic Research, I(1):69–79, January

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N1_1967_E.pdf.

Cowan:1967:LVC

[8] C. I. Cowan. On-line visual correction and make-up systems — I: Hardware. The Journal of Typographic Research, I(1):80–89, January

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N1_1967_E.pdf.

Fabrizio:1967:RFS

[9] Ralph Fabrizio, Ira Kaplan, and Gilbert Teal. Readability as a function of the straightness of right-hand margins. The Journal of Typographic Research, I(1):90–95, January 1967. ISSN 0449-329X. URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V1N1_1967_E.pdf.

Malkiel:1967:SULa

[10] Yakov Malkiel. Secondary uses of letters in language. [Part I]. The Journal of Typographic Research, I(1):96–110, January 1967. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V1N1_1967_E.pdf. Reprinted from Romance Philology, 19 (1965).

Anonymous:1967:Aa

[11] Anonymous. The authors. The Journal of Typographic Research, I(1): 111–113, January 1967. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V1N1_1967_E.pdf.

Harrison:1967:CTT

[12] Randall Harrison and Clyde D. J. Morris. Communication theory and typographic research. The Journal of Typographic Research, I(2):115–124,

April 1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V1N2_1967_E.pdf.

Payne:1967:RTM

[13] Donald R. Payne. Readability of typewritten material; proportional versus standard spacing. The Journal of Typographic Research, I(2): 125–136, April 1967. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V1N2_1967_E.pdf.

Frutiger:1967:OBS

[14] Adrian Frutiger. OCR-B: a standardized character for optical recognition. The Journal of Typographic Research, I(2):137–146, April

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N2_1967_E.pdf.

Wheelock:1967:IVD

[15] Warren H. Wheelock and Nicholas J. Silvaroli. An investigation of visual discrimination training for beginning readers. The Journal of Typographic Research, I(2):147–156, April 1967. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V1N2_1967_E.pdf.

Tanselle:1967:TRB

[16] G. Thomas Tanselle. Typographic research and bibliography. The Journal of Typographic Research, I(2):157–163, April 1967. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V1N2_1967_E.pdf.

Wakefield:1967:PLD

[17] R. J. Wakefield. Print layout and design with a computer CRT system. The Journal of Typographic Research, I(2):165–168, April

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N2_1967_E.pdf.

Malkiel:1967:SULb

[18] Yakov Malkiel. Secondary uses of letters in language. [Part II]. The Journal of Typographic Research, I(2):169–190, April 1967. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V1N2_1967_E.pdf.

Downing:1967:CMP

[19] John Downing. Commentary: Methodological problems in research on simplified alphabets and regularized writing-systems. The Journal of Typographic Research, I(2):191–201, April 1967. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V1N2_1967_E.pdf. With a reply by Edward Fry.

Anonymous:1967:C

[20] Anonymous. Correspondence. The Journal of Typographic Research, I (2):202–203, April 1967. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V1N2_1967_E.pdf.

Anonymous:1967:BR

[21] Anonymous. Book reviews. The Journal of Typographic Research, I(2): 204–209, April 1967. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V1N2_1967_E.pdf.

Anonymous:1967:AJAa

[22] Anonymous. Abstracts of journal articles in French and German. The Journal of Typographic Research, I(2):210–213, April

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N2_1967_E.pdf.

Anonymous:1967:Ab

[23] Anonymous. The authors. The Journal of Typographic Research, I(2): 214–215, April 1967. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V1N2_1967_E.pdf.

Pierce:1967:PIA

[24] James Smith Pierce. Pictographs, ideograms, and alphabets in the work of Paul Klee. The Journal of Typographic Research, I(3):219–244, July

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N3_1967_E.pdf.

Huss:1967:CLT

[25] Richard E. Huss. A chronological list of typesetting machines and ancillary equipment, 1822–1925. The Journal of Typographic Research, I (3):245–274, July 1967. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V1N3_1967_E.pdf. See com-

ments [73].

Hailstone:1967:SED

[26] M. Hailstone and Jeremy J. Foster. Studies of the efficiency of drug labeling. The Journal of Typographic Research, I(3):275–284, July

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N3_1967_E.pdf.

Frutiger:1967:TIS

[27] Adrian Frutiger. Typography with the IBM Selectric Composer. The Journal of Typographic Research, I(3):285–292, July 1967. ISSN 0449-

329X. URL https://s3-us-west-2.amazonaws.com/visiblelanguage/

pdf/V1N3_1967_E.pdf. See comments [47, 85].

Weaver:1967:CP

[28] Mike Weaver. Concrete poetry. The Journal of Typographic Research, I (3):293–326, July 1967. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V1N3_1967_E.pdf.

Wendt:1967:RPR

[29] Dirk Wendt. Research in progress: Research on readability and congeniality of print. The Journal of Typographic Research, I(3):327, July

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N3_1967_E.pdf.

Middleton:1967:ER

[30] R. Hunter Middleton. Exhibition review. The Journal of Typographic Research, I(3):328–330, July 1967. ISSN 0449-329X. URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V1N3_1967_E.pdf.

Anonymous:1967:BRR

[31] Anonymous. Book review: R. C. Alston, A Bibliography of the English Language from the Invention of Printing to the Year 1800, Volume VIII: Treatises on Short-Hand. Printed for the author by E. J. Arnold & Son Limited, Leeds, England; 1966. xiii + 152 pp. £5. 1Os. The Journal of Typographic Research, I(3):331–335, July 1967. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V1N3_1967_E.pdf.

Anonymous:1967:AJAb

[32] Anonymous. Abstracts of journal articles in French and German. The Journal of Typographic Research, I(3):336–338, July

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N3_1967_E.pdf.

Anonymous:1967:Ac

[33] Anonymous. The authors. The Journal of Typographic Research, I (3):339, July 1967. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V1N3_1967_E.pdf.

Mathews:1967:TFC

[34] M. V. Mathews, Carol Lochbaum, and Judith A. Moss. Three fonts of computer-drawn letters. The Journal of Typographic Research, I(4): ??, October 1967. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V1N4_1967_E.pdf.

Bass:1967:DCN

[35] Rudi Bass. The development of CBS News 36. The Journal of Typographic Research, I(4):357–372, October 1967. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V1N4_1967_E.pdf.

Baudin:1967:TER

[36] Fernand Baudin. Typography: Evolution and revolution. The Journal of Typographic Research, I(4):373–386, October 1967. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V1N4_1967_E.pdf. See commentary [70, 84].

Wright:1967:VGE

[37] Edward Wright and J. Collins. Vertical group exercises in graphic design. The Journal of Typographic Research, I(4):??, October

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N4_1967_E.pdf.

Tomohiko:1967:JC

[38] Horie Tomohiko. Japanese calligraphy. The Journal of Typographic Research, I(4):??, October 1967. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V1N4_1967_E.pdf.

Tansell:1967:ITB

[39] G. Thomas Tansell. Identification of typefaces in bibliographical description. The Journal of Typographic Research, I(4):??, October

1967. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V1N4_1967_E.pdf.

Wendt:1968:SDT

[40] Dirk Wendt. Semantic differentials of typefaces as a method of congeniality research. The Journal of Typographic Research, II(1):3–25, January

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N1_1968_E.pdf.

Gurtler:1968:DE

[41] André Gürtler. The design of Egyptian 505. The Journal of Typographic Research, II(1):27–42, January 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V2N1_1968_E.pdf.

Hinds:1968:WCT

[42] Lillian R. Hinds and William G. Dodds. Words in color: Two experimental studies. The Journal of Typographic Research, II(1):43–52, January

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N1_1968_E.pdf.

Jones:1968:RRC

[43] F. Kenneth Jones. A research report on colour story reading. The Journal of Typographic Research, II(1):53–58, January 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V2N1_1968_E.pdf.

Evers:1968:AUC

[44] C. H. Evers. Adjustments to unjustified composition on the Rotterdamsch Nieuwsblad. The Journal of Typographic Research, II(1):59–74, January

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N1_1968_E.pdf.

Anonymous:1968:AGJ

[45] Anonymous. Author’s guide to The Journal of Typographic Research. The Journal of Typographic Research, II(1):75–100, January

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N1_1968_E.pdf.

Baudin:1968:RAAa

[46] Fernand Baudin. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, II(1):101– 103, January 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N1_1968_E.pdf.

Schmoller:1968:E

[47] Hans Schmoller. To the editor. The Journal of Typographic Research, II (1):103, January 1968. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V2N1_1968_E.pdf. Critical

comment on [27]. See comments [85, 154].

Anonymous:1968:Aa

[48] Anonymous. The authors. The Journal of Typographic Research, II(1): 104, January 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N1_1968_E.pdf.

Smith:1968:TRM

[49] Glendon C. Smith. Type reading machines for the blind. The Journal of Typographic Research, II(2):107–125, April 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V2N2_1968_E.pdf.

Enstrom:1968:RHR

[50] E. A. Enstrom. Reading–handwriting research. The Journal of Typographic Research, II(2):126, April 1968. ISSN 0449-329X. URL https://

s3-us-west-2.amazonaws.com/visiblelanguage/pdf/V2N2_1968_E.

pdf. Excerpted with kind permission from The Reading Teacher, Volume 21, Number 6 (March 1968); 1968 by International Reading Association, Newark, Delaware 19711.

Click:1968:RRN

[51] J. W. Click and Guido H. Stempel III. Reader response to newspaper front-page format. The Journal of Typographic Research, II(2):127–142,

April 1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V2N2_1968_E.pdf.

Horowitz:1968:WFA

[52] Leonard M. Horowitz, Margaret A. White, and Douglas W. Atwood. Words fragments as aids to recall: the organization of a word. The Journal of Typographic Research, II(2):143–156, April 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V2N2_1968_E.pdf.

Turnbull:1968:RAA

[53] A. T. Turnbull and David E. Carter. Readership of advertisements with all-display typography. The Journal of Typographic Research, II (2):157–169, April 1968. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V2N2_1968_E.pdf.

Anonymous:1968:JEE

[54] Anonymous. Journal editor for experimental design. The Journal of Typographic Research, II(2):170, April 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V2N2_1968_E.pdf. Pieter Brattinga h as been named Editor for Experimental Design for The Journal of Typographic Research.

Shera:1968:DM

[55] Jesse H. Shera and Conrad H. Rawski. The diagram is the message. The Journal of Typographic Research, II(2):171–188, April

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N2_1968_E.pdf.

Coleman:1968:BRS

[56] Carroll Coleman. Book review: The Spiral Press through Four Decades, An Exhibition of Books and Ephemera, with a Commentary by Joseph Blumenthal. New York: Pierpont Morgan Library, 1966. 100 pages. Clothbound edition out of print; paperbound edition, $2.85. The Journal of Typographic Research, II(2):189–193, April 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V2N2_1968_E.pdf.

Taks:1968:E

[57] J. J. Taks. To the Editor. The Journal of Typographic Research, II (2):194, April 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N2_1968_E.pdf.

Baudin:1968:RAAb

[58] Fernand Baudin. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, II(2): 195–197, April 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N2_1968_E.pdf.

Anonymous:1968:Ab

[59] Anonymous. The authors. The Journal of Typographic Research, II(2): 197–198, April 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N2_1968_E.pdf.

Anonymous:1968:PA

[60] Anonymous. The Penrose Annual 1968. The Journal of Typographic Research, II(2):199, April 1968. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V2N2_1968_E.pdf.

Bonsiepe:1968:MQO

[61] Gui Bonsiepe. A method of quantifying order in typographic design. The Journal of Typographic Research, II(3):203–220, July

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N3_1968_E.pdf.

Mountford:1968:WA

[62] John Mountford. ‘Writing’ and ‘alphabet’. The Journal of Typographic Research, II(3):221–232, July 1968. ISSN 0449-329X. URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V2N3_1968_E.pdf.

Telingater:1968:SAG

[63] Solomon B. Telingater. The standardization of alphabetic graphemes. The Journal of Typographic Research, II(3):233–240, July 1968.

ISSN 0449-329X.  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N3_1968_E.pdf.

Slothouber:1968:TA

[64] Jan Slothouber and William Graatsma. Typographic anamorphosis. The Journal of Typographic Research, II(3):241–244, July 1968.

ISSN 0449-329X.  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N3_1968_E.pdf.

Brown:1968:TSM

[65] Dwight D. Brown. Towards a standard for measuring the accuracy of any computer-hyphenation program. The Journal of Typographic Research, II(3):245–258, July 1968. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V2N3_1968_E.pdf.

vanderMarck:1968:PRP

[66] Jan van der Marck. Pictures to be read/poetry to be seen. The Journal of Typographic Research, II(3):259–270, July 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V2N3_1968_E.pdf.

Zachrisson:1968:TLR

[67] Bror Zachrisson. The A.Typ.I. legibility research committee. The Journal of Typographic Research, II(3):271–276, July 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V2N3_1968_E.pdf.

Anonymous:1968:LM

[68] Anonymous. List of members. The Journal of Typographic Research, II(3):277, July 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N3_1968_E.pdf.

Anonymous:1968:TXC

[69] Anonymous. A.TYP.I XIth Congress; ICOGRADA 1968 Congress. The Journal of Typographic Research, II(3):278, July 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V2N3_1968_E.pdf.

Foster:1968:CPR

[70] Jeremy J. Foster. Commentary: Psychological research into legibility. The Journal of Typographic Research, II(3):279–282, July

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N3_1968_E.pdf. See [36, 84].

Welsh:1968:BRC

[71] George S. Welsh. Book reviews: I. Coppersmith, Stanley (ed.). Frontiers of Psychological Research: Readings from Scientific American. San Francisco: Freeman and Co., 1966. 322 pages. $10.00 ($4.95 in paper).

2. Gregory, R. L., Eye and Brain: The Psychology of Seeing. New York: McGraw-Hill, J 966. 254 pages. $2.45 (paper). 3. Vernon, M. D. (ed.). Experiments in Visual Perception: Selected Readings. Baltimore: Penguin Books, 1966. 443 pages. $1.95 (paper). The Journal of Typographic Research, II(3):283–286, July 1968. ISSN 0449-329X. URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V2N3_1968_E.pdf.

Anonymous:1968:BR

[72] Anonymous. Books received. The Journal of Typographic Research, II (3):287–291, July 1968. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V2N3_1968_E.pdf.

Hutt:1968:C

[73] G. Allen Hutt. Correspondence. The Journal of Typographic Research, II(3):292, July 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N3_1968_E.pdf. See [25].

Baudin:1968:RAAc

[74] Fernand Baudin. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, II(3): 293–295, July 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N3_1968_E.pdf.

Anonymous:1968:Ac

[75] Anonymous. The authors. The Journal of Typographic Research, II (3):296, July 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N3_1968_E.pdf.

Mergler:1968:OAC

[76] H. W. Mergler and P. M. Vargo. One approach to computer assisted letter design. The Journal of Typographic Research, II(4):299–322, October

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N4_1968_E.pdf.

Moorsel:1968:TL

[77] L. Leering van Moorsel. The typography of El Lissitzky. The Journal of Typographic Research, II(4):323–340, October 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V2N4_1968_E.pdf.

Seybold:1968:EVC

[78] John W. Seybold. Esthetic values in computerized photocomposition. The Journal of Typographic Research, II(4):341–350, October

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N4_1968_E.pdf.

Zapf:1968:CLD

[79] Hermann Zapf. Changes in letterforms due to technical developments. The Journal of Typographic Research, II(4):351–368, October

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N4_1968_E.pdf.

Emans:1968:CCA

[80] Robert Emans. Context clues as an aid to the reader. The Journal of Typographic Research, II(4):369–373, October 1968. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V2N4_1968_E.pdf.

Flarsheim:1968:RP

[81] Carl Flarsheim, Warren Infield, Alan Levine, and Adrienne C. McCoy. Research in progress. The Journal of Typographic Research, II(4):374– 378, October 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N4_1968_E.pdf.

Hoch:1968:CLW

[82] Ernest Hoch. Commentary: Linguistics, “writing,” and typography. The Journal of Typographic Research, II(4):379–382, October

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N4_1968_E.pdf.

Meyer:1968:BRS

[83] Peter Meyer and Henry Hawley. Book reviews: I. Stephen Bann (Ed.). Concrete Poetry. An international anthology. London: London Magazine Editions No. 13, 1967. 199 pages. 30s. 2. Eugene Wildman (ed.). The Chicago Review Anthology of Concretism. Chicago: Swallow Press, 1967. 15 7 pages. $2.00. 3. Emmett William (Ed.). An Anthology of Concrete Poetry. New York: Something Else Press, 196 7. X + 342 pages. $10.00 (cloth), $3.00 (paper). John Russell Taylor. The Art Nouveau Book in Britain. Cambridge, Mass.: M.I.T. Press, 1966, 175 pp. $12.95. The Journal of Typographic Research, II(4):383–391, October

1968. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V2N4_1968_E.pdf. See letters [105].

Baudin:1968:C

[84] Fernand Baudin. Correspondence. The Journal of Typographic Research, II(4):397, October 1968. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V2N4_1968_E.pdf. See [36,

70].

Gurtler:1968:C

[85] André Gürtler. Correspondence. The Journal of Typographic Research, II(4):397–398, October 1968. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V2N4_1968_E.pdf. See

[47].

Spear:1968:C

[86] Frank Spear, Jack Stauffacher, Daniel Ford, John Willett, and Howard Paine. Correspondence. The Journal of Typographic Research, II(4): 398–401, October 1968. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V2N4_1968_E.pdf.

Baudin:1968:RAAd

[87] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, II(4):402–404, October 1968. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V2N4_1968_E.pdf.

Anonymous:1968:IVI

[88] Anonymous. Index to Volume II. The Journal of Typographic Research, II(4):405–407, October 1968. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V2N4_1968_E.pdf.

Anonymous:1968:Ad

[89] Anonymous. The authors. The Journal of Typographic Research, II(4): 408, October 1968. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V2N4_1968_E.pdf.

Meiss:1969:FAT

[90] Millard Meiss. The first alphabetical treatises in the Renaissance. The Journal of Typographic Research, III(1):3–30, January

1969. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N1_1969_E.pdf.

Munson:1969:CRH

[91] John H. Munson. Computer recognition of hand-printed text. The Journal of Typographic Research, III(1):31–61, January 1969.

ISSN 0449-329X.  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N1_1969_E.pdf.

Crosby:1969:RMP

[92] R. M. N. Crosby. Recognizing the marks on paper. The Journal of Typographic Research, III(1):63–78, January 1969. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V3N1_1969_E.pdf.

Shurtleff:1969:RLL

[93] Donald Shurtleff. Relative legibility of Leroy and Lincoln/MITRE fonts on television. The Journal of Typographic Research, III(1):79–90, Jan-

uary 1969. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V3N1_1969_E.pdf.

Lindley:1969:TEE

[94] William R. Lindley. Typographic education. England: The Working Party on Typographic Teaching. USA: Typography in journalism school curricula. The Journal of Typographic Research, III(1):91–106, January

1969. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N1_1969_E.pdf.

Baudin:1969:RAAa

[95] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, III(1):107–109, January 1969. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V3N1_1969_E.pdf.

Anonymous:1969:Aa

[96] Anonymous. The authors. The Journal of Typographic Research, III (1):110, January 1969. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V3N1_1969_E.pdf.

Anonymous:1969:BNJ

[97] Anonymous. Back numbers of the Journal. The Journal of Typographic Research, III(1):111–112, January 1969. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V3N1_1969_E.pdf.

Wrolstad:1969:LRN

[98] Merald E. Wrolstad. Letterform research needs definition and direction: A report from the Editor. The Journal of Typographic Research, III(2): 115–126, April 1969. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V3N2_1969_E.pdf.

Bartz:1969:TVP

[99] Barbara S. Bartz. Type variation and the problem of cartographic type legibility — Part One: Cartographic typography as a medium for communication; the cartographic view of legibility. The Journal of Typographic Research, III(2):127–144, April 1969. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V3N2_1969_E.pdf. See correspondence [154].

Kolers:1969:CLR

[100] Paul A. Kolers. Clues to a letter’s recognition: Implications for the design of characters. The Journal of Typographic Research, III(2):145–168, April

1969. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N2_1969_E.pdf.

DiMeo:1969:VMS

[101] Katherine P. DiMeo. Visual-motor skills: Response characteristics and prereading behavior. The Journal of Typographic Research, III(2):169– 182, April 1969. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V3N2_1969_E.pdf.

Rice:1969:SCT

[102] Stanley Rice. A standard code for typographic character identification. The Journal of Typographic Research, III(2):183–192, April

1969. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N2_1969_E.pdf.

Roth:1969:ETM

[103] Joel A. Roth. Excerpt: Typography that makes the reader work. The Journal of Typographic Research, III(2):193–196, April 1969.

ISSN 0449-329X.  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N2_1969_E.pdf.

Baudin:1969:BRH

[104] Fernand Baudin, Colin Banks, and Ruari McLean. Book reviews: H. D. L. Vervliet, Sixteenth-century Printing Types of the Low Countries. Jan Tschichold, Asymmetric Typography. David Diringer, The Alphabet. The Journal of Typographic Research, III(2):197–208, April

1969. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N2_1969_E.pdf.

HamiltonFinlay:1969:C

[105] Ian Hamilton Finlay and Peter Mayer. Correspondence. The Journal of Typographic Research, III(2):209–212, April 1969. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V3N2_1969_E.pdf. See [83].

Baudin:1969:RAAb

[106] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, III(2):213–215, April 1969. ISSN 0449-329X. URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V3N2_1969_E.pdf.

Anonymous:1969:Ab

[107] Anonymous. The authors. The Journal of Typographic Research, III (2):216, April 1969. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V3N2_1969_E.pdf.

Bruns:1969:MTL

[108] Gerald L. Bruns. Mallarmé: The transcendence of language and the aesthetics of the book. The Journal of Typographic Research, III(3): 219–240, July 1969. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V3N3_1969_E.pdf.

Wendt:1969:X

[109] Dirk Wendt. O or 0? The Journal of Typographic Research, III(3): 241–248, July 1969. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V3N3_1969_E.pdf. See discus-

sion [189, 244].

Vartabedian:1969:PFG

[110] Allen G. Vartabedian. A proposed fontstyle for the graphic representation of the oh and zero. The Journal of Typographic Research, III (3):249–258, July 1969. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V3N3_1969_E.pdf. See cor-

respondence [154, 189].

Tanselle:1969:UTD

[111] G. Thomas Tanselle. The use of type damage as evidence in bibliographical description. The Journal of Typographic Research, III(3):259–276,

July 1969. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V3N3_1969_E.pdf.

Burns:1969:RGA

[112] J. R. Burns. A report generator approach to automated page composition. The Journal of Typographic Research, III(3):277–286, July

1969. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N3_1969_E.pdf.

Anonymous:1969:LA

[113] Anonymous. Letterforms in the arts. The Journal of Typographic Research, III(3):287–292, July 1969. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V3N3_1969_E.pdf.

Reynolds:1969:CMM

[114] Lloyd Reynolds and R. J. Wakefield. Comment: Marshall McLuhan and italic handwriting; comment: Publishing by computer. The Journal of Typographic Research, III(3):293–300, July 1969. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V3N3_1969_E.pdf.

Anonymous:1969:ER

[115] Anonymous. Exhibition review. The Journal of Typographic Research, III

(3):301–305, July 1969. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N3_1969_E.pdf.

Banks:1969:BRE

[116] Colin Banks. Book review: Erik Dal, Scandinavian Bookmaking in the Twentieth Century. Copenhagen: Christian Ejlers Forlag (also published Urbana: The University of Illinois Press), 1968. 128pp. $5.50. The Journal of Typographic Research, III(3):306–308, July 1969.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V3N3_1969_E.pdf.

Baudin:1969:RAAc

[117] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research,

III(3):309–311, July 1969. URL https://s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V3N3_1969_E.pdf.

Anonymous:1969:Ac

[118] Anonymous. The authors. The Journal of Typographic Research,

III(3):312, July 1969. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N3_1969_E.pdf.

Kaldor:1969:GRG

[119] Ivan L. Kaldor. The genesis of Russian Grazhdanskii Shrift or civil type — Part I. The Journal of Typographic Research, III(4):315–344, October

1969. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N4_1969_E.pdf. See also Part II [146].

Turbayne:1969:VLV

[120] Colin Murray Turbayne. Visual language from the verbal model. The Journal of Typographic Research, III(4):345–370, October 1969.

ISSN 0449-329X.  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3N4_1969_E.pdf.

Ovink:1969:FTD

[121] G. W. Ovink. Fashion in type design. The Journal of Typographic Research, III(4):371–377, October 1969. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V3N4_1969_E.pdf.

Anonymous:1969:CJS

[122] Anonymous. Changes in journal subscription policy. The Journal of Typographic Research, III(4):378, October 1969. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V3N4_1969_E.pdf.

Brattinga:1969:HSN

[123] Pieter Brattinga. The house-style of The Netherlands PTT. The Journal of Typographic Research, III(4):379–386, October 1969. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V3N4_1969_E.pdf.

Bartz:1969:SAC

[124] Barbara S. Bartz. Search: an approach to cartographic type legibility measurement. The Journal of Typographic Research, III(4):387–398, Oc-

tober 1969. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V3N4_1969_E.pdf. A summary of this re-

search is contained in Barbara S. Ba rtz, Type Variation and the Problem of Cartographic Type Legibility (Madison: Ph.D. Dissertation, University of Wisconsin, 1969). See correspondence [154].

Anonymous:1969:RS

[125] Anonymous. Response sheet. The Journal of Typographic Research, III (4):399, October 1969. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V3N4_1969_E.pdf.

Anonymous:1969:PAV

[126] Anonymous. The Penrose Annual 1969, Volume 62. The Journal of Typographic Research, III(4):400, October 1969. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V3N4_1969_E.pdf.

Baudin:1969:RAAd

[127] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, III(4):401–403, October 1969. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V3N4_1969_E.pdf.

Anonymous:1969:Ad

[128] Anonymous. The authors. The Journal of Typographic Research, III (4):404, October 1969. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V3N4_1969_E.pdf.

Anonymous:1969:IVI

[129] Anonymous. Index to Volume III. The Journal of Typographic Research, III(4):405–407, October 1969. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V3N4_1969_E.pdf.

Anonymous:1970:GIa

[130] Anonymous. General information. The Journal of Typographic Research, IV(1):3–6, Winter 1970. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Lees:1970:IDP

[131] John Lees and Melvin Farman. An investigation of the design and performance of traffic control devices. The Journal of Typographic Research, IV(1):7–38, Winter 1970. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Scorsone:1970:LDC

[132] Joesph S. Scorsone. Ligature design for contemporary technology. The Journal of Typographic Research, IV(1):39–50, Winter

1970. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N1_1970_E.pdf.

Crouwel:1970:TDC

[133] Wim Crouwel. Type design for the computer age. The Journal of Typographic Research, IV(1):51–59, Winter 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N1_1970_E.pdf.

Becker:1970:RPT

[134] D. Becker, J. Heinrich, R. von Sichowsky, and Dirk Wendt. Reader preferences for typeface and leading. The Journal of Typographic Research, IV(1):61–66, Winter 1970. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Montague:1970:DIT

[135] Arleigh Montague. Designing the initial teaching alphabet in five typefaces. The Journal of Typographic Research, IV(1):67–72, Winter

1970. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N1_1970_E.pdf.

Brown:1970:SRM

[136] W. S. Brown. Speed-reading made easy. The Journal of Typographic Research, IV(1):73–75, Winter 1970. ISSN 0449-329X. URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Anonymous:1970:LRA

[137] Anonymous. Legibility research abstracts. The Journal of Typographic Research, IV(1):76, Winter 1970. ISSN 0449-329X. URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Ong:1970:CVP

[138] Walter J. Ong. Comment: Voice, print, and culture. The Journal of Typographic Research, IV(1):77–83, Winter 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N1_1970_E.pdf.

Parker:1970:WI

[139] Brant Parker and Johnny Hart. The Wizard of Id. The Journal of Typographic Research, IV(1):84, Winter 1970. ISSN 0449-329X.

URL https://en.wikipedia.org/wiki/Brant_Parker;   https:/
/en.wikipedia.org/wiki/Johnny_Hart;   https://s3-us-west-2.
amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Noordzij:1970:RJD

[140] Gerrit Noordzij. Reading the journal: a discussion of starting points in the issue of April 1969. The Journal of Typographic Research, IV(1): 85–90, Winter 1970. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Baudin:1970:RAAa

[141] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, IV(1):91–93, Winter 1970. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Anonymous:1970:BA

[142] Anonymous. [Book] announcements. The Journal of Typographic Research, IV(1):94, Winter 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Anonymous:1970:Aa

[143] Anonymous. The authors. The Journal of Typographic Research, IV (1):96, Winter 1970. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Anonymous:1970:GIb

[144] Anonymous. General information. The Journal of Typographic Research, IV(2):99–102, Spring 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N1_1970_E.pdf.

Goodman:1970:PUR

[145] Kenneth S. Goodman. Psycholinguistic universals in the reading process. The Journal of Typographic Research, IV(2):103–110, Spring

1970. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N2_1970_E.pdf.

Kaldor:1970:GRG

[146] Ivan L. Kaldor. The genesis of the Russian Grazhdanskii Shrift or civil type — Part II. The Journal of Typographic Research, IV(2): 111–138, Spring 1970. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V4N2_1970_E.pdf. See also

Part I [119].

Foster:1970:DCF

[147] Jeremy J. Foster. Directional consistency in form identification. The Journal of Typographic Research, IV(2):139–145, Spring 1970.

ISSN 0449-329X.  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N2_1970_E.pdf.

Burnhill:1970:CSW

[148] Peter Burnhill. The case for a standard word space. The Journal of Typographic Research, IV(2):146, Spring 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N2_1970_E.pdf.

Bartz:1970:EUS

[149] Barbara S. Bartz. Experimental use of the search task in an analysis of type legibility in cartography. The Journal of Typographic Research, IV (2):147–167, Spring 1970. ISSN 0449-329X. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V4N2_1970_E.pdf.

Anonymous:1970:GBC

[150] Anonymous. Goudy, Baskerville, Caslon, Bodoni. The Journal of Typographic Research, IV(2):168, Spring 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V4N2_1970_E.pdf. Cartoon from The New Yorker Magazine, 1964.

Pulos:1970:CDE

[151] Arthur J. Pulos. Comment: Design education. The Journal of Typographic Research, IV(2):169–172, Spring 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N2_1970_E.pdf.

Kindersley:1970:BRJ

[152] David Kindersley. Book reviews: Jan Tschichold, Treasury of Alphabets and Lettering. New York: Reinhold Publishing Corporation, 1966. (Originally published in Germany, Meisterbuch der Schriff. Ravensburg: Otto Maier Verlag.) 234 pages (175 pages of illustration). $16.50. Tom Gourdie, A Guide to Better Handwriting. New York: The Viking Press, 1967 (London: Studio Vista, 1967). 96 pages. $1.95. Dana Atchley, ABC Design. New York: George Wittenborn, 1965. 30 pages of illustration, some hand-written text. $9.00. L’Harl Copeland, Design of the Roman Letters. New York: Philosophical Library, 1966. ix + 66 pages, illustrated. $3.75. Eric Lindegren, ABC of Lettering and Printing Types. New York: Museum Books. Produced by Eric Lindegren Gratisk Studio, Askim, Sweden. Volume A: 150 pages of illustrations, 4 pages of text, 1964. $12.50. Volume B: 293 pages of illustrations, 30 pages of text, 1965. $15.00. Volume C: 133 pages of illustrations with text, 1965. $15.00. The Journal of Typographic Research, IV(2):173–176, Spring 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N2_1970_E.pdf.

Baudin:1970:BRH

[153] Fernand Baudin. Book reviews: Herbert Spencer, The Visible Word. London: Lund Humphries in association with the Royal College of Art, 1969. 107pp. 50s (New York: Hastings House, $7.95). Ruari McLean, Magazine Design. London: Oxford University Press, 1969. 354 pp. £5.5.0. The Journal of Typographic Research, IV(2):176–178, Spring

1970. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N2_1970_E.pdf.

Zapf:1970:C

[154] Hermann Zapf, Allen G. Vartabedian, Adrian Frutiger, J. A. M. Haddon, and G. W. Ovink. Correspondence. The Journal of Typographic Research, IV(2):179–188, Spring 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N2_1970_E.pdf.

Baudin:1970:RAAb

[155] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, IV(2):189–191, Spring 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N2_1970_E.pdf.

Anonymous:1970:Ab

[156] Anonymous. The authors. The Journal of Typographic Research, IV (2):192, Spring 1970. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V4N2_1970_E.pdf.

Anonymous:1970:GIc

[157] Anonymous. General information. The Journal of Typographic Research, IV(3):195–198, Summer 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N3_1970_E.pdf.

Arnheim:1970:WTP

[158] Rudolf Arnheim. Words in their place. The Journal of Typographic Research, IV(3):199–212, Summer 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N3_1970_E.pdf.

Noordzij:1970:BSC

[159] Gerrit Noordzij. Broken scripts and the classification of typefaces. The Journal of Typographic Research, IV(3):213–240, Summer

1970. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N3_1970_E.pdf. See follow-up work [186] and

comments [189].

Patteson:1970:SIA

[160] Roy K. Patteson, Jr. The Siloam Inscription and alphabetic origins. The Journal of Typographic Research, IV(3):241–257, Summer

1970. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N3_1970_E.pdf.

Senator:1970:S

[161] Cassiodorus Senator. On scribes. The Journal of Typographic Research, IV(3):258, Summer 1970. ISSN 0449-329X. URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V4N3_1970_E.pdf.

Reprinted from An Introduction to Divine and Human Readings, Chapter XXX; translated with an introduction and notes by Leslie Webber Jones (New York: Columbia University Press, 1946).

Hunt:1970:TRR

[162] Allen Hunt. Times Roman: a re-assessment. The Journal of Typographic Research, IV(3):259–270, Summer 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N3_1970_E.pdf.

Anonymous:1970:PAN

[163] Anonymous. Proposed American National Standard: Presentation of alphameric characters for information processing. The Journal of Typographic Research, IV(3):271–280, Summer 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N3_1970_E.pdf.

Brattinga:1970:BRH

[164] Pieter Brattinga and Lawrence W. Wallis. Book reviews: Herbert Spencer. Pioneers of Modern Typography. London: Lund Humphries, 1969, 162 pp. 84/-. Arthur H. Phillips. Computer Peripherals and Typesetting. London: Her Majesty’s Stationery Office, 1968, 665 pp. 160/- (Cloth). The Journal of Typographic Research, IV(3):281–284, Summer

1970. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N3_1970_E.pdf.

Baudin:1970:RAAc

[165] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, IV(3):285–287, Summer 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N3_1970_E.pdf.

Anonymous:1970:Ac

[166] Anonymous. The authors. The Journal of Typographic Research, IV(3): 288, Summer 1970. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V4N3_1970_E.pdf.

Anonymous:1970:GId

[167] Anonymous. General information. The Journal of Typographic Research, IV(4):291–294, Autumn 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N4_1970_E.pdf.

Mountford:1970:SPC

[168] John Mountford. Some psycholinguistic components of initial standard literacy. The Journal of Typographic Research, IV(4):295–306, Autumn

1970. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N4_1970_E.pdf.

Boyle:1970:EGH

[169] Leonard E. Boyle. The emergence of Gothic handwriting. The Journal of Typographic Research, IV(4):307–316, Autumn 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V4N4_1970_E.pdf. See letter [254].

Marck:1970:STU

[170] Louis Marck. Some thoughts of the use and disuse of diacritics. The Journal of Typographic Research, IV(4):317–326, Autumn

1970. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N4_1970_E.pdf. See comments [207].

Frutiger:1970:LPT

[171] Adrian Frutiger. Letterforms in photo-typography. The Journal of Typographic Research, IV(4):327–335, Autumn 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N4_1970_E.pdf.

Wormald:1970:RMM

[172] Francis Wormald. Reading in the medieval monastery. The Journal of Typographic Research, IV(4):336, Autumn 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V4N4_1970_E.pdf. Excerpted from the author’s The Monastic Library in The Year 1200: a Background Survey, II (New York: Metropolitan Museum of Art, 1970), 170.

Molyneux:1970:IQH

[173] Lindsay Molyneux. An index of the quality of a hyphenation algorithm. The Journal of Typographic Research, IV(4):337–341, Autumn

1970. ISSN 0449-329X. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V4N4_1970_E.pdf.

Kindersley:1970:LA

[174] David Kindersley. Letterforms in the arts. The Journal of Typographic Research, IV(4):342–352, Autumn 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N4_1970_E.pdf.

Burnhill:1970:TEH

[175] Peter Burnhill. Typographic education: Headings in text. The Journal of Typographic Research, IV(4):353–365, Autumn 1970. ISSN 0449-329X.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V4N4_1970_E.pdf.

Anonymous:1970:SE

[176] Anonymous. Shaw enough. The Journal of Typographic Research, IV(4): 366–367, Autumn 1970. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V4N4_1970_E.pdf. Commen-

tary on George Bernard Shaw’s proposed spelling reform of English, and on his Shavian phonetic alphabet for English. Reprinted from The Sunday Times (London), 8 March 1970.

Beyer:1970:C

[177] Ralph Beyer, Peter Burnhill, Nicolete Gray, James Mosley, Michael Twyman, Berthold Wolpe, The Governors of the St.Bride Foundation, and David Ford. Correspondence. The Journal of Typographic Research, IV(4):367–368, Autumn 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N4_1970_E.pdf.

Anonymous:1970:BR

[178] Anonymous. Books received. The Journal of Typographic Research, IV (4):369–376, Autumn 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N4_1970_E.pdf.

Baudin:1970:RAAd

[179] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. The Journal of Typographic Research, IV(4):377–379, Autumn 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N4_1970_E.pdf.

Anonymous:1970:Ad

[180] Anonymous. The authors. The Journal of Typographic Research, IV(4): 380, Autumn 1970. ISSN 0449-329X. URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V4N4_1970_E.pdf.

Anonymous:1970:IVI

[181] Anonymous. Index to Volume IV. The Journal of Typographic Research, IV(4):381–384, Autumn 1970. ISSN 0449-329X. URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V4N4_1970_E.pdf.

Wrolstad:1971:VLJ

[182] Merald E. Wrolstad. Visible Language: The journal for research on the visual media of language expression. a report from the Editor. Visible Language, V(1):5–12, Winter 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V5N1_1971_E.pdf.

Osley:1971:CAC

[183] A. S. Osley. Calligraphy — an aid to cartography? Visible Language, V(1):13–32, Winter 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N1_1971_E.pdf.

Bass:1971:DV

[184] Rudi Bass. The development of Vidifont. Visible Language, V(1): 33–48, Winter 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N1_1971_E.pdf.

Crofts:1971:CMO

[185] Marjorie Crofts. Creating a Mundurukú orthography. Visible Language, V(1):49–58, Winter 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N1_1971_E.pdf.

Tracy:1971:TDC

[186] Walter Tracy. Type design classification. Visible Language, V(1): 59–66, Winter 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N1_1971_E.pdf. See also previous work [159].

Carpenter:1971:CSB

[187] Edmund Carpenter. Comment: Not since Babel. Visible Language, V (1):67–74, Winter 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N1_1971_E.pdf.

Lotz:1971:CRS

[188] John Lotz. Comment: The role of script in describing the languages of the world. Visible Language, V(1):75–81, Winter 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V5N1_1971_E.pdf.

Nesbitt:1971:C

[189] Alexander Nesbitt, Gerrit Noordzij, and John A. Schappler. Correspondence. Visible Language, V(1):82–89, Winter 1971. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V5N1_1971_E.pdf. Discussion of [159] and [109, 110, 154, 244].

Anonymous:1971:Aa

[190] Anonymous. The authors. Visible Language, V(1):90, Winter 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V5N1_1971_E.pdf.

Baudin:1971:RAAa

[191] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, V(1):92, Winter 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V5N1_1971_E.pdf.

Anonymous:1971:BNJ

[192] Anonymous. Back numbers of The Journal of Typographic Research. Visible Language, V(1):93–96, Winter 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V5N1_1971_E.pdf.

Cyrus:1971:LFS

[193] Virginia J. Cyrus. Linguistic features of scribal spacing. Visible Language, V(2):101–110, Spring 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V5N2_1971_E.pdf.

OHara:1971:UHI

[194] F. M. O’Hara, Jr. Use of the hyphen to indicate divided words. Visible Language, V(2):111–124, Spring 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V5N2_1971_E.pdf. See com-

ments [230].

Troxel:1971:ARP

[195] Donald E. Troxel. Automated reading of the printed page. Visible Language, V(2):125–144, Spring 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V5N2_1971_E.pdf.

Cleland:1971:VSR

[196] Donald L. Cleland. Vocalism in silent reading. Visible Language, V (2):145–157, Spring 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N2_1971_E.pdf.

Pollak:1971:DFC

[197] Michael Pollak. Durability of Fifteenth-century type. Visible Language, V(2):159–188, Spring 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N2_1971_E.pdf.

Baudin:1971:RAAb

[198] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, V(2): 189–191, Spring 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N2_1971_E.pdf.

Anonymous:1971:Ab

[199] Anonymous. The authors. Visible Language, V(2):192, Spring 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V5N2_1971_E.pdf.

Anonymous:1971:GIc

[200] Anonymous. General information. Visible Language, V(3):195– 196, Summer 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N3_1971_E.pdf.

Marcus:1971:PCP

[201] Aaron Marcus. A prototype computerized page-design system. Visible Language, V(3):197–220, Summer 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V5N3_1971_E.pdf.

Gordon:1971:LNL

[202] Arthur E. Gordon. Letter names of the Latin alphabet. Visible Language, V(3):221–228, Summer 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V5N3_1971_E.pdf.

Yasuhara:1971:STH

[203] Makoto Yasuhara. Steps toward handwriting analysis and recognition. Visible Language, V(3):229–248, Summer 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V5N3_1971_E.pdf.

Fontein:1971:CCZ

[204] Jan Fontein and Money L. Hickman. Calligraphy of Ch’an and Zen monks. Visible Language, V(3):249–264, Summer 1971. CODEN VSL-GAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V5N3_1971_E.pdf.

Hartley:1971:EUT

[205] James Hartley and Peter Burnhill. Experiments with unjustified text. Visible Language, V(3):265–278, Summer 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V5N3_1971_E.pdf.

Banks:1971:BRN

[206] Colin Banks. Book review: E. Nicolette Gray, Lettering as Drawing: The Moving Line [volume 1] and Contour and Silhouette [volume 2]. London: Oxford University Press, 1970. 96 pages each with many black and white illustrations. 75 new pence. Paperbacks. Visible Language, V (3):279–281, Summer 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N3_1971_E.pdf.

Champney:1971:C

[207] Freeman Champney and Louis Marck. Correspondence. Visible Language, V(3):282–283, Summer 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V5N3_1971_E.pdf. See [170].

Louis Marck also comments on the too-similar appearance of ‘b‘ and ‘h’ in Monotype Garamond (Series 156 Italic).

Baudin:1971:RAAc

[208] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, V(3):285– 287, Summer 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N3_1971_E.pdf.

Anonymous:1971:Ac

[209] Anonymous. The authors. Visible Language, V(3):288, Summer 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V5N3_1971_E.pdf.

Anonymous:1971:GId

[210] Anonymous. General information. Visible Language, V(4):291– 292, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N4_1971_E.pdf.

Coe:1971:AMW

[211] Michael D. Coe. Ancient Maya writing and calligraphy. Visible Language, V(4):293–307, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V5N4_1971_E.pdf.

GT:1971:DNJ

[212] Group Typo. Design of a new Japanese typeface: Typos. Visible Language, V(4):309–317, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V5N4_1971_E.pdf.

Basu:1971:LMA

[213] Hella Basu. Letterforms as a medium for artistic expression. Visible Language, V(4):319–330, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V5N4_1971_E.pdf.

Anderson:1971:CHC

[214] Donald M. Anderson. Cresci and his capital alphabet. Visible Language, V(4):331–352, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V5N4_1971_E.pdf.

Robinson:1971:WSI

[215] David Owen Robinson, Michael Abbamonte, and Selby H. Evans. Why serifs are important: the perception of small print. Visible Language, V (4):353–359, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N4_1971_E.pdf.

Bauries:1971:RBG

[216] Fred Bauries. Research in brief: Graphic trends in consumer magazine advertising, 1938–1968. Visible Language, V(4):360–364, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V5N4_1971_E.pdf.

Smith:1971:RBG

[217] John M. Smith and Maxwell E. McCombs. Research in brief: The graphics of prose. Visible Language, V(4):365–369, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V5N4_1971_E.pdf.

Banks:1971:BRE

[218] Colin Banks. Book review: Edward Johnstone, Formal Penmanship. London: Lund Humphries, 1971. 160 pages of text, 20 pages of plates. £5.80. Visible Language, V(4):370–372, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V5N4_1971_E.pdf. See [254]: Johnstone should be Johnston.

vanKrimpen:1971:C

[219] Huib van Krimpen, The Educational Committee of A. Typ. I., and Peter Burnhill. Correspondence. Visible Language, V(4):373–376, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V5N4_1971_E.pdf. Comments on the financial and copyright relations between type designers and typefounders, on the teaching of writing and lettering, and disapproval of the journal name change.

Baudin:1971:RAAd

[220] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, V(4):377– 379, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N4_1971_E.pdf.

Anonymous:1971:Ad

[221] Anonymous. The authors. Visible Language, V(4):380, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V5N4_1971_E.pdf.

Anonymous:1971:IVV

[222] Anonymous. Index to Volume V. Visible Language, V(4):381– 384, Autumn 1971. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V5N4_1971_E.pdf.

Anonymous:1972:GIa

[223] Anonymous. General information. Visible Language, VI(1):3–4, Winter 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V6N1_1972_E.pdf.

Powell:1972:OSS

[224] Marvin A. Powell, Jr. The origin of the sexagesimal system: Interaction of language and writing. Visible Language, VI(1):5–18, Winter 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V6N1_1972_E.pdf.

Seaman:1972:DVP

[225] David W. Seaman. The development of visual poetry in France. Visible Language, VI(1):19–44, Winter 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N1_1972_E.pdf.

Furby:1972:DWP

[226] Lita Furby. Development of word perception and problem solving strategies. Visible Language, VI(1):45–58, Winter 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V6N1_1972_E.pdf.

Wells:1972:WBR

[227] James Wells. The work of Bruce Rogers. Visible Language, VI(1): 59–78, Winter 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N1_1972_E.pdf.

Lotz:1972:SBS

[228] John Lotz. A select bibliography on script and language. Visible Language, VI(1):79–80, Winter 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V6N1_1972_E.pdf.

Stauffacher:1972:HAT

[229] Jack W. Stauffacher. Homage to Alberto Tallone, 1898–1960. Visible Language, VI(1):81–89, Winter 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N1_1972_E.pdf.

Tanselle:1972:C

[230] G. Thomas Tanselle. Correspondence. Visible Language, VI(1): 90–91, Winter 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N1_1972_E.pdf. See [194, 254].

Baudin:1972:RAAa

[231] Fernand Baudin, Dirk Wendt, and Tony Évora. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, VI(1):92–95, Winter 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N1_1972_E.pdf.

Anonymous:1972:Aa

[232] Anonymous. The authors. Visible Language, VI(1):96, Winter 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V6N1_1972_E.pdf.

Wrolstad:1972:ER

[233] Merald E. Wrolstad. Editor’s report. Visible Language, VI(2):99– 108, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N2_1972_E.pdf. See comments [313].

Solt:1972:TVC

[234] Mary Ellen Solt. Typography and the visual concrete poem. Visible Language, VI(2):109–122, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N2_1972_E.pdf.

David:1972:PDL

[235] Robert E. David. Proposal for a diagrammatic language for design. Visible Language, VI(2):123–137, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N2_1972_E.pdf.

Lamaitre:1972:ACM

[236] ?. Lamaître. Albert Camus and the men of the stone. Visible Language, VI(2):138, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N2_1972_E.pdf.

Stern:1972:OPE

[237] Henry R. Stern. The orthographic practices of Elias Molee. Visible Language, VI(2):139–149, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N2_1972_E.pdf. See letter

[313].

Scott:1972:GPC

[238] Myrtle Scott. Grapheme–phoneme correspondence in beginning reading of disadvantaged five-year-olds. Visible Language, VI(2):151–165, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V6N2_1972_E.pdf. See comments and corrections [301].

Anonymous:1972:IYB

[239] Anonymous. [1972 — international year of the book]. Visible Language, VI(2):166, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N2_1972_E.pdf.

Wendt:1972:EIU

[240] Dirk Wendt and Hans Weckerle. On effects of indentation and underlining in reference work. Visible Language, VI(2):167–171, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V6N2_1972_E.pdf.

Coleman:1972:BBC

[241] A. D. Coleman. Bonnie Baldwin Collier’s delightful Thesis Box. Visible Language, VI(2):172, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/v6n2_1972_e.pdf.

vanKrimpen:1972:BRA

[242] Huib van Krimpen and Fernand Baudin. Book reviews: Albert Kapr. Schriftkunst. Geschichte, Anatomie und Schønheit der lateinischen Buchstaben. Dresden: VEB Verlag der Kunst, 1971. 11 3/4 × 8 1/2 inches, 468 pages, 465 illustrations, 321 type specimens. Price 70 MDN (approx. $18.00). W. P. Jaspert, W. T. Berry, A F. Johnson, The Encyclopaedia of Typefaces, 4th ed. (London: Blandford Press, 1970). Visible Language, VI(2):173–181, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N2_1972_E.pdf.

Anonymous:1972:INI

[243] Anonymous. ICOGRAPHIC: a New International Design Quarterly. Visible Language, VI(2):181, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N2_1972_E.pdf.

Nadler:1972:C

[244] Morton Nadler, Dirk Wendt, and Trilokesh Mukherjee. Correspondence. Visible Language, VI(2):182–187, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N2_1972_E.pdf. See [109,

189, 219].

Baudin:1972:RAAb

[245] Fernand Baudin, Dirk Wendt, and Tony Évora. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, VI(2):188–191, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N2_1972_E.pdf.

Anonymous:1972:Ab

[246] Anonymous. The authors. Visible Language, VI(2):192, Spring 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V6N2_1972_E.pdf.

Anonymous:1972:GIc

[247] Anonymous. General information. Visible Language, VI(3):195– 196, Summer 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N3_1972_E.pdf.

Steiner:1972:AB

[248] George Steiner. After the book? Visible Language, VI(3):197– 210, Summer 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N3_1972_E.pdf. See replies [249].

Brain:1972:FRG

[249] John V. Brain, Norman S. Fiering, John Freeman, and Leland E. Warren. Four replies to George Steiner’s address to the Ferguson Seminar in Publishing. Visible Language, VI(3):211–228, Summer 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V6N3_1972_E.pdf. See [248].

Greene:1972:MMO

[250] Gordon K. Greene. From mistress to master: The origins of polyphonic music and as a visible language. Visible Language, VI(3): 229–253, Summer 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N3_1972_E.pdf.

Ohlgren:1972:VLO

[251] Thomas H. Ohlgren. Visual language in the Old English Cadmonian Genesis. Visible Language, VI(3):253–276, Summer 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V6N3_1972_E.pdf.

Anonymous:1972:CRC

[252] Anonymous. On the cover: a rubbing of a Claudian bronze tablet. Visible Language, VI(3):277–278, Summer 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N3_1972_E.pdf.

Anonymous:1972:EDP

[253] Anonymous. Experimental design pages. Visible Language, VI(3): 280, Summer 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N3_1972_E.pdf.

Gilman:1972:C

[254] E. W. Gilman, Alf K. Ebsen, Peter Burnhill, Roy Brewer, Colin Banks, Mark Ash II, and Harry Duncan. Correspondence. Visible Language, VI (3):281–284, Summer 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N3_1972_E.pdf. See [230, 218, 169].

Baudin:1972:RAAc

[255] Fernand Baudin, Dirk Wendt, and Tony Évora. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, VI(3):285–287, Summer 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N3_1972_E.pdf.

Anonymous:1972:Ac

[256] Anonymous. The authors. Visible Language, VI(3):288, Summer 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V6N3_1972_E.pdf.

Gough:1972:OSR

[257] Philip B. Gough. One second of reading. Visible Language, VI(4): 291–320, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N4_1972_E.pdf.

Marcus:1972:TEC

[258] Susan Marcus. The typographic element in cubism, 1911–1915: Its formal and semantic implications. Visible Language, VI(4):321– 340, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N4_1972_E.pdf.

Ovink:1972:CRT

[259] G. W. Ovink. Changing responsibilities of the typographic designer. Visible Language, VI(4):341–354, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N4_1972_E.pdf.

Bedno:1972:PDV

[260] Ed Bedno. Program for developing visual symbols. Visible Language, VI (4):355–363, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N4_1972_E.pdf.

Anonymous:1972:DVS

[261] Anonymous. Design a visual symbol for ‘visible language’ — win a ten-year subscription to Visible Language. Visible Language, VI (4):364, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N4_1972_E.pdf.

Bourdon:1972:CWA

[262] David Bourdon. Cover: Words about Ed Ruscha. Visible Language, VI (4):365–368, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N4_1972_E.pdf.

Foster:1972:BRF

[263] Jeremy J. Foster. Book reviews: Frank Smith, Understanding Reading: a Psycholinguistic Analysis of Reading and Learning to Read, New York: Holt, Rinehart & Winston, 1971. xiv + 239 pages, $4.95 paper. John Gilliland, Readability. London: University of London Press, 1972. 127 pages. 80p. Visible Language, VI(4):369–373, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V6N4_1972_E.pdf.

Greenfield:1972:TEW

[264] Patricia M. Greenfield and Jerome S. Bruner. The training embodied in written language. Visible Language, VI(4):374, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V6N4_1972_E.pdf. Excerpt from the authors’ Work with the Wolog, Psychology Today (July 1971), pages 78–79.

Anonymous:1972:GId

[265] Anonymous. General information. Visible Language, VI(4):375– 376, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N4_1972_E.pdf.

Baudin:1972:RAAd

[266] Fernand Baudin, Dirk Wendt, and Tony Évora. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, VI(4):377–379, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V6N4_1972_E.pdf.

Anonymous:1972:Ad

[267] Anonymous. The authors. Visible Language, VI(4):380, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V6N4_1972_E.pdf.

Anonymous:1972:IVV

[268] Anonymous. Index to Volume VI. Visible Language, VI(4):381– 384, Autumn 1972. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V6N4_1972_E.pdf.

Anonymous:1973:WBM

[269] Anonymous. Have we bitten off more than we can chew? A reappraisal of Visible Language for its 25th issue. Visible Language, VII (1):5–18, Winter 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N1_1973_E.pdf. Commentary on the journal

by numerous letter writers.

Olson:1973:GLP

[270] Richard A. Olson. Greek letterforms on the Parthian drachms. Visible Language, VII(1):19–40, Winter 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N1_1973_E.pdf.

Tritt:1973:LCS

[271] Carleton S. Tritt. The language of capitalization in Shakespeare’s first folio. Visible Language, VII(1):41–50, Winter 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V7N1_1973_E.pdf.

Poggenpohl:1973:VLE

[272] Sharon H. Poggenpohl. Visible Language: an experimental course. Visible Language, VII(1):51–61, Winter 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N1_1973_E.pdf.

Proust:1973:ROE

[273] Marcel Proust. “Reading, in its original essence… ”. Visible Language, VII(1):62, Winter 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N1_1973_E.pdf. From the author’s On Read-

ing (translated and edited by Jean Autret and William Burford). New York: Macmillan, 1971; p. 31.

Timko:1973:DTT

[274] Henry G. Timko. The discrimination of three types of graphic stimuli. Visible Language, VII(1):63–72, Winter 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N1_1973_E.pdf.

Hochuli:1973:BRE

[275] Jost Hochuli. Book review: Edward M. Catich’s The Origin of the Serif Davenport, Iowa: The Catfish Press, St. Ambrose College, 1968. 8 1/2 × 11 1/4 inches, xii + 310 pages. $24.00. Visible Language, VII (1):73–91, Winter 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N1_1973_E.pdf. See letter [322].

Baudin:1973:RAAa

[276] Fernand Baudin, Dirk Wendt, and Tony Évora. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, VII(1):93–95, Winter 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N1_1973_E.pdf.

Anonymous:1973:Aa

[277] Anonymous. The authors. Visible Language, VII(1):96, Winter 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V7N1_1973_E.pdf.

Anonymous:1973:GIb

[278] Anonymous. General information. Visible Language, VII(2):99– 100, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N2_1973_E.pdf.

Hewes:1973:EFR

[279] Gordon W. Hewes. An explicit formulation of the relationship between tool-using, tool-making, and emergence of language. Visible Language, VII(2):101–127, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V7N2_1973_E.pdf.

Walker:1973:QB

[280] Martin Walker. Qwerty birthday. Visible Language, VII(2):128, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V7N2_1973_E.pdf. A hundred years ago last week Christopher Latham Sholes sold the idea of a production typewriter to Remington, the American gun and sewing machine worker.

Friedman:1973:VIE

[281] Daniel Friedman. A view: Introductory education in typography. Visible Language, VII(2):129–144, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N2_1973_E.pdf.

Downing:1973:LAE

[282] John Downing. Is literacy acquisition easier in some languages than in others? Visible Language, VII(2):144–154, Spring 1973. CODEN VSL-GAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V7N2_1973_E.pdf.

Barber:1973:FEW

[283] E. J. W. Barber. Formal economy of written signs. Visible Language, VII (2):155–166, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N2_1973_E.pdf.

Hartley:1973:RBS

[284] James Hartley, Susan J. Timson, and Peter Burnhill. Research in brief: Subjective preference and retrieval of information from reference materials. Visible Language, VII(2):167–170, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V7N2_1973_E.pdf. Adapted from the introduction to the author’s Archives in the Ancient World, Cambridge, Mass.: Harvard University Press.

Posner:1973:EAA

[285] Ernest Posner. Excerpt: Archives in the Ancient World. Visible Language, VII(2):171–178, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V7N2_1973_E.pdf.

Moran:1973:CFR

[286] James Moran. Comment: The future role of the printing historical society. Visible Language, VII(2):179–180, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V7N2_1973_E.pdf.

Standard:1973:BRO

[287] Paul Standard. Book review: A. S. Osley: LUMINARIO. An Introduction to the Italian Writing-Books of the 16th and 17th Centuries. xiii + 173 pp. with 116 illustrations; plus checklist of 105 first edns; bibl. of 113 items, & full index; clothbound sm. folio with paper wrapper. Nieuwkoop/Netherlands: Miland Publishers, 1972. Price £21.50 or $56.50. Visible Language, VII(2):181–188, Spring 1973. CODEN VSL-GAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V7N2_1973_E.pdf.

Praz:1973:SC

[288] Mario Praz. A secret of calligraphy. Visible Language, VII(2):189, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V7N2_1973_E.pdf. Extract from the author’s Mnemosyne: The Parallel between Literature and the Visual Arts. Princeton: Princeton University Press, 1970; pp. 25–27.

Baudin:1973:RAAb

[289] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, VII(2): 190–191, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N2_1973_E.pdf.

Anonymous:1973:Ab

[290] Anonymous. The authors. Visible Language, VII(2):192, Spring 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V7N2_1973_E.pdf.

Anonymous:1973:APP

[291] Anonymous. Announcement: Mit press to publish Visible Language. Visible Language, VII(3):195, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N3_1973_E.pdf.

Anonymous:1973:GIc

[292] Anonymous. General information. Visible Language, VII(3):197– 198, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N3_1973_E.pdf.

Harary:1973:T

[293] Frank Harary. Typographs. Visible Language, VII(3):199–208, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V7N3_1973_E.pdf. This article was received for publication on January 11, 1970, at which time the name of this journal was as in reference [2] (Journal of Typographic Research). It was circulated in preprint form and inspired the insightful and interesting paper: J. A. Bondy, The “graph theory” of the Greek alphabet. Graph theory and applications (Y. Alavi, et al., eds.), Berlin: Springer, 1972. Pp. 43–54. This article was prepared with the assistance of Linda Bidelman.

Blesser:1973:CRB

[294] B. Blesser, R. Shillman, C. Cox, T. Kuklinsky, J. Ventura, and M. Eden. Character recognition based on phenomenological attributes. Visible Language, VII(3):209–223, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N3_1973_E.pdf.

Golomb:1973:FBF

[295] Solomon W. Golomb. Forth and back and forth. Visible Language, VII (3):224, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N3_1973_E.pdf. Excerpted from an article by

the author in the Harvard Bulletin, March 1972.

Douglass:1973:RBB

[296] Malcolm P. Douglass. Reading between and beyond the lines. Visible Language, VII(3):225–234, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N3_1973_E.pdf. Adapted

from the author’s keynote address to the Fortieth Annual Claremont Reading Conference (Claremont, Calif. 91711) in February 1973.

Viehmeyer:1973:GLC

[297] L. Allen Viehmeyer. Gothic letterforms and Codex Vindobonensis. Visible Language, VII(3):235–246, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N3_1973_E.pdf.

Cushman:1973:CST

[298] Donald R. Cushman. The cue summation theory tested with meaningful verbal information. Visible Language, VII(3):247–260, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V7N3_1973_E.pdf.

Venturi:1973:ELV

[299] Robert Venturi, Denise Scott Brown, and Steven Izenour. Excerpt: Learning from Las Vegas. Visible Language, VII(3):261–270, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V7N3_1973_E.pdf. Excerpted from the authors’ Learning from Las Vegas (Cambridge, Mass.: MIT Press, 1972), xviii + 188 pages, $25.

Neruda:1973:OT

[300] Pablo Neruda. Ode to typography. Visible Language, VII(3):271– 275, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N3_1973_E.pdf. Translated from the Spanish

by Carlos Lozano; it appeared in the Chicago Review, Vol. 17, No. 1 (1964).

Downing:1973:C

[301] John Downing. Correspondence. Visible Language, VII(3):276– 277, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N3_1973_E.pdf. See [238].

Baudin:1973:RAAc

[302] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, VII(3): 278–279, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N3_1973_E.pdf.

Anonymous:1973:Ac

[303] Anonymous. The authors. Visible Language, VII(3):280, Summer 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V7N3_1973_E.pdf.

Anonymous:1973:GId

[304] Anonymous. General information. Visible Language, VII(4):283– 284, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N4_1973_E.pdf.

Hoffman:1973:VSI

[305] William E. Hoffman. Vision, sign, and inference. Visible Language, VII (4):285–309, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N4_1973_E.pdf.

Coleman:1973:G

[306] A. Coleman. Graphicacy. Visible Language, VII(4):310, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V7N4_1973_E.pdf. Excerpted from the author’s note in the December 1972 Cartographic Journal, Journal of the British Cartographic Society.

Kindersley:1973:SC

[307] David Kindersley. Space craft. Visible Language, VII(4):311– 324, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N4_1973_E.pdf.

Lehman:1973:HLM

[308] C. L. Lehman. Handwriting legibility: a method of objective evaluation. Visible Language, VII(4):325–344, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V7N4_1973_E.pdf.

Pollack:1973:PDI

[309] Michael Pollack. The Prophetess Deborah and the invention of printing. Visible Language, VII(4):345–350, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V7N4_1973_E.pdf.

Wright:1973:RBU

[310] Patricia Wright. Research in brief: Understanding tabular displays. Visible Language, VII(4):351–359, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N4_1973_E.pdf.

Conway:1973:CSG

[311] Patricia Conway. Cover: Subway graffiti in New York City. Visible Language, VII(4):360–362, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V7N4_1973_E.pdf. From an

article by the author, Subway Graffiti: The Message from Underground. Print, XXVII (May/June 1973), 25–32.

Banks:1973:BRA

[312] Colin Banks. Book review: Allen Hutt. Fournier: The Compleat Typographer. London: Frederick Muller Ltd, 1973 £4.50. Visible Language, VII (4):363–364, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N4_1973_E.pdf. The author of this book,

Allen Hutt, died in August, 1973.

Haugen:1973:C

[313] Einar Haugen, Louis Marck, and David Abercrombie. Correspondence. Visible Language, VII(4):365–369, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V7N4_1973_E.pdf. See

[237, 233].

Baudin:1973:RAAd

[314] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, VII(4): 370–371, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N4_1973_E.pdf.

Anonymous:1973:Ad

[315] Anonymous. The authors. Visible Language, VII(4):372, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V7N4_1973_E.pdf.

Anonymous:1973:IVV

[316] Anonymous. Index to Volume VI. Visible Language, VII(4):373– 376, Autumn 1973. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V7N4_1973_E.pdf.

Anonymous:1974:GIa

[317] Anonymous. General information. Visible Language, VIII(1):3–4, Winter 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V8N1_1974_E.pdf.

Herrick:1974:TAS

[318] Earl M. Herrick. A taxonomy of alphabets and scripts. Visible Language, VIII(1):5–32, Winter 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N1_1974_E.pdf.

Haugen:1974:RSS

[319] Einar Haugen. The rune stones of Spirit Pond, Maine. Visible Language, VIII(1):33–64, Winter 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N1_1974_E.pdf.

Turner:1974:EGM

[320] E. G. Turner. Excerpt: Greek manuscripts of the Ancient World. Visible Language, VIII(1):65–82, Winter 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V8N1_1974_E.pdf.

Reineck:1974:DBW

[321] Gay Beste Reineck. Design in brief: Word/forms. Visible Language, VIII (1):85–88, Winter 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N1_1974_E.pdf.

Palladino:1974:C

[322] Robert J. Palladino. Correspondence. Visible Language, VIII(1): 89–93, Winter 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N1_1974_E.pdf. See [275].

Baudin:1974:RAAa

[323] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, VIII(1):

94–95, Winter 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N1_1974_E.pdf.

Anonymous:1974:Aa

[324] Anonymous. The authors. Visible Language, VIII(1):96, Winter 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V8N1_1974_E.pdf.

Anonymous:1974:GIb

[325] Anonymous. General information. Visible Language, VIII(2):99– 100, Spring 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N2_1974_E.pdf.

Barganz:1974:POR

[326] Robert A. Barganz. Phonological and orthographic relationships of reading performance. Visible Language, VIII(2):101–122, Spring 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V8N2_1974_E.pdf.

Bormuth:1974:LPE

[327] John R. Bormuth. Literacy policy and the emerging technology of readability. Visible Language, VIII(2):123–135, Spring 1974. CODEN VSL-GAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V8N2_1974_E.pdf.

Anonymous:1974:BAG

[328] Anonymous. [Boston area gravestone]. Visible Language, VIII(2): 136, Spring 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N2_1974_E.pdf.

Duval:1974:IOP

[329] Francis Duval and Ivan Rigby. Inscriptions of our past. Visible Language, VIII(2):137–150, Spring 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V8N2_1974_E.pdf.

Shillman:1974:BCR

[330] R. Shillman, C. Cox, T. Kuklinski, J. Ventura, M. Eden, and B. Blesser. A bibliography in character recognition: Techniques for describing characters. Visible Language, VIII(2):151–166, Spring 1974. CODEN VSL-GAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V8N2_1974_E.pdf.

Raben:1974:HCL

[331] Joseph Raben. The humanist in the computer lab. Visible Language, VIII (2):167–177, Spring 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N2_1974_E.pdf.

Sutherland:1974:PSS

[332] Eric Sutherland. [photograph of Sentence Structures]. Visible Language, VIII(2):178, Spring 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N2_1974_E.pdf.

Lettieri:1974:EWC

[333] Larry Lettieri. Excerpt: Which computer printer when? Visible Language, VIII(2):179–185, Spring 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V8N2_1974_E.pdf.

Banks:1974:BRR

[334] Colin Banks. Book review: Ruari McLean. Victorian Book Design and Colour Printing. London: Faber & Faber, 1973. £15. Visible Language, VIII(2):187–188, Spring 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V8N2_1974_E.pdf.

Baudin:1974:RAAb

[335] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, VIII(2): 189–191, Spring 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N2_1974_E.pdf.

Anonymous:1974:Ab

[336] Anonymous. The authors. Visible Language, VIII(2):192, Spring 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V8N2_1974_E.pdf.

Anonymous:1974:GIc

[337] Anonymous. General information. Visible Language, VIII(3):195– 196, Summer 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N3_1974_E.pdf.

Beiman:1974:CPS

[338] Abbie W. Beiman. Concrete poetry: a study in metaphor. Visible Language, VIII(3):197–223, Summer 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V8N3_1974_E.pdf.

Longuet-Higgins:1974:LM

[339] H. C. Longuet-Higgins. The language of music. Visible Language, VIII (3):224, Summer 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N3_1974_E.pdf. Excerpted from the author’s

“The Language of Music,” The Times Literary Supplement, November 20, 1970, pp. 1351–1352.

Frye:1974:RB

[340] Northrop Frye. The renaissance of books. Visible Language, VIII(3): 225–240, Summer 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N3_1974_E.pdf.

Pearson:1974:WRL

[341] P. David Pearson and Michael L. Kamil. Word recognition latencies as a function of form class, stem length, and affix length. Visible Language, VIII(3):241–246, Summer 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V8N3_1974_E.pdf.

Gray:1974:LS

[342] Nicolete Gray. Lettering and society. Visible Language, VIII(3):247– 260, Summer 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N3_1974_E.pdf.

Crouwel:1974:PEL

[343] Wim Crouwel. A proposition for education in letterforms and handwriting. Visible Language, VIII(3):261–264, Summer 1974. CODEN VSL-GAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V8N3_1974_E.pdf.

Ehri:1974:RBP

[344] Linnea C. Ehri and Lee S. Wilce. Research in brief: Printed intonation cues and reading in children. Visible Language, VIII(3):265–274, Summer 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V8N3_1974_E.pdf.

Clay:1974:RBO

[345] Marie M. Clay. Research in brief: Orientation to the spatial characteristics of the open book. Visible Language, VIII(3):275–282, Summer 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V8N3_1974_E.pdf.

Wendt:1974:KBA

[346] Dirk Wendt and Fernand Baudin. Kurzfassung der Beiträge [abstracts of journal articles in German and French]. Visible Language, VIII(3): 283–285, Summer 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N3_1974_E.pdf.

Anonymous:1974:Ac

[347] Anonymous. The authors. Visible Language, VIII(3):287, Summer 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V8N3_1974_E.pdf.

Anonymous:1974:GId

[348] Anonymous. General information. Visible Language, VIII(4):291– 292, Autumn 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N4_1974_E.pdf.

Gelb:1974:RWD

[349] I. J. Gelb. Records, writing, and decipherment. Visible Language, VIII (4):293–318, Autumn 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N4_1974_E.pdf.

Germain:1974:FSI

[350] Edward Germain. Four surrealist images. Visible Language, VIII(4): 319–332, Autumn 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N4_1974_E.pdf.

Marcus:1974:IVS

[351] Aaron Marcus. An introduction to the visual syntax of concrete poetry. Visible Language, VIII(4):333–360, Autumn 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V8N4_1974_E.pdf. See

comments [369, 389].

damen:1974:ETT

[352] herman damen. excerpt: towards a three-dimensional pouitry. Visible Language, VIII(4):361–368, Autumn 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V8N4_1974_E.pdf.

Kostka:1974:AG

[353] Robert Kostka. Aspects of graffiti. Visible Language, VIII(4):369– 375, Autumn 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N4_1974_E.pdf.

Baudin:1974:RAAd

[354] Fernand Baudin, Dirk Wendt, and Felix Beltran. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, VIII(4):376–378, Autumn 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V8N4_1974_E.pdf.

Anonymous:1974:Ad

[355] Anonymous. The authors. Visible Language, VIII(4):380, Autumn 1974. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V8N4_1974_E.pdf.

Anonymous:1974:IVV

[356] Anonymous. Index to Volume VIII. Visible Language, VIII(4):382– 384, Autumn 1974. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V8N4_1974_E.pdf.

Anonymous:1975:GIa

[357] Anonymous. General information. Visible Language, IX(1):3–4, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V9N1_1975_E.pdf.

Anonymous:1975:BC

[358] Anonymous. Brief communications. Visible Language, IX(1):5–6, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V9N1_1975_E.pdf.

Venezky:1975:CRL

[359] Richard L. Venezky. The curious role of letter names in reading instruction. Visible Language, IX(1):7–23, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V9N1_1975_E.pdf. See

comments [398].

Steig:1975:CDB

[360] William Steig. C D B! Visible Language, IX(1):24, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V9N1_1975_E.pdf. Reprint of letter cartoons from the author’s book of the same title, New York: Windmill Books, Inc., Simon & Schuster, Inc. (1968).

King:1975:RDW

[361] Terrance J. King. Radial design in Wallace Stevens. Visible Language, IX(1):25–46, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N1_1975_E.pdf.

Dole:1975:CWD

[362] William Dole and Gerald Nordland. The collages of William Dole. Visible Language, IX(1):47–56, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V9N1_1975_E.pdf.

Hauser:1975:PRL

[363] Robert A. Hauser. Photographic restoration of letterforms. Visible Language, IX(1):57–66, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V9N1_1975_E.pdf.

Groff:1975:RBS

[364] Patrick Groff. Research in brief: Shapes as cues to word recognition. Visible Language, IX(1):67–71, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V9N1_1975_E.pdf.

Charyn:1975:AFL

[365] Jerome Charyn. Azazian is a frenetic language …. Visible Language, IX(1):72, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N1_1975_E.pdf. Except from the author’s

book, Eisenhower, My Eisenhower, New York: Holt, Rinehart & Winston.

Doblin:1975:ESA

[366] Jay Doblin, Inder Agrawal, Marianna Porter, and Robert Peterson. Excerpt: Simplifying the ABC’s. Visible Language, IX(1):73–86, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V9N1_1975_E.pdf. Excerpted from an article in the January/February 1974 issue of Industrial Design magazine.

Nesbitt:1975:CDL

[367] Alexander Nesbitt. Comment: The designer and language. Visible Language, IX(1):87–89, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V9N1_1975_E.pdf.

sandberg:1975:AT

[368] willem sandberg. art and typography. Visible Language, IX(1):90, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V9N1_1975_E.pdf.

Mayer:1975:Ca

[369] Peter Mayer, James Mosley, and Hermann Zapf. Correspondence. Visible Language, IX(1):91–92, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V9N1_1975_E.pdf. Commen-

tary on the phrase “26 (more or less) soldiers (of lead).”, plus Zapf’s disapproval of the cover, and threat to resign from the Advisory Board. See response to Zapf [379].

Baudin:1975:RAAa

[370] Fernand Baudin, Dirk Wendt, and Felix Beltran. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, IX(1):93–95, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V9N1_1975_E.pdf.

Anonymous:1975:Aa

[371] Anonymous. The authors. Visible Language, IX(1):96, Winter 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V9N1_1975_E.pdf.

Anonymous:1975:GIb

[372] Anonymous. General information. Visible Language, IX(2):99– 100, Spring 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N2_1975_E.pdf.

Conley:1975:VSP

[373] Tom Conley. Verbal shape in the poetry of Villon and Marot. Visible Language, IX(2):101–122, Spring 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V9N2_1975_E.pdf.

Antonsen:1975:IWS

[374] Elmer H. Antonsen. The inscription of the Whetstone from Strøm. Visible Language, IX(2):123–132, Spring 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V9N2_1975_E.pdf.

Herrick:1975:LAB

[375] Earl M. Herrick. Letters with alternative basic shapes. Visible Language, IX(2):133–144, Spring 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N2_1975_E.pdf.

Suen:1975:HEB

[376] Ching Y. Suen. Handwriting education — a bibliography of contemporary publications. Visible Language, IX(2):145–158, Spring 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V9N2_1975_E.pdf.

ODonnell:1975:MRC

[377] Thomas D. O’Donnell. Maurice Roche: Crâne, carne. Visible Language, IX(2):159–172, Spring 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N2_1975_E.pdf.

AIGA:1975:DPP

[378] American Institute of Graphic Arts. The development of passenger/ pedestrian oriented symbols for use in transportation-related facilities. Visible Language, IX(2):173–185, Spring 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V9N2_1975_E.pdf.

Mayer:1975:Cb

[379] Peter Mayer and Ken Komai. Correspondence. Visible Language, IX (2):186–188, Spring 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N2_1975_E.pdf. See [351, 369].

Baudin:1975:RAAb

[380] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, IX(2): 189–191, Spring 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N2_1975_E.pdf.

Anonymous:1975:Ab

[381] Anonymous. The authors. Visible Language, IX(2):192, Spring 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V9N2_1975_E.pdf.

Anonymous:1975:GIc

[382] Anonymous. General information. Visible Language, IX(3):195– 196, Summer 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N3_1975_E.pdf.

Steinberg:1975:RBS

[383] Danny D. Steinberg and Miho T. Steinberg. Reading before speaking. Visible Language, IX(3):197–224, Summer 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V9N3_1975_E.pdf.

Moxley:1975:AWS

[384] Roy A. Moxley, Jr. Acquisition of writing skills. Visible Language, IX (3):225–248, Summer 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N3_1975_E.pdf.

Mendel:1975:LTI

[385] Mark Mendel. Line transmitter installation — a poem in the environment. Visible Language, IX(3):249–262, Summer 1975. CODEN VSL-GAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-

us-west-2.amazonaws.com/visiblelanguage/pdf/V9N3_1975_E.pdf.

Kristeller:1975:MRR

[386] Paul Oskar Kristeller. Methods of research in renaissance manuscripts. Visible Language, IX(3):263–275, Summer 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V9N3_1975_E.pdf.

Anno:1975:EAA

[387] Mitsumasa Anno. Excerpt from Anno’s Alphabet: An Adventure in Imagination. Visible Language, IX(3):276, Summer 1975. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V9N3_1975_E.pdf. Photographs from the author’s book, New York: Thomas Y. Crowell Company.

Mountford:1975:CMA

[388] John Mountford. Comment: The medial aspect of language: a linguistic framework for literacy. Visible Language, IX(3):277–281, Summer 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V9N3_1975_E.pdf.

Kinniburg:1975:C

[389] Ian A. G. Kinniburg. Correspondence. Visible Language, IX(3):282– 285, Summer 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N3_1975_E.pdf. See [351].

Baudin:1975:RAAc

[390] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, IX(3):285– 287, Summer 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N3_1975_E.pdf.

Anonymous:1975:Ac

[391] Anonymous. The authors. Visible Language, IX(3):288, Summer 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V9N3_1975_E.pdf.

Anonymous:1975:GId

[392] Anonymous. General information. Visible Language, IX(4):291– 292, Autumn 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N4_1975_E.pdf.

Watt:1975:WPC

[393] W. C. Watt. What is the proper characterization of the alphabet? I. Desiderata. Visible Language, IX(4):293–327, Autumn 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V9N4_1975_E.pdf.

Anonymous:1975:RC

[394] Anonymous. The research connection. Visible Language, IX(4): 328, Autumn 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N4_1975_E.pdf.

Ascher:1975:QVL

[395] Marcia Ascher and Robert Ascher. The quipu as a visible language. Visible Language, IX(4):329–356, Autumn 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V9N4_1975_E.pdf.

Robson:1975:OWW

[396] Ernest M. Robson. An orthographic way of writing English prosody. Visible Language, IX(4):357–372, Autumn 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V9N4_1975_E.pdf.

Feagler:1975:RTN

[397] Dick Feagler. RTA — the transit network? Visible Language, IX(4): 373–374, Autumn 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N4_1975_E.pdf. Comment on the logos of the

Regional Transit Authority and the Radio Corporation of America.

Arnheim:1975:C

[398] Rudolf Arnheim and Peter Mayer. Correspondence. Visible Language, IX (4):375–377, Autumn 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N4_1975_E.pdf. Arnheim comments on the

orientation of roman capitals, and Mayer on letternames [359].

Baudin:1975:RAAd

[399] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, IX(4):378– 379, Autumn 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N4_1975_E.pdf.

Anonymous:1975:Ad

[400] Anonymous. The authors. Visible Language, IX(4):380, Autumn 1975. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V9N4_1975_E.pdf.

Anonymous:1975:IVI

[401] Anonymous. Index to Volume IX. Visible Language, IX(4):381– 384, Autumn 1975. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V9N4_1975_E.pdf.

Anonymous:1976:GIa

[402] Anonymous. General information. Visible Language, X(1):3–4, Winter 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V10N1_1976_E.pdf.

Wrolstad:1976:MVL

[403] Merald E. Wrolstad. A manifesto for visible language. Visible Language, X(1):5–40, Winter 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N1_1976_E.pdf. See comments [433].

Gaskell:1976:NLR

[404] Phillip Gaskell. Nomenclature of the letterforms of roman type. Visible Language, X(1):41–51, Winter 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V10N1_1976_E.pdf. See com-

ments [433].

Oldenburg:1976:AGC

[405] Claes Oldenburg. Alphabet goodhumor — cloth study, 1972–73. Visible Language, X(1):52, Winter 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N1_1976_E.pdf. Photo-

graph of art work.

White:1976:ASA

[406] John J. White. The argument for a semiotic approach to shape writing: The case of Italian futurist typography. Visible Language, X (1):53–86, Winter 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N1_1976_E.pdf.

vanKrimpen:1976:BRJ

[407] Huib van Krimpen. Book review: Jan Tschichold on Typography: Jan Tschichold, Ansgewählte Anfsätze über Fragen der Gestalt des Buches und der Typographie (Selected Essays on Book Design and Typography). Basel: Birkhäuser Verlag, 1975; ISBN 3-7643-0711-0, 23 Swiss Francs. Visible Language, X(1):87–90, Winter 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N1_1976_E.pdf.

Baudin:1976:RAAa

[408] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, X(1): 91–93, Winter 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N1_1976_E.pdf.

Anonymous:1976:Aa

[409] Anonymous. The authors. Visible Language, X(1):95, Winter 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V10N1_1976_E.pdf.

Anonymous:1976:BS

[410] Anonymous. Burma Shave. Visible Language, X(1):96, Winter 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V10N1_1976_E.pdf. Photograph of widely-seen advertising sign from Frank Rowsome, Jr., The Verse by the Side of the Road: The Story of the Burma-Shave Signs and Jingles, The Stephen Greene Press (Brattleboro, Vermont 05310).

Anonymous:1976:GIb

[411] Anonymous. General information. Visible Language, X(2):99–100, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V10N2_1976_E.pdf.

Wanat:1976:RR

[412] Stanley F. Wanat. Reading readiness. Visible Language, X(2):101– 127, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N2_1976_E.pdf.

Zenker:1976:S

[413] Helmut Zenker. Sunday. Visible Language, X(2):128, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V10N2_1976_E.pdf. Excerpted from Alan Riddell, Typewriter Art, London: London Magazine Editions, 1975.

Nelson:1976:JMC

[414] T. M. Nelson and C. J. Ladan. Judgment of meaningfulness of Chinese characters by English-speaking observers. Visible Language, X(2): 129–143, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N2_1976_E.pdf.

Segalen:1976:CCP

[415] Victor Segalen. [Chinese characters and poem]. Visible Language, X (2):144, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N2_1976_E.pdf. Excerpt from the author’s

Stele du chemin de l’ame, The Greenwood Press (San Francisco, CA 94133), 1976.

Orth:1976:LND

[416] Bernard Orth, Hans Weckerle, and Dirk Wendt. Legibility of numerals displayed in a 4 × 7 dot matrix and seven-segment digits. Visible Language, X(2):145–155, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N2_1976_E.pdf.

Parker:1976:HFC

[417] George A. Parker. How to find characters per pica for caps. Visible Language, X(2):156, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V10N2_1976_E.pdf.

Ranta:1976:PPG

[418] Jerrald Ranta. Palindromes, poems and geometric form. Visible Language, X(2):157–172, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V10N2_1976_E.pdf. Reprinted

from the author’s article in College English, XXXXVI (October 1974), 161–172, National Council of Teachers of English.

Marcus:1976:CPP

[419] Aaron Marcus and Joe Rothrock. A course in poetry and printing. Visible Language, X(2):173–182, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N2_1976_E.pdf.

Banks:1976:BRF

[420] Colin Banks. Book review: F. E. Pardoe, John Baskerville. London: Muller, 1976; ISBN 0-584-10354-9; £9.50. Visible Language, X(2): 183–188, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N2_1976_E.pdf.

Wrolstad:1976:EN

[421] Merald E. Wrolstad. Editor’s note. Visible Language, X(2):188, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V10N2_1976_E.pdf.

Baudin:1976:RAAb

[422] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, X(2): 189–190, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N2_1976_E.pdf.

Anonymous:1976:Ab

[423] Anonymous. The authors. Visible Language, X(2):192, Spring 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V10N2_1976_E.pdf.

Anonymous:1976:GIc

[424] Anonymous. General information. Visible Language, X(3):195– 196, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N3_1976_E.pdf.

Littlewood:1976:ISB

[425] A. R. Littlewood. An ‘ikon of the soul‘: The Byzantine letter. Visible Language, X(3):197–226, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N3_1976_E.pdf.

Gomringer:1976:PMS

[426] Eugen Gomringer. Poetry as a means for the structuring of a social environment. Visible Language, X(3):227–241, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V10N3_1976_E.pdf. Translated by Mark E. Cory.

Panian:1976:U

[427] Ed Panian. Untitled. Visible Language, X(3):242, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V10N3_1976_E.pdf. Signatories of the Declaration of Independence.

Weiss:1976:S

[428] Jack Weiss. Statements 76. Visible Language, X(3):243–246, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V10N3_1976_E.pdf.

Otto:1976:OHE

[429] Wayne Otto and Cathy Stallard. One hundred essential sight words. Visible Language, X(3):247–252, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N3_1976_E.pdf.

Perrin:1976:DIK

[430] Stephen G. Perrin. Declaration of Independence kit. Visible Language, X (3):253–256, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N3_1976_E.pdf. Excerpted permission from

the author’s Rutabaga in Eight Languages.

Johnson:1976:MPH

[431] Dale D. Johnson and Richard L. Venezky. Models for predicting how adults pronounce vowel digraph spellings in unfamiliar words. Visible Language, X(3):257–268, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N3_1976_E.pdf.

Greengold:1976:C

[432] Jane Greengold and Chris Tanz. “Correspondences”. Visible Language, X (3):269–276, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N3_1976_E.pdf.

Hewes:1976:C

[433] Gordon W. Hewes, Alan Lomax, Martyn Hitchcock, and Philip Gaskell. Correspondence. Visible Language, X(3):277–283, Summer 1976. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V10N3_1976_E.pdf. See [403, 404, 490].

Wrolstad:1976:SOM

[434] Merald E. Wrolstad. [Statement of ownership, management and circulation of Visible Language]. Visible Language, X(3):283, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V10N3_1976_E.pdf.

Baudin:1976:RAAc

[435] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, X(3):284–287, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N3_1976_E.pdf.

Anonymous:1976:Ac

[436] Anonymous. The authors. Visible Language, X(3):288, Summer 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V10N3_1976_E.pdf.

Anonymous:1976:GId

[437] Anonymous. General information. Visible Language, X(4):291– 292, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N4_1976_E.pdf.

Venezky:1976:ISI

[438] Richard L. Venezky. An introduction to this special issue. Visible Language, X(4):293–294, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V10N4_1976_E.pdf.

Jaquith:1976:DAP

[439] James Jaquith. Digraphia in advertising: The public as guinea pig. Visible Language, X(4):295–308, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V10N4_1976_E.pdf. See com-

ments [490].

Secrist:1976:IEO

[440] Robert H. Secrist. Internalization of English orthographic patterns. Visible Language, X(4):309–322, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N4_1976_E.pdf.

Kerek:1976:PRS

[441] Andrew Kerek. The phonological relevance of spelling pronunciation. Visible Language, X(4):323–338, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N4_1976_E.pdf.

Gates:1976:MTL

[442] Arthur I. Gates and Esther H. Chase. Methods and theories of learning to spell tested by studies of deaf children. Visible Language, X(4): 339–350, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N4_1976_E.pdf.

Venezky:1976:NHE

[443] Richard L. Venezky. Notes on the history of English spelling. Visible Language, X(4):351–365, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N4_1976_E.pdf.

Webster:1976:EAS

[444] Noah Webster. Excerpt from The American Spelling Book. Visible Language, X(4):366–376, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N4_1976_E.pdf.

Baudin:1976:RAAd

[445] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, X(4):377–380, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V10N4_1976_E.pdf.

Anonymous:1976:Ad

[446] Anonymous. The authors. Visible Language, X(4):380, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V10N4_1976_E.pdf.

Anonymous:1976:IVX

[447] Anonymous. Index to Volume X. Visible Language, X(4):381– 384, Autumn 1976. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V10N4_1976_E.pdf.

Anonymous:1977:GIa

[448] Anonymous. General information. Visible Language, XI(1):3–4, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N1_1977_E.pdf.

Rawski:1977:SSS

[449] Conrad H. Rawski. The scientific study of subject literatures. Visible Language, XI(1):5–23, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N1_1977_E.pdf.

Anonymous:1977:SR

[450] Anonymous. Stereogrammic relations. Visible Language, XI(1): 24, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N1_1977_E.pdf.

Gurtler:1977:CGF

[451] André Gürtler and Christian Mengelt. Cyrillic Gothic: Formal modifications in the design of a Russian sans-serif typeface. Visible Language, XI(1):25–36, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N1_1977_E.pdf.

Walker:1977:CWS

[452] Laurence Walker. Comprehension of writing and spontaneous speech. Visible Language, XI(1):37–51, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N1_1977_E.pdf.

Anonymous:1977:VLS

[453] Anonymous. Visible language as speech written down. Visible Language, XI(1):52, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N1_1977_E.pdf.

Walklin:1977:LAC

[454] Carol Walklin. Letters, art, and children. Visible Language, XI(1): 53–62, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N1_1977_E.pdf.

Quirk:1977:SNW

[455] Randolph Quirk. Setting new word records. Visible Language, XI (1):63–74, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N1_1977_E.pdf.

Bonavia:1977:CWW

[456] David Bonavia. China’s war of words. Visible Language, XI(1): 75–78, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N1_1977_E.pdf. Reprinted from Far Eastern

Economic Review, March 4, 1977, pp. 20–21.

Banks:1977:BRC

[457] Colin Banks. Book review: Catalogue of the Edward Clark Library (with typographical notes by Harry Carter and an essay on the Printing of Illustrations by Frank P. Restall). P. J. W. Kilpatrick, general editor. Edinburgh: privately printed for Napier College of Commerce and Technology, Lothian Regional Council, 1976. Two volumes, 685 pages, 74 illustrations. ISBN 0-902703-12-9. Visible Language, XI(1):79–86, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N1_1977_E.pdf.

Baudin:1977:RAAa

[458] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XI(1):87–89, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N1_1977_E.pdf.

Anonymous:1977:Aa

[459] Anonymous. The authors. Visible Language, XI(1):90, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N1_1977_E.pdf.

Anonymous:1977:CAI

[460] Anonymous. Cumulative author index for Volumes I–X. Visible Language, XI(1):91–96, Winter 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V11N1_1977_E.pdf.

Anonymous:1977:DC

[461] Anonymous. Diagram of contents. Visible Language, XI(2):2–3, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N2_1977_E.pdf.

Marcus:1977:EM

[462] Aaron Marcus. At the edge of meaning. Visible Language, XI(2): 4–21, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N2_1977_E.pdf.

Franke:1977:OCP

[463] Herbert W. Franke. Observations concerning practical visual languages. Visible Language, XI(2):22–32, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N2_1977_E.pdf.

Howe:1977:SR

[464] Nelson Howe. To the sincere reader. Visible Language, XI(2):32, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N2_1977_E.pdf.

Ockerse:1977:D

[465] Thomas Ockerse. Documentracings. Visible Language, XI(2):33– 37, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N2_1977_E.pdf.

Knowlton:1977:CPG

[466] Ken Knowlton and Leon Harmon. Computer-produced grey scales. Visible Language, XI(2):38–49, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N2_1977_E.pdf.

Kostelanetz:1977:N

[467] Richard Kostelanetz. On numbers. Visible Language, XI(2):50–51, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N2_1977_E.pdf.

Price:1977:S

[468] Jonathan Price and Joel Katz. Signs. Visible Language, XI(2): 52–67, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N2_1977_E.pdf.

Cumming:1977:SSS

[469] Robert Cumming. “Sentence structures” and “a structure”. Visible Language, XI(2):68–69, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V11N2_1977_E.pdf.

Finlay:1977:S

[470] Ian Hamilton Finlay. ‘Sundial’. Visible Language, XI(2):70, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N2_1977_E.pdf.

Sky:1977:PW

[471] Alison Sky. The public word. Visible Language, XI(2):71–80, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N2_1977_E.pdf.

Baudin:1977:RTE

[472] Fernand Baudin. Reflection on the theme: At the edge of meaning. Visible Language, XI(2):81–92, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N2_1977_E.pdf.

Friedman:1977:SBD

[473] Daniel Friedman. A study in basic design and meaning. Visible Language, XI(2):93–101, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N2_1977_E.pdf.

Kenedy:1977:WDS

[474] R. C. Kenedy. Wortgebilde Durch Spiel und Kombinatorik: Or, why Duchamps loved words. Visible Language, XI(2):102–127, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N2_1977_E.pdf.

Anonymous:1977:AB

[475] Anonymous. Abstract/biography. Visible Language, XI(2):128– 134, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N2_1977_E.pdf.

Baudin:1977:RAAb

[476] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XI(2):135–139, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N2_1977_E.pdf.

Anonymous:1977:DN

[477] Anonymous. Design notes. Visible Language, XI(2):140, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N2_1977_E.pdf.

Anonymous:1977:PN

[478] Anonymous. Production notes. Visible Language, XI(2):141, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N2_1977_E.pdf.

Wrolstad:1977:BM

[479] Merald E. Wrolstad. [back matter]. Visible Language, XI(2):142– 143, Spring 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N2_1977_E.pdf.

Anonymous:1977:GIc

[480] Anonymous. General information. Visible Language, XI(3):243– 245, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N3_1977_E.pdf.

Hellerstein:1977:PCG

[481] Nina S. Hellerstein. Paul Claudel and Guillaume Apollinaire as visual poets: Idéogrammes occidentaux and Calligrammes. Visible Language, XI (3):245–270, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N3_1977_E.pdf.

McCarthy:1977:CLA

[482] Lenore McCarthy. A child learns the alphabet. Visible Language, XI (3):271–284, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N3_1977_E.pdf.

Allington:1977:EGS

[483] Richard L. Allington and Michael Strange. Effects of grapheme substitutions in connected text upon reading behaviors. Visible Language, XI (3):285–297, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N3_1977_E.pdf.

Anonymous:1977:DPF

[484] Anonymous. Decorative page from Festschrift of Originals given to and in honor of Paul A. Bennett by The Typophiles, New York City. Visible Language, XI(3):298, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N3_1977_E.pdf.

Bank:1977:CAB

[485] Arnold Bank. Calligraphy of Arnold Bank. Visible Language, XI(3): 299–306, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N3_1977_E.pdf.

Hatherly:1977:RR

[486] Ana Hatherly. The reinvention of reading. Visible Language, XI(3): 307–320, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N3_1977_E.pdf.

Nesbitt:1977:DLI

[487] Alexander Nesbitt. The designer and language II. Visible Language, XI (3):321–324, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N3_1977_E.pdf.

Wrolstad:1977:SOM

[488] Merald E. Wrolstad. [Statement of ownership, management and circulation of Visible Language]. Visible Language, XI(3):324, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N3_1977_E.pdf.

Bigelow:1977:BRD

[489] Charles A. Bigelow. Book review: David Kindersley. Optical Letter Spacing for New Printing Systems. London: Wynkyn de Worde Society, 1976. 4-1/2 × 11 inches. 40 pp., 45 figs. ISBN 0-85331-360-1. (Distributed by Lund Humphries Publishers Ltd.). Visible Language, XI(3): 325–329, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N3_1977_E.pdf.

Herrick:1977:C

[490] Earl M. Herrick and Laurence Urdang. Correspondence. Visible Language, XI(3):330–332, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V11N3_1977_E.pdf. See [433,

439].

Baudin:1977:RAAc

[491] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XI(3):333–335, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N3_1977_E.pdf.

Anonymous:1977:Ac

[492] Anonymous. The authors. Visible Language, XI(3):336, Summer 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N3_1977_E.pdf.

Barthes:1977:LRR

[493] Roland Barthes. Letter to Randolph Runyon. Visible Language, XI (4):338, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N4_1977_E.pdf.

Ungar:1977:BES

[494] Steven Ungar. Beyond the empire of signs. Visible Language, XI (4):338, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N4_1977_E.pdf.

Barthes:1977:ELS

[495] Roland Barthes. Excerpt: L’Empire des Signes. Visible Language, XI (4):339–340, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N4_1977_E.pdf.

Caplan:1977:NLB

[496] Jay Caplan. Nothing but language: on Barthes’s Empire of Signs. Visible Language, XI(4):341–362, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N4_1977_E.pdf.

Conley:1977:BES

[497] Tom Conley. Barthes’s Excès: The silent apostrophe of S/Z. Visible Language, XI(4):355–385, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N4_1977_E.pdf.

Gallop:1977:B

[498] Jane Gallop. “b s”. Visible Language, XI(4):364–387, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N4_1977_E.pdf.

Runyon:1977:FAD

[499] Randolph Runyon. Fragments of an amorous discourse: Canon in ubis. Visible Language, XI(4):387–389, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V11N4_1977_E.pdf.

Ungar:1977:WLB

[500] Steven Ungar. From writing to the letter: Barthes and alphabetese. Visible Language, XI(4):390–400, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N4_1977_E.pdf.

Baudin:1977:RAAd

[501] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XI(4):430, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V11N4_1977_E.pdf.

Anonymous:1977:GId

[502] Anonymous. General information. Visible Language, XI(4):430, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N4_1977_E.pdf.

Anonymous:1977:Ad

[503] Anonymous. The authors. Visible Language, XI(4):432, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V11N4_1977_E.pdf.

Anonymous:1977:IVX

[504] Anonymous. Index to Volume XI. Visible Language, XI(4):432– ??, Autumn 1977. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V11N4_1977_E.pdf.

Massaro:1978:SMR

[505] Dominic W. Massaro. A stage model of reading and listening. Visible Language, XII(1):3–26, Winter 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V12N1_1978_E.pdf.

Pastore:1978:PAC

[506] Richard E. Pastore. Phonemes and alphanumeric characters: Possible components of parallel human communications systems. Visible Language, XII(1):27–42, Winter 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V12N1_1978_E.pdf.

Frith:1978:PMP

[507] Uta Frith. From print to meaning and from print to sound, or how to read without knowing how to spell. Visible Language, XII(1):43–54, Winter 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N1_1978_E.pdf.

Baron:1978:USS

[508] Jonathan Baron and June Hodge. Using spelling–sound correspondences without trying to learn them. Visible Language, XII(1):55–70, Winter 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N1_1978_E.pdf.

Martin:1978:VRD

[509] James G. Martin, Richard H. Meltzer, and Carol B. Mills. Visual rhythms: Dynamic text display for learning to read a second language. Visible Language, XII(1):71–80, Winter 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V12N1_1978_E.pdf.

Levy:1978:SAD

[510] Betty Ann Levy. Speech analysis during sentence processing: Reading and listening. Visible Language, XII(1):81–102, Winter 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N1_1978_E.pdf.

Anonymous:1978:B

[511] Anonymous. Biography. Visible Language, XII(1):103, Winter 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N1_1978_E.pdf.

Anonymous:1978:Aa

[512] Anonymous. Acknowledgements. Visible Language, XII(1):104, Winter 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N1_1978_E.pdf.

Anonymous:1978:RAA

[513] Anonymous. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XII(1):105–110, Winter 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N1_1978_E.pdf.

Wrolstad:1978:GI

[514] Merald E. Wrolstad. General information. Visible Language, XII(1): 111–112, Winter 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N1_1978_E.pdf.

Bormuth:1978:VVL

[515] John Bormuth. Value and volume of literacy. Visible Language, XII (2):118–161, Spring 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N2_1978_E.pdf.

Arnheim:1978:SAG

[516] Rudolf Arnheim. Spatial aspects of graphological expression. Visible Language, XII(2):162–170, Spring 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V12N2_1978_E.pdf.

Cowie:1978:BCA

[517] Frederick J. Cowie. Boniface (c. 675–754): Archbishop, legate, and Postmaster General. Visible Language, XII(2):171–182, Spring 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N2_1978_E.pdf.

Hartley:1978:ELL

[518] James Hartley, Peter Burnhill, and Lindsey Davis. The effects of line length and paragraph denotation on retrieval of information from prose text. Visible Language, XII(2):183–194, Spring 1978. CODEN VSLGAO.

ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V12N2_1978_E.pdf.

Duffelmeyer:1978:VCP

[519] Frederick Duffelmeyer. The vowel cluster: Pronunciation preferences of proficient and non-proficient adult readers. Visible Language, XII (2):195–200, Spring 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N2_1978_E.pdf.

Foster:1978:LLR

[520] Jeremy Foster. Locating legibility research: a guide for the graphic designer. Visible Language, XII(2):201–205, Spring 1978. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N2_1978_E.pdf.

Weber:1978:BRW

[521] Rose-Marie Weber. Book review: What did I write? Marie M. Clay. What Did I Write? Auckland: Heinemann Educational Books, 1975. 7 1/2 × 9 inches. 78 pp. Visible Language, XII(2):206–209, Spring 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N2_1978_E.pdf.

Baudin:1978:AFD

[522] Fernand Baudin, Dirk Wendt, and Ana Fisch. Abstracts: Français; Deutsch; Español. Visible Language, XII(2):211–219, Spring 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N2_1978_E.pdf.

Anonymous:1978:Ab

[523] Anonymous. Authors. Visible Language, XII(2):220–221, Spring 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N2_1978_E.pdf.

Anonymous:1978:GI

[524] Anonymous. General information. Visible Language, XII(2):222, Spring 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N2_1978_E.pdf.

Wrolstad:1978:SOM

[525] Merald E. Wrolstad. [Statement of ownership, management and circulation of Visible Language]. Visible Language, XII(2):222, Spring 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N2_1978_E.pdf.

Anonymous:1978:C

[526] Anonymous. Colophon. Visible Language, XII(2):222, Spring 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N2_1978_E.pdf.

Hedges:1978:CWM

[527] Inez Hedges. Cinematic writing of Maurice Roche. Visible Language, XII(3):??, Summer 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N3_1978_E.pdf.

Hubert:1978:PVP

[528] Renée Riese Hubert. Ponge: Visual and poetic writing. Visible Language, XII(3):??, Summer 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N3_1978_E.pdf.

Andermatt:1978:WLL

[529] Verna Andermatt. Writing the letter: Lower-case of hélène cixous. Visible Language, XII(3):??, Summer 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V12N3_1978_E.pdf.

McKenna:1978:DHP

[530] A. J. McKenna. Derrida and his precursors. Visible Language, XII (3):??, Summer 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N3_1978_E.pdf.

Conley:1978:S

[531] Tom Conley. Serres (É). Visible Language, XII(3):??, Summer 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N3_1978_E.pdf.

Lydon:1978:LLM

[532] Mary Lydon. Love letters: Michel Butor and visible language. Visible Language, XII(3):??, Summer 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V12N3_1978_E.pdf.

Ames:1978:MLD

[533] S. S. Ames. Mallarmé’s letters in Duras. Visible Language, XII (3):??, Summer 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N3_1978_E.pdf.

Rodriguez:1978:RWV

[534] L. S. Rodriguez. Readable/writable/visible. Visible Language, XII (3):??, Summer 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N3_1978_E.pdf.

Anonymous:1978:ABa

[535] Anonymous. Advisory Board. Visible Language, XII(4):388, Autumn 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N4_1978_E.pdf.

Wrolstad:1978:GIS

[536] Merald E. Wrolstad. General information. Subscription rates. Reprints and back copies. Manuscripts. Letters to the editor. Advertising. Visible Language, XII(4):389, Autumn 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V12N4_1978_E.pdf.

Marcel:1978:URE

[537] Tony Marcel. Unconscious reading: Experiments on people who do not know they are reading. Visible Language, XII(4):391–404, Autumn 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V12N4_1978_E.pdf. See correction in figure layout [564].

Cosmos:1978:TVS

[538] Spencer Cosmos. Toward a visual stylistics: Assent and denial in Chaucer. Visible Language, XII(4):406–427, Autumn 1978. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N4_1978_E.pdf.

Cox:1978:GCP

[539] Charles H. Cox III, Barry A. Blesser, and Murray Eden. Graphical context of printed characters. Visible Language, XII(4):428–447, Autumn 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V12N4_1978_E.pdf.

Mosenthal:1978:WRR

[540] Peter Mosenthal, Sean Walmsley, and Richard Allington. Word recognition reconsidered: Toward a multi-context model. Visible Language, XII (4):448–468, Autumn 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N4_1978_E.pdf.

Baudin:1978:RAAd

[541] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XII(4):469–474, Autumn 1978. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V12N4_1978_E.pdf.

Wrolstad:1978:AB

[542] Merald E. Wrolstad. Advisory Board. Visible Language, XII(4): 475, Autumn 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N4_1978_E.pdf.

Anonymous:1978:IVX

[543] Anonymous. Index to Volume XII. Visible Language, XII(4):476– 479, Autumn 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N4_1978_E.pdf.

Anonymous:1978:CNB

[544] Anonymous. Credits. Note. Back numbers. Visible Language, XII (4):480, Autumn 1978. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V12N4_1978_E.pdf.

Lehman:1979:TLC

[545] Charles Lehman. Teaching and learning the craft of handwriting. Visible Language, XIII(1):4–15, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V13N1_1979_E.pdf.

Graves:1979:LCS

[546] Donald H. Graves. Let children show us how to help them write. Visible Language, XIII(1):16–28, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V13N1_1979_E.pdf. See [612].

Macleod:1979:DAT

[547] Iain Macleod and Peter Procter. A dynamic approach to teaching handwriting skills. Visible Language, XIII(1):29–42, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N1_1979_E.pdf. See [612].

Rhodes:1979:CDN

[548] Chris Rhodes. Considerations for the design of a new pen range. Visible Language, XIII(1):43–48, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V13N1_1979_E.pdf.

Briem:1979:WHF

[549] Gunnlauger S. E. Briem. Wanted: Handwriting to fit pens. Visible Language, XIII(1):49–62, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V13N1_1979_E.pdf. See [612].

Gray:1979:TNH

[550] Nicolete Gray. Towards a new handwriting adapted to the ballpoint pen. Visible Language, XIII(1):63–69, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V13N1_1979_E.pdf. See [612].

Osley:1979:CRH

[551] A. S. Osley. Cannons of Renaissance handwriting. Visible Language, XIII(1):70–94, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N1_1979_E.pdf. See [612].

Baudin:1979:RAAa

[552] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XIII(1):95–99, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V13N1_1979_E.pdf.

Anonymous:1979:Aa

[553] Anonymous. The authors. Visible Language, XIII(1):100–101, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N1_1979_E.pdf.

Anonymous:1979:GIa

[554] Anonymous. General information. Visible Language, XIII(1):102– 103, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N1_1979_E.pdf.

Wrolstad:1979:SOM

[555] Merald E. Wrolstad. [Statement of ownership, management and circulation of Visible Language]. Visible Language, XIII(1):103, Winter 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N1_1979_E.pdf.

Anonymous:1979:GIb

[556] Anonymous. General information. Visible Language, XIII(2):107, Spring 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N2_1979_E.pdf.

Wendt:1979:EAI

[557] Dirk Wendt. An experimental approach to the improvement of the typographic design of textbooks. Visible Language, XIII(2):108–133, Spring 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N2_1979_E.pdf.

Unger:1979:DT

[558] Gerard Unger. The design of a typeface. Visible Language, XIII(2): 134–149, Spring 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N2_1979_E.pdf.

Brann:1979:MDP

[559] Noel L. Brann. A monastic dilemma posed by the invention of printing: the context of De laude Scriptorum Manualium by Abbot Johann Trithemius (1462–1516). Visible Language, XIII(2):150–167, Spring 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N2_1979_E.pdf.

Marmolin:1979:MRP

[560] Hans Marmolin, Lars-Göran Nilsson, and Hans Smedshammar. Mediated reading process of the partially sighted. Visible Language, XIII (2):168–183, Spring 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N2_1979_E.pdf.

Larcher:1979:DWL

[561] Jean Larcher. A different way of looking at typography. Visible Language, XIII(2):184–191, Spring 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V13N2_1979_E.pdf.

Wendt:1979:KBAa

[562] Dirk Wendt, Fernand Baudin, and Ana Fisch. Kurzfassung der Beiträge [Abstracts of journal articles in German, French, and Spanish]. Visible Language, XIII(2):192–197, Spring 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V13N2_1979_E.pdf.

Anonymous:1979:Ab

[563] Anonymous. The authors. Visible Language, XIII(2):198, Spring 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N2_1979_E.pdf.

Anonymous:1979:ERT

[564] Anonymous. Errata: Re: Tony Marcel, “Unconscious Reading: Experiments on People Who Do Not Know That They Are Reading,” XII 4 (Autumn 1978), pp. 319–404. Visible Language, XIII(2):199–200, Spring 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N2_1979_E.pdf. See [537].

Anonymous:1979:GIc

[565] Anonymous. General information. Visible Language, XIII(3):207, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-

5529 (electronic).  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N3_1979_E.pdf.

Wing:1979:ISI

[566] Alan M. Wing. Introduction to this special issue [on behavioural studies of the handwriting skill]. Visible Language, XIII(3):208, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N3_1979_E.pdf.

Woodworth:1979:FHU

[567] R. S. Woodworth. Facility of handwriting using different movements. Visible Language, XIII(3):209–217, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V13N3_1979_E.pdf.

Tuelings:1979:CAA

[568] Hans-Leo H. M. Tuelings and Arnold J. W. M. Thomassen. Computeraided analysis of handwriting movements. Visible Language, XIII(3): 218–231, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N3_1979_E.pdf.

Watson:1979:CCS

[569] R. S. Watson and P. J. Pobgee. A computer to check signatures. Visible Language, XIII(3):232–238, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V13N3_1979_E.pdf.

Ansell:1979:HCF

[570] Michael Ansell. Handwriting classification of forsenic science. Visible Language, XIII(3):239–251, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V13N3_1979_E.pdf.

Hollerbach:1979:CMH

[571] John Hollerbach. A competence model of handwriting. Visible Language, XIII(3):252–264, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V13N3_1979_E.pdf.

Ellis:1979:SP

[572] Andrew W. Ellis. Slips of the pen. Visible Language, XIII(3):265– 282, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N3_1979_E.pdf.

Wing:1979:VHC

[573] Alan M. Wing. Variability of handwritten characters. Visible Language, XIII(3):283–298, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V13N3_1979_E.pdf.

Thomassen:1979:DDP

[574] Arnold J. W. M. Thomassen and Hans-Leo H. M. Tuelings. The development of directional preference in writing movements. Visible Language, XIII(3):299–313, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V13N3_1979_E.pdf.

Sovik:1979:SCP

[575] Nils Søvik. Some constructional parameters related to children’s copying performance. Visible Language, XIII(3):314–330, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N3_1979_E.pdf.

Kao:1979:HE

[576] Henry S. R. Kao. Handwriting ergonomics. Visible Language, XIII(3): 331–339, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N3_1979_E.pdf.

Wendt:1979:KBAb

[577] Dirk Wendt, Fernand Baudin, and Ana Fisch. Kurzfassung der Beiträge [Abstracts of journal articles in German, French, and Spanish]. Visible Language, XIII(3):341–345, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V13N3_1979_E.pdf.

Anonymous:1979:Ac

[578] Anonymous. The authors. Visible Language, XIII(3):346–347, Summer 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N3_1979_E.pdf.

Anonymous:1979:Ad

[579] Anonymous. Abstracts. Visible Language, XIII(4):350–351, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N4_1979_E.pdf.

Poggenpohl:1979:GDP

[580] Sharon Poggenpohl. Graphic design: a practice in search of theory. Visible Language, XIII(4):352–357, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V13N4_1979_E.pdf.

Ockerse:1979:SGD

[581] Thomas Ockerse and Hans C. van Dijk. Semiotics and graphic design education. Visible Language, XIII(4):358–378, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N4_1979_E.pdf.

Stuhr:1979:LFT

[582] David Stuhr. A language of form: The two-dimensional isometric system. Visible Language, XIII(4):379–409, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V13N4_1979_E.pdf.

Manning:1979:NVD

[583] Robert Manning. Notes on the visual differential theory. Visible Language, XIII(4):410–427, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V13N4_1979_E.pdf.

Reynolds:1979:GIR

[584] Linda Reynolds. The graphic information research unit: Background and recent research. Visible Language, XIII(4):428–448, Autumn 1979.

CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N4_1979_E.pdf.

Bigelow:1979:NAT

[585] Charles A. Bigelow. Native American texts. Visible Language, XIII (4):450, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

White:1979:WSD

[586] Constance White. Writing system design: Finding the optimal written translation of the English language. Visible Language, XIII(4): 451, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

Spencer:1979:IDB

[587] Herbert Spencer, Linda Reynolds, and Brian Coe. Image degradation and background noise/legibility. Visible Language, XIII(4):452, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N4_1979_E.pdf.

Reynolds:1979:LRV

[588] Linda Reynolds, Herbert Spencer, and George Glaze. Legibility and readability of Viewdata displays. Visible Language, XIII(4):453, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N4_1979_E.pdf.

Skaggs:1979:TMV

[589] Steven Skaggs. Type manipulation with video technology. Visible Language, XIII(4):454, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V13N4_1979_E.pdf.

Tanaka:1979:IWD

[590] Tadao Tanaka. Investigation of the word design in the katakana form of Japanese. Visible Language, XIII(4):455, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N4_1979_E.pdf.

Kuyper:1979:IHG

[591] Jerry L. Kuyper. Investigation/hand gestures. Visible Language, XIII (4):456, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

Kuyper:1979:ZSS

[592] Jerry L. Kuyper. Zoo symbol sign. Visible Language, XIII(4): 457, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

Remington:1979:SG

[593] R. Roger Remington. The sign game. Visible Language, XIII(4): 458–459, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

Marcus:1979:VGI

[594] Aaron Marcus. Visualizing global interdependencies. Visible Language, XIII(4):460–461, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V13N4_1979_E.pdf.

Remington:1979:HGD

[595] R. Roger Remington. History of graphic design. Visible Language, XIII (4):462, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

Anonymous:1979:DTA

[596] Anonymous. Design thesis abstract. Visible Language, XIII(4): 463, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

Anonymous:1979:CPS

[597] Anonymous. Class project survey. Visible Language, XIII(4):464, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-

5529 (electronic).  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

Anonymous:1979:Ca

[598] Anonymous. Contributors. Visible Language, XIII(4):465–466, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N4_1979_E.pdf.

Baudin:1979:RAAd

[599] Fernand Baudin, Ana Fisch, and Dirk Wendt. Résumé des articles [abstracts of journal articles in French, Spanish, and German]. Visible Language, XIII(4):467–470, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V13N4_1979_E.pdf.

Anonymous:1979:DN

[600] Anonymous. Design notes. Visible Language, XIII(4):470, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N4_1979_E.pdf.

Anonymous:1979:Cb

[601] Anonymous. Colophon. Visible Language, XIII(4):470, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V13N4_1979_E.pdf.

Anonymous:1979:GId

[602] Anonymous. General information. Visible Language, XIII(4):471, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-

5529 (electronic).  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

Anonymous:1979:IVX

[603] Anonymous. Index [to Volume XIII]. Visible Language, XIII(4):473– 475, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

Wrolstad:1979:AB

[604] Merald E. Wrolstad. Advisory Board. Visible Language, XIII(4): 476, Autumn 1979. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V13N4_1979_E.pdf.

Wrolstad:1980:ABa

[605] Merald E. Wrolstad. Advisory Board. Visible Language, XIV (1):2, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N1_1980_E.pdf.

Anonymous:1980:GIa

[606] Anonymous. General information. Visible Language, XIV(1):3–4, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N1_1980_E.pdf.

Shimron:1980:DVI

[607] Joseph Shimron and David Navon. The distribution of visual information in the vertical dimension of Roman and Hebrew letters. Visible Language, XIV(1):5–12, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N1_1980_E.pdf.

Meares:1980:FGB

[608] Olive Meares. Figure/ground, brightness contrast, and reading disabilities. Visible Language, XIV(1):13–29, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V14N1_1980_E.pdf.

Wood:1980:VLP

[609] Richard E. Wood. Visible language policy — bilingualism and multinlingualism on postage stamps. Visible Language, XIV(1):30–51, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N1_1980_E.pdf.

James:1980:SAE

[610] Jean M. James. Signs in Ancient Egypt: Another look at the relation of figure to hieroglyph. Visible Language, XIV(1):52–61, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N1_1980_E.pdf.

Hartley:1980:SCT

[611] James Hartley. Spatial cues in text: Some comments on the paper by Frase & Schwartz (1980). Visible Language, XIV(1):62–79, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V14N1_1980_E.pdf. See corrections to layout [628].

Ebsen:1980:C

[612] Alf K. Ebsen. Correspondence. Visible Language, XIV(1):80–83, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V14N1_1980_E.pdf. Comments on several papers in volume 13 number 1. See replies [613, 614, 615, 616, 617].

Briem:1980:CNL

[613] Gunnlauger S. E. Briem. Correspondence: No, late roman cursive wasn’t a worm. Visible Language, XIV(1):84, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N1_1980_E.pdf. See [612].

Gray:1980:RE

[614] Nicolete Gray. Reply [to Ebsen]. Visible Language, XIV(1):85–86, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N1_1980_E.pdf. See [612].

Macleod:1980:RE

[615] Iain Macleod and Peter Procter. Reply [to Ebsen]. Visible Language, XIV(1):86–88, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N1_1980_E.pdf. See [612].

Graves:1980:RE

[616] Donald H. Graves. Reply [to Ebsen]. Visible Language, XIV(1): 89, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N1_1980_E.pdf. See [612].

Osley:1980:RE

[617] A. S. Osley. Reply [to Ebsen]. Visible Language, XIV(1):89–90, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N1_1980_E.pdf. See [612].

Flach:1980:RM

[618] V. H. Flach. Righting’s merror. Visible Language, XIV(1):91, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N1_1980_E.pdf.

Baudin:1980:RAAa

[619] Fernand Baudin and Dirk Wendt. Résumé des articles [abstracts of journal articles in French and German]. Visible Language, XIV(1): 92–93, Winter 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N1_1980_E.pdf.

Anonymous:1980:GIb

[620] Anonymous. General information. Visible Language, XIV(2):98, Spring 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N2_1980_E.pdf.

Stratton:1980:WCL

[621] Jon Stratton. Writing and the concept of law in Ancient Greece. Visible Language, XIV(2):99–121, Spring 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N2_1980_E.pdf.

Smith:1980:DCE

[622] Philip T. Smith. In defense of conservatism in English orthography. Visible Language, XIV(2):122–136, Spring 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N2_1980_E.pdf.

Sakiey:1980:SFC

[623] Elizabeth Sakiey, Edward Fry, Albert Goss, and Barry Loigman. A syllable frequency count. Visible Language, XIV(2):137–150, Spring 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N2_1980_E.pdf.

Wright:1980:STD

[624] Patricia Wright. Strategy and tactics in the design of forms. Visible Language, XIV(2):151–193, Spring 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N2_1980_E.pdf.

Lyttelton:1980:BS

[625] Humphrey Lyttelton. ‘Beware of the scribes’. Visible Language, XIV (2):194–203, Spring 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N2_1980_E.pdf. This article originally ap-

peared in No. 102 of The journal of the Society for Italic Handwriting from which it is reprinted with permission.

Baudin:1980:RAAb

[626] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XIV(2):204–206, Spring 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N2_1980_E.pdf.

Anonymous:1980:Ab

[627] Anonymous. The authors. Visible Language, XIV(2):206–207, Spring 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N2_1980_E.pdf.

Hartley:1980:C

[628] James Hartley. Correspondence. Visible Language, XIV(2):208, Spring 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N2_1980_E.pdf. See [611].

Wrolstad:1980:SOM

[629] Merald E. Wrolstad. [Statement of ownership, management and circulation of Visible Language]. Visible Language, XIV(2):208, Spring 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N2_1980_E.pdf.

Woodruff:1980:CSRa

[630] Lori Woodruff and Susan Warren. Covers: a stele for Roland Barthes. Visible Language, XIV(3):c1, c2, Summer 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N3_1980_E.pdf.

Wrolstad:1980:ABc

[631] Merald E. Wrolstad. Advisory Board. Visible Language, XIV(3): 210, Summer 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N3_1980_E.pdf.

Lydon:1980:VLF

[632] Mary Lydon. Visible language: Freud’s imprint. Visible Language, XIV (3):211–218, Summer 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N3_1980_E.pdf.

McKenna:1980:LIF

[633] Andrew J. McKenna. Lex icon: Freud and Rimbaud. Visible Language, XIV(3):219–240, Summer 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V14N3_1980_E.pdf.

Ames:1980:KBO

[634] Sanford S. Ames. Killer bees: an ontology in abeyance. Visible Language, XIV(3):241–249, Summer 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V14N3_1980_E.pdf.

Bauer:1980:RGT

[635] George H. Bauer. Robbe-Grillet on target Or interrogation by the numbers. Visible Language, XIV(3):250–272, Summer 1980. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N3_1980_E.pdf.

Gallop:1980:FIC

[636] Jane Gallop. Freud’s invisible chiasmus, or you can’t judge a book by its cover. Visible Language, XIV(3):273–283, Summer 1980. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N3_1980_E.pdf.

Bartman:1980:DDB

[637] Susanna Bartman. Defoe’s daydream: Becoming Moll Flanders. Visible Language, XIV(3):283–305, Summer 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N3_1980_E.pdf.

Skoller:1980:FLC

[638] Eleanor Honig Skoller. Franked letters: Crossing the bar. Visible Language, XIV(3):306–319, Summer 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N3_1980_E.pdf.

Woodruff:1980:CSRb

[639] Lori Woodruff and Susan Warren. The contract: a stele for Roland Barthes. Visible Language, XIV(3):320–327, Summer 1980. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V14N3_1980_E.pdf. This essay first appeared in Cream City Review, VI, 1980. With Roland Barthes: In Memoriam by Mary Lydon.

Baudin:1980:RAAc

[640] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XIV(3):328–332, Summer 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N3_1980_E.pdf.

Anonymous:1980:Ac

[641] Anonymous. The authors. Visible Language, XIV(3):333, Summer 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N3_1980_E.pdf.

Wrolstad:1980:ABd

[642] Merald E. Wrolstad. Advisory Board. Visible Language, XIV(4): 338, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N4_1980_E.pdf.

Anonymous:1980:GId

[643] Anonymous. General information. Visible Language, XIV(4):339, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-

5529 (electronic).  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N4_1980_E.pdf.

Wason:1980:I

[644] Peter Wason. Introduction. Visible Language, XIV(4):340, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N4_1980_E.pdf.

Young:1980:ACG

[645] Richard E. Young. Arts, crafts, gifts, and knacks: Some disharmonies in the new rhetoric. Visible Language, XIV(4):341–350, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N4_1980_E.pdf.

Wason:1980:CCW

[646] Peter C. Wason. Conformity and commitment in writing. Visible Language, XIV(4):351–363, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N4_1980_E.pdf.

Galbriath:1980:ECG

[647] David Galbriath. The effect of conflicting goals on writing: a case study. Visible Language, XIV(4):364–375, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V14N4_1980_E.pdf.

Stack:1980:WCA

[648] Richard Stack. Writing as conversation. Addendum: Letter to a novelist. Visible Language, XIV(4):376–382, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V14N4_1980_E.pdf.

Lowenthal:1980:MLR

[649] David Lowenthal. Mixing levels of revision. Visible Language, XIV(4): 383–387, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N4_1980_E.pdf.

Hayes:1980:WPS

[650] John R. Hayes and Linda S. Flower. Writing as problem solving. Visible Language, XIV(4):388–399, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N4_1980_E.pdf.

Bracewell:1980:WCA

[651] Robert J. Bracewell. Writing as a cognitive activity. Visible Language, XIV(4):400–422, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V14N4_1980_E.pdf.

Anonymous:1980:Q

[652] Anonymous. [quotations]. Visible Language, XIV(4):423, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N4_1980_E.pdf.

Baudin:1980:RAAd

[653] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XIV(4):424–427, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V14N4_1980_E.pdf.

Anonymous:1980:Ad

[654] Anonymous. The authors. Visible Language, XIV(4):428–429, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N4_1980_E.pdf.

Elbow:1980:W

[655] Peter Elbow. [writing]. Visible Language, XIV(4):429, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V14N4_1980_E.pdf.

Anonymous:1980:IVX

[656] Anonymous. Index to Volume XIV. Visible Language, XIV(4):430– 432, Autumn 1980. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V14N4_1980_E.pdf.

Wrolstad:1981:ABa

[657] Merald E. Wrolstad. Advisory Board. Visible Language, XV(1):2, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N1_1981_E.pdf.

Anonymous:1981:GIa

[658] Anonymous. General information. Visible Language, XV(1):3, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N1_1981_E.pdf.

Hartley:1981:E

[659] James Hartley and Peter Burnhill. Editorial. Visible Language, XV (1):4, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N1_1981_E.pdf.

Twyman:1981:TW

[660] Michael Twyman. Typography without words. Visible Language, XV (1):5–12, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N1_1981_E.pdf.

Hartley:1981:ECL

[661] James Hartley and Mark Trueman. The effects of changes in layout and changes in wording on preferences for instructional text. Visible Language, XV(1):13–31, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V15N1_1981_E.pdf.

Jewett:1981:MLW

[662] Don L. Jewett. Multi-level writing in theory and practice. Visible Language, XV(1):32–40, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V15N1_1981_E.pdf.

Shebilske:1981:TSC

[663] Wayne L. Shebilske and John A. Rotondo. Typographical and spatial cues that facilitate learning from textbooks. Visible Language, XV (1):41–54, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N1_1981_E.pdf.

Jonassen:1981:IMD

[664] David Scott Jonassen. Information mapping: a description, rationale, and comparison with programmed instruction. Visible Language, XV (1):55–66, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N1_1981_E.pdf.

Norton:1981:PNT

[665] Linda S. Norton. Patterned note-taking: an evaluation. Visible Language, XV(1):67–85, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N1_1981_E.pdf.

Sloboda:1981:USM

[666] John Sloboda. The uses of space in music notation. Visible Language, XV(1):86–110, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N1_1981_E.pdf.

Anonymous:1981:SB

[667] Anonymous. A selected bibliography. Visible Language, XV(1):111– 113, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N1_1981_E.pdf.

Anonymous:1981:Aa

[668] Anonymous. The authors. Visible Language, XV(1):114–115, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N1_1981_E.pdf.

Anonymous:1981:C

[669] Anonymous. Correspondence. Visible Language, XV(1):115, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N1_1981_E.pdf.

Baudin:1981:RAAa

[670] Fernand Baudin, Dirk Wendt, and Ana Fisch. Résumé des articles [abstracts of journal articles in French, German, and Spanish]. Visible Language, XV(1):116–120, Winter 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V15N1_1981_E.pdf.

Wrolstad:1981:ABb

[671] Merald E. Wrolstad. Advisory Board. Visible Language, XV(2): 122, Spring 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N2_1981_E.pdf.

Anonymous:1981:GIb

[672] Anonymous. General information. Visible Language, XV(2):123, Spring 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N2_1981_E.pdf.

Rayner:1981:NGE

[673] Keith Rayner. A note from the Guest Editor. Visible Language, XV (2):124, Spring 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N2_1981_E.pdf.

Rayner:1981:VCW

[674] Keith Rayner. Visual cues in word recognition and reading: Introduction. Visible Language, XV(2):125–128, Spring 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V15N2_1981_E.pdf.

Morrison:1981:VFE

[675] Robert E. Morrison and Albrecht-Werner Inhoff. Visual factors and eye movements in reading. Visible Language, XV(2):129–146, Spring 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N2_1981_E.pdf.

Haber:1981:VCR

[676] Ralph Norman Haber and Lyn R. Haber. Visual components of the reading process. Visible Language, XV(2):147–182, Spring 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N2_1981_E.pdf.

Brady:1981:TCT

[677] Michael Brady. Toward a computational theory of early visual processing in reading. Visible Language, XV(2):183–215, Spring 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N2_1981_E.pdf.

Anonymous:1981:Ab

[678] Anonymous. The authors. Visible Language, XV(2):216, Spring 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N2_1981_E.pdf.

Ehrlich:1981:CWR

[679] Susan F. Ehrlich. Children’s word recognition in prose context. Visible Language, XV(3):219–244, Summer 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V15N3_1981_E.pdf.

Coltheart:1981:DRT

[680] Max Coltheart. Disorders of reading and their implications for models of normal reading. Visible Language, XV(3):245–286, Summer 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N3_1981_E.pdf.

Well:1981:WPR

[681] Arnold D. Well and Alexander Pollatsek. Word processing in reading: a commentary on the papers. Visible Language, XV(3):287–308, Summer 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N3_1981_E.pdf.

Anonymous:1981:Ac

[682] Anonymous. The authors. Visible Language, XV(3):309–310, Summer 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N3_1981_E.pdf.

Anonymous:1981:GIc

[683] Anonymous. General information. Visible Language, XV(3):311, Summer 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N3_1981_E.pdf.

Wrolstad:1981:ABc

[684] Merald E. Wrolstad. Advisory Board. Visible Language, XV(3): 312, Summer 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N3_1981_E.pdf.

Wrolstad:1981:ABd

[685] Merald E. Wrolstad. Advisory Board. Visible Language, XV(4): 318, Autumn 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N4_1981_E.pdf.

Powell:1981:EF

[686] Marvin A. Powell. Editor’s foreword. Visible Language, XV(4): 320, Autumn 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N4_1981_E.pdf.

Schmandt-Besserat:1981:TTR

[687] Denise Schmandt-Besserat. From tokens to tablets: a re-evaluation of the so-called “numerical tablets”. Visible Language, XV(4):321–344, Autumn 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N4_1981_E.pdf.

Green:1981:CIC

[688] M. W. Green. The construction and implementation of the cuneiform system. Visible Language, XV(4):345–372, Autumn 1981. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N4_1981_E.pdf.

Justus:1981:VSC

[689] Carol Justus. Visual sentences in cuneiform Hittite. Visible Language, XV(4):373–408, Autumn 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V15N4_1981_E.pdf.

Sack:1981:TSC

[690] Ronald H. Sack. The temple scribe in Chaldean Uruk. Visible Language, XV(4):409–418, Autumn 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V15N4_1981_E.pdf.

Powell:1981:TPH

[691] Marvin A. Powell. Three problems in the history of cuneiform writing: Origins, direction of script, and literacy. Visible Language, XV(4): 419–440, Autumn 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N4_1981_E.pdf.

Anonymous:1981:Ad

[692] Anonymous. The authors. Visible Language, XV(4):441, Autumn 1981. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V15N4_1981_E.pdf.

Anonymous:1981:IVX

[693] Anonymous. Index for Volume XV. Visible Language, XV(4):442– 444, Autumn 1981. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V15N4_1981_E.pdf.

Wrolstad:1982:ABa

[694] Merald E. Wrolstad. Advisory Board. Visible Language, XVI (1):2, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N1_1982_E.pdf.

Knuth:1982:CMF

[695] Donald E. Knuth. The concept of a Meta-Font. Visible Language, XVI (1):3–27, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N1_1982_E.pdf. See comments [737, 738].

French translation by M. R. Delorme in Communication et Langages no. 55, (1983), 40–53; reprinted in Typographie et Informatique, proceedings of INRIA conference held 21–25 January 1985, organized by Jacques André and Patrick Sallio (Rennes, France: INRIA/IRISA – CCETT, 1985), 119–132.

Skousen:1982:ESP

[696] Royal Skousen. English spelling and phonemic representation. Visible Language, XVI(1):28–38, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V16N1_1982_E.pdf.

Kelley:1982:CNM

[697] David H. Kelley. Costume and name in Mesoamerica. Visible Language, XVI(1):39–48, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N1_1982_E.pdf.

Weidemann:1982:BDT

[698] Kurt Weidemann. Biblica: Designing a typeface for the Bible. Visible Language, XVI(1):49–61, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V16N1_1982_E.pdf. Excerpted

from Deutscher Drucker (Satztechnische Beilage), nr. 34. 22. X. 1981.

Sekey:1982:TDS

[699] Andrew Sekey and Jerome Tietz. Text display by ‘saccadic scrolling’. Visible Language, XVI(1):63–77, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V16N1_1982_E.pdf.

Greene:1982:LW

[700] Maxine Greene. Literacy for what? Visible Language, XVI(1): 78–87, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N1_1982_E.pdf.

Wendt:1982:ILT

[701] Dirk Wendt. Improving the legibility of textbooks: Effects of wording and typographic design. Visible Language, XVI(1):88–93, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N1_1982_E.pdf.

Baudin:1982:BRO

[702] Fernand Baudin. Book review: A. S. Osley. Scribes and Sources: Handbook of the Chancery Hand in the Sixteenth Century. London: Faber & Faber, Ltd., 1980. 272 pp. 22 b/w illustrations. £16.00. Visible Language, XVI(1):94–102, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V16N1_1982_E.pdf.

Anonymous:1982:Aa

[703] Anonymous. The authors. Visible Language, XVI(1):103–104, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N1_1982_E.pdf.

Anonymous:1982:GIa

[704] Anonymous. General information. Visible Language, XVI(1):104, Winter 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N1_1982_E.pdf.

Wrolstad:1982:ABb

[705] Merald E. Wrolstad. Advisory Board. Visible Language, XVI(2): 106, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N2_1982_E.pdf.

Anonymous:1982:GIb

[706] Anonymous. General information. Visible Language, XVI(2):107, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N2_1982_E.pdf.

Popp:1982:LCP

[707] Helen M. Popp. Literacy and competency: The proceedings of a conference hosted by the Reading Laboratory at Harvard Graduate School of Education. Visible Language, XVI(2):109–110, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N2_1982_E.pdf.

Ylvisaker:1982:I

[708] Paul Ylvisaker. Introductions. Visible Language, XVI(2):111, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N2_1982_E.pdf.

Featherstone:1982:I

[709] Jay Featherstone. Introductions. Visible Language, XVI(2):112, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N2_1982_E.pdf.

Venezky:1982:OPD

[710] Richard L. Venezky. The origins of the present-day chasm between adult literacy needs and school literacy instruction. Visible Language, XVI (2):113–127, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N2_1982_E.pdf. See discussion [711].

Featherstone:1982:D

[711] Jay Featherstone, David Olson, Carman St. John Hunter, Richard Venezky, and Audience Participants. Discussion. Visible Language, XVI (2):128–136, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N2_1982_E.pdf. See [710].

StJohnHunter:1982:LES

[712] Carman St. John Hunter. Literacy for empowerment and social change. Visible Language, XVI(2):137–143, Spring 1982. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V16N2_1982_E.pdf. See discussion [713].

Olson:1982:D

[713] David Olson, Carman St. John Hunter, Richard Venezky, Jay Featherstone, and Audience Participants. Discussion. Visible Language, XVI (2):144–150, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N2_1982_E.pdf. See [712].

Olson:1982:WSW

[714] David R. Olson. What is said and what is meant in speech and writing. Visible Language, XVI(2):151–161, Spring 1982. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V16N2_1982_E.pdf. See discussion [715].

StJohnHunter:1982:D

[715] Carman St. John Hunter, David Olson, Richard Venezky, Jay Featherstone, Jeanne Chall, and Audience Participants. Discussion. Visible Language, XVI(2):162–174, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V16N2_1982_E.pdf. See [714].

Gray:1982:UAR

[716] William S. Gray. Universal ability to read and write! Visible Language, XVI(2):175, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N2_1982_E.pdf. Excerpt from the author’s

The Teaching of Reading and Writing: an International Survey. UN-ESCO Monographs on Fundamental Education, X. Paris: UNESCO, 1956, p. 9.

Anonymous:1982:CBT

[717] Anonymous. [colloquial book text]. Visible Language, XVI(2):176, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N2_1982_E.pdf.

Bowdre:1982:EDP

[718] Paul H. Bowdre, Jr. Eye dialect as a problem in graphics. Visible Language, XVI(2):177–183, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V16N2_1982_E.pdf.

Thogmartin:1982:MDB

[719] Clyde Thogmartin. Mr. Dooley’s brogue:the literary dialect of Finley Peter Dunne. Visible Language, XVI(2):184–198, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N2_1982_E.pdf.

Anonymous:1982:Ab

[720] Anonymous. The authors. Visible Language, XVI(2):199–200, Spring 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N2_1982_E.pdf.

Wrolstad:1982:ABc

[721] Merald E. Wrolstad. Advisory Board. Visible Language, XVI(3): 202, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N3_1982_E.pdf.

Anonymous:1982:GIc

[722] Anonymous. General information. Visible Language, XVI(3):203, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-

5529 (electronic).  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N3_1982_E.pdf.

Skemp:1982:M

[723] Richard R. Skemp. [on mathematics]. Visible Language, XVI(3): 204, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N3_1982_E.pdf.

Adda:1982:DMS

[724] Josette Adda. Difficulties with mathematical symbolism: Synonymy and homonymy. Visible Language, XVI(3):205–214, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N3_1982_E.pdf.

Buxton:1982:ERS

[725] Laurie G. Buxton. Emotional responses to symbolism. Visible Language, XVI(3):215–220, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V16N3_1982_E.pdf.

Goldin:1982:MLP

[726] Gerald A. Goldin. Mathematical language and problem solving. Visible Language, XVI(3):221–238, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V16N3_1982_E.pdf.

Higginson:1982:SIM

[727] William Higginson. Symbols, icons, and mathematical understanding. Visible Language, XVI(3):239–248, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V16N3_1982_E.pdf.

James:1982:TR

[728] Nick James and John Mason. Towards recording. Visible Language, XVI (3):249–258, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N3_1982_E.pdf.

Lay:1982:MIA

[729] L. Clark Lay. Mental images and arithmetical symbols. Visible Language, XVI(3):259–274, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V16N3_1982_E.pdf.

Lowenthal:1982:LAT

[730] Francis Lowenthal. Language acquisition through mathematical symbolism. Visible Language, XVI(3):275–280, Summer 1982. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N3_1982_E.pdf.

Skemp:1982:CMS

[731] Richard R. Skemp. Communicating mathematics: Surface structures and deep structures. Visible Language, XVI(3):281–288, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N3_1982_E.pdf.

Woodrow:1982:MS

[732] Derek Woodrow. Mathematical symbolism. Visible Language, XVI(3): 289–302, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N3_1982_E.pdf.

Anonymous:1982:Ac

[733] Anonymous. The authors. Visible Language, XVI(3):303–304, Summer 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N3_1982_E.pdf.

Wrolstad:1982:ABd

[734] Merald E. Wrolstad. Advisory Board. Visible Language, XVI(4): 306, Autumn 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N4_1982_E.pdf.

Anonymous:1982:GId

[735] Anonymous. General information. Visible Language, XVI(4):307, Autumn 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-

5529 (electronic).  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N4_1982_E.pdf.

Wrolstad:1982:EN

[736] Merald E. Wrolstad. Editor’s note. Visible Language, XVI(4): 308, Autumn 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N4_1982_E.pdf.

Hofstadter:1982:MFM

[737] Douglas R. Hofstadter. Meta-font, metamathematics, and metaphysics: Comments on Donald Knuth’s article “The Concept of a Meta-Font”. Visible Language, XVI(4):309–338, Autumn 1982. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N4_1982_E.pdf. See [695, 780].

Baudin:1982:OCT

[738] Fernand Baudin, Charles Bigelow, Henri-Paul Bronsard, David Ford, Gary Gore, Ed Fisher, Jr., W. P. Jaspert, Peter Karow, Albert Kapr, Alexander Nesbitt, John Schappler, Edward Rondthaler, Walter Tracy, Gerard Unger, Hermann Zapf, and Donald E. Knuth. Other comments by type designers and type directors. Visible Language, XVI(4): 339–359, Autumn 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N4_1982_E.pdf. See [695].

HaasDyson:1982:EVL

[739] Anne Haas Dyson. The emergence of visible language: Interrelationships: Drawing and early writing. Visible Language, XVI(4):360–381, Autumn 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N4_1982_E.pdf.

Bruce:1982:VCC

[740] Margaret Bruce and Jeremy J. Foster. The visibility of colored characters on colored backgrounds in Viewdata displays. Visible Language, XVI(4): 382–390, Autumn 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N4_1982_E.pdf.

Green:1982:WJT

[741] T. R. G. Green and S. J. Payne. The wooly jumper: Typographical problems of concurrency in information display. Visible Language, XVI (4):391–403, Autumn 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N4_1982_E.pdf.

Wrolstad:1982:SOM

[742] Merald E. Wrolstad. [Statement of ownership, management and circulation of Visible Language]. Visible Language, XVI(4):403, Autumn 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N4_1982_E.pdf.

Anonymous:1982:Ad

[743] Anonymous. The authors. Visible Language, XVI(4):404, Autumn 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V16N4_1982_E.pdf.

Anonymous:1982:IVX

[744] Anonymous. Index to Volume XVI. Visible Language, XVI(4):406– 408, Autumn 1982. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V16N4_1982_E.pdf.

Briem:1983:OR

[745] Gunnlauger S. E. Briem. Opening remarks. Visible Language, XVII (1):4–7, Winter 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N1_1983_E.pdf.

Briem:1983:WPY

[746] Gunnlauger S. E. Briem. ‘What parts of your work give you the most trouble?’. Visible Language, XVII(1):8–112, Winter 1983. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N1_1983_E.pdf.

Wrolstad:1983:ABa

[747] Merald E. Wrolstad. Advisory Board. Visible Language, XVII(2): 114, Spring 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N2_1983_E.pdf.

Anonymous:1983:GIb

[748] Anonymous. General information. Visible Language, XVII(2):115, Spring 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N2_1983_E.pdf.

Kline:1983:IRH

[749] Charles Robert Kline, Jr. Introduction: The renascence of die Hermeneute. Visible Language, XVII(2):117–119, Spring 1983. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N2_1983_E.pdf.

Kinneavy:1983:RWP

[750] James L. Kinneavy. The relation of the whole to the part in interpretation theory and in the composing process. Visible Language, XVII (2):120–145, Spring 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N2_1983_E.pdf.

Hyde:1983:HPA

[751] Michael J. Hyde. The hermeneutic phenomenon and the authenticity of discourse. Visible Language, XVII(2):156–162, Spring 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N2_1983_E.pdf.

Kline:1983:RWR

[752] Charles R. Kline, Jr. and Roland K. Huff. Reading, writing: Radix. Visible Language, XVII(2):163–176, Spring 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V17N2_1983_E.pdf.

Witte:1983:TSW

[753] Stephen P. Witte. Topical structure and writing quality: Some possible text-based explanations of readers’ judgments of student writing. Visible Language, XVII(2):177–205, Spring 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V17N2_1983_E.pdf.

Marino:1983:HC

[754] Adrian Marino. Hermeneutics as criticism. Visible Language, XVII (2):206–215, Spring 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N2_1983_E.pdf.

Anonymous:1983:Ab

[755] Anonymous. The authors. Visible Language, XVII(2):216, Spring 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N2_1983_E.pdf.

Anonymous:1983:GIc

[756] Anonymous. General information. Visible Language, XVII(3):3, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N3_1983_E.pdf.

Wrolstad:1983:ABc

[757] Merald E. Wrolstad. Advisory Board. Visible Language, XVII (3):4, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N3_1983_E.pdf.

Foster:1983:EF

[758] Stephen C. Foster. Editor’s foreword. Visible Language, XVII(3): 5–6, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N3_1983_E.pdf.

Foster:1983:LPV

[759] Stephen C. Foster. Letterisme: a point of view. Visible Language, XVII (3):7–12, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N3_1983_E.pdf.

Curtay:1983:C

[760] Jean-Paul Curtay. Chronology [1944–1982]. Visible Language, XVII (3):13–17, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N3_1983_E.pdf.

Seaman:1983:LSR

[761] David W. Seaman. Letterisme — a stream that runs own course. Visible Language, XVII(3):18–25, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V17N3_1983_E.pdf.

Curtay:1983:SWA

[762] Jean-Paul Curtay. Super-writing 1983 — America 1683. Visible Language, XVII(3):26–47, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V17N3_1983_E.pdf.

Devaux:1983:ALC

[763] Frédérique Devaux. Approaching letterist cinema. Visible Language, XVII(3):48–56, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V17N3_1983_E.pdf. Translated

by David W. Seaman.

Zurbrugg:1983:LLI

[764] Nicholas Zurbrugg. The limitations of letterisme: an interview with Henri Chopin. Visible Language, XVII(3):57–69, Summer 1983. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N3_1983_E.pdf.

Seaman:1983:STT

[765] David W. Seaman. Selected theoretical texts from letterists. Visible Language, XVII(3):70–83, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V17N3_1983_E.pdf.

Ferrua:1983:RLB

[766] Pietro Ferrua. Researching letterisme and bibliography. Visible Language, XVII(3):84–93, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V17N3_1983_E.pdf.

Anonymous:1983:EC

[767] Anonymous. Exhibition checklist. Visible Language, XVII(3):94– 110, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N3_1983_E.pdf.

Anonymous:1983:Ac

[768] Anonymous. The authors. Visible Language, XVII(3):111–112, Summer 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N3_1983_E.pdf.

Wrolstad:1983:ABd

[769] Merald E. Wrolstad. Advisory Board. Visible Language, XVII(4): 330, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N4_1983_E.pdf.

Anonymous:1983:GId

[770] Anonymous. General information. Visible Language, XVII(4):331, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-

5529 (electronic).  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N4_1983_E.pdf.

Anonymous:1983:ACa

[771] Anonymous. [about the cover]. Visible Language, XVII(4):332, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N4_1983_E.pdf.

Myers:1983:HEE

[772] Prue Wallis Myers. Handwriting in English education. Visible Language, XVII(4):333–356, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V17N4_1983_E.pdf.

Tsao:1983:IDC

[773] Yao-Chung Tsao and Tsai-Guey Wang. Information distribution in Chinese characters. Visible Language, XVII(4):357–364, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N4_1983_E.pdf.

Bar-Lev:1983:HH

[774] Zev Bar-Lev. Hebrew hieroglyphics. Visible Language, XVII(4):365– 379, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N4_1983_E.pdf.

Perlman:1983:GAT

[775] Gary Perlman and Thomas D. Erickson. Graphical abstractions of technical documents. Visible Language, XVII(4):380–389, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N4_1983_E.pdf.

Weber:1983:BS

[776] Rose-Marie Weber. Behind the slash. Visible Language, XVII(4): 390–395, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N4_1983_E.pdf.

Groff:1983:SWL

[777] Patrick Groff. Significance of word length. Visible Language, XVII(4): 396–398, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N4_1983_E.pdf.

Moskin:1983:FBE

[778] J. Robert Moskin. The future for books in the electronic era: a report on a seminar jointly sponsored by the Jerusalem International Book Fair and the Aspen Institute for Humanistic Studies. Visible Language, XVII (4):399–407, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N4_1983_E.pdf.

McNelly:1983:FDV

[779] Willis E. McNelly. Floppy discs vs. coffee stains. Visible Language, XVII (4):408–409, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N4_1983_E.pdf.

Sampson:1983:RTO

[780] Geoffrey Sampson. Is roman type an open-ended system? A response to Douglas Hofstadter. Visible Language, XVII(4):410–412, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V17N4_1983_E.pdf. See reply [737, 781].

Hofstadter:1983:R

[781] Douglas Hofstadter. [reply]. Visible Language, XVII(4):413–416, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N4_1983_E.pdf. See [737, 780].

Knuth:1983:C

[782] Donald E. Knuth and William C. Stokoe. Correspondence. Visible Language, XVII(4):417–418, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V17N4_1983_E.pdf. See [737,

695].

Anonymous:1983:Ad

[783] Anonymous. The authors. Visible Language, XVII(4):419–420, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N4_1983_E.pdf.

Wrolstad:1983:MAB

[784] Merald E. Wrolstad. [In memoriam: Advisory Board members G. Willem Ovink and Bror Zachrisson]. Visible Language, XVII(4):420, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V17N4_1983_E.pdf.

Anonymous:1983:IVX

[785] Anonymous. Index to Volume XVII. Visible Language, XVII(4): 422–424, Autumn 1983. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V17N4_1983_E.pdf.

Wrolstad:1984:ABa

[786] Merald E. Wrolstad. Advisory Board. Visible Language, XVIII (1):2, Winter 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N1_1984_E.pdf.

Anonymous:1984:GIa

[787] Anonymous. General information. Visible Language, XVIII(1):3, Winter 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N1_1984_E.pdf.

Anonymous:1984:PYA

[788] Anonymous. Photograph of young Ashley Smith. Visible Language, XVIII(1):4, Winter 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N1_1984_E.pdf.

Yaden:1984:RRM

[789] David B. Yaden, Jr. Reading research in metalinguistic awareness: Findings, problems, and classroom applications. Visible Language, XVIII (1):5–47, Winter 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N1_1984_E.pdf.

Schmandt-Besserat:1984:BN

[790] Denise Schmandt-Besserat. Before numerals. Visible Language, XVIII (1):48–60, Winter 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N1_1984_E.pdf.

Keenan:1984:ECL

[791] Stacey A. Keenan. Effects of chunking and line length on reading efficiency. Visible Language, XVIII(1):61–80, Winter 1984. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N1_1984_E.pdf.

Baudin:1984:VET

[792] Fernand Baudin. The visual editing of texts. Visible Language, XVIII (1):81–86, Winter 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N1_1984_E.pdf.

Avrin:1984:HMO

[793] Leila Avrin. Hebrew micrography: One thousand years of art in script. Visible Language, XVIII(1):87–95, Winter 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V18N1_1984_E.pdf.

Anonymous:1984:Aa

[794] Anonymous. The authors. Visible Language, XVIII(1):96, Winter 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N1_1983_E.pdf.

Wrolstad:1984:ABb

[795] Merald E. Wrolstad. Advisory Board. Visible Language, XVIII (2):98, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N2_1984_E.pdf.

Wright:1984:ISI

[796] Patricia Wright. Introduction to the special issue: Some effects of communication medium on visible language. Visible Language, XVIII (2):98–101, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N2_1984_E.pdf.

Walker:1984:HTC

[797] Sue Walker. How typewriters changed correspondence. Visible Language, XVIII(2):102–117, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V18N2_1984_E.pdf.

Baron:1984:CMC

[798] Naomi S. Baron. Computer mediated communication as a force in language change. Visible Language, XVIII(2):118–141, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N2_1984_E.pdf.

Bryant:1984:RLC

[799] Philip Bryant. Reading library catalogues and indexes. Visible Language, XVIII(2):142–153, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V18N2_1984_E.pdf.

Norrish:1984:MTP

[800] Pat Norrish. Moving tables from paper to CRT screen. Visible Language, XVIII(2):154–170, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V18N2_1984_E.pdf.

Pullinger:1984:DPC

[801] David J. Pullinger. Design and presentation of Computer Human Factors journal on the BLEND system. Visible Language, XVIII(2): 171–185, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N2_1984_E.pdf.

Wright:1984:IRR

[802] P. Wright and A. Lickorish. Investigating referee’s requirements in an electronic medium. Visible Language, XVIII(2):186–205, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N2_1984_E.pdf.

Anonymous:1984:GIb

[803] Anonymous. General information. Visible Language, XVIII(2): 206, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N2_1984_E.pdf.

Anonymous:1984:Ab

[804] Anonymous. The authors. Visible Language, XVIII(2):207, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N2_1984_E.pdf.

Wrolstad:1984:SOM

[805] Merald E. Wrolstad. [Statement of ownership, management and circulation of Visible Language]. Visible Language, XVIII(2):208, Spring 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N2_1984_E.pdf.

Wrolstad:1984:ABc

[806] Merald E. Wrolstad. Advisory Board. Visible Language, XVIII(3): 210, Summer 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N3_1984_E.pdf.

Anonymous:1984:GIc

[807] Anonymous. General information. Visible Language, XVIII(3): 211, Summer 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N3_1984_E.pdf.

Seeley:1984:SIA

[808] Chris Seeley. Special issue: Aspects of the Japanese writing system. Visible Language, XVIII(3):213–218, Summer 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V18N3_1984_E.pdf.

Backhouse:1984:AGS

[809] A. E. Backhouse. Aspects of the graphological structure of Japanese. Visible Language, XVIII(3):219–228, Summer 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V18N3_1984_E.pdf.

Twine:1984:APJ

[810] Nannette Twine. The adoption of punctuation in Japanese script. Visible Language, XVIII(3):229–237, Summer 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V18N3_1984_E.pdf.

Unger:1984:JOC

[811] J. Marshall Unger. Japanese orthography in the computer age. Visible Language, XVIII(3):238–253, Summer 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V18N3_1984_E.pdf.

Unger:1984:JB

[812] J. Marshall Unger. Japanese Braille. Visible Language, XVIII(3): 254–266, Summer 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N3_1984_E.pdf.

Seeley:1984:JSS

[813] Chris Seeley. Japanese script since 1900. Visible Language, XVIII(3): 267–301, Summer 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N3_1984_E.pdf.

Anonymous:1984:Ac

[814] Anonymous. The authors. Visible Language, XVIII(3):303, Summer 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N3_1984_E.pdf.

Wrolstad:1984:AB

[815] Merald E. Wrolstad. Advisory Board. Visible Language, XVIII(4): 306, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N4_1984_E.pdf.

Anonymous:1984:GI

[816] Anonymous. General information. Visible Language, XVIII(4): 307, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N4_1984_E.pdf.

Massaro:1984:SIP

[817] Dominic Massaro. Special issue: Psychological processes in reading. edited by Dominic Massaro, University of California, Santa Cruz and Cesare Cornoldi, University of Padova, Italy. Visible Language, XVIII (4):309, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N4_1984_E.pdf.

Baddeley:1984:RWM

[818] Alan Baddeley. Reading and working memory. Visible Language, XVIII (4):311–322, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N4_1984_E.pdf.

Massaro:1984:RAK

[819] Dominic W. Massaro. Reading ability and knowledge of orthographic structure. Visible Language, XVIII(4):323–332, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N4_1984_E.pdf.

Simion:1984:ADC

[820] Francesca Simion, Beatrice Benelli, and Franca Tarantini. Is activation of different codes related to age and stimulus material? Visible Language, XVIII(4):333–341, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V18N4_1984_E.pdf.

Burani:1984:MSL

[821] Cristina Burani, Dario Salmaso, and Alfonso Caramazza. Morphological structures and lexical access. Visible Language, XVIII(4):342– 352, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N4_1984_E.pdf.

FloresdArcais:1984:LKW

[822] Giavanni B. Flores d’Arcais. Lexical knowledge and word recognition: Children’s reading of function words. Visible Language, XVIII (4):353–371, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N4_1984_E.pdf.

Cavedon:1984:SVS

[823] Adele Cavedon, Cesare Cornoldi, and Rossana DeBeni. Structural vs. semantic coding in reading of isolated works by deaf children. Visible Language, XVIII(4):372–381, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V18N4_1984_E.pdf.

Masterson:1984:SDD

[824] Jacqueline Masterson. Surface dyslexia and developmental disorders of reading. Visible Language, XVIII(4):382–390, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N4_1984_E.pdf.

Blachowicz:1984:RRC

[825] Camille L. Z. Blachowicz. Reading and remembering: a constructivist perspective on reading comprehension and its disorders. Visible Language, XVIII(4):391–403, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V18N4_1984_E.pdf.

Anonymous:1984:Ad

[826] Anonymous. The authors. Visible Language, XVIII(4):404–405, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V18N4_1984_E.pdf.

Anonymous:1984:IVX

[827] Anonymous. Index to Volume XVIII. Visible Language, XVIII(4): 406–408, Autumn 1984. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V18N4_1984_E.pdf.

Wrolstad:1985:ABa

[828] Merald E. Wrolstad. Advisory Board. Visible Language, XIX (1):2, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_E.pdf.

Anonymous:1985:GIa

[829] Anonymous. General information. Visible Language, XIX(1):3, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N1_1985_E.pdf.

Anonymous:1985:PMF

[830] Anonymous. The printer’s mark of Friedrich Peypus. Nuremberg, active 1512–1534. Visible Language, XIX(1):4, Winter 1985. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N1_1985_E.pdf.

Bigelow:1985:I

[831] Charles Bigelow. Introduction. Visible Language, XIX(1):5–10, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N1_1985_E.pdf.

Dreyfus:1985:TPT

[832] John Dreyfus. A turning point in type design. Visible Language, XIX (1):11–22, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_E.pdf.

Zapf:1985:FTT

[833] Hermann Zapf. Future tendencies in type design: The scientific approach to letterforms. Visible Language, XIX(1):23–34, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N1_1985_E.pdf.

Knuth:1985:LLM

[834] Donald E. Knuth. Lessons learned from Metafont. Visible Language, XIX(1):35–53, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_E.pdf.

Anonymous:1985:PMW

[835] Anonymous. The printer’s mark of Wechel. 1495–1554. Visible Language, XIX(1):54, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_E.pdf.

LopesCardozo:1985:SKW

[836] Lida Lopes Cardozo. Stonecuttings from Kindersley’s workshop. Visible Language, XIX(1):55–59, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N1_1985_E.pdf.

Stauffacher:1985:TPK

[837] Jack Stauffacher. The Transylvanian phoenix: The Kis–Janson types in the digital era. Visible Language, XIX(1):60–76, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N1_1985_E.pdf.

Carter:1985:GMR

[838] Matthew Carter. Galliard: a modern revival of types of Robert Granjon. Visible Language, XIX(1):77–98, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N1_1985_E.pdf.

Drost:1985:PD

[839] Henk Drost. Punchcutting demonstration. Visible Language, XIX(1): 99–106, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_E.pdf.

Nelson:1985:MMM

[840] Stan Nelson. Mould making, matrix fitting, and hand casting. Visible Language, XIX(1):107–120, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N1_1985_E.pdf.

Anonymous:1985:PMH

[841] Anonymous. The printer’s mark of The Hackiana Press. Leyden, 1671. Visible Language, XIX(1):121, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N1_1985_E.pdf.

Anonymous:1985:PMAa

[842] Anonymous. The printer’s mark of Antoine De Harsy. Lyon, 1550. Visible Language, XIX(1):122, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N1_1985_E.pdf.

Gurtler:1985:FRM

[843] André Gürtler and Christian Mengelt. Fundamental research methods and form innovations in type design compared to technological developments in type production. Visible Language, XIX(1):123–147, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N1_1985_E.pdf.

Anonymous:1985:PMG

[844] Anonymous. The printer’s mark of The Guinta Press. Antwerp. Visible Language, XIX(1):148, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N1_1985_E.pdf.

Gottschall:1985:SAT

[845] Edward Gottschall. The state of the art in typeface design protection. Visible Language, XIX(1):149–155, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V19N1_1985_E.pdf.

Anonymous:1985:PMAb

[846] Anonymous. The printer’s mark of Aldus Manutius. Venice, 1450–1515. Visible Language, XIX(1):156, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N1_1985_E.pdf.

Baudin:1985:AFA

[847] Fernand Baudin, Inga Wennik, and Margaret Tsai. Abstracts — French; abstracts — German; abstracts — Japanese. Visible Language, XIX (1):157–163, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_E.pdf.

Anonymous:1985:Aa

[848] Anonymous. The authors. Visible Language, XIX(1):164–166, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N1_1985_E.pdf.

Anonymous:1985:C

[849] Anonymous. Colophon. Visible Language, XIX(1):167–168, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N1_1985_E.pdf.

Poggenpohl:1985:CGB

[850] Sharon Helmer Poggenpohl. Computer graphics: Bridging the gap. Visible Language, XIX(2):173–177, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N2_1985_E.pdf.

Owen:1985:GDC

[851] Charles Owen, Roger Remington, Michael Twyman, and Sharon Poggenpohl. Graphic design: Computer graphics: What do they mean? How do they fit? Visible Language, XIX(2):178–225, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N2_1985_E.pdf.

Shamonsky:1985:SGG

[852] Dorothy Shamonsky. Scripting graphic with graphics. Visible Language, XIX(2):226–242, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V19N2_1985_E.pdf.

MillerRubin:1985:SCA

[853] Morissa Miller Rubin. Spatial context as an aid to page layout: a system for planning and sketching. Visible Language, XIX(2):243–250, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N2_1985_E.pdf.

Nishimura:1985:DID

[854] Yoshiki Nishimura and Keiichi Sato. Dynamic information display. Visible Language, XIX(2):251–271, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N2_1985_E.pdf.

Allmendinger:1985:CDE

[855] Leif Allmendinger and Mihai Nadin. Computers in design education: a case study. Visible Language, XIX(2):272–287, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N2_1985_E.pdf.

Jones:1985:GDT

[856] Mary Jones. Graphic design: Towards digital applications. Visible Language, XIX(2):288–296, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N2_1985_E.pdf.

Anonymous:1985:Ba

[857] Anonymous. Bibliography. Visible Language, XIX(2):297, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N2_1985_E.pdf.

Anonymous:1985:ATF

[858] Anonymous. Abstracts in translation [French, German]. Visible Language, XIX(2):298–301, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V19N2_1985_E.pdf.

Anonymous:1985:Bb

[859] Anonymous. Biographies. Visible Language, XIX(2):302–304, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N2_1985_E.pdf.

Wrolstad:1985:ABb

[860] Merald E. Wrolstad. Advisory Board. Visible Language, XIX(2): 304–305, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N2_1985_E.pdf.

Anonymous:1985:GIb

[861] Anonymous. General information. Visible Language, XIX(2):305, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N2_1985_E.pdf.

Anonymous:1985:BI

[862] Anonymous. Back issues. Visible Language, XIX(2):306, Spring 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N2_1985_E.pdf.

Wrolstad:1985:ABc

[863] Merald E. Wrolstad. Advisory Board. Visible Language, XIX(3): 308, Summer 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N3_1985_E.pdf.

Anonymous:1985:GIc

[864] Anonymous. General information. Visible Language, XIX(3):309– 310, Summer 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N3_1985_E.pdf.

Nickerson:1985:ALT

[865] Raymond S. Nickerson. Adult literacy and technology. Visible Language, XIX(3):311–355, Summer 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V19N3_1985_E.pdf.

Harker:1985:NIL

[866] W. John Harker. The new imperative in literary criticism. Visible Language, XIX(3):356–372, Summer 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N3_1985_E.pdf.

Viglionese:1985:IFW

[867] Paschal C. Viglionese. The inner functioning of words: Iconicity in poetic language. Visible Language, XIX(3):373–386, Summer 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N3_1985_E.pdf.

Nickerson:1985:MEC

[868] Jeff Nickerson. The mind’s eye and the CRT terminal: Towards a diagrammatic interface. Visible Language, XIX(3):387–400, Summer 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N3_1985_E.pdf.

Anonymous:1985:Ac

[869] Anonymous. The authors. Visible Language, XIX(3):401, Summer 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N3_1985_E.pdf.

Anonymous:1985:BIV

[870] Anonymous. Back issues of Visible Language in print. Visible Language, XIX(3):402, Summer 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N3_1985_E.pdf.

Wrolstad:1985:ABd

[871] Merald E. Wrolstad. Advisory Board. Visible Language, XIX(4): 404, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N4_1985_E.pdf.

Preckshot:1985:ISI

[872] Judith E. Preckshot. Introduction to the special issue: The book, inside and out. Visible Language, XIX(4):405–409, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N4_1985_E.pdf.

Edson:1985:VLC

[873] Laurie Edson. Visible language in contemporary culture. Visible Language, XIX(4):410–425, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V19N4_1985_E.pdf.

Winspur:1985:PPP

[874] Steven Winspur. Poetry, portrait, poetrait. Visible Language, XIX(4): 426–438, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N4_1985_E.pdf.

James:1985:NSS

[875] Carol Plyley James. “No, says the signified”: The ‘logical status’ of words in painting. Visible Language, XIX(4):439–461, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N4_1985_E.pdf.

Christin:1985:VBC

[876] Anne-Marie Christin. A visionary book: Charles Nodier’s L’Histoire du Roi de Bohême et de ses sept châteaux. Visible Language, XIX(4): 462–483, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N4_1985_E.pdf.

Conley:1985:SBC

[877] Tom Conley. Samuel Beckett: Color, letter, and line. Visible Language, XIX(4):484–498, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V19N4_1985_E.pdf.

Preckshot:1985:PAP

[878] Judith Preckshot. Press art: Poets and their printing machines. Visible Language, XIX(4):499–518, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V19N4_1985_E.pdf.

RieseHubert:1985:RVR

[879] Renée Riese Hubert. Readable — visible: Reflections on the illustrated book. Visible Language, XIX(4):519–538, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N4_1985_E.pdf.

Anonymous:1985:Ad

[880] Anonymous. The authors. Visible Language, XIX(4):539–540, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N4_1985_E.pdf.

Anonymous:1985:GId

[881] Anonymous. General information. Visible Language, XIX(4):541– 542, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N4_1985_E.pdf.

Wrolstad:1985:SOM

[882] Merald E. Wrolstad. [Statement of ownership, management and circulation of Visible Language]. Visible Language, XIX(4):542, Autumn 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V19N4_1985_E.pdf.

Wrolstad:1986:ABa

[883] Merald E. Wrolstad. Advisory Board. Visible Language, XX(1):2, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N1_1986_E.pdf.

Anonymous:1986:GIa

[884] Anonymous. General information. Visible Language, XX(1):3, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N1_1986_E.pdf.

Higgins:1986:F

[885] Dick Higgins. Foreword. Visible Language, XX(1):5–7, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N1_1986_E.pdf.

Ernst:1986:FPT

[886] Ulrich Ernst. The figured poem: Towards a definition of genre. Visible Language, XX(1):8–27, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V20N1_1986_E.pdf.

Higgins:1986:CBO

[887] Dick Higgins. The corpus of British and other English-language pattern poetry. Visible Language, XX(1):28–51, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V20N1_1986_E.pdf.

Hatherly:1986:RPS

[888] Ana Hatherly. Reading paths in Spanish and Portuguese baroque labyrinths. Visible Language, XX(1):52–64, Winter 1986. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N1_1986_E.pdf.

Rypson:1986:LP

[889] Piotr Rypson. The labyrinth poem. Visible Language, XX(1):65– 95, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N1_1986_E.pdf.

Franke:1986:CPT

[890] Herbert Franke. Chinese patterned texts. Visible Language, XX(1): 96–108, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N1_1986_E.pdf.

Jha:1986:SCW

[891] Kalāñath Jhā. Sanskrit Citrakāvyas and the Western pattern poem: a critical appraisal. Visible Language, XX(1):109–120, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N1_1986_E.pdf.

Adler:1986:PTS

[892] Jeremy Adler. Pastoral typography: Sigmund von Birken and the “picture-rhymes” of Johann Helwig. Visible Language, XX(1):121– 135, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N1_1986_E.pdf.

Otto:1986:GWL

[893] Karl F. Otto, Jr. Georg Weber’s Lebens-Früchte (1649). Visible Language, XX(1):136–145, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V20N1_1986_E.pdf.

Anonymous:1986:Aa

[894] Anonymous. The authors. Visible Language, XX(1):146–147, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N1_1986_E.pdf.

Anonymous:1986:BIVa

[895] Anonymous. Back issues of Visible Language still in print. Visible Language, XX(1):148, Winter 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V20N1_1986_E.pdf.

Wrolstad:1986:ABb

[896] Merald E. Wrolstad. Advisory Board. Visible Language, XX(2): 150, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N2_1986_E.pdf.

Anonymous:1986:GIb

[897] Anonymous. General information. Visible Language, XX(2):151, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N2_1986_E.pdf.

Steinberg:1986:IPP

[898] Erwin R. Steinberg. Introduction: Promoting plain English. Visible Language, XX(2):153–154, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V20N2_1986_E.pdf.

Bowen:1986:AVA

[899] Betsy A. Bowen, Thomas M. Duffy, and Erwin R. Steinberg. Analyzing the various approaches of plain language laws. Visible Language, XX (2):155–165, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N2_1986_E.pdf.

Williams:1986:PER

[900] Joseph M. Williams. Plain English: The remaining problems. Visible Language, XX(2):166–173, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V20N2_1986_E.pdf.

Huckin:1986:PLP

[901] Thomas N. Huckin, Elizabeth H. Curtin, and Debra Graham. Prescriptive linguistics and plain English: The case of “whiz-deletions”. Visible Language, XX(2):174–187, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V20N2_1986_E.pdf.

Miller:1986:CCS

[902] Lance A. Miller. Computers for composition: a stage model approach to helping. Visible Language, XX(2):188–218, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N2_1986_E.pdf.

Jereb:1986:PEP

[903] Barry Jereb. Plain English on the plant floor. Visible Language, XX (2):219–225, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N2_1986_E.pdf.

Steinberg:1986:PID

[904] Erwin R. Steinberg. A program for improving documentation. Visible Language, XX(2):226–232, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V20N2_1986_E.pdf.

Agnew:1986:WBU

[905] G. Reed Agnew. Writing backwards: The use of visual models in writing. Visible Language, XX(2):233–241, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V20N2_1986_E.pdf.

Anonymous:1986:Ab

[906] Anonymous. The authors. Visible Language, XX(2):242–243, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N2_1986_E.pdf.

Anonymous:1986:BIVb

[907] Anonymous. Back issues of Visible Language still in print. Visible Language, XX(2):??, Spring 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V20N2_1986_E.pdf.

Anonymous:1986:GIc

[908] Anonymous. General information. Visible Language, XX(3):246, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N3_1986_E.pdf.

Wrolstad:1986:ABc

[909] Merald E. Wrolstad. Advisory Board. Visible Language, XX(3): 247, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N3_1986_E.pdf.

Olson:1986:ISI

[910] David Olson and Derrick de Kerckove. Introduction: [to special issue]. Visible Language, XX(3):248–249, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V20N3_1986_E.pdf.

Schmandt-Besserat:1986:TFI

[911] Denise Schmandt-Besserat. Tokens: Facts and interpretations. Visible Language, XX(3):250–273, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V20N3_1986_E.pdf.

deKerckhove:1986:ALB

[912] Derrick de Kerckhove. Alphabetic literacy and brain processes. Visible Language, XX(3):274–293, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V20N3_1986_E.pdf.

Stock:1986:TRE

[913] Brian Stock. Texts, readers, and enacted narratives. Visible Language, XX(3):394–301, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V20N3_1986_E.pdf.

Olson:1986:ITI

[914] David R. Olson. Interpreting texts and interpreting nature: The effects of literacy on hermeneutics and epistemology. Visible Language, XX(3): 302–317, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N3_1986_E.pdf.

Goody:1986:WRR

[915] Jack Goody. Writing, religion, and revolt in Bahia. Visible Language, XX (3):318–343, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N3_1986_E.pdf.

diNorcia:1986:SBF

[916] Vincent di Norcia. Of stone, books, and freedom. Visible Language, XX (3):344–354, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N3_1986_E.pdf.

Marvin:1986:IMM

[917] Carolyn Marvin. Innis, McLuhan, and Marx. Visible Language, XX(3): 355–359, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N3_1986_E.pdf.

Anonymous:1986:Ac

[918] Anonymous. The authors. Visible Language, XX(3):360–361, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N3_1986_E.pdf.

Anonymous:1986:BIVc

[919] Anonymous. Back issues of Visible Language still in print. Visible Language, XX(3):362, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V20N3_1986_E.pdf.

Anonymous:1986:NPM

[920] Anonymous. Notes on preparation of a manuscript for Visible Language. Visible Language, XX(3):363, Summer 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V20N3_1986_E.pdf.

Wrolstad:1986:ABd

[921] Merald E. Wrolstad. Advisory Board. Visible Language, XX(4): 366, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N4_1986_E.pdf.

Anonymous:1986:GId

[922] Anonymous. General information. Visible Language, XX(4):367, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N4_1986_E.pdf.

Kerr:1986:ITT

[923] Stephen T. Kerr. Instructional text: The transition from page to screen. Visible Language, XX(4):368–392, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V20N4_1986_E.pdf.

Clement:1986:ISC

[924] Richard W. Clement. Italian Sixteenth-Century Italian writing books and the scribal reality of Verona. Visible Language, XX(4):393– 412, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N4_1986_E.pdf.

Weber:1986:VSS

[925] Rose-Marie Weber. Variations in spelling and the special case of colloquial contractions. Visible Language, XX(4):413–426, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N4_1986_E.pdf.

DeJulio:1986:PDM

[926] Maryanne De Julio. Proximité du murmure: Dupin and Ubac collaboratel. Visible Language, XX(4):427–436, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V20N4_1986_E.pdf.

Ormrod:1986:DBG

[927] Jeanne Ellis Ormrod. Differences between good and poor spellers in reading style and short-term memory. Visible Language, XX(4): 437–447, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N4_1986_E.pdf.

Moriarty:1986:LLS

[928] Sandra E. Moriarty. Line lengths and starch scores. Visible Language, XX (4):448–455, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N4_1986_E.pdf.

Anonymous:1986:Ad

[929] Anonymous. The authors. Visible Language, XX(4):456–457, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V20N4_1986_E.pdf.

Anonymous:1986:IVX

[930] Anonymous. Index to Volume XX. Visible Language, XX(4):458– 459, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V20N4_1986_E.pdf.

Anonymous:1986:BIVd

[931] Anonymous. Back issues of Visible Language still in print. Visible Language, XX(4):460, Autumn 1986. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V20N4_1986_E.pdf.

Briem:1987:VTM

[932] Gunnlaugur S. E. Briem, Richard L. Venezky, Wim Crouwel, Michael Twyman, Robert Norton, John Mountford, Colin Banks, Krzysztof Lenk, Dominic W. Massaro, Fernand Baudin, and Sharon Helmer Poggenpohl. A visible tribute to Merald Wrolstad. Visible Language, XXI (1):3–15, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N1_1987_E.pdf.

Hodgson:1987:F

[933] Richard Hodgson and Ralph Sarkonak. Foreword. Visible Language, XXI (1):16–17, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N1_1987_E.pdf.

Hodgson:1987:GCL

[934] Richard Hodgson and Ralph Sarkonak. Graphic collisions: Languages in contact. Visible Language, XXI(1):18–41, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N1_1987_E.pdf.

Crawford:1987:BT

[935] Alistair Crawford. Bilingual typography. Visible Language, XXI(1): 42–65, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N1_1987_E.pdf.

Schogt:1987:FLD

[936] Henry Schogt. Foreign loanwords in Dutch: Integration and adaptation. Visible Language, XXI(1):66–87, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V21N1_1987_E.pdf.

Saint-Jacques:1987:RAJ

[937] Bernard Saint-Jacques. The Roman alphabet in the Japanese writing system. Visible Language, XXI(1):88–105, Winter 1987. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V21N1_1987_E.pdf. See also later study [1508].

Wrenn:1987:OMG

[938] Phyllis Wrenn. Ortho and morpho-graphic transcoding of Acadian ‘Franglais’. Visible Language, XXI(1):106–129, Winter 1987. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N1_1987_E.pdf.

Flores:1987:CLW

[939] Lauro Flores. Converging languages in a world of conflicts: Codeswitching in Chicano poetry. Visible Language, XXI(1):130–152, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N1_1987_E.pdf.

Anonymous:1987:Aa

[940] Anonymous. The authors. Visible Language, XXI(1):153–154, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N1_1987_E.pdf.

Anonymous:1987:SFR

[941] Anonymous. Suggestions for further reading. Visible Language, XXI (1):155–156, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N1_1987_E.pdf.

Poggenpohl:1987:ABa

[942] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, XXI (1):157, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N1_1987_E.pdf.

Anonymous:1987:GIa

[943] Anonymous. General information. Visible Language, XXI(1):158, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N1_1987_E.pdf.

Anonymous:1987:BIa

[944] Anonymous. Back issues. Visible Language, XXI(1):159, Winter 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N1_1987_E.pdf.

Monaghan:1987:I

[945] E. Jennifer Monaghan. Introduction. Visible Language, XXI(2): 161–165, Spring 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N2_1987_E.pdf.

Monaghan:1987:RWC

[946] E. Jennifer Monaghan. Readers writing: The curriculum of the writing schools of Eighteenth Century Boston. Visible Language, XXI(2): 167–213, Spring 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N2_1987_E.pdf.

Hodges:1987:ASI

[947] Richard E. Hodges. American spelling instruction: Retrospect and prospect. Visible Language, XXI(2):215–234, Spring 1987. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N2_1987_E.pdf.

vonBrachtDonsky:1987:WP

[948] Barbara von Bracht Donsky. Writing as praxis 1900–1959. Visible Language, XXI(2):237–251, Spring 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V21N2_1987_E.pdf.

Campbell:1987:FGW

[949] Anne Campbell. Fourth grade writing achievement and instruction, 1974–1984: NAEP’s report card. Visible Language, XXI(2): 253–268, Spring 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N2_1987_E.pdf.

Mavrogenes:1987:YCT

[950] Nancy A. Mavrogenes. Young children then and now: Recent research on emergent literacy. Visible Language, XXI(2):271–297, Spring 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N2_1987_E.pdf.

Anonymous:1987:Ab

[951] Anonymous. The authors. Visible Language, XXI(2):298–299, Spring 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N2_1987_E.pdf.

Poggenpohl:1987:AB

[952] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, XXI (2):300, Spring 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N2_1987_E.pdf.

Anonymous:1987:GIc

[953] Anonymous. General information. Visible Language, XXI(3–4):308, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N3_1987_E.pdf.

Anonymous:1987:DN

[954] Anonymous. Design notes. Visible Language, XXI(3–4):309, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N3_1987_E.pdf.

Anonymous:1987:C

[955] Anonymous. Colophon. Visible Language, XXI(3–4):310, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N3_1987_E.pdf.

Allen:1987:P

[956] Roy F. Allen, Stephen C. Foster, and Estera Milman. Preface. Visible Language, XXI(3–4):311, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V21N2_1987_E.pdf.

Anonymous:1987:Ac

[957] Anonymous. Abstracts. Visible Language, XXI(3–4):312, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N2_1987_E.pdf.

Foster:1987:PT

[958] Stephen C. Foster. The prerequisite text. Visible Language, XXI(3– 4):313–333, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V21N3_1987_E.pdf.

Milman:1987:TMA

[959] Estera Milman. The text and the myth of the avant-garde. Visible Language, XXI(3–4):335–363, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V21N3_1987_E.pdf.

Benson:1987:TCA

[960] Timothy O. Benson. The text and the coming of age of the avant-garde in Germany. Visible Language, XXI(3–4):365–411, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N3_1987_E.pdf.

Guenther:1987:BD

[961] Peter Guenther. Berlin dada. Visible Language, XXI(3–4):413–451, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N3_1987_E.pdf.

Rumold:1987:DTP

[962] Rainer Rumold. The Dadaist text: Politics, aesthetics, and alternative cultures? Visible Language, XXI(3–4):453–489, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N3_1987_E.pdf.

Anonymous:1987:Ad

[963] Anonymous. Abstracts. Visible Language, XXI(3–4):491–495, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N2_1987_E.pdf.

Anonymous:1987:IVX

[964] Anonymous. Index to Volume XXI. Visible Language, XXI(3–4):497–501, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N2_1987_E.pdf.

Allen:1987:TAG

[965] Roy Allen. The tradition of the avant-garde. Visible Language, XXI (3–4):494–??, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V21N3_1987_E.pdf.

Greenberg:1987:BTA

[966] Allen Greenberg. Between text and audience: a path to the future. Visible Language, XXI(3–4):454–??, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V21N3_1987_E.pdf.

Anonymous:1987:CC

[967] Anonymous. Catalogue checklist. Visible Language, XXI(3–4):416–??, Summer/Autumn 1987. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V21N2_1987_E.pdf.

Poggenpohl:1988:ABa

[968] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, XXII (1):4, Winter 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N1_1988_E.pdf.

Campbell:1988:IGE

[969] Ruth Campbell. Introduction by Guest Editor. Visible Language, XXII (1):5–7, Winter 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N1_1988_E.pdf.

Massaro:1988:VLS

[970] Dominic W. Massaro, Michael M. Cohen, and Laura A. Thompson. Visible language in speech perception: Lipreading and reading. Visible Language, XXII(1):8–31, Winter 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V22N1_1988_E.pdf.

Campbell:1988:TLM

[971] Ruth Campbell. Tracing lip movements: Making speech visible. Visible Language, XXII(1):32–57, Winter 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V22N1_1988_E.pdf.

Dodd:1988:CME

[972] Barbara Dodd, Michael Oerlemans, and Ray Robinson. Cross-model effects in repetition priming: a comparison of lipread graphic and heard stimuli. Visible Language, XXII(1):58–77, Winter 1988. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V22N1_1988_E.pdf.

Vinter:1988:PFM

[973] Annie Vinter. Perception of facial movements in early infancy: Some reflections in relation to speech perception. Visible Language, XXII (1):78–111, Winter 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N1_1988_E.pdf.

Storey:1988:RSD

[974] Darryl Storey and Martin Roberts. Reading the speech of digital lips: Motives and method for audio-visual speech synthesis. Visible Language, XXII(1):112–127, Winter 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V22N1_1988_E.pdf.

Graves:1988:STS

[975] Roger E. Graves and Susan M. Potter. Speaking with two sides of the mouth. Visible Language, XXII(1):128–137, Winter 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V22N1_1988_E.pdf.

Poggenpohl:1988:VLA

[976] Sharon Helmer Poggenpohl. Visible Language advisors, research interests, and upcoming issues. Visible Language, XXII(1):138–144, Winter 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V22N1_1988_E.pdf.

Kintgen:1988:LL

[977] Eugene R. Kintgen. Literacy literacy. Visible Language, XXII(2–3):149– 168, Spring/Summer 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N23_1988_E.pdf.

Bradford:1988:SWP

[978] Richard Bradford. Speech and writing in poetry and its criticism. Visible Language, XXII(2–3):169–194, Spring/Summer 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V22N23_1988_E.pdf.

Lange:1988:VVT

[979] Martha Scotford Lange. Verbal and visual translation of Mayakovsky’s and Lissitsky’s for reading out loud. Visible Language, XXII(2–3):195– 222, Spring/Summer 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N23_1988_E.pdf.

Bojko:1988:RLC

[980] Szymon Bojko and Krzysztof Lenk. For reading out loud in context. Visible Language, XXII(2–3):223–231, Spring/Summer 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V22N23_1988_E.pdf.

Vande:1988:MRM

[981] William J. Vande and Allen Shoemaker. Metadiscourse and the recall of modality markers. Visible Language, XXII(2–3):232–272, Spring/Summer 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N23_1988_E.pdf.

Garofalo:1988:TCA

[982] Karen M. Garofalo. Typographic cues as an aid to learning from textbooks. Visible Language, XXII(2–3):273–297, Spring/Summer 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V22N23_1988_E.pdf. See response [1020] and reply [1021]..

Cavalier:1988:MVT

[983] Todd Cavalier. Meditation: Visual transition as a bridge between form and meaning. Visible Language, XXII(2–3):298–329, Spring/Summer 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V22N23_1988_E.pdf.

Baudin:1988:BRG

[984] Fernand Baudin. Book review: György Haiman, Nicholas Kis: a Hungarian Punch-Cutter and Printer 1650–1702, The Greenwood Press, 1983. 452 pages, 8 color plates, 173 figures. Visible Language, XXII(2–3):330– 337, Spring/Summer 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

Gray:1988:BRW

[985] Nicolette Gray. Book review: Wolf Von Eckhardt, Please write. How to improve your hand-writing for business and pleasure in ten quick and easy lessons, Anthenum, New York, 1988. Visible Language, XXII(2– 3):338–340, Spring/Summer 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

Saper:1988:ITM

[986] Craig Saper. Instant theory: Making thinking popular. Visible Language, XXII(4):371–398, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V22N4_1988_E.pdf.

Ulmer:1988:HTH

[987] Gregory L. Ulmer. Handbook for a theory hobby: The hobby-horse is the sawhorse of theory. Visible Language, XXII(4):399–422, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V22N4_1988_E.pdf.

Ray:1988:AVT

[988] Robert B. Ray. The ABC of visual theory. Visible Language, XXII(4): 423–448, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N4_1988_E.pdf.

Sparling:1988:DP

[989] Bonnie Sparling. Decoder process. Visible Language, XXII(4):449– 454, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N4_1988_E.pdf.

Laiwan:1988:IS

[990] Laiwan. The imperialism of syntax. Visible Language, XXII(4):455– 458, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N4_1988_E.pdf.

Ronell:1988:CA

[991] Avital Ronell. Condensed article. Visible Language, XXII(4):459– 482, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N4_1988_E.pdf.

Ragland-Sullivan:1988:WR

[992] Ellie Ragland-Sullivan. A writing of the real. Visible Language, XXII (4):483–495, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N4_1988_E.pdf.

Anonymous:1988:Ad

[993] Anonymous. The authors. Visible Language, XXII(4):486–496, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V22N4_1988_E.pdf.

Anonymous:1988:VXI

[994] Anonymous. Volume XXII index. Visible Language, XXII(4):498– 500, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N4_1988_E.pdf.

Poggenpohl:1988:ABd

[995] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, XXII (4):501, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N4_1988_E.pdf.

Anonymous:1988:JI

[996] Anonymous. Journal information. Visible Language, XXII(4):502– 503, Autumn 1988. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V22N4_1988_E.pdf.

Bradford:1989:VPE

[997] Richard Bradford. The visual poem in the Eighteenth Century. Visible Language, XXIII(1):??, January 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V23N1_1989_E.pdf.

York:1989:MAU

[998] R. A. York. Mallarmé and apollinaire: The unpunctured text. Visible Language, XXIII(1):??, January 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V23N1_1989_E.pdf.

Webster:1989:WFO

[999] Michael Webster. Words-in-freedom and the oral tradition. Visible Language, XXIII(1):??, January 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V23N1_1989_E.pdf.

Berry:1989:VFF

[1000] Eleanor Berry. Visual form in free verse. Visible Language, XXIII (1):??, January 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N1_1989_E.pdf.

Markham:1989:WPR

[1001] E. A. Markham. Which poem am i reading? Visible Language, XXIII (1):??, January 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N1_1989_E.pdf.

Waterman:1989:SAL

[1002] Andrew Waterman. Soundings along the lines. Visible Language, XXIII (1):??, January 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N1_1989_E.pdf.

Gandelman:1989:WII

[1003] Claude Gandelman. By way of introduction: Inscriptions as subversion. Visible Language, XXIII(2–3):??, May 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V23N23_1989_E.pdf.

Barasch:1989:SOP

[1004] Moshe Barasch. Some oriental pseudo-inscriptions in renaissance art. Visible Language, XXIII(2–3):??, May 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V23N23_1989_E.pdf.

Martin:1989:OWO

[1005] Louis Martin. The order of words and the order of things in painting. Visible Language, XXIII(2–3):??, May 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V23N23_1989_E.pdf.

Lussier:1989:CDD

[1006] Mark Lussier. The contra-diction of design: Blake’s illustrations to Gray’s “Ode on the Death of a Favourite Cat”. Visible Language, XXIII(2–3):??, May 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N23_1989_E.pdf.

Roque:1989:MWI

[1007] George Roque. Magritte’s words and images. Visible Language, XXIII (2–3):??, May 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N23_1989_E.pdf.

Hubert:1989:BFL

[1008] Renée Riese Hubert. The books of fernand léger: Illustration and inscription. Visible Language, XXIII(2–3):??, May 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V23N23_1989_E.pdf.

Levinger:1989:JJP

[1009] Esther Levinger. Jasper Johns’ painted words. Visible Language, XXIII (2–3):??, May 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N23_1989_E.pdf.

Gandelman:1989:VAM

[1010] Claude Gandelman. On the verbal art of a modern painter: the work of Jules Kirschenbaum. Visible Language, XXIII(2–3):??, May 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V23N23_1989_E.pdf.

Ortquist:1989:MCR

[1011] Leslie Ortquist. Magritte’s captivity in Robbe-Grillet’s La Belle Captive: The subjugation of the image by the word. Visible Language, XXIII(2–3):??, May 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N23_1989_E.pdf.

Poggenpohl:1989:ABd

[1012] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, XXIII (4):318, Autumn 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N4_1989_E.pdf.

Berger:1989:DTS

[1013] Paul Eric Berger. Documentation: Time, space and the blackboard. Visible Language, XXIII(4):321–325, Autumn 1989. CODEN VSLGAO.

ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V23N4_1989_E.pdf.

Moilanen:1989:EIH

[1014] Carolyn Moilanen and Charles Lehman. The effects of italic handwriting on legibility: The methods and findings of a three-year study. Visible Language, XXIII(4):327–352, October 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V23N4_1989_E.pdf.

Bear:1989:WBR

[1015] Donald R. Bear. Why beginning reading must be word-by-word: Disfluent oral reading and orthographic development. Visible Language, XXIII (4):353–367, October 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N4_1989_E.pdf.

Riedinger:1989:TTT

[1016] Edward A. Riedinger. The tales typography tells. Visible Language, XXIII(4):369–374, October 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V23N4_1989_E.pdf.

Kadalenos:1989:INR

[1017] Emma Kadalenos. Image and narrativity: Robbe-Grillet’s La Belle Captive. Visible Language, XXIII(4):375–392, October 1989. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V23N4_1989_E.pdf.

Twyman:1989:BRH

[1018] Michael Twyman. Book review: Herman Zapf & His Design Philosophy. Visible Language, XXIII(4):393–400, Autumn 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V23N4_1989_E.pdf.

Singleton:1989:MA

[1019] Barbara Cunliffe Singleton. Margin artist. Visible Language, XXIII(4): 401–405, Autumn 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N4_1989_E.pdf.

Hartley:1989:STC

[1020] James Hartley. Spatial and typographic cues in text: a response to Garofalo. Visible Language, XXIII(4):407–411, Autumn 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V23N4_1989_E.pdf. See [982, 1021].

Garofalo:1989:GRJ

[1021] Karen M. Garofalo. Garofalo responds to James Hartley. Visible Language, XXIII(4):411, Autumn 1989. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V23N4_1989_E.pdf. See [982,

1020].

Nesbitt:1989:RC

[1022] Alexander Nesbitt, Fernand Baudin, and Adam Blatner. Reader comments. Visible Language, XXIII(4):412–415, Autumn 1989. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V23N4_1989_E.pdf. See [?] [Book review in Volume XXII, 2–3].

Anonymous:1989:IVX

[1023] Anonymous. Index to Volume XXIII. Visible Language, XXIII(4): 416–418, Autumn 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N4_1989_E.pdf.

Anonymous:1989:GId

[1024] Anonymous. General information. Visible Language, XXIII(4):419– 420, Autumn 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N4_1989_E.pdf.

Anonymous:1989:CPR

[1025] Anonymous. Call for papers: RIDT-91: the second international workshop on raster imaging and digital typography. Visible Language, XXIII (4):420, Autumn 1989. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V23N4_1989_E.pdf.

Poggenpohl:1990:ABa

[1026] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, XXIV (1):2, Winter 1990. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V24N1_1990_E.pdf.

Schmandt-Besserat:1990:EF

[1027] Denise Schmandt-Besserat. Editor’s foreword. Visible Language, XXIV (1):5, Winter 1990. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V24N1_1990_E.pdf.

Reilly:1990:OM

[1028] F. Kent Reilly III and Brian Stross. An overview of Mesoamerica. Visible Language, XXIV(1):6–11, Winter 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V24N1_1990_E.pdf.

Reilly:1990:CRF

[1029] F. Kent Reilly III. Cosmos and rulership: The function of Olmec-style symbols in formative period Mesoamerica. Visible Language, XXIV (1):12–37, Winter 1990. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V24N1_1990_E.pdf.

Stross:1990:MWC

[1030] Brian Stross. Mesoamerican writing at the crossroads: The Late Formative. Visible Language, XXIV(1):38–61, Winter 1990. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V24N1_1990_E.pdf. See comments [1082].

Fields:1990:DMH

[1031] Virginia M. Fields. Deciphering Maya hieroglyphic writing: The state of the art. Visible Language, XXIV(1):62–73, Winter 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V24N1_1990_E.pdf.

Troike:1990:PHP

[1032] Nancy P. Troike. Pre-Hispanic pictoral communication: The codex system of the Mixtec of Oaxaca, Mexico. Visible Language, XXIV (1):74–87, Winter 1990. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V24N1_1990_E.pdf.

Justeson:1990:ETM

[1033] John S. Justeson and Peter Mathews. Evolutionary trends in Mesoamerican hieroglyphic writing. Visible Language, XXIV(1):88–132, Winter 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V24N1_1990_E.pdf.

Anonymous:1990:WGR

[1034] Anonymous. The William S. Gray Research Collection in Reading. Visible Language, XXIV(1):133, Winter 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V24N1_1990_E.pdf.

Anonymous:1990:NPM

[1035] Anonymous. Notes on the preparation of manuscripts. Visible Language, XXIV(1):134–135, Winter 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V24N1_1990_E.pdf.

Anonymous:1990:USI

[1036] Anonymous. Upcoming special issues of Visible Language. Visible Language, XXIV(1):136, Winter 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V24N1_1990_E.pdf.

Poggenpohl:1990:AB

[1037] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, XXIV (2):134, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V24N2_1990_E.pdf.

Baker:1990:TYB

[1038] Jan Baker. [for a thousand years before the alphabet …]. Visible Language, XXIV(2):136, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V24N2_1990_E.pdf.

Sassoon:1990:HHM

[1039] Rosemary Sassoon. Handwriting — how much do we know about it? Visible Language, XXIV(2):137–143, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V24N2_1990_E.pdf.

Sassoon:1990:WEI

[1040] John Sassoon. Who on earth invented the alphabet? Visible Language, XXIV(2):144–163, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V24N2_1990_E.pdf.

Sedeyn:1990:QDH

[1041] Marie-Jeanne Sedeyn. Questioned documents: The human trace as a body flow. Visible Language, XXIV(2):164–175, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V24N2_1990_E.pdf.

Twyman:1990:PTN

[1042] Michael Twyman and Susan Walker. Preliminary thoughts on nomenclature for teachers of handwriting. Visible Language, XXIV(2):176– 193, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V24N2_1990_E.pdf.

Patkin:1990:AH

[1043] Michael Patkin. Anatomy of the hand. Visible Language, XXIV(2): 194–197, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V24N2_1990_E.pdf.

Sassoon:1990:WC

[1044] Rosemary Sassoon. Writer’s cramp. Visible Language, XXIV(2): 198–213, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V24N2_1990_E.pdf.

Ericsson:1990:GSD

[1045] Kjerstin Ericsson. Graphic skills as a diagnostic tool for working with the elderly. Visible Language, XXIV(2):214–226, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V24N2_1990_E.pdf.

Baker:1990:VN

[1046] Jan Baker. Visual notations. Visible Language, XXIV(2):227, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V24N2_1990_E.pdf.

Anonymous:1990:C

[1047] Anonymous. Colophon. Visible Language, XXIV(2):228, Spring 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V24N2_1990_E.pdf.

Kapr:1990:CBP

[1048] Albert Kapr. Concerning the beginning of printing in 15th century strassburg. Visible Language, XXIV(3–4):??, July 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V24N34_1990_E.pdf.

Zapf:1990:CAD

[1049] Hermann Zapf. Is creativity in alphabet design still wanted? Visible Language, XXIV(3–4):??, July 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V24N34_1990_E.pdf.

Bonsiepe:1990:IDG

[1050] Gui Bonsiepe. Interface design • graphics • language: Interpretations of human user interface. Visible Language, XXIV(3–4):??, July 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V24N34_1990_E.pdf.

Nemeth:1990:WDB

[1051] Christopher Nemeth. World design: Broadening the bandwidth communication in the 90’s. Visible Language, XXIV(3–4):??, July 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V24N34_1990_E.pdf.

Burnett:1990:CVS

[1052] Kathleen Burnett. Communication with visual sound: Herbert Bayer and the design of type. Visible Language, XXIV(3–4):??, July 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V24N34_1990_E.pdf.

Stiff:1990:BWU

[1053] Paul Stiff. From the bookshelves: What the user tells the designer. Visible Language, XXIV(3–4):??, July 1990. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V24N34_1990_E.pdf.

Poggenpohl:1991:ABa

[1054] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 25 (1):2, Winter 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N1_1991_E.pdf.

Benson:1991:LER

[1055] Philippa Jane Benson. Leading-edge research or lost cause: The search for interscriptual Stroop effects. Visible Language, 25(1):4–17, Winter 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V25N1_1991_E.pdf.

Brown:1991:LAP

[1056] R. A. Brown. Literary assessments in polyscriptal societies: Chinese character literacy in Korea and Japan. Visible Language, 25(1): 18–39, Winter 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N1_1991_E.pdf.

Bear:1991:CFO

[1057] Donald R. Bear. Copying fluency and orthographic development. Visible Language, 25(1):40–53, Winter 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V25N1_1991_E.pdf.

Cruttenden:1991:IC

[1058] Alan Cruttenden. Intonation and the comma. Visible Language, 25 (1):54–73, Winter 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N1_1991_E.pdf.

Bever:1991:SPT

[1059] Thomas G. Bever, Steven Jandreau, Rebecca Burwell, Ron Kaplan, and Annie Zaenen. Spacing printed text to isolate major phrases improves readability. Visible Language, 25(1):74–87, Winter 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V25N1_1991_E.pdf.

Polkinhorn:1991:BFI

[1060] Harry Polkinhorn. Bastard in the family: The impact of cubofuturist book art on structural linguistics. Visible Language, 25(1): 88–109, Winter 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N1_1991_E.pdf.

Anonymous:1991:ANa

[1061] Anonymous. Author notes. Visible Language, 25(1):110, Winter 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V25N1_1991_E.pdf.

Anonymous:1991:NPM

[1062] Anonymous. Notes on the preparation of manuscripts. Visible Language, 25(1):111–112, Winter 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N1_1991_E.pdf.

Hubert:1991:IAB

[1063] Renée Riese Hubert. Introduction to the artists’ books. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N23_1991_E.pdf.

Polkinhorn:1991:BAB

[1064] Harry Polkinhorn. From book to anti-book. Visible Language, 25 (2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N23_1991_E.pdf.

Lydon:1991:BTH

[1065] Mary Lydon. The book as the Trojan Horse of art: Walter Hamady, the Perishable Press Limited and Gabberjabbs 1–6. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N23_1991_E.pdf.

Perloff:1991:ITA

[1066] Marjorie Perloff. “Inner tension/in attention”: Steve McCaffery’s book art. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V25N23_1991_E.pdf.

Deguy:1991:DDD

[1067] Michel Deguy. Deguy/Dorny Dorny/Deguy. Visible Language, 25 (2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N23_1991_E.pdf.

Hoyem:1991:WTC

[1068] Andrew Hoyem. Working together: Collaboration in the book arts. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V25N23_1991_E.pdf.

Zelevansky:1991:CMM

[1069] Paul Zelevansky. The computer made me do it: Computers and books. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V25N23_1991_E.pdf.

Drucker:1991:TMP

[1070] Johanna Drucker. Typographic manipulation of the poetic text in the early twentieth-century avant-garde. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V25N23_1991_E.pdf.

Anzalone:1991:CTO

[1071] John Anzalone and Ruth Copans. Covering the text: the object of bookbinding. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO.

ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V25N23_1991_E.pdf.

Haskell:1991:RMB

[1072] Eric T. Haskell. Reading the multimedia book: the case of les fleurs du mal. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V25N23_1991_E.pdf.

Prinz:1991:NBN

[1073] Jessica Prinz. The “non-book”: New dimensions in the contemporary artist’s book. Visible Language, 25(2–3):??, April 1991. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V25N23_1991_E.pdf.

Sayre:1991:IAB

[1074] Henry M. Sayre. Ida applebroog and the book as a performance. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V25N23_1991_E.pdf.

Saillard:1991:BEM

[1075] Martine Saillard. A book exhibit at the musée pompidou. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V25N23_1991_E.pdf.

Shipe:1991:LAB

[1076] Timothy Shipe. The librarian and the artist’s book: Notes on the subversive art of cataloging. Visible Language, 25(2–3):??, April 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V25N23_1991_E.pdf.

Poggenpohl:1991:ABd

[1077] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 25 (4):336, Autumn 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N4_1991_E.pdf.

Hartley:1991:PWC

[1078] James Hartley. Psychology, writing and computers: a review of research. Visible Language, 25(4):338–375, Autumn 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V25N4_1991_E.pdf.

DeKay:1991:EDM

[1079] Michael L. DeKay and Jennifer J. Freyd. The effects of drawing method on the discriminability of characters. Visible Language, 25(4): 376–414, Autumn 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N4_1991_E.pdf.

Smudde:1991:BRL

[1080] Peter M. Smudde. Book review: Language and Typography. Cal Swann. New York: Van Nostrand, 1991. 94 pages, large format, amply illustrated, many photos, black and white. $19.95. Visible Language, 25(4): 415–420, Autumn 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N4_1991_E.pdf.

Anonymous:1991:BR

[1081] Anonymous. Books received. Visible Language, 25(4):420–423, Autumn 1991. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V25N4_1991_E.pdf.

Zapf:1991:RC

[1082] Hermann Zapf, Sharon Helmer Poggenpohl, Hans von Winning, Brian Stross, and Muriel Underwood. Reader comments. Visible Language, 25 (4):424–427, Autumn 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N4_1991_E.pdf. See [1030].

Anonymous:1991:IV

[1083] Anonymous. Index to volume 25. Visible Language, 25(4):428– 430, Autumn 1991. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V25N4_1991_E.pdf.

Milman:1992:HPT

[1084] Estera Milman. Historical precedents, trans-historical strategies, and the myth of democratization. Visible Language, 26(1–2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V26N12_1992_E.pdf.

Foster:1992:HDS

[1085] Stephen C. Foster. Historical design and social purpose: a note on the relationship of Fluxus to modernism. Visible Language, 26(1– 2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V26N12_1992_E.pdf.

Smith:1992:PFU

[1086] Owen Smith. Proto-Fluxus in the United States 1959–1961: The establishment of a like-minded community of artists. Visible Language, 26 (1–2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V26N12_1992_E.pdf.

Snyder:1992:JCD

[1087] Ellsworth Snyder. John Cage discusses Fluxus. Visible Language, 26 (1–2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V26N12_1992_E.pdf.

Allen:1992:FL

[1088] Roy F. Allen. Fluxus and literature. Visible Language, 26(1–2): ??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V26N12_1992_E.pdf.

Saper:1992:FII

[1089] Craig Saper. Fluxacademy: From intermedia to interactive education. Visible Language, 26(1–2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V26N12_1992_E.pdf.

Milman:1992:RSS

[1090] Estera Milman. Road shows, street events, and Fluxus people; a conversation with Alison Knowles. Visible Language, 26(1–2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V26N12_1992_E.pdf.

Anderson:1992:OSC

[1091] Eric Anderson, Stephen C. Foster, and Estera Milman. On open structures and the crisis of meaning, a dialogue. Visible Language, 26(1–2): 133–142, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V26N12_1992_E.pdf.

Higgins:1992:TSC

[1092] Dick Higgins. Two sides of a coin: Fluxus and the Something Else Press. Visible Language, 26(1–2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V26N12_1992_E.pdf.

Friedman:1992:FGC

[1093] Ken Friedman and James Lewes. Fluxus: Global community, human dimensions. Visible Language, 26(1–2):??, January 1992. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V26N12_1992_E.pdf.

Melton:1992:NSR

[1094] Hollis Melton. Notes on SoHo and a reminiscence. Visible Language, 26(1–2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V26N12_1992_E.pdf.

Milman:1992:CFC

[1095] Estera Milman. Circle of friends: a conversation with Alice Hutchins. Visible Language, 26(1–2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V26N12_1992_E.pdf.

Frank:1992:FFN

[1096] Peter Frank. Fluxus fallout: New York in the wake of the new sensibility. Visible Language, 26(1–2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V26N12_1992_E.pdf.

Partridge:1992:FIA

[1097] Michael Partridge and Joan Huntley. FluxBase: an interactive art exhibition. Visible Language, 26(1–2):??, January 1992. CODEN VSLGAO.

ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V26N12_1992_E.pdf.

Hanhardt:1992:CTW

[1098] John G. Hanhardt and Peter Moore. Dé-collage and television: Wolf Vostell in New York, 1963–64. Visible Language, 26(1–2):??, January 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V26N12_1992_E.pdf.

Poggenpohl:1992:FRD

[1099] Sharon Helmer Poggenpohl and Dietmar R. Winkler. The frame of reference: Diagrams as tools for worldmaking. Visible Language, 26 (3–4):??, July 1992. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V26N34_1992_E.pdf.

Lenk:1992:SEI

[1100] Krzysztof Lenk and Paul Kahn. To show and explain: The information graphics of Stevin and Comenius. Visible Language, 26(3–4):??, July 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V26N34_1992_E.pdf.

McArthur:1992:SFP

[1101] Douglas McArthur. Sign function and potential of the printed word. Visible Language, 26(3–4):??, July 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V26N34_1992_E.pdf.

Chizlett:1992:DLS

[1102] Clive Chizlett. Damned lies. And statistics. Otto Neurath and Soviet propaganda in the 1930s. Visible Language, 26(3–4):??, July 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V26N34_1992_E.pdf. See negative commentary [1139, 1143].

Sims-Knight:1992:PPH

[1103] Judith E. Sims-Knight. To picture or not to picture: How to decide. Visible Language, 26(3–4):??, July 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V26N34_1992_E.pdf.

Storkerson:1992:EIG

[1104] Peter Storkerson. Explicit and implicit graphs: Changing the frame. Visible Language, 26(3–4):??, July 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V26N34_1992_E.pdf.

Cohen:1992:BZT

[1105] Michael Cohen. Blush and zebrackets: Two schemes for typographical representation of nested associativity. Visible Language, 26(3–4):??, July 1992. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V26N34_1992_E.pdf.

Hodgson:1993:SDP

[1106] Richard Hodgson and Ralph Sarkonak. Seeing in depth: the practice of bilingual writing. Visible Language, 27(1–2):??, January 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V27N12_1993_E.pdf.

Mackey:1993:LDB

[1107] William Mackey. Literary diglossia, biculturalism and cosmopolitanism in literature. Visible Language, 27(1–2):??, January 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V27N12_1993_E.pdf.

Cooper:1993:BBC

[1108] Jerrold Cooper. Bilingual Babel: Cuneiform texts in two or more languages from Ancient Mesopotamia and beyond. Visible Language, 27 (1–2):??, January 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N12_1993_E.pdf.

Picard:1993:JAC

[1109] Daniel Picard. Jackhammers and alarm clocks: Perceptions in stereo. Visible Language, 27(1–2):??, January 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V27N12_1993_E.pdf.

Hewson:1993:BET

[1110] Lance Hewson. The bilingual edition in translation studies. Visible Language, 27(1–2):??, January 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V27N12_1993_E.pdf.

Lubell:1993:BHT

[1111] Stephen Lubell. Bilingualism in the Hebrew text. Visible Language, 27 (1–2):??, January 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N12_1993_E.pdf.

Grutman:1993:MVS

[1112] Rainier Grutman. Mono versus stereo: Bilingualism’s double face. Visible Language, 27(1–2):??, January 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V27N12_1993_E.pdf.

Wrenn:1993:CAP

[1113] Phyllis Wrenn. A case for acadian — the politics of style. Visible Language, 27(1–2):??, January 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V27N12_1993_E.pdf.

Nassar:1993:TEM

[1114] Joseph Nassar. Transformations in exile: The multilingual exploits of Nabokov’s Pnin and Kinbote. Visible Language, 27(1–2):??, January 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V27N12_1993_E.pdf.

Poggenpohl:1993:ABc

[1115] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 27 (3):274, July 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N3_1993_E.pdf.

Poggenpohl:1993:E

[1116] Sharon Helmer Poggenpohl. Editorial. Visible Language, 27(3): 275–277, July 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N3_1993_E.pdf.

Salen:1993:STR

[1117] Katie Salen. Speaking in text: The resonance of syntactic difference in text interpretation. Visible Language, 27(3):278–301, July 1993.

CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V27N3_1993_E.pdf.

Hailstone:1993:TTT

[1118] Max Hailstone. “Te Tiriti” (the treaty). Visible Language, 27(3): 302–321, July 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N3_1993_E.pdf.

Davis:1993:SPF

[1119] Christopher Davis. Saving pictures from the flood: Using visual art in creative writing workshops. Visible Language, 27(3):322–335, July 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V27N3_1993_E.pdf.

Gilreath:1993:GCT

[1120] Charles T. Gilreath. Graphic cueing of text: The typographic and diagraphic dimensions. Visible Language, 27(3):336–361, July 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V27N3_1993_E.pdf.

Black:1993:PRM

[1121] Alison Black and Darren Watts. Proof-reading monospaced and proportionally spaced typefaces: Should we check typewritten or typeset. Visible Language, 27(3):362–377, July 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V27N3_1993_E.pdf.

Nemeth:1993:BRC

[1122] Christopher Nemeth. Book review: Computers as Theater, Brenda Laurel. New York: Addison-Wesley, Publishers, 1992 211 pages, cloth, 8 full color pages, $19.50 ISBN 0-201-51048-0. Visible Language, 27 (3):378–380, July 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N3_1993_E.pdf.

Poggenpohl:1993:BRM

[1123] Sharon Helmer Poggenpohl. Book review: Modern Typography, an essay in critical history. Robin Kinross, London: Hyphen Press, 1992 206 pages, paper, illustrations, $30.00 ISBN 0-907259-05-7. Visible Language, 27(3):381–383, July 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N3_1993_E.pdf.

Poggenpohl:1993:ESC

[1124] Sharon Helmer Poggenpohl. Editorial, subscriptions, claims, reproduction. Visible Language, 27(3):384, July 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V27N3_1993_E.pdf.

Poggenpohl:1993:AB

[1125] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 27 (4):386, Autumn 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N4_1993_E.pdf.

Polkinhorn:1993:VPI

[1126] Harry Polkinhorn. Visual poetry: an introduction. Visible Language, 27 (4):389–393, Autumn 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N4_1993_E.pdf.

Menezes:1993:BVP

[1127] Philadelpho Menezes. Brazilian visual poetry. Visible Language, 27(4): 394–409, Autumn 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N4_1993_E.pdf.

Gutierrez:1993:CVP

[1128] Pedro Juan Gutierrez. Cuban visual poetry. Visible Language, 27(4): 410–421, Autumn 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N4_1993_E.pdf.

Minarelli:1993:IVP

[1129] Enzo Minarelli. Italian visual poetry. Visible Language, 27(4):422– 435, Autumn 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N4_1993_E.pdf.

Espinosa:1993:MVP

[1130] César Espinosa. Mexican visual poetry. Visible Language, 27(4): 436–443, Autumn 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N4_1993_E.pdf.

Aguiar:1993:PVP

[1131] Fernando Aguiar. Portuguese visual poetry. Visible Language, 27(4): 444–466, Autumn 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N4_1993_E.pdf.

Padin:1993:UVP

[1132] Clemette Padin. Uruguayan visual poetry. Visible Language, 27(4): 468–480, Autumn 1993. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V27N4_1993_E.pdf.

Polkinhorn:1993:VPU

[1133] Harry Polkinhorn. Visual poetry from the United States. Visible Language, 27(4):482–493, Autumn 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V27N4_1993_E.pdf.

Anonymous:1993:VI

[1134] Anonymous. Volume 27 index. Visible Language, 27(4):494–396, Autumn 1993. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V27N4_1993_E.pdf.

Poggenpohl:1994:ABa

[1135] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 28 (1):2, Winter 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N1_1994_E.pdf.

Haussamen:1994:FES

[1136] Brock Haussamen. The future of the English sentence. Visible Language, 28(1):4–25, Winter 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N1_1994_E.pdf.

Yule:1994:PFR

[1137] Valerie Yule. Problems that face research in the design of spelling. Visible Language, 28(1):26–46, Winter 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V28N1_1994_E.pdf.

Roth:1994:UBH

[1138] Susan King Roth. The unconsidered ballot: How design effects voting behavior. Visible Language, 28(1):48–67, Winter 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V28N1_1994_E.pdf.

Kinross:1994:BEI

[1139] Robin Kinross. Blind eyes, innuendo and the politics of design: a reply to Clive Chizlett. Visible Language, 28(1):68–79, Winter 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V28N1_1994_E.pdf. See [1102, 1143].

Remington:1994:BRS

[1140] R. Roger Remington. Book review: SPIRALS 91. Thomas Ockerse, editor. Providence, Rhode Island: Rhode Island School of Design. Large format, five softbound books in a slipcase, printed in two colors, many illustrations. $125.00 . Visible Language, 28(1):80–82, Winter 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V28N1_1994_E.pdf.

Poggenpohl:1994:BRB

[1141] Sharon Helmer Poggenpohl. Book review: Black Riders, The Visible Language of Modernism. Jerome McGann. Princeton, New Jersey: Princeton University Press, 1993. $35.00 cloth, $12.95 paper. ISBN 0-691-01544-9. Visible Language, 28(1):83–85, Winter 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V28N1_1994_E.pdf.

Banks:1994:BRS

[1142] Colin Banks. Book review: Stop Stealing Sheep and Find Out How Type Works. Erik Spiekermann and E. M. Ginger. Mountain View, California: Adobe Press. $19.50 and $24.95 (Canadian) ISBN 0-672-48543-5. Visible Language, 28(1):86–88, Winter 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V28N1_1994_E.pdf.

Twyman:1994:LRC

[1143] Michael Twyman and Daniel Picard. Letters: Regarding Clive Chizlett’s article Damned Lies. And Statistics. Visible Language, 28(1): 89–91, Winter 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N1_1994_E.pdf. See [1102, 1139].

Anonymous:1994:BR

[1144] Anonymous. Books received. Visible Language, 28(1):92–96, Winter 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V28N1_1994_E.pdf.

Poggenpohl:1994:ABb

[1145] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 28 (2):98, Spring 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N2_1994_E.pdf.

Close:1994:RO

[1146] Eleanor O. Close. Recollect orality. Visible Language, 28(2):100– 109, Spring 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N2_1994_E.pdf.

Meyer:1994:CYS

[1147] Charles F. Meyer, Robert A. Morris, and Ed Blanchman. Can you see whose speech is overlapping? Visible Language, 28(2):110–133, Spring 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V28N2_1994_E.pdf.

Grow:1994:WPV

[1148] Gerald Grow. The writing problems of visual thinkers. Visible Language, 28(2):134–161, Spring 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N2_1994_E.pdf.

Keppler:1994:NPL

[1149] Joseph F. Keppler. The news as a post-literary spectacle. Visible Language, 28(2):162–171, Spring 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V28N2_1994_E.pdf.

Poggenpohl:1994:MTB

[1150] Sharon Helmer Poggenpohl. More than a book review of The Electronic Word. Visible Language, 28(2):172–193, Spring 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V28N2_1994_E.pdf.

Poggenpohl:1994:ABc

[1151] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 28 (3):194, Summer 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N3_1994_E.pdf.

Blauvelt:1994:FDB

[1152] Andrew Blauvelt. Foreword: Disciplinary bodies: the resistance to theory and the cut of the critic. Visible Language, 28(3):196–202, Summer 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V28N3_1994_E.pdf.

Blauvelt:1994:OGD

[1153] Andrew Blauvelt. An opening: Graphic design’s discursive spaces. Visible Language, 28(3):204–216, Summer 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V28N3_1994_E.pdf.

Bush:1994:TLG

[1154] Anne Bush. Through the looking glass: Territories of the historiographic gaze. Visible Language, 28(3):218–231, Summer 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V28N3_1994_E.pdf.

Margolin:1994:NPG

[1155] Victor Margolin. Narrative problems of graphic design history. Visible Language, 28(3):232–243, Summer 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V28N3_1994_E.pdf.

Baker:1994:PGD

[1156] Steve Baker. A poetics of graphic design? Visible Language, 28(3): 244–259, Summer 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N3_1994_E.pdf.

Mermoz:1994:MHS

[1157] Gérard Mermoz. Masks on hire: In search of typographic histories. Visible Language, 28(3):260–285, Summer 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V28N3_1994_E.pdf.

Anonymous:1994:JIa

[1158] Anonymous. Journal information. Visible Language, 28(3):286– 287, Summer 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N3_1994_E.pdf.

Anonymous:1994:PID

[1159] Anonymous. PhD Institute of Design. Visible Language, 28(3): 288, Summer 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N3_1994_E.pdf.

Poggenpohl:1994:ABd

[1160] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 28 (4):286, Autumn 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N4_1994_E.pdf.

Blauvelt:1994:FPP

[1161] Andrew Blauvelt. Foreword: The personal is political: The social practices of graphic design. Visible Language, 28(4):289–295, Autumn 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V28N4_1994_E.pdf.

Smith:1994:CLP

[1162] Marilyn Crafton Smith. Culture is the limit: Pushing the boundaries of graphic design criticism and practice. Visible Language, 28(4): 297–315, Autumn 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N4_1994_E.pdf.

vanToorn:1994:DR

[1163] Jan van Toorn. Design and reflexivity. Visible Language, 28(4):317– 325, Autumn 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N4_1994_E.pdf.

McKee:1994:SH

[1164] Stuart McKee. Simulated histories. Visible Language, 28(4):327– 343, Autumn 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N4_1994_E.pdf.

Lupton:1994:DGD

[1165] Ellen Lupton and J. Abbott Miller. Deconstruction and graphic design: History meets theory. Visible Language, 28(4):345–365, Autumn 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V28N4_1994_E.pdf.

Scotford:1994:MHV

[1166] Martha Scotford. Messy history vs. neat history: Toward an expanded view of women in graphic design. Visible Language, 28(4): 367–387, Autumn 1994. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N4_1994_E.pdf.

Anonymous:1994:VI

[1167] Anonymous. Volume 28 index. Visible Language, 28(4):392–393, Autumn 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V28N4_1994_E.pdf.

Anonymous:1994:JIb

[1168] Anonymous. Journal information. Visible Language, 28(4):394, Autumn 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V28N4_1994_E.pdf.

Poggenpohl:1995:ABa

[1169] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 29 (1):2, Winter 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N1_1995_E.pdf.

Blauvelt:1995:FST

[1170] Andrew Blauvelt. Foreword: Surface tensions: Between explanation and understanding. Visible Language, 29(1):4–10, Winter 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V29N1_1995_E.pdf.

Sellers:1995:HLB

[1171] Susan Sellers. How long has this been going on? Harpers Bazaar, Funny Face and the construction of the modernist woman. Visible Language, 29(1):12–34, Winter 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N1_1995_E.pdf.

Williamson:1995:EHI

[1172] Jack Williamson. Embodiments of human identity: Detecting and interpreting hidden narratives in twentieth-century design history. Visible Language, 29(1):36–70, Winter 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V29N1_1995_E.pdf.

Triggs:1995:ASR

[1173] Teal Triggs. Alphabet soup: Reading British fanzines. Visible Language, 29(1):72–87, Winter 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N1_1995_E.pdf.

Butler:1995:NDT

[1174] Frances Butler. New demotic typography: The search for new indices. Visible Language, 29(1):88–111, Winter 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V29N1_1995_E.pdf.

Anonymous:1995:JIa

[1175] Anonymous. Journal information. Visible Language, 29(1):112, Winter 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V29N1_1995_E.pdf.

Poggenpohl:1995:ABb

[1176] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 29 (2):114, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N2_1995_E.pdf.

Breder:1995:LOV

[1177] Hans Breder and Herman Rapaport. The luminous object: Video art and video theory: an introduction. Visible Language, 29(2):116–121, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V29N2_1995_E.pdf.

Zurbrugg:1995:NJP

[1178] Nicholas Zurbrugg. Nam June Paik: an interview. Visible Language, 29 (2):122–137, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N2_1995_E.pdf.

Hanhardt:1995:FIE

[1179] John G. Hanhardt. Film image — electronic image: The construction of abstraction, 1960–1990. Visible Language, 29(2):138–159, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V29N2_1995_E.pdf.

Rapaport:1995:TLD

[1180] Herman Rapaport. Time and light: David Garcia, an interview. Visible Language, 29(2):160–179, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V29N2_1995_E.pdf.

Sherk:1995:LLN

[1181] Bonnie Sherk. A living library: New model for global electronic interactivity and networking in the garden. Visible Language, 29(2): 180–185, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N2_1995_E.pdf.

Rapaport:1995:TUD

[1182] Herman Rapaport. Television and the unconscious, Donald Kuspit: an interview. Visible Language, 29(2):186–197, Spring 1995.

CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V29N2_1995_E.pdf.

Kostelanetz:1995:LV

[1183] Richard Kostelanetz. Literary video. Visible Language, 29(2):198– 203, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N2_1995_E.pdf.

Rapaport:1995:CAK

[1184] Herman Rapaport. Carole Anne Klonarides: an interview. Visible Language, 29(2):204–213, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V29N2_1995_E.pdf.

Zurbrugg:1995:JCV

[1185] Nicholas Zurbrugg. Jameson’s complaint: Video-art and the intertextual “time-wall”. Visible Language, 29(2):214–237, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V29N2_1995_E.pdf.

Rapaport:1995:LE

[1186] Herman Rapaport. The liminal eye. Visible Language, 29(2):238– 253, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N2_1995_E.pdf.

Anonymous:1995:JIb

[1187] Anonymous. Journal information. Visible Language, 29(2):254– 255, Spring 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N2_1995_E.pdf.

Poggenpohl:1995:BVD

[1188] Sharon Helmer Poggenpohl. Between visual and digital tokens: a look at the abstraction of money. Visible Language, 29(3–4):??, September 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V29N34_1995_E.pdf. See discussion of Dwiggins’ fabrications [1188].

Barrett:1995:FM

[1189] Dawn Barrett. Flying money. Visible Language, 29(3–4):??, September 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V29N34_1995_E.pdf.

Barrett:1995:MEM

[1190] Dawn Barrett. “Modest enquiry” and major innovation: Franklin’s early American currency. Visible Language, 29(3–4):??, September 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V29N34_1995_E.pdf.

Smith:1995:JGB

[1191] Sandra Smith. J. S. G. Boggs: Life size and in color. Visible Language, 29 (3–4):??, September 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N34_1995_E.pdf.

Gonzalez:1995:IJG

[1192] Manuel Gonzalez. Interview with J. S. G. Boggs. Visible Language, 29 (3–4):??, September 1995. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V29N34_1995_E.pdf.

Shell:1995:PGA

[1193] Marc Shell. Paper, gold and art as representation and exchange. Visible Language, 29(3–4):??, September 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V29N34_1995_E.pdf.

Nemeth:1995:FMC

[1194] Christopher Nemeth. Funny money, coupons, scrip, chips and other quasi-official media of exchange. Visible Language, 29(3–4):??, September 1995. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V29N34_1995_E.pdf.

Poggenpohl:1996:ABa

[1195] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 30 (1):4, January 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N1_1996_E.pdf.

DeFrancis:1996:HEC

[1196] John DeFrancis. How efficient is the Chinese writing system? Visible Language, 30(1):6–44, January 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V30N1_1996_E.pdf.

Smith:1996:VWJ

[1197] Janet Shibamoto Smith and David L. Schmidt. Variability in written Japanese: Towards a sociolinguistics of script choice. Visible Language, 30(1):46–71, January 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N1_1996_E.pdf.

Elkins:1996:BPP

[1198] James Elkins. Between picture and proposition: Torturing paintings in Wittgenstein’s Tractatus. Visible Language, 30(1):72–95, January 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V30N1_1996_E.pdf.

Anonymous:1996:ECB

[1199] Anonymous. Editorial correspondence. Business correspondence. Visible Language, 30(1):96, January 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V30N1_1996_E.pdf.

Kac:1996:I

[1200] Eduardo Kac. Introduction. Visible Language, 30(2):98–101, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V30N2_1996_E.pdf.

Rosenberg:1996:IDS

[1201] Jim Rosenberg. The interactive diagram sentence: Hypertext as a medium of thought. Visible Language, 30(2):102–116, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V30N2_1996_E.pdf.

Bootz:1996:PM

[1202] Philippe Bootz. Poetic machinations. Visible Language, 30(2): 118–137, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N2_1996_E.pdf.

deMeloeCastro:1996:V

[1203] E. M. de Melo e Castro. Videopoetry. Visible Language, 30(2): 138–149, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N2_1996_E.pdf.

Vallias:1996:WUD

[1204] Andras Vallias. We have not understood Descartes. Visible Language, 30(2):150–157, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N2_1996_E.pdf.

Gyori:1996:VP

[1205] Ladislao Pablo Györi. Virtual poetry. Visible Language, 30(2): 158–163, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N2_1996_E.pdf.

Cayley:1996:BCP

[1206] John Cayley. Beyond codexspace: Potentialities of literary cybertext. Visible Language, 30(2):164–183, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V30N2_1996_E.pdf.

Kac:1996:H

[1207] Eduardo Kac. Holopoetry. Visible Language, 30(2):184–212, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V30N2_1996_E.pdf.

Vos:1996:NMP

[1208] Eric Vos. New media poetry — theory and strategies. Visible Language, 30(2):214–233, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N2_1996_E.pdf.

Kac:1996:SW

[1209] Eduardo Kac. Selected webliography. Visible Language, 30(2): 234–237, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N2_1996_E.pdf.

Poggenpohl:1996:ABb

[1210] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 30 (2):238, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N2_1996_E.pdf.

Anonymous:1996:JIb

[1211] Anonymous. Journal information. Visible Language, 30(2):239, May 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V30N2_1996_E.pdf.

Kostelanetz:1996:PFS

[1212] Richard Kostelanetz. A poetry-film storyboard: Transformations. Visible Language, 30(3):244–245, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V30N3_1996_E.pdf.

Sless:1996:BIP

[1213] David Sless. Better information presentation: Satisfying customers? Visible Language, 30(3):246–267, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V30N3_1996_E.pdf.

Walker:1996:DBB

[1214] Sue Walker, Viv Edwards, and Ruth Blacksell. Designing bilingual books for children. Visible Language, 30(3):268–283, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V30N3_1996_E.pdf.

Nissani:1996:AAW

[1215] Moti Nissani. The apprenticeship approach to writing instruction. Visible Language, 30(3):284–313, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V30N3_1996_E.pdf.

Storkerson:1996:JTL

[1216] Peter Storkerson. Jan Tschichold and the language of modernism. Visible Language, 30(3):314–339, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V30N3_1996_E.pdf.

Banks:1996:BRK

[1217] Colin Banks. Book reviews: Ken Garland, Mr. Beck’s Underground Map: a History. Middlesex: Capital Transport Publishing: 1995, illustrated, limited edition, £10.95. Visible Language, 30(3):340–345, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V30N3_1996_E.pdf.

Blatner:1996:BRJ

[1218] Adam Blatner. Book review: Johanna Drucker, The Alphabetic Labyrinth: The Letters in History and Imagination. New York: Thames & Hudson: 1995. 322 pages, illustrated, $45.00. Visible Language, 30(3): 346–352, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N3_1996_E.pdf.

Anonymous:1996:VPE

[1219] Anonymous. Virtuoso penmanship: examples from the early 1700s. Visible Language, 30(3):353, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V30N3_1996_E.pdf.

Anonymous:1996:BR

[1220] Anonymous. Books received. Visible Language, 30(3):354–361, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N3_1996_E.pdf.

Anonymous:1996:NPM

[1221] Anonymous. Notes on the preparation of a manuscript. Visible Language, 30(3):362–363, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V30N3_1996_E.pdf.

Anonymous:1996:IV

[1222] Anonymous. Index to volume 30. Visible Language, 30(3):364– 365, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N3_1996_E.pdf.

Anonymous:1996:JIc

[1223] Anonymous. Journal information. Visible Language, 30(3):366– 367, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V30N3_1996_E.pdf.

Gibbons:1996:LE

[1224] Charles Gibbons, Paul Shaw, and Sharon Helmer Poggenpohl. Letters to the Editor. Visible Language, 30(3):368, September 1996. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V30N3_1996_E.pdf. See [1188].

Poggenpohl:1997:ABa

[1225] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 31 (1):2, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V31N1_1997_E.pdf.

Miller:1997:WAG

[1226] Errol Miller. What avant-garde may be. Visible Language, 31(1): 4–5, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V31N1_1997_E.pdf.

Baron:1997:TLWa

[1227] Naomi S. Baron. Thinking, learning and the written word. Visible Language, 31(1):6–37, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V31N1_1997_E.pdf.

Kanawati:1997:HCL

[1228] Dianne G. Kanawati. How can I be literate: Counting the ways. Visible Language, 31(1):38–51, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V31N1_1997_E.pdf.

Haussamen:1997:PPD

[1229] Brock Haussamen. Puns, public discourse and postmodernism. Visible Language, 31(1):52–61, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V31N1_1997_E.pdf.

Miller:1997:HPH

[1230] Errol Miller. How poetry happens. Visible Language, 31(1):62–63, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V31N1_1997_E.pdf.

Hayes:1997:BAU

[1231] Kevin J. Hayes. The book in American Utopia literature, 1883–1917. Visible Language, 31(1):64–84, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V31N1_1997_E.pdf.

Miller:1997:AE

[1232] Errol Miller. Attempting to explain. Visible Language, 31(1):86– 87, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V31N1_1997_E.pdf.

McGuinne:1997:SD

[1233] Dermot McGuinne. Simply a dot. Visible Language, 31(1):88– 107, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V31N1_1997_E.pdf.

Miller:1997:P

[1234] Errol Miller. Poetry. Visible Language, 31(1):108–109, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V31N1_1997_E.pdf.

Skelly:1997:DN

[1235] Barbara Louise Skelly. Designer’s note. Visible Language, 31(1): 110, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V31N1_1997_E.pdf.

Burnhill:1997:E

[1236] Peter Burnhill. To the Editor. Visible Language, 31(1):111, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V31N1_1997_E.pdf.

Anonymous:1997:ECBa

[1237] Anonymous. Editorial correspondence. Business correspondence. Visible Language, 31(1):112, Winter 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V31N1_1997_E.pdf.

Poggenpohl:1997:ABb

[1238] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 31 (2):114, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V31N2_1997_E.pdf.

Winkler:1997:LEG

[1239] Dietmar Winkler. Loss of an empire/gaining another? Visible Language, 31(2):116–125, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V31N2_1997_E.pdf.

Storkerson:1997:HAM

[1240] Peter Storkerson and Janine Wong. Hypertext & the art of memory. Visible Language, 31(2):126–157, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V31N2_1997_E.pdf.

Stone:1997:ESC

[1241] Greg Stone. Exploring the special communications experiences of online education. Visible Language, 31(2):158–181, May 1997. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V31N2_1997_E.pdf.

Rasheed:1997:CMI

[1242] Thomas Rasheed and Leif Allmendinger. A conceptual model of interactive exhibits for African–American children. Visible Language, 31 (2):182–199, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V31N2_1997_E.pdf.

Twarog:1997:IIA

[1243] Carl Twarog. Inclusive interaction: Ability enhancing multimedia design. Visible Language, 31(2):200–213, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V31N2_1997_E.pdf.

Gromala:1997:VAS

[1244] Diana J. Gromala. Virtual avatars: Subjectivity in virtual environments. Visible Language, 31(2):214–229, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V31N2_1997_E.pdf.

Poggenpohl:1997:BRM

[1245] Sharon Helmer Poggenpohl. Book review: M. Christine Boyer 1996. Cybercities, Visual Perception in the Age of Electronic Communication. New York: Princeton Architectural Press, 1996. 245 pages, cloth, $19.95. ISBN 1-56898-048-5. Visible Language, 31(2):231–233, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V31N2_1997_E.pdf.

Poggenpohl:1997:BRA

[1246] Sharon Helmer Poggenpohl. Book review: Alberto Manguel, A History of Reading, New York: Viking, 1996. 372 pages, cloth, illustrated, $26.95. ISBN 0-670-84302-4. Visible Language, 31(2):234–236, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V31N2_1997_E.pdf.

Anonymous:1997:BR

[1247] Anonymous. Books received. Visible Language, 31(2):237–241, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V31N2_1997_E.pdf.

Anonymous:1997:NPM

[1248] Anonymous. Notes on the preparation of a manuscript. Visible Language, 31(2):249–251, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V31N2_1997_E.pdf.

Anonymous:1997:JIb

[1249] Anonymous. Journal information. Visible Language, 31(2):252–253, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V31N2_1997_E.pdf.

Anonymous:1997:WAW

[1250] Anonymous. Website announcement [www.id.iit.edu/visiblelanguage]. Visible Language, 31(2):255, May 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V31N2_1997_E.pdf.

Poggenpohl:1997:AB

[1251] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 31 (3):258, November 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3132_1997_E.pdf.

Salen:1997:DRS

[1252] Katie Salen and Sharyn O’Mara. Dis[appearances]: Representational strategies and operational needs in codexspace and screenspace. Visible Language, 31(3):260–285, November 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V31N3_1997_E.pdf.

Wallraff:1997:EGT

[1253] Martin Wallraff. Early Greek typography in Milan: a historical note on a new Greek typeface. Visible Language, 31(3):286–299, November 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V31N3_1997_E.pdf.

Gibson:1997:TCA

[1254] Michael Gibson. Teaching critical analytical methods in the digital typography classroom. Visible Language, 31(3):300–325, November 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V31N3_1997_E.pdf.

Wendt:1997:EHA

[1255] Dirk Wendt, Wiebke Groggel, and Georg Gutschmidt. On the effectiveness of highlighting ads in telephone directories by color. Visible Language, 31(3):326–337, November 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V31N3_1997_E.pdf.

Kostelanetz:1997:BRS

[1256] Richard Kostelanetz. Book review: Sloppy “scholarship”: The Century of Artists’ Books by Johanna Drucker. New York: Granary Books, 1995. ISBN 1-887123-01-6. 377 pages, hardbound, illustrated, $35. Visible Language, 31(3):338–342, November 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V3132_1997_E.pdf.

Banks:1997:BRE

[1257] Colin Banks. Book review: Encyclopedia of the Book, Geoffrey Ashall Glaister. London: The British Library, cloth. £65, paperback £35. United States: Oak Knoll Press, $75. ISBN 1-884718-15-9. 550 pages, 260 × 180mms. Visible Language, 31(3):343–344, November 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V3132_1997_E.pdf.

Poggenpohl:1997:BRP

[1258] Sharon Helmer Poggenpohl. Book review: Practice with philosophy: Design Writing Research, Ellen Lupton and J. Abbott Miller. New York: Kiosk, 1996. ISBN 1-56898-047-7. 211 pages, hardbound, illustrated, some in color, $45. Visible Language, 31(3):345–348, November 1997. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V3132_1997_E.pdf.

Anonymous:1997:IV

[1259] Anonymous. Index to volume 31. Visible Language, 31(3):349– 350, November 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3132_1997_E.pdf.

Anonymous:1997:JIc

[1260] Anonymous. Journal information. Visible Language, 31(3):351– 352, November 1997. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V3132_1997_E.pdf.

Poggenpohl:1998:ABa

[1261] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 32 (1):2, January 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N1_1998_E.pdf.

Poggenpohl:1998:TSE

[1262] Sharon Helmer Poggenpohl. Twenty-six not-so-easy pieces. Visible Language, 32(1):5–32, January 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V32N1_1998_E.pdf.

Miller:1998:RH

[1263] Errol Miller. Ripples on the horizon. Visible Language, 32(1):33, January 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V32N1_1998_E.pdf.

Baron:1998:WAE

[1264] Naomi S. Baron. Writing in the age of email: The impact of ideology versus technology. Visible Language, 32(1):35–53, January 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V32N1_1998_E.pdf.

Miller:1998:SD

[1265] Errol Miller. The seventh day. Visible Language, 32(1):54–55, January 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V32N1_1998_E.pdf.

Balhorn:1998:PRN

[1266] Mark Balhorn. Paper representations of the non-standard voice. Visible Language, 32(1):56–74, January 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V32N1_1998_E.pdf.

Miller:1998:SR

[1267] Errol Miller. Self-realization. Visible Language, 32(1):76–77, January 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V32N1_1998_E.pdf.

Golec:1998:CBR

[1268] Michael Golec. Cloth-bound reverie. Visible Language, 32(1):78– 92, January 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N1_1998_E.pdf.

Miller:1998:CP

[1269] Errol Miller. The crowbar principle. Visible Language, 32(1):93– 94, January 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N1_1998_E.pdf.

Anonymous:1998:JIa

[1270] Anonymous. Journal information. Visible Language, 32(1):95– 96, January 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N1_1998_E.pdf.

Poggenpohl:1998:ABb

[1271] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 32 (2):98, May 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N2_1998_E.pdf.

Karow:1998:ECD

[1272] Peter Karow. Extending control of digital typography. Visible Language, 32(2):100–127, May 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N2_1998_E.pdf.

Lucke:1998:CDB

[1273] Karsten Lücke. Customized digital books on demand: Issues in the creation of a flexible document format. Visible Language, 32(2): 128–149, May 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N2_1998_E.pdf.

Dyson:1998:ELL

[1274] Mary C. Dyson and Gary J. Kipping. The effects of line length and method of movement on patterns of reading from screen. Visible Language, 32(2):150–181, May 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V32N2_1998_E.pdf.

Winkler:1998:BRW

[1275] Dietmar Winkler. Book review: Writing and Its Use. An International Handbook of International Research Volumes 1 and 2. Hartmut Gunter and Otto Ludwig, editors. Compiled in collaboration with J. Baurmann, F. Coulmas, K. Ehlich, P. Eisenberg, H. W. Giese, H. Gluck, K. B. Gunther, U. Knoop, B. Pompino-Marschall, E. Scheerer and R. Weingarten. Berlin: Walter de Gruyter, 1996. ISBN 3-11-011129-2. Volume 1: 902 pages, hardbound, illustrated, some in color, $532.00. ISBN 3-11-014744-0. Volume 2: 863 pages, hardbound, illustrated, approximately 50 figures in black and white, $482.00. Visible Language, 32(2):184–188, May 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V32N2_1998_E.pdf.

Poggenpohl:1998:BRU

[1276] Sharon Helmer Poggenpohl. Book review: User-Centered Graphic Design: Mass Communication and Social Change. Jorge Frascara. London: Taylor and Francis, 1997. ISBN 0-7484-0672-7. 147 pages, softbound, illustrated, one color, $44.95. Visible Language, 32(2):188–190, May 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V32N2_1998_E.pdf.

Poggenpohl:1998:BRF

[1277] Sharon Helmer Poggenpohl. Book review: The Footnote*: *A curious history, Anthony Grafton. Cambridge: Harvard University Press, 1997. ISBN 0-674-90215-7. 242 pages, hardbound, $22.95. Visible Language, 32(2):190–191, May 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N2_1998_E.pdf.

Anonymous:1998:JIb

[1278] Anonymous. Journal information. Visible Language, 32(2):192–193, May 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V32N2_1998_E.pdf.

Poggenpohl:1998:ABc

[1279] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 32 (3):194, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N3_1998_E.pdf.

Poggenpohl:1998:VRI

[1280] Sharon Helmer Poggenpohl. Visual rhetoric: an introduction. Visible Language, 32(3):197–199, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V32N3_1998_E.pdf.

Poggenpohl:1998:DDR

[1281] Sharon Helmer Poggenpohl. Doubly damned, rhetorical and visual. Visible Language, 32(3):200–233, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V32N3_1998_E.pdf.

Brizuela:1998:TOR

[1282] Alejandro Brizuela. Tale of the origin: a rhetorical visual analysis of a Mexican mythical tale. Visible Language, 32(3):234–255, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V32N3_1998_E.pdf.

NavarroTapia:1998:PCS

[1283] Claudia Navarro Tapia. Pre-Columbian stamps: Pintaderas. Visible Language, 32(3):256–263, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V32N3_1998_E.pdf.

GonzalezdeCosio:1998:RL

[1284] Mária González de Cosío. Rhetoric in logotypes. Visible Language, 32(3): 264–279, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N3_1998_E.pdf.

Salazar:1998:PNC

[1285] Martha Salazar. The peso: National currency as rhetoric. Visible Language, 32(3):280–293, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V32N3_1998_E.pdf.

Anonymous:1998:SRB

[1286] Anonymous. Subscription rates. Back copies. Copyright information. Visible Language, 32(3):294, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V32N3_1998_E.pdf.

Anonymous:1998:ECB

[1287] Anonymous. Editorial correspondence. Business correspondence. Visible Language, 32(3):295, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V32N3_1998_E.pdf.

Anonymous:1998:IV

[1288] Anonymous. Index to volume 32. Visible Language, 32(3):296– 297, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V32N3_1998_E.pdf.

Gibbons:1998:ND

[1289] Charles Gibbons and John Nordyke. A note on the design. Visible Language, 32(3):302, September 1998. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V32N3_1998_E.pdf.

Kostelanetz:1999:RDH

[1290] Richard Kostelanetz. Remembering Dick Higgins (1938–1998). Visible Language, 33(1):3, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V33N1_1999_E.pdf.

Poggenpohl:1999:ABa

[1291] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 33 (1):4, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V33N1_1999_E.pdf.

Kostelanetz:1999:TYV

[1292] Richard Kostelanetz. Thirty years of visible writing: a memoir. Visible Language, 33(1):6–41, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V33N1_1999_E.pdf.

Beckman:1999:ACT

[1293] Laurel Beth Beckman. ASCII classroom, a text dependent investigation of the studio art classroom. Visible Language, 33(1):42–51, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N1_1999_E.pdf.

McLeer:1999:ALA

[1294] Brigid McLeer. Axis — a line about which a body, contextualizing photo-text work. Visible Language, 33(1):52–73, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N1_1999_E.pdf.

Frascara:1999:CEO

[1295] Jorge Frascara. Cognition, emotion and other inescapable dimensions of human experience. Visible Language, 33(1):74–89, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N1_1999_E.pdf.

Poggenpohl:1999:BRDa

[1296] Sharon Helmer Poggenpohl. Book review: Design Culture: an Anthology of Writing from the AIGA journal of Graphic Design, Steven Heller and Marie Finamore, editors. New York: Allworth Press, 1997. ISBN 1-80559-71-4. 304 pages. softbound, $19.95. Visible Language, 33(1):91, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N1_1999_E.pdf.

Poggenpohl:1999:BRG

[1297] Sharon Helmer Poggenpohl. Book review: Graphic Design Sources, Kenneth V. Hiebert, New Haven: Yale University Press, 1998. ISBN 0-300-07461-1. 214 pages, softbound, illustrated: 140 full color and 440 one color, §45.00. Visible Language, 33(1):91, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N1_1999_E.pdf.

Poggenpohl:1999:BRDb

[1298] Sharon Helmer Poggenpohl. Book review: Designing books practice and theory, Jost Hochuli and Robin Kinross. London: Hyphen Books, 1996. ISBN 0-9072759-08-1. 168 pager, hardbound, illustrated in two colors, $40.00. Visible Language, 33(1):92–93, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V33N1_1999_E.pdf.

Poggenpohl:1999:BRDc

[1299] Sharon Helmer Poggenpohl. Book review: Dimensional Typography: Case Studies on the Shape of Letters in Virtual Environments, J. Abbott Miller. Princeton: Kiosk Report (Princeton Architectural Press), 1996. ISBN 1-56898-089-2. 6o pages, softbound, illustrated in various colors, $11.95. Visible Language, 33(1):94, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V33N1_1999_E.pdf.

Poggenpohl:1999:BRH

[1300] Sharon Helmer Poggenpohl. Book review: A History of Communication Study: a Biographical Approach, Everett M. Rogers. New York: The Free Press, 1997. ISBN 0-684-84001-4.. 576 pages, softbound, $18.95. Visible Language, 33(1):94, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V33N1_1999_E.pdf.

Poggenpohl:1999:BRT

[1301] Sharon Helmer Poggenpohl. Book review: Typographis Polyglotta: A Comparative Study in Multilingual Typesetting, George Sadek and Maxim Zhukov. New York: The Cooper Union, 1991. 40 pages, softbound, $11.95. Visible Language, 33(1):95, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V33N1_1999_E.pdf.

Anonymous:1999:JIa

[1302] Anonymous. Journal information. Visible Language, 33(1):96, January 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N1_1999_E.pdf.

Weiss:1999:SWH

[1303] Irving Weiss. She was here a moment ago. Visible Language, 33 (2):101, May 1999. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V33N2_1999_E.pdf.

Herrick:1999:TDT

[1304] Earl Herrick. Toward disambiguating the term “Roman”. Visible Language, 33(2):102–127, May 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V33N2_1999_E.pdf.

Broglio:1999:BZ

[1305] Ron Broglio. Becoming-zoa. Visible Language, 33(2):128–149, May 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N2_1999_E.pdf.

Hallensleben:1999:WAA

[1306] Markus Hallensleben. The work of art in the age of digital reproduction: On the relationships between early twentieth century avant-garde movements and new media. Visible Language, 33(2):150–171, May 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N2_1999_E.pdf.

Presno:1999:SG

[1307] Caroline Presno. Screen gem. Visible Language, 33(2):173, May 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N2_1999_E.pdf.

Banks:1999:WBC

[1308] Colin Banks. What’s been cooking in the type kitchen? A report on the ATypI conference MultiType98. Visible Language, 33(2):174–182, May 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N2_1999_E.pdf.

Weiss:1999:X

[1309] Irving Weiss. S. Visible Language, 33(2):183, May 1999. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N2_1999_E.pdf.

Banks:1999:BRB

[1310] Colin Banks. Book review: The British Library Guide to Printing, History and Technique, Michael Twyman, 1998. ISBN 0-7123-4588-4. Paperback, illustrated, £9.95. Visible Language, 33(2):184–186, May 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N2_1999_E.pdf.

Poggenpohl:1999:BRP

[1311] Sharon Helmer Poggenpohl. Book review: Pentagram Book Five. Fifty case histories in architecture, graphics, and industrial design from the international consultancy, Pentagram, New York: The Monacelli Press, 1999. ISBN 1-58093-003-4 494 pages, hardbound, illustrated in full color, $90.00. Visible Language, 33(2):187–189, May 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V33N2_1999_E.pdf.

Poggenpohl:1999:ABb

[1312] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 33 (2):192, May 1999. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V33N2_1999_E.pdf.

Poggenpohl:1999:AB

[1313] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 33 (3):194, October 1999. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V33N3_1999_E.pdf.

Kim:1999:ERR

[1314] Min-Soo Kim. ‘An eccentric reversible reaction‘: Yi Sang’s experimental poetry in the 1930s and its meaning to contemporary design. Visible Language, 33(3):196–235, October 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V33N3_1999_E.pdf. See cor-

rection [1380].

Gluth:1999:RSV

[1315] Stuart Gluth. Roxane, a study in visual factors effecting legibility. Visible Language, 33(3):236–253, October 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V33N3_1999_E.pdf.

Gibbons:1999:ZRD

[1316] Charles Gibbons. Zealand, reflections on developing a typeface. Visible Language, 33(3):254–283, October 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V33N3_1999_E.pdf.

Ferguson:1999:CDA

[1317] Lynne Ferguson. CurioCity, developing an ‘active learning’ game. Visible Language, 33(3):284–307, October 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V33N3_1999_E.pdf.

Poggenpohl:1999:BRF

[1318] Sharon Helmer Poggenpohl. Book review: The Fluxus Reader, Ken Friedman, Editor. West Sussux, United Kingdom: Academy Editions, 1998 ISBN 0-471-97858-2, 308 pages, no illustrations, $29.95. Visible Language, 33(3):309–312, October 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V33N3_1999_E.pdf.

Poggenpohl:1999:BRI

[1319] Sharon Helmer Poggenpohl. Book review: Imagining Language, an Anthology, Jed Rasula and Steve McCaffery, Editors. Cambridge: MIT Press, 1998. ISBN 0-262-18186-X, 618 pages, hardbound, illustrated, $55.00. Visible Language, 33(3):313–315, October 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N3_1999_E.pdf.

Anonymous:1999:BRN

[1320] Anonymous. Book review: A New Dictionary of the Avant-Gardes, Richard Kostelanetz, New York: Schirmer/Macmillan. Expected publication in 2000. Visible Language, 33(3):316–317, October 1999. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V33N3_1999_E.pdf.

Anonymous:1999:IV

[1321] Anonymous. Index to volume 33. Visible Language, 33(3):318– 319, October 1999. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V33N3_1999_E.pdf.

Anonymous:1999:JI

[1322] Anonymous. Journal information. Visible Language, 33(3):319– 320, October 1999. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V33N3_1999_E.pdf.

Poggenpohl:2000:WSIa

[1323] Sharon Helmer Poggenpohl. Words in space: an introduction. Visible Language, 34(1):4–7, January 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V34N1_2000_E.pdf.

Poggenpohl:2000:PWK

[1324] Sharon Helmer Poggenpohl and Sang-Soo Ahn. Preserving words: The Korean Tripitaka. Visible Language, 34(1):8–13, January 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V34N1_2000_E.pdf.

Hayes:2000:BIC

[1325] Kevin Hayes. Bookcover as intertitle in the cinema of Jean-Luc Godard. Visible Language, 34(1):14–29, January 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V34N1_2000_E.pdf.

Owens:2000:RCW

[1326] Mark Owens. Reading the city: Writing and the construction of urban space in Jem Cohen’s Lost Book Found. Visible Language, 34(1): 30–55, January 2000. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V34N1_2000_E.pdf.

Wollensak:2000:VP

[1327] Andrea Wollensak. Visualizing place. Visible Language, 34(1):56– 75, January 2000. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V34N1_2000_E.pdf.

Patel:2000:BRP

[1328] Mookesh Patel. Book review: The Printed Bengali Character and its evolution, Fiona G. E. Ross. Surrey, United Kingdom: Curzon Press, 1999 ISBN 0-7007-1135-X. 244 pages, hardbound, illustrated: 110 one color. Visible Language, 34(1):77–81, January 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V34N1_2000_E.pdf.

Poggenpohl:2000:BRD

[1329] Sharon Helmer Poggenpohl. Book review: The Domain of Images, James Elkins. Ithaca, New York: Cornell University Press, 1999. ISBN 0-8014-3559-5, 282 pages, illustrated one color, cloth, $45.00. Visible Language, 34(1):82–86, January 2000. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V34N1_2000_E.pdf.

Rochon:2000:BRA

[1330] Alain Rochon and Sylvie Pouliot. Book review: ALIRE and DOC(K)S, MOTS-VOIR, AKENATON DOC(K)S, and the authors. October 1997. ISSN DOC(K)S 0396-3004. ISSN ALIRE 1260-8750. 258 pages, softbound, illustrated, one color, CD-ROM included email: akenaton docks@sitec.fr. Visible Language, 34(1):87–90, January 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V34N1_2000_E.pdf.

Poggenpohl:2000:BRH

[1331] Sharon Helmer Poggenpohl. Book review: The History of Counting, Denise Schmandt-Besserat. Michael Hays, illustrator. New York: Wm. Morrow & Company, 1999. ISBN 0-688-141188-8. 48 pages, cloth,ful1 color illustration, $17.00. Visible Language, 34(1):91, January 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V34N1_2000_E.pdf.

Anonymous:2000:JIa

[1332] Anonymous. Journal information. Visible Language, 34(1):92– 94, January 2000. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V34N1_2000_E.pdf.

Poggenpohl:2000:WSIb

[1333] Sharon Helmer Poggenpohl. Words in space: an introduction. Visible Language, 34(2):100–103, May 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V34N2_2000_E.pdf.

Shep:2000:REU

[1334] Sydney Shep. The restaurant at this end of the universe: Edible typography in New Zealand. Visible Language, 34(2):104–141, May 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V34N2_2000_E.pdf.

Rogal:2000:SBM

[1335] Maria Rogal. South of the border: Down Mexico way. Visible Language, 34(2):142–161, May 2000. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V34N2_2000_E.pdf.

Hitchcock:2000:WSB

[1336] Lucinda Hitchcock. Word space/book space/poetic space, experiments in transformation. Visible Language, 34(2):162–197, May 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V34N2_2000_E.pdf.

Poggenpohl:2000:RWS

[1337] Sharon Helmer Poggenpohl. Reflections on words in space. Visible Language, 34(2):198–218, May 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V34N2_2000_E.pdf.

Anonymous:2000:JIb

[1338] Anonymous. Journal information. Visible Language, 34(2):220–222, May 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V34N2_2000_E.pdf.

Poggenpohl:2000:AB

[1339] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 34 (3):226, September 2000. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V34N3_2000_E.pdf.

Wilkerson:2000:GPT

[1340] Kyoko Takahasi Wilkerson and Douglas Wilkerson. The gloss as poetics: Transcending the didactic. Visible Language, 34(3):228–263, September 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V34N3_2000_E.pdf.

Golec:2000:TID

[1341] Michael Golec. A typography of impoverishment: D. C. McMurtrie’s reception of European modernist typography and an American economic depression. Visible Language, 34(3):264–279, September 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V34N3_2000_E.pdf.

Seki:2000:ULI

[1342] Yusaku Seki. Using lists to improve text access: The role of layout in reading. Visible Language, 34(3):280–295, September 2000.

CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V34N3_2000_E.pdf.

Hartley:2000:PLT

[1343] James Hartley and Matthew Johnson. Portrait or landscape? Typographical layouts for patient information leaflets. Visible Language, 34(3):296–309, September 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V34N3_2000_E.pdf.

Banks:2000:MBR

[1344] Colin Banks. Making books: a review and critical commentary: Design in British Publishing Since 1945, Alan Bartram. London: British Library and Oak Knoll Press, ISBN 1-884718-93-0, 160 pages, 10 × 11 inches, illustrated, $39.95. Visible Language, 34(3):310–317, September 2000. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V34N3_2000_E.pdf.

Anonymous:2000:IV

[1345] Anonymous. Index for volume 34. Visible Language, 34(3):318– 319, September 2000. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V34N3_2000_E.pdf.

Anonymous:2000:JIc

[1346] Anonymous. Journal information. Visible Language, 34(3):320, September 2000. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V34N3_2000_E.pdf.

Poggenpohl:2001:ABa

[1347] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 35 (1):2, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N1_2001_E.pdf.

Poggenpohl:2001:P

[1348] Sharon Helmer Poggenpohl. Preface. Visible Language, 35(1):4, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V35N1_2001_E.pdf.

Minarelli:2001:FVI

[1349] Enzo Minarelli. Foreword: Voicimage: an international collection of essays about sound and image in contemporary poetical experimentalism. Visible Language, 35(1):5, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V35N1_2001_E.pdf.

Costa:2001:WPS

[1350] Mario Costa. The word of poetry, sounds of the voice and technology. Visible Language, 35(1):6–11, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V35N1_2001_E.pdf.

Polkinhorn:2001:THS

[1351] Harry Polkinhorn. True heritage: The sound image in experimental poetry. Visible Language, 35(1):12–19, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V35N1_2001_E.pdf.

Zurbrugg:2001:TPA

[1352] Nicholas Zurbrugg. Technology, polypoetry and the aura of polyperformance. Visible Language, 35(1):20–35, January 2001. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V35N1_2001_E.pdf.

Zurbrugg:2001:VLA

[1353] Nicholas Zurbrugg. Visible language, audible language, inarticulable language and the “supplementary signifier”. Visible Language, 35(1): 36–47, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N1_2001_E.pdf.

Szkarosi:2001:SCH

[1354] Endre Szkárosi. A soundscape of contemporary Hungarian poetry. Visible Language, 35(1):48–63, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V35N1_2001_E.pdf.

Menezes:2001:EPB

[1355] Philadelpho Menezes. Experimental poetics based on sound poetry today. Visible Language, 35(1):64–75, January 2001. CODEN VSLGAO.

ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V35N1_2001_E.pdf.

Cost:2001:EPB

[1356] Lis Cost. Experimental poetry in Barcelona during the 1990s. Visible Language, 35(1):76–91, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V35N1_2001_E.pdf.

Scholz:2001:RBS

[1357] Christian Scholz. Relations between sound poetry and visual poetry, the path from the optophonetic poem to the multime. Visible Language, 35 (1):92–103, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N1_2001_E.pdf.

Minarelli:2001:SBV

[1358] Enzo Minarelli. The singing blackbird, voice, images, technology in polypoetry. Visible Language, 35(1):104–115, January 2001. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V35N1_2001_E.pdf.

Minarelli:2001:MP

[1359] Enzo Minarelli. The manifesto of polypoetry. Visible Language, 35(1): 116–125, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N1_2001_E.pdf.

Anonymous:2001:JIa

[1360] Anonymous. Journal information. Visible Language, 35(1):126– 128, January 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N1_2001_E.pdf.

Poggenpohl:2001:ABb

[1361] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 35 (2):130, May 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N2_2001_E.pdf.

Salen:2001:SMT

[1362] Katie Salen. Surrogate multiplicities: Typography in the age of invisibility. Visible Language, 35(2):132–153, May 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V35N2_2001_E.pdf.

Maryniak:2001:ER

[1363] Nadia Maryniak. Ethnographic reflections. Visible Language, 35(2): 154–163, May 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N2_2001_E.pdf.

Frascara:2001:DWT

[1364] Jorge Frascara. Diagramming as a way of thinking ecologically. Visible Language, 35(2):164–177, May 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V35N2_2001_E.pdf.

Hsu:2001:EML

[1365] Sheng-Hsiung Hsu and Kuo-Chen Huang. Effects of minimal legible size characters on Chinese word recognition. Visible Language, 35 (2):178–191, May 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N2_2001_E.pdf.

Frascara:2001:BRI

[1366] Jorge Frascara. Book review: Interface / An Approach to Design, Gui Bonsiepe. Maastricht: Jan Van Eyck Akademie, 1999. ISBN 90-6617-212-6 168 pages, softbound, illustrated, one color. Visible Language, 35 (2):192–205, May 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N2_2001_E.pdf.

Burke:2001:LER

[1367] ?. Burke and Earl M. Herrick. Letter to the Editor: a response to Earl M. Herrick’s article “Toward Disambiguating the Term “Roman” in Visible Language 33.2, by Dr. Burke. Visible Language, 35(2):206–214, May 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V35N2_2001_E.pdf.

Anonymous:2001:CPCa

[1368] Anonymous. Call for papers: Communication design research for new media: a special issue of Visible Language for 2002. Visible Language, 35(2):216–219, May 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N2_2001_E.pdf.

Anonymous:2001:JIb

[1369] Anonymous. Journal information. Visible Language, 35(2):220–223, May 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V35N2_2001_E.pdf.

Poggenpohl:2001:ABc

[1370] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 35 (3):228, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N3_2001_E.pdf.

Winkler:2001:LLL

[1371] Dietmar Winkler. Limits of language, limits of world. Visible Language, 35(3):230–243, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V35N3_2001_E.pdf.

Loschky:2001:STP

[1372] Lester Loschky. Some things that pictures are good for: an information processing perspective. Visible Language, 35(3):244–265, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V35N3_2001_E.pdf.

McClain:2001:CVT

[1373] Matthew McClain. Critical viewing of television. Visible Language, 35 (3):266–277, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N3_2001_E.pdf.

Gunji:2001:NME

[1374] Jennifer Gunji. New media, experience and Japanese way of tea (Chado). Visible Language, 35(3):278–291, October 2001. CODEN VSLGAO.

ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V35N3_2001_E.pdf.

Anonymous:2001:BRW

[1375] Anonymous. Books received: Writing with language and by hand: Rethinking Writing, Roy Harris. Bloomington: Indiana University Press, 2000. 270 pages, hardbound, $39.95 ISBN 0-253-33776-3. Handwriting of the Twentieth Century, Rosemary Sassoon. London: Routledge, 1999. 208 pages, illustrated, one color, softbound, $27.99. ISBN 0-415-17882-7. Books and Pages, Polish Avant-garde and Artists’ Books in the 20th Century, Piotr Rypson. Warsaw: Center for Contemporary Art, 2000. 168 pages, illustrated, full color, hard paperback, $35.00. Available from P 0 Box 796, 00-950 Warsaw, Poland, ISBN 200083-88277-36-7. Designalltag — Symbols, logos, Identities, Ruedi Ruegg. Zurich: Designalltag, 1999. 76 pages, illustrated, one color, paperback, free upon request. ISBN 3-9520297-1-8. dotlinepixel — Thoughts on Cross-Media Design, Michele Iannuzzi and Richard Smith. Switzerland: Gabriele Capelli Editore. 44 pages, illustrated, some in color, hardbound, ISBN 88-87469-00-8. Emotional Digital — a Sourcebook of Contemporary Typographics, Alexander Branczyk, Jutta Nachtwey et al., editors. London: Thames & Hudson, 1999. 312 pages, illustrated, full color, hardbound, $50.00, ISBN 0-500-01925-8.. The Form of News, Kevin G. Bornhurst and John Nerone. New York: The Guilford Press, 2001. 326 pages, illustrated, one color, hardbound, $35.00, ISBN 1-57230-637-8, …. Visible Language, 35(3): 292–297, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N3_2001_E.pdf.

Stevens:2001:BRD

[1376] Graham Stevens. Book review: Design & typography: Type in Use — effective typography for electronic publishing, Alex W. White. New York: W. W. Norton, 208 pages, $26.95 ISBN 0-393-73034-4. Visible Language, 35(3):294–296, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V35N3_2001_E.pdf.

Hartley:2001:BRB

[1377] James Hartley. Book review: On Book Design, Richard Hendel, New Haven: Yale University Press, 1998, 210 pages, illustrated, one color, hardbound, $35.00. ISBN 300-07570-7. Visible Language, 35(3):298– 302, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N3_2001_E.pdf.

Miles:2001:BRP

[1378] John Miles. Book review: Printing Type Designs — a new history from Gutenberg to 2000, Dundan Glen. Akros Publications, 2001 ISBN 0-86142-110-8. [Obtainable direct from Duncan Glen, 33 Lady Nairn Avenue, Kirkcaldy, Fife, United Kingdom]. Visible Language, 35(3): 304–306, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N3_2001_E.pdf.

Poggenpohl:2001:BRN

[1379] Sharon Helmer Poggenpohl. Book review: New media: Alphabet to Email — How written English evolved and where it’s heading, Naomi S. Baron. London: Routledge, 2000, 316 pages, $18.95 ISBN 0-415-18685-4. Visible Language, 35(3):308–311, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V35N3_2001_E.pdf.

Lew:2001:CYS

[1380] Walter K. Lew. Correction: to Yi Sang poems. Visible Language, 35 (3):312–313, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N3_2001_E.pdf. See [1314].

Anonymous:2001:IV

[1381] Anonymous. Index for Volume 35. Visible Language, 35(3):314– 315, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N3_2001_E.pdf.

Anonymous:2001:JIc

[1382] Anonymous. Journal information. Visible Language, 35(3):316– 317, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V35N3_2001_E.pdf.

Anonymous:2001:FID

[1383] Anonymous. Forthcoming issues: Design research bibliographies; cultural dimensions of visual communication; search and research in communication design. Visible Language, 35(3):318, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V35N3_2001_E.pdf.

Anonymous:2001:CPCb

[1384] Anonymous. Call for papers: Cultural dimensions of visual communication. Visible Language, 35(3):319, October 2001. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V35N3_2001_E.pdf.

Poggenpohl:2002:ABa

[1385] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 36 (1):2, January 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N1_2002_E.pdf.

Strickler:2002:VDI

[1386] Zoe Strickler and Patricia Neafsey. Visual design of interactive software for older adults: Preventing drug interactions in older adults. Visible Language, 36(1):4–28, January 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V36N1_2002_E.pdf.

Ingemann:2002:MPE

[1387] Bruno Ingemann. The Mirage Project: an experimental qualitative reception study. Visible Language, 36(1):29–47, January 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V36N1_2002_E.pdf.

Golec:2002:ALM

[1388] Michael Golec. Alix Lambert’s ‘The Mark of Cain’. Visible Language, 36(1):48–61, January 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N1_2002_E.pdf.

Justus:2002:BRA

[1389] Carol F. Justus. Book review: Ancestor of the West. Writing, Reasoning, and Religion in Mesopotamia, Elam, and Greece, Jean Bottéro, Clarisse Herrenschmidt and Jean-Pierre Vernant. With a Foreword by François Zabbal. Teresa Lavender Fagan, translator. Chicago: University of Chicago Press, 2000. 194 pages, Illustrated, Hardbound, $25.00, ISBN 0-226-06715-7. Visible Language, 36(1):62–74, January 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V36N1_2002_E.pdf.

Peters:2002:BRF

[1390] Marsha Peters. Book review: The Form of The News: A History, Kevin G. Barnhurst and John Nerone. New York: Guilford Press, 2000. Hardcover, 326 pages, Illustrated, one color, $35. ISBN 1-52730-637-8. Visible Language, 36(1):76–83, January 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V36N1_2002_E.pdf.

Banks:2002:BRZ

[1391] Colin Banks. Book review: Zero: Hans Schleger — a Life of Design, Pat Schleger. New York: Princeton Architectural Press, 2000. 248 × 210mm, illustrated, hardbound, 272 pages, $45. ISBN 1-56898-273-9. Visible Language, 36(1):84–88, January 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V36N1_2002_E.pdf.

Anonymous:2002:CPV

[1392] Anonymous. Call for papers: Visual literacy and the dimensions of consciousness. Visible Language, 36(1):91–93, January 2002. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V36N1_2002_E.pdf.

Anonymous:2002:JIa

[1393] Anonymous. Journal information. Visible Language, 36(1):94– 95, January 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N1_2002_E.pdf.

Poggenpohl:2002:ABb

[1394] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 36 (2):98, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N2_2002_E.pdf.

Anonymous:2002:A

[1395] Anonymous. Acknowledgements. Visible Language, 36(2):100, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V36N2_2002_E.pdf.

Poggenpohl:2002:PCI

[1396] Sharon Helmer Poggenpohl. Preface: Cultivating an interest in design research. Visible Language, 36(2):101, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V36N2_2002_E.pdf.

Chayutsahakij:2002:ADR

[1397] Praima Chayutsahakij. Annotated design research bibliography process overview. Visible Language, 36(2):102–117, May 2002. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V36N2_2002_E.pdf.

Teeravarunyou:2002:PBP

[1398] Sakol Teeravarunyou and Carlos Teixeira. Perspectives on building a philosophy of design. Visible Language, 36(2):119–135, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V36N2_2002_E.pdf.

Anonymous:2002:APT

[1399] Anonymous. Annotated philosophy and theory of design list. Visible Language, 36(2):136–155, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V36N2_2002_E.pdf.

Jeamsinkul:2002:PBF

[1400] Chujit Jeamsinkul and Napawan Sawasdichai. Perspectives on building a foundation for design research. Visible Language, 36(2):156–173, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V36N2_2002_E.pdf.

Anonymous:2002:APM

[1401] Anonymous. Annotated principles and methods of design research list. Visible Language, 36(2):174–193, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V36N2_2002_E.pdf.

Chayutsahakij:2002:PBD

[1402] Praima Chayutsahakij. Perspectives on building a discourse between design theory and practice. Visible Language, 36(2):194–211, May 2002.

CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V36N2_2002_E.pdf.

Anonymous:2002:ATP

[1403] Anonymous. Annotated theory and practice in design list. Visible Language, 36(2):212–235, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V36N2_2002_E.pdf.

Anonymous:2002:AAL

[1404] Anonymous. Alphabetical author list. Visible Language, 36(2): 236–237, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N2_2002_E.pdf.

Anonymous:2002:GEB

[1405] Anonymous. Guest Editors biography. Visible Language, 36(2): 238–239, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N2_2002_E.pdf.

Anonymous:2002:JIb

[1406] Anonymous. Journal information. Visible Language, 36(2):240, May 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V36N2_2002_E.pdf.

Poggenpohl:2002:ABc

[1407] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 36 (3):242, September 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N3_2002_E.pdf.

Anonymous:2002:CCB

[1408] Anonymous. Commemoration: Colin Banks (1932–2002). Visible Language, 36(3):244, September 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V36N3_2002_E.pdf.

Poggenpohl:2002:CID

[1409] Sharon Helmer Poggenpohl. Cultivating an interest in design research. Visible Language, 36(3):246–252, September 2002. CODEN VSLGAO.

ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V36N3_2002_E.pdf.

Jeamsinkul:2002:MUM

[1410] Chujit Jeamsinkul and Sharon Poggenpohl. Methodology for uncovering motion affordance in interactive media. Visible Language, 36(3): 254–280, September 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N3_2002_E.pdf.

GonzalezdeCosio:2002:MME

[1411] Mária González de Cosío and Mary C. Dyson. Methods for manipulating electronic documents in relation to information retrieval. Visible Language, 36(3):282–306, September 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V36N3_2002_E.pdf.

Woolley:2002:CCE

[1412] S. I. Woolley, T. R. Davis, N. J. Flowers, J. Pinilla-Dutoit, A. Livingstone, and T. N. Arvanitis. Communicating cuneiform: The evolution of a multimedia cuneiform database. Visible Language, 36(3): 308–324, September 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N3_2002_E.pdf.

Dyson:2002:TCS

[1413] Mary C. Dyson and Judy Gregory. Typographic cueing on screen. Visible Language, 36(3):326–346, September 2002. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V36N3_2002_E.pdf.

Anonymous:2002:IV

[1414] Anonymous. Index for Volume 36. Visible Language, 36(3):347– 348, September 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N3_2002_E.pdf.

Anonymous:2002:JIc

[1415] Anonymous. Journal information. Visible Language, 36(3):350– 351, September 2002. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V36N3_2002_E.pdf.

Poggenpohl:2003:DTA

[1416] Sharon Helmer Poggenpohl. Developing theory, analysis and effective research communication for design. Visible Language, 37(1):5–12, January 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N1_2003_E.pdf.

Storkerson:2003:DTC

[1417] Peter Storkerson. Designing theory in communication. Visible Language, 37(1):13–42, January 2003. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V37N1_2003_E.pdf.

Mealing:2003:VAT

[1418] Stuart Mealing. Value-added text: Where graphic design meets paralinguistics. Visible Language, 37(1):43–58, January 2003. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N1_2003_E.pdf.

Sawasdichai:2003:UAF

[1419] Napawan Sawasdichai and Sharon Poggenpohl. User analysis framework: Thoughts on user’s cognitive factors for information design on Web-based media. Visible Language, 37(1):59–92, January 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V37N1_2003_E.pdf.

Nemeth:2003:GRN

[1420] Christopher Nemeth. Get real: The need for effective design research. Visible Language, 37(1):93–110, January 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V37N1_2003_E.pdf.

Anonymous:2003:JIa

[1421] Anonymous. Journal information. Visible Language, 37(1):111– 112, January 2003. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V37N1_2003_E.pdf.

Poggenpohl:2003:ABa

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V37N2_2003_E.pdf.

Poggenpohl:2003:CDV

[1423] Sharon Helmer Poggenpohl. Cultural dimensions of visual communication: an introduction. Visible Language, 37(2):116–121, May 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N2_2003_E.pdf.

Scotford:2003:FES

[1424] Martha Scotford. Function and expression: Student typographic work in English and Indian languages. Visible Language, 37(2):124–145, May 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N2_2003_E.pdf.

Search:2003:MDI

[1425] Patricia Search. The metastructural dynamics of interactive electronic design. Visible Language, 37(2):146–165, May 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V37N2_2003_E.pdf.

Kim:2003:MGG

[1426] Min-Soo Kim. Mapping a graphic genome: a cross-cultural comparison between Korean and Japanese designers. Visible Language, 37 (2):168–205, May 2003. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V37N2_2003_E.pdf.

Ciochetto:2003:TSE

[1427] Lynn Ciochetto. Toilet signage as effective communication. Visible Language, 37(2):208–221, May 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V37N2_2003_E.pdf.

Martinson:2003:GAN

[1428] Barbara Martinson and Sauman Chu. Graphic assimilation: New immigrants and social identity. Visible Language, 37(2):222–233, May 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N2_2003_E.pdf.

Poggenpohl:2003:BRL

[1429] Sharon Helmer Poggenpohl. Book review: Language Culture Type: International Type Design in the Age of Unicode, John D. Berry, editor. New York: Association Typographique Internationale (ATypI), 2002. Hardbound, 374 pages, extensively illustrated, some signatures in color, $60 Designed by Maxim Zhukov. ISBN 1-932026-01-0. Visible Language, 37 (2):233–236, May 2003. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V37N2_2003_E.pdf.

Anonymous:2003:JIb

[1430] Anonymous. Journal information. Visible Language, 37(2):239, May 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N2_2003_E.pdf.

Poggenpohl:2003:AB

[1431] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 37 (3):242, September 2003. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V37N3_2003_E.pdf.

Golec:2003:IPR

[1432] Michael Golec and Aron Vinegar. Instruction and provocation, or relearning from Las Vegas. Visible Language, 37(3):244–245, September 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N3_2003_E.pdf.

Bhatt:2003:AAC

[1433] Ritu Bhatt. Aesthetic or AnAesthetic: The competing symbols of Las Vegas Strip. Visible Language, 37(3):248–265, September 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N3_2003_E.pdf.

Golec:2003:DID

[1434] Michael Golec. “Doing it deadpan:” Venturi, Scott Brown and Izenour’s Learning from Las Vegas. Visible Language, 37(3):266–287, September 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N3_2003_E.pdf.

Vinegar:2003:SOB

[1435] Aron Vinegar. Skepticism and the ordinary — from Burnt Norton to Las Vegas. Visible Language, 37(3):288–311, September 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N3_2003_E.pdf.

Whiteley:2003:LVA

[1436] Nigel Whiteley. Learning from Las Vegas… and Los Angeles and Reyner Banham. Visible Language, 37(3):314–331, September 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V37N3_2003_E.pdf.

Upton:2003:STW

[1437] Dell Upton. Signs taken for wonders. Visible Language, 37(3):332– 351, September 2003. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V37N3_2003_E.pdf.

Anonymous:2003:JIc

[1438] Anonymous. Journal information. Visible Language, 37(3):352, September 2003. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V37N3_2003_E.pdf.

Hamann:2004:SMS

[1439] Byron Hamann. Seeing and the Mixtec screenfolds. Visible Language, 38(1):??, January 2004. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V38N1_2004_E.pdf.

Marcus:2004:VLA

[1440] Aaron Marcus and Valentina Johanna Baumgartner. A visible language analysis of user-interface design components and culture dimensions. Visible Language, 38(1):??, January 2004. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V38N1_2004_E.pdf.

Poggenpohl:2004:ABb

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V38N2_2004_E.pdf.

Poggenpohl:2004:I

[1442] Sharon Helmer Poggenpohl. Introduction. Visible Language, 38(2): 132–137, May 2004. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V38N2_2004_E.pdf.

Poggenpohl:2004:PCD

[1443] Sharon Helmer Poggenpohl. Practicing collaboration in design. Visible Language, 38(2):138–157, May 2004. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V38N2_2004_E.pdf.

Remington:2004:CSC

[1444] R. Roger Remington. A case study in collaboration: Looking back at the national graphic design archive. Visible Language, 38(2):158–167, May 2004. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V38N2_2004_E.pdf.

Melican:2004:USF

[1445] Jay Melican. User studies: Finding a place in design practice and education. Visible Language, 38(2):168–196, May 2004. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V38N2_2004_E.pdf.

Conley:2004:WDM

[1446] Chris Conley. Where are the design methodologists? Visible Language, 38(2):196–217, May 2004. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V38N2_2004_E.pdf.

Sato:2004:PDR

[1447] Keiichi Sato. Perspectives of design research: Collective views for forming the foundation of design research. Visible Language, 38(2):218–237, May 2004. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V38N2_2004_E.pdf.

Anonymous:2004:AN

[1448] Anonymous. Author notes. Visible Language, 38(2):238–239, May 2004. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V38N2_2004_E.pdf.

Poggenpohl:2004:EC

[1449] Sharon Helmer Poggenpohl. Editorial correspondence. Visible Language, 38(2):240, May 2004. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V38N2_2004_E.pdf.

Poggenpohl:2004:ABc

[1450] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 38 (3):242, September 2004. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V38N3_2004_E.pdf.

Reid:2004:TRF

[1451] Larry D. Reid, Meta L. Reid, and Audrey Bennett. Towards a readerfriendly font. Visible Language, 38(3):246–259, September 2004. CODEN

VSLGAO. ISSN 0022-2224. URL https://s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V38N3_2004_E.pdf.

Peachey:2004:LHW

[1452] Ian Peachey. Left-handedness: a writing handicap? Visible Language, 38(3):262–287, September 2004. CODEN VSLGAO. ISSN 0022-2224.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V38N3_2004_E.pdf.

Throop:2004:TPH

[1453] Liz C. Throop. Thinking on paper: Hindu–Arabic numerals in European typography. Visible Language, 38(3):290–303, September 2004. CODEN

VSLGAO. ISSN 0022-2224. URL https://s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V38N3_2004_E.pdf.

Abulhab:2004:MTS

[1454] Saad Abulhab. The Mutamathil type style: towards free, technologicallyfriendly ‘Arabetic’ types. Visible Language, 38(3):306–335, September 2004. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V38N3_2004_E.pdf. See comments [1470] and reply [1471]. See also

Anonymous:2004:JI

[1455] Anonymous. Journal information. Visible Language, 38(3):336, September 2004. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V38N3_2004_E.pdf.

Poggenpohl:2005:ABa

[1456] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 39 (1):1, January 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N1_2005_E.pdf.

McVarish:2005:ETE

[1457] Emily McVarish. Existential textuality: Engagement in the form of a letterpress book, John Crombie’s So. Visible Language, 39(1):4–32, January 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V39N1_2005_E.pdf.

Ruecker:2005:MLD

[1458] Stan Ruecker, Eric Homich, and Stéfan Sinclair. Multi-level document visualization. Visible Language, 39(1):33–41, January 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V39N1_2005_E.pdf.

Vukic:2005:RDP

[1459] Fedja Vukiç. Re-design in public space. Visible Language, 39(1): 42–44, January 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N1_2005_E.pdf.

Cehovin:2005:RDP

[1460] Eduard Cehovin. Redesign in public space: The work. Visible Language, 39(1):45–63, January 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N1_2005_E.pdf.

Baron:2005:BRC

[1461] Naomi Baron. Book review: Cybertalk at work and at play: Cyberpl@y: Communicating Online, Brenda Danet; Oxford and New York: Berg, 2001. ISBN 1-85973-424-3. 418 pages, illustrated, some in full color, softbound, $29.95. The Mobile Connection: The Cell Phone’s Impact on Society, Rich Ling. San Francisco: Morgan Kaufmann, 2004. ISBN 1-55860-936-9. 244 pages, softbound, $34.95. Netinguistics: Language, Discourse, and Ideology in Internet, Santiago Posteguillo, Castelló de la Plana: Publicaciones de la Universitat Jaume I, 2003. 180 pages, softbound, 13 EUR. Visible Language, 39(1):64–83, January 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V39N1_2005_E.pdf.

Hahn:2005:BRA

[1462] Young ae Hahn. Book review: Activity Centered Design: An Ecological Approach to Designing Smart Tools and Usable Systems, Geri Gay and Helene Hembrooke. Cambridge, Massachusetts: The MIT Press, 2004. ISBN 0-262-07248-3. 135 pages, cloth, illustrated, $30.00. Visible Language, 39(1):84–85, January 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V39N1_2005_E.pdf.

Anonymous:2005:BR

[1463] Anonymous. Books received. Visible Language, 39(1):86–92, January 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V39N1_2005_E.pdf.

Anonymous:2005:JIa

[1464] Anonymous. Journal information. Visible Language, 39(1):93– 94, January 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N1_2005_E.pdf.

Poggenpohl:2005:ABb

[1465] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 39 (2):97, May 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N2_2005_E.pdf.

Meirelles:2005:DVF

[1466] Isabel Meirelles. Dynamic visual formation. Visible Language, 39 (2):100–120, May 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N2_2005_E.pdf.

Nanavati:2005:OLL

[1467] Anuj A. Nanavati and Randolph G. Bias. Optimal line length in reading — a literature review. Visible Language, 39(2):121–145, May 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V39N2_2005_E.pdf.

Rane:2005:RDS

[1468] Mandar S. Rane. Rationalizing design sensitivity. Visible Language, 39 (2):146–167, May 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N2_2005_E.pdf.

Snart:2005:UWB

[1469] Jason Snart. UnReading William Blake’s marginalia. Visible Language, 39(2):168–193, May 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N2_2005_E.pdf.

Chahine:2005:LER

[1470] Nadine Chahine. Letters to the Editor: a reply to “The Mutamathil Type Style”. Visible Language, 39(2):194–201, May 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V39N2_2005_E.pdf. See [1454, 1471].

Abulhab:2005:R

[1471] Saad Abulhab. Response. Visible Language, 39(2):202–207, May 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V39N2_2005_E.pdf. See [1454, 1470].

Anonymous:2005:JIb

[1472] Anonymous. Journal information. Visible Language, 39(2):208, May 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V39N2_2005_E.pdf.

Poggenpohl:2005:ABc

[1473] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 39 (3):209, September 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N3_2005_E.pdf.

Friedman:2005:I

[1474] Ken Friedman and Owen Smith. Introduction. Visible Language, 39(3): 212–217, September 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N3_2005_E.pdf.

Smith:2005:TLA

[1475] Owen F. Smith. Teaching and learning about Fluxus: thoughts, observations and suggestions from the front lines. Visible Language, 39(3): 218–235, September 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N3_2005_E.pdf.

Clavez:2005:FRP

[1476] Bertand Clavez. Fluxus — reference or paradigm for young contemporary artists? Visible Language, 39(3):236–247, September 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V39N3_2005_E.pdf.

Higgins:2005:FO

[1477] Hannah Higgins. Fluxkids (overview). Visible Language, 39(3):248– 277, September 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N3_2005_E.pdf.

Bloom:2005:FFB

[1478] Ina Bloom. Fluxus futures, Ben Vautier’s signature acts and the historiography of the avant-garde. Visible Language, 39(3):278–307, September 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V39N3_2005_E.pdf.

Friedman:2005:HHL

[1479] Ken Friedman and Owen F. Smith. History, historiography and legacy. Visible Language, 39(3):308–317, September 2005. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V39N3_2005_E.pdf.

Anonymous:2005:IV

[1480] Anonymous. Index to Volume 39. Visible Language, 39(3):318– 319, September 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N3_2005_E.pdf.

Anonymous:2005:JIc

[1481] Anonymous. Journal information. Visible Language, 39(3):320, September 2005. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V39N3_2005_E.pdf.

Poggenpohl:2006:ABa

[1482] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 40 (1):2, January 2006. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V40N1_2006_E.pdf.

Friedman:2006:DL

[1483] Ken Friedman and Owen F. Smith. The dialectics of legacy. Visible Language, 40(1):4–11, January 2006. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V40N1_2006_E.pdf.

Klefstad:2006:WFC

[1484] Ann Klefstad. What has Fluxus created? Visible Language, 40(1): 12–27, January 2006. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V40N1_2006_E.pdf.

Moren:2006:ISM

[1485] Lisa Moren. Introductions: Signatures, music, computers, paranoia, smells, danger & the sky. Visible Language, 40(1):28–43, January 2006. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V40N1_2006_E.pdf.

Anonymous:2006:AS

[1486] Anonymous. Artists’ statements. Visible Language, 40(1):44–65, January 2006. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V40N1_2006_E.pdf.

Pearce:2006:GAA

[1487] Celia Pearce. Games as art: The aesthetics of play. Visible Language, 40(1):66–89, January 2006. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V40N1_2006_E.pdf.

Friedman:2006:LF

[1488] Ken Friedman. The literature of Fluxus. Visible Language, 40(1): 90–113, January 2006. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V40N1_2006_E.pdf.

Friedman:2006:FB

[1489] Ken Friedman and Owen F. Smith. A Fluxus bibliography. Visible Language, 40(1):114–127, January 2006. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V40N1_2006_E.pdf.

Anonymous:2006:JIa

[1490] Anonymous. Journal information. Visible Language, 40(1):128, January 2006. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V40N1_2006_E.pdf.

Poggenpohl:2006:ABb

[1491] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 40 (2):130, May 2006. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V40N2_2006_E.pdf.

Zuern:2006:SSS

[1492] John Zuern. System, suspension, seduction: Anne Bush’s critical design practice. Visible Language, 40(2):132–154, May 2006. CODEN VSL-

GAO. ISSN 0022-2224. URL http://visiblelanguage.herokuapp.
com/issue/182;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V40N2_2006_E.pdf.

Shelton:2006:BSP

[1493] Barrie Shelton and Emiko Okayama. Between script and pictures in Japan. Visible Language, 40(2):155–176, May 2006. CODEN VSL-

GAO. ISSN 0022-2224. URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V40N2_2006_E.pdf.

Zender:2006:AID

[1494] Mike Zender. Advancing icon design for global nonverbal communication: or what does the word bow mean? Visible Language, 40(2):177–206, May 2006. CODEN VSLGAO. ISSN 0022-2224. URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V40N2_2006_E.pdf.

Swanson:2006:TDA

[1495] Eric Swanson, Stacie Sabady, and Chris Yin. Teaching design: Analysis from three different analytical perspectives. Visible Language, 40(2):207–237, May 2006. CODEN VSLGAO. ISSN 0022-2224.

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V40N2_2006_E.pdf.

Anonymous:2006:JIb

[1496] Anonymous. Journal information. Visible Language, 40(2):238–240, May 2006. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V40N2_2006_E.pdf.

Poggenpohl:2006:ABc

[1497] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 40 (3):244, September 2006. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V40N3_2006_E.pdf.

Reynolds:2006:CRL

[1498] Linda Reynolds, Sue Walker, and Allison Duncan. Children’s responses to line spacing in early reading books or “holes to tell which line you’re on”. Visible Language, 40(3):246–267, September 2006. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V40N3_2006_E.pdf.

Moura:2006:AMI

[1499] Heloisa Moura. Analyzing multimodal interaction within a classroom setting. Visible Language, 40(3):270–291, September 2006. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V40N3_2006_E.pdf.

Abulhab:2006:TBA

[1500] Saad D. Abulhab. Typography behind Arabetic calligraphy veil. Visible Language, 40(3):294–307, September 2006. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V40N3_2006_E.pdf. See also

[1454, 1550].

Chiang:2006:CMO

[1501] William Chiang. Comparison of Maya and oracle bone scripts. Visible Language, 40(3):310–333, September 2006. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V40N3_2006_E.pdf.

Anonymous:2006:IV

[1502] Anonymous. Index for Volume 40. Visible Language, 40(3):334– 335, September 2006. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V40N3_2006_E.pdf.

Anonymous:2006:JIc

[1503] Anonymous. Journal information. Visible Language, 40(3):336, September 2006. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V40N3_2006_E.pdf.

Poggenpohl:2007:ABa

[1504] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 41 (1):2, January 2007. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V41N1_2007_E.pdf.

Ruecker:2007:DRP

[1505] Stan Ruecker, Lisa M. Given, Elizabeth Sadler, Andrea Ruskin, and Heather Simpson. Design of a rich-prospect browsing interface for seniors. Visible Language, 41(1):4–22, January 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V41N1_2007_E.pdf.

Zender:2007:VLE

[1506] Mike Zender and Keith A. Crutcher. Visual language for the expression of scientific concepts. Visible Language, 41(1):23–49, January 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N1_2007_E.pdf.

Ruecker:2007:BEB

[1507] Stan Ruecker and Kirsten C. Uszkalo. Binding the electronic book: Design features for bibliophiles. Visible Language, 41(1):50–69, January 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N1_2007_E.pdf.

Backhaus:2007:AAP

[1508] Peter Backhaus. Alphabet ante portas: How English text invades Japanese public space. Visible Language, 41(1):70–87, January 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/

V41N1_2007_E.pdf. See also earlier study [937].

Poggenpohl:2007:BRJa

[1509] Sharon Helmer Poggenpohl. Book review: Janet Abrams and Peter Hall, Editors, Else/Where: Mapping New Cartographies of Networks and Territories, Minneapolis: University of Minnesota Press, 2006. ISBN 0-9729696-2-4, 320 pages, softbound, full color illustrations. Visible Language, 41(1):88–89, January 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V41N1_2007_E.pdf.

Poggenpohl:2007:BRJb

[1510] Sharon Helmer Poggenpohl. Book review: John Maeda, The Laws of Simplicity: Design, Technology, Business, Life, Cambridge, MA: MIT Press, 2006. ISBN 0-262-13472-1. 108 pages, hardbound, $20.00. Visible Language, 41(1):90–91, January 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V41N1_2007_E.pdf.

Poggenpohl:2007:BRJc

[1511] Sharon Helmer Poggenpohl. Book review: Jonas Löwgren and Erik Stolterman, Thoughtful Interaction Design: a Design Perspective on Information Technology, MIT Press, 2004. ISBN 0-262-12271-5, 198 pages, hardbound, black and white illustrations, $35.00. Visible Language, 41 (1):92–93, January 2007. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V41N1_2007_E.pdf.

Poggenpohl:2007:BRJd

[1512] Sharon Helmer Poggenpohl. Book review: Jay David Bolter and Diane Gromala, Windows and Mirrors: Interaction Design, Digital Art and the Myth of Transparency Cambridge, MA: MIT Press, 2003. ISBN 0-262-02545-0, 182 pages, softbound, black and white illustrations, $17.95 . Visible Language, 41(1):94–95, January 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V41N1_2007_E.pdf.

Anonymous:2007:JIa

[1513] Anonymous. Journal information. Visible Language, 41(1):96, January 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N1_2007_E.pdf.

Sless:2007:DP

[1514] David Sless. Designing philosophy. Visible Language, 41(2), May 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N2_2007_E.pdf.

Poggenpohl:2007:ABb

[1515] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 41 (2):98, May 2007. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V41N2_2007_E.pdf.

Siess:2007:DP

[1516] David Siess. Designing philosophy. Visible Language, 41(2):101–126, May 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N2_2007_E.pdf.

Kenix:2007:HIN

[1517] Linda Jean Kenix. The homogenized imagery of non-profit organizations on the Internet. Visible Language, 41(2):127–161, May 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N2_2007_E.pdf.

Yu:2007:RVH

[1518] Lawrence Chun wai Yu. Relating the visual and the headline in Chinese print advertisements. Visible Language, 41(2):163–189, May 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N2_2007_E.pdf.

Anonymous:2007:JIb

[1519] Anonymous. Journal information. Visible Language, 41(2):190, May 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N2_2007_E.pdf.

Poggenpohl:2007:ABc

[1520] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 41 (3):194, September 2007. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V41N3_2007_E.pdf.

Westendorp:2007:VMU

[1521] Piet Westendorp and Karel van der Waarde. Visual metaphors in user instructions. Visible Language, 41(3):196–203, September 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N3_2007_E.pdf.

Meirelles:2007:UMD

[1522] Isabel Meirelles. The use of metaphors in dietary visual displays around the world. Visible Language, 41(3):204–219, September 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N3_2007_E.pdf.

Mitchel:2007:RTC

[1523] Marilyn Mitchel and Peter van Sommers. Representations of time in computer interface design. Visible Language, 41(3):220–245, September 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N3_2007_E.pdf.

Jones:2007:IUT

[1524] Phil Jones. Italicization and understanding texts through metaphoric projections of movement. Visible Language, 41(3):246–265, September 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N3_2007_E.pdf.

Wang:2007:MGA

[1525] Regina W. Y. Wang and Chun Cheng Hsu. The method of graphic abstraction in visual metaphor. Visible Language, 41(3):266–279, September 2007. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V41N3_2007_E.pdf.

Davis:2007:SD

[1526] Stephen Boyd Davis. A schema for depiction. Visible Language, 41(3): 280–300, September 2007. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V41N3_2007_E.pdf.

Anonymous:2007:JIc

[1527] Anonymous. Journal information. Visible Language, 41(3):301– 302, September 2007. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V41N3_2007_E.pdf.

Anonymous:2007:IVL

[1528] Anonymous. Index for Visible Language volume 41. Visible Language, 41(3):302–303, September 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V41N3_2007_E.pdf.

Anonymous:2007:C

[1529] Anonymous. Colophon. Visible Language, 41(3):304, September 2007. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V41N3_2007_E.pdf.

Poggenpohl:2008:ABa

[1530] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 42 (1):2, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N1_2008_E.pdf.

Armstrong:2008:AGL

[1531] David Scott Armstrong and Patrick Mahon. After the grave: Language and materiality. Visible Language, 42(1):4–13, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N1_2008_E.pdf.

Armstrong:2008:SPL

[1532] David Scott Armstrong. Sfumato, or, print: Like a vanishing point grown over by its picture plane. Visible Language, 42(1):14–27, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N1_2008_E.pdf.

Mahon:2008:XBE

[1533] Patrick Mahon. Xu Bing, Ed Pien and Gu Xiong: Lost and found in translation. Visible Language, 42(1):28–43, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N1_2008_E.pdf.

Takahashi:2008:ADF

[1534] Tess Takahashi. After the death of film: Writing the natural world in the digital age. Visible Language, 42(1):44–69, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N1_2008_E.pdf.

Anonymous:2008:AP

[1535] Anonymous. Artist’s projects. Visible Language, 42(1):70–108, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N1_2008_E.pdf.

Thib:2008:IT

[1536] Jeannie Thib. “Image” and “text”. Visible Language, 42(1):70– 75, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N1_2008_E.pdf.

Brennan:2008:OSF

[1537] Blair Brennan. Only a sudden flaming word. Visible Language, 42 (1):76–81, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N1_2008_E.pdf.

Gardner:2008:SCN

[1538] Joscelyn Gardner. Subverting a Caribbean ‘natural’ history. Visible Language, 42(1):82–85, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V42N1_2008_E.pdf.

Balfour:2008:WM

[1539] Barbara Balfour. A writer’s manual. Visible Language, 42(1):86– 89, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N1_2008_E.pdf.

Merritt:2008:AAC

[1540] David Merritt. Allmusic aspires to the condition of image. Visible Language, 42(1):90–93, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V42N1_2008_E.pdf.

Mahon:2008:EDL

[1541] Patrick Mahon. Excerpts from the drawn like money series. Visible Language, 42(1):94–97, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V42N1_2008_E.pdf.

Armstrong:2008:TTT

[1542] David Scott Armstrong. Turning, turner, turned. Visible Language, 42 (1):98–101, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N1_2008_E.pdf.

Lexier:2008:RS

[1543] Micah Lexier. Revelation series. Visible Language, 42(1):102– 108, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N1_2008_E.pdf.

Anonymous:2008:AN

[1544] Anonymous. Author notes. Visible Language, 42(1):109–110, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N1_2008_E.pdf.

Anonymous:2008:JIa

[1545] Anonymous. Journal information. Visible Language, 42(1):111, January 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N1_2008_E.pdf.

Poggenpohl:2008:AB

[1546] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 42 (2):115, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N2_2008_E.pdf.

Bessa:2008:SSS

[1547] Pedro Bessa. Skittish skirts and scanty silhouettes: The tribulations of gender in modern signage. Visible Language, 42(2):119–141, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N2_2008_E.pdf.

Noel:2008:LIF

[1548] Guillerminna Noël. Language impairment, family interaction and the design of a game. Visible Language, 42(2):143–157, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N2_2008_E.pdf.

Lee:2008:ONI

[1549] Jae Young Lee. Otto Neurath’s Isotype and the rhetoric of neutrality. Visible Language, 42(2):159–180, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V42N2_2008_E.pdf.

Abulhab:2008:AAT

[1550] Saad D. Abulhab. Anatomy of an Arabetic type design. Visible Language, 42(2):181–193, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N2_2008_E.pdf. See also [1454, 1500].

Anonymous:2008:BRA

[1551] Anonymous. Book review: Acting with Technology, Activity Theory and Interaction Design, Victor Kaptelinin and Bonnie A. Nardi. Cambridge, MA: MIT Press, 2006. ISBN-13 978-0-262-11298-7. Hardbound, 333 pages, black and white with diagrams, $35.00. Visible Language, 42(2):195, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N2_2008_E.pdf.

Anonymous:2008:BRE

[1552] Anonymous. Book review: Evocative Objects, Things We Think With, Sherry Turkle, editor. Cambridge, MA: MIT Press, 2007. ISBN 0-262-20168-2. Hard bound, 385 pages, black and white, illustrated, $24.95. Visible Language, 42(2):195, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V42N2_2008_E.pdf.

Anonymous:2008:BRL

[1553] Anonymous. Book review: Laws of Seeing, Wolfgang Metzger. Cambridge, MA: MIT Press, 2007. ISBN 0-262-13467-5. Hardbound, 203 pages, black and white, Illustrated, $48.00. Visible Language, 42 (2):196, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N2_2008_E.pdf.

Anonymous:2008:BRN

[1554] Anonymous. Book review: New Typographic Design, Roger Faxcett-Tang. New Haven, CT: Yale University Press, 2007. ISBN 10: 0-300-11775-2. Large format, 192 pages, full color, heavily illustrated $35.00. Visible Language, 42(2):196, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V42N2_2008_E.pdf.

Anonymous:2008:BRP

[1555] Anonymous. Book review: Print Is Dead, Books in Our Digital Age, Jeff Gomez. New York: Macmillan, 2008. ISBN 13: 978-0-230-52716-4. Hardbound, 221 pages, $24.95. Visible Language, 42(2):197, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N2_2008_E.pdf.

Anonymous:2008:BRU

[1556] Anonymous. Book review: Underground Maps After Beck, Maxwell J. Roberts. Harrow, UK: Capital Transport Publishing, 2005. Hardbound, large format, 112 pages with full color illustrations. ISBN 1-85414-286-0. Visible Language, 42(2):197–198, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V42N2_2008_E.pdf.

Anonymous:2008:BRW

[1557] Anonymous. Book review: Words to be Looked At, Language in 1960s Art, Liz Kotz. Cambridge, MA: MIT Press, 2007. Hardbound, 333 pages, black and white, illustrated, $29.95. Visible Language, 42(2):198, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N2_2008_E.pdf.

Anonymous:2008:JIb

[1558] Anonymous. Journal information. Visible Language, 42(2):201–202, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N2_2008_E.pdf.

Anonymous:2008:CPS

[1559] Anonymous. Call for papers: Special issue: Communication design failures: Function & interpretation scrutinized. Visible Language, 42 (2):205–206, May 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N2_2008_E.pdf.

Poggenpohl:2008:DLD

[1560] Sharon Helmer Poggenpohl. Design literacy, discourse and communities of practice. Visible Language, 42(3):213–236, September 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V42N3_2008_E.pdf.

Poon:2008:PFP

[1561] Janie Chun Nei Poon. Palimpsest: The future of the past. Visible Language, 42(3):237–264, September 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V42N3_2008_E.pdf.

Fathulla:2008:UDP

[1562] Kamaran Fathulla. Understanding diagrams: a pointer to the development of diagramming software. Visible Language, 42(3):265– 284, September 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N3_2008_E.pdf.

Djurek:2008:NWM

[1563] Nikola Djurek. The new way of making fonts with DTL Font Master. Visible Language, 42(3):285–300, September 2008. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V42N3_2008_E.pdf.

Anonymous:2008:JIc

[1564] Anonymous. Journal information. Visible Language, 42(3):301– 302, September 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N3_2008_E.pdf.

Anonymous:2008:IV

[1565] Anonymous. Index to Volume 42. Visible Language, 42(3):303, September 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N3_2008_E.pdf.

Anonymous:2008:DP

[1566] Anonymous. Distribution of pages. Visible Language, 42(3):304, September 2008. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V42N3_2008_E.pdf.

Poggenpohl:2009:ABa

[1567] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 43 (1):2, January 2009. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V43N1_2009_E.pdf.

Winkler:2009:VCV

[1568] Dietmar Winkler. Visual culture and visual communication in the context of globalization. Visible Language, 43(1):4–43, January 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V43N1_2009_E.pdf.

Moura:2009:TDC

[1569] Heloisa Moura, Dale Fahnstrom, Greg Prygrocki, and T. J. McLeish. Thinkeringspace: Designing for collaboration (around the book and beyond). Visible Language, 43(1):44–59, January 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V43N1_2009_E.pdf.

Gainor:2009:MBU

[1570] Rhiannon Gainor, Stéfan Sinclair, Stan Ruecker, Matt Patey, and Sandra Gabriele. A Mandala browser user study: Visualizing XML versions of Shakespeare’s plays. Visible Language, 43(1):60–85, January 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V43N1_2009_E.pdf.

Anonymous:2009:BRD

[1571] Anonymous. Book review: The Designer, half a century of change in image, training and techniques, Rosemary Sassoon. Bristol, UK: Intellect Books, 2008. ISBN 978-1-84150-195-6. Softbound, 144 pages, some illustrations, $30.00. Visible Language, 43(1):86–87, January 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V43N1_2009_E.pdf.

Anonymous:2009:BRV

[1572] Anonymous. Book review: Visual Thinking for Design, Colin Ware. New York: Morgan Kaufmann, 2008. ISBN 978-0-12-370896-0. Softbound, 197 pages, color illustration, $39.95. Visible Language, 43(1):87, January 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V43N1_2009_E.pdf.

Anonymous:2009:BRW

[1573] Anonymous. Book review: When Writing Met Art — from symbol to story Denise Schmandt-Besserat. Austin, TX: University of Texas Press, 2007. ISBN 0-292-71334-7. Hardbound, 134 pages, black and white illustration, $45.00. Visible Language, 43(1):88, January 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V43N1_2009_E.pdf.

Anonymous:2009:JIa

[1574] Anonymous. Journal information. Visible Language, 43(1):90– 91, January 2009. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V43N1_2009_E.pdf.

Anonymous:2009:SIV

[1575] Anonymous. Special issues of Visible Language since 2005. Visible Language, 43(1):92–93, January 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V43N1_2009_E.pdf.

Anonymous:2009:UDS

[1576] Anonymous. Upcoming double special issue: Glide: Global interaction in design. Visible Language, 43(1):94, January 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V43N1_2009_E.pdf.

Winkler:2009:CF

[1577] Dietmar R. Winkler and Sharon Poggenpohl. Celebrating failure. Visible Language, 43(2–3):??, July 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V43N23_2009_E.pdf.

Doherty:2009:RSA

[1578] Mike Doherty. ‘Realist’ stakeholder analysis in design. Visible Language, 43(2–3):??, July 2009. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V43N23_2009_E.pdf.

Singer:2009:PCF

[1579] Len D. Singer. Product communication, form, failure and safety. Visible Language, 43(2–3):??, July 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V43N23_2009_E.pdf.

Barnes:2009:MIF

[1580] Carolyn Barnes, Simone Taffe, and Lucy Miceli. Multiple information failure: a case of different investments in form and content in graphic design. Visible Language, 43(2–3):??, July 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V43N23_2009_E.pdf.

Roesler:2009:LTM

[1581] Axel Roesler. Lessons from Three Mile Island: Visual design in a high-stakes environment. Visible Language, 43(2–3):??, July 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V43N23_2009_E.pdf.

Brown:2009:PPB

[1582] Stephen Brown. Paper prototypes and beyond. Visible Language, 43(2–3):??, July 2009. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V43N23_2009_E.pdf.

Yee:2009:RRF

[1583] Joyce Yee, Matthew Lieveslay, and Louise Taylor. Recognizing riskof-failure in communication design projects. Visible Language, 43 (2–3):??, July 2009. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V43N23_2009_E.pdf.

Winkler:2009:FII

[1584] Dietmar Winkler. Failure? Isn’t it time to slay the design-dragon? Visible Language, 43(2–3):??, July 2009. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V43N23_2009_E.pdf.

Poggenpohl:2010:ABa

[1585] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 44 (1):2, January 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N1_2010_E.pdf.

Storkerson:2010:ASG

[1586] Peter Storkerson. Antinomies of semiotics in graphic design. Visible Language, 44(1):6–36, January 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V44N1_2010_E.pdf.

vanderWaarde:2010:VCM

[1587] Karel van der Waarde. Visual communication for medicines: Malignant assumptions and benign design? Visible Language, 44(1):40–69, January 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V44N1_2010_E.pdf.

McDonald:2010:FMC

[1588] Ann McDonald. Failure to manage constant change. Visible Language, 44(1):72–101, January 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N1_2010_E.pdf.

Lee:2010:IBE

[1589] Chae Ho Lee. Ibn Battuta Mall: Edutaining the world? Visible Language, 44(1):104–125, January 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V44N1_2010_E.pdf.

Winkler:2010:CWW

[1590] Sharon Poggenpohl and Dietmar R. Winkler. What have we learned from communication design failure? Visible Language, 44(1):128–139, January 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V44N1_2010_E.pdf.

Anonymous:2010:JIa

[1591] Anonymous. Journal information. Visible Language, 44(1):140– 141, January 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N1_2010_E.pdf.

Poggenpohl:2010:ABb

[1592] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 44 (2):146, May 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N2_2010_E.pdf.

Bennett:2010:GID

[1593] Audrey Grace Bennett. Global interaction in design. Visible Language, 44(2):149–159, May 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N2_2010_E.pdf.

McKee:2010:HPC

[1594] Stuart McKee. How print culture came to be indigenous. Visible Language, 44(2):161–186, May 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V44N2_2010_E.pdf.

Buck-Coleman:2010:NCC

[1595] Audra Buck-Coleman. Navigating cross-cultures, curriculum and confrontation. Visible Language, 44(2):187–206, May 2010. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V44N2_2010_E.pdf.

Blair:2010:BBP

[1596] Adream Blair. Beyond borders: Participatory design research and the changing role of design. Visible Language, 44(2):207–218, May 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V44N2_2010_E.pdf.

Moldenhauer:2010:VCG

[1597] Judith A. Moldenhauer. Virtual conferencing in global design education: Dreams and realities. Visible Language, 44(2):219–238, May 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V44N2_2010_E.pdf.

Lawson:2010:NSC

[1598] Cynthia Lawson. The New School collaborates: Organization and communication in immersive international field programs with artisan communities. Visible Language, 44(2):239–265, May 2010. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V44N2_2010_E.pdf.

Anonymous:2010:BRE

[1599] Anonymous. Book review: The Economics of Attention, Style and Substance in the Age of Information, Richard A. Lanham. Chicago, IL: University of Chicago Press, 2006. ISBN 10: 0-226-46867-4. Paperbound, 312 pages, black and white illustrations $I8.00. Visible Language, 44 (2):266–267, May 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N2_2010_E.pdf.

Anonymous:2010:BRGa

[1600] Anonymous. Book review: Graphic Design Translated, A Visual Directory of Terms for Global Design, Peter J. Wolf. Beverly, MA: Rockport Publishers, 2010. ISBN 978-1-59253-595-8. Hardbound, 432 pages, full color illustrations. $50.00. Visible Language, 44(2):267–268, May 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V44N2_2010_E.pdf.

Anonymous:2010:BRGb

[1601] Anonymous. Book review: The Grid Book, Hannah B. Higgins. Cambridge, MA: MIT Press, 2009. ISBN 978-0-262-51240-4. Paperback, 300 pages, black and white illustrations. $25.00. Visible Language, 44 (2):268–269, May 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N2_2010_E.pdf.

Anonymous:2010:JIb

[1602] Anonymous. Journal information. Visible Language, 44(2):270–272, May 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V44N2_2010_E.pdf.

Poggenpohl:2010:ABc

[1603] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 44 (3):276, September 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N3_2010_E.pdf.

Fong:2010:AOI

[1604] Michèle Wong Kung Fong. Audience/online information interactions: New research in learning preferences. Visible Language, 44(3):279– 303, September 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N3_2010_E.pdf.

Wang:2010:AIR

[1605] Hsiu-Feng Wang. The appropriateness of icon representations for Taiwanese computer users. Visible Language, 44(3):305–329, September 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V44N3_2010_E.pdf.

Mitchell:2010:DAS

[1606] Marilyn Mitchell. The development of automobile speedometer dials: a balance of ergonomics and style, regulation and power. Visible Language, 44(3):331–366, September 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V44N3_2010_E.pdf.

Winkler:2010:HFF

[1607] Dietmar R. Winkler. Helvetica, the film and the face in context. Visible Language, 44(3):367–378, September 2010. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V44N3_2010_E.pdf.

Anonymous:2010:IV

[1608] Anonymous. Index for Volume 44. Visible Language, 44(3):379– 381, September 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N3_2010_E.pdf.

Anonymous:2010:JIc

[1609] Anonymous. Journal information. Visible Language, 44(3):382– 384, September 2010. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V44N3_2010_E.pdf.

Toner:2011:SP

[1610] Anne Toner. Seeing punctuation. Visible Language, 45(1–2):??, August 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V45N12_2011_E.pdf.

Durrenmatt:2011:IVB

[1611] Jacques Dürrenmatt. From invisibility to visibility and backwards: Punctuation in comics. Visible Language, 45(1–2):??, August 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V45N12_2011_E.pdf.

Baron:2011:NSU

[1612] Naomi S. Baron and Rich Ling. Necessary smileys & useless periods: Redefining punctuation in electronically mediated communication. Visible Language, 45(1–2):??, August 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V45N12_2011_E.pdf.

Hall:2011:SNY

[1613] Nigel Hall and Sue Sing. Seven- to nine-year-olds’ understandings of speech marks: Some issues and problems. Visible Language, 45 (1–2):??, August 2011. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V45N12_2011_E.pdf.

Blackburn:2011:EMS

[1614] Nick Blackburn. Early modern ‘speech’ marks. Visible Language, 45 (1–2):??, August 2011. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V45N12_2011_E.pdf.

Lennard:2011:IVP

[1615] John Lennard. In/visible punctuation. Visible Language, 45(1– 2):??, August 2011. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V45N12_2011_E.pdf.

Luna:2011:MSB

[1616] Paul Luna. Marks, spaces and boundaries: Punctuation (and other effects) in the typography of dictionaries. Visible Language, 45(1– 2):??, August 2011. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V45N12_2011_E.pdf.

Poggenpohl:2011:ABc

[1617] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 45 (3):166, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V45N3_2011_E.pdf.

Dobson:2011:IVP

[1618] Teresa Dobson, Piotr Michura, Stan Ruecker, Monica Brown, and Omar Rodriguez. Interactive visualizations of plot in fiction. Visible Language, 45(3):169–191, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V45N3_2011_E.pdf.

Dyson:2011:DDS

[1619] Mary C. Dyson. Do designers show categorical perception of typefaces? Visible Language, 45(3):193–220, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V45N3_2011_E.pdf.

Wong:2011:CCE

[1620] Ho Lan Helena Wong. Critique: a communicative event in design education: a qualitative research on Western faculty and Asian students. Visible Language, 45(3):221–247, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V45N3_2011_E.pdf.

Anonymous:2011:BRA

[1621] Sharon Helmer Poggenpohl. Book review: The Architecture of Patterns, Paul Andersen and David Salomon. New York, NY: W. W. Norton & Company, 2010. ISBN 978-0-393-73293-1. Softbound, 144 pages, full color illustrations, $24.95. Visible Language, 45(3):250–251, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V45N3_2011_E.pdf.

Anonymous:2011:BRS

[1622] Sharon Helmer Poggenpohl. Book review: The Shallows: What the Internet is Doing to Our Brains, Nicholas Carr. New York, NY: W. W. Norton & Company, 2010. ISBN 978-0-393-07222-8, Hardbound, 276 pages, black and white, $26.95. Visible Language, 45(3):252–253, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V45N3_2011_E.pdf.

Storkerson:2011:BRD

[1623] Peter Storkerson. Book review: Design Integrations: Research and Collaboration, Sharon Poggenpohl and Keiichi Sato, Editors. Bristol, UK: Intellect, 2009. ISBN 978-1-84150-240-3. Softbound, 306 pages, black and white illustrations, $40.00. Visible Language, 45(3):254–257, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V45N3_2011_E.pdf.

Poggenpohl:2011:BRL

[1624] Sharon Helmer Poggenpohl. Book review: Limited Language: Rewriting Design, Responding to a feedback culture, Colin Davies and Monika Parrinder, Editors. Basel, SW: Birkhäuser Verlag AG, 2010. ISBN 2-00-993486-5. Softbound, 288 pages, full color illustrations, $26.00. Visible Language, 45(3):258–260, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V45N3_2011_E.pdf.

Poggenpohl:2011:BRH

[1625] Sharon Helmer Poggenpohl. Book review: Helvetica and the New York City Subway System, Paul Shaw. Cambridge, MA: MIT Press, 2011. ISBN 978-0-262-01548-6. Hardbound, 132 pages, full color illustration, $39.95. Visible Language, 45(3):261–263, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V45N3_2011_E.pdf.

Poggenpohl:2011:BRO

[1626] Sharon Helmer Poggenpohl. Book review: Open Design Now: Why Design Cannot Remain Exclusive, Bas van Abel, Lucus Evers, Roel Klaassen and Peter Troxler. Amsterdam, NL: BIS Publishers, 2011. ISBN 978-90-6369-259-9. Softbound, 320 pages, full color illustrations, $39.00. Visible Language, 45(3):264–267, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V45N3_2011_E.pdf.

Anonymous:2011:IV

[1627] Anonymous. Index to Volume 45. Visible Language, 45(3):268– 269, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V45N3_2011_E.pdf.

Anonymous:2011:JIc

[1628] Anonymous. Journal information. Visible Language, 45(3):270, December 2011. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-

5529 (electronic).  URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V45N3_2011_E.pdf.

Poggenpohl:2012:EFD

[1629] Sharon Poggenpohl. Envisioning a future design education, introduction. Visible Language, 46(1–2):??, May 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V46N12_2012_E.pdf.

vanderWaarde:2012:CDE

[1630] Karel van der Waarde and Maurits Vroombout. Communication design education: Could nine reflections be sufficient? Visible Language, 46(1–2):??, May 2012. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V46N12_2012_E.pdf.

Frascara:2012:WMD

[1631] Jorge Frascara and Guillermina Noël. What’s missing in design education today? Visible Language, 46(1–2):??, May 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V46N12_2012_E.pdf.

Sless:2012:DDE

[1632] David Sless. Design or ‘design’ — envisioning a future design education. Visible Language, 46(1–2):??, May 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V46N12_2012_E.pdf.

Bennett:2012:GDG

[1633] Audrey Bennett. Good design is good social change: Envisioning an age of accountability in communication design. Visible Language, 46 (1–2):??, May 2012. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V46N12_2012_E.pdf.

Ockerse:2012:LCD

[1634] Thomas Ockerse. Learn from the core — design from the core. Visible Language, 46(1–2):??, May 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V46N12_2012_E.pdf.

Myers:2012:HHN

[1635] Chris Myers. Handsomely, handsomely now! 5 impromptus for the early part of the century. Visible Language, 46(1–2):??, May 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V46N12_2012_E.pdf.

Davis:2012:LGE

[1636] Meredith Davis. Leveraging graduate education for a more relevant future. Visible Language, 46(1–2):??, May 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V46N12_2012_E.pdf.

Ruecker:2012:PPC

[1637] Stan Ruecker. The perennial and the particular challenges of design education. Visible Language, 46(1–2):??, May 2012. CODEN VSLGAO.

ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V46N12_2012_E.pdf.

Friedman:2012:MDE

[1638] Ken Friedman. Models of design: Envisioning a future design education. Visible Language, 46(1–2):??, May 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V46N12_2012_E.pdf.

Poggenpohl:2012:ABc

[1639] Sharon Helmer Poggenpohl. Advisory Board. Visible Language, 46 (3):178, September 2012. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V46N3_2012_E.pdf.

Walker:2012:DDC

[1640] Sue Walker. Describing the design of children’s books: an analytical approach. Visible Language, 46(3):180–199, September 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V46N3_2012_E.pdf.

Beier:2012:HMB

[1641] Sofie Beier. How my brain stopped reading. Visible Language, 46(3): 200–205, September 2012. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V46N3_2012_E.pdf.

Koch:2012:ETD

[1642] Beth E. Koch. Emotions in typographic design: an empirical examination. Visible Language, 46(3):206–227, September 2012. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V46N3_2012_E.pdf.

Wang:2012:CMO

[1643] Regina W. Y. Wang and Chiung-Fen Wang. Composition methodology of optical illusion design. Visible Language, 46(3):228–245, September 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V46N3_2012_E.pdf.

Black:2012:DCI

[1644] Alison Black and Karen L. Stanbridge. Documents as ‘critical incidents’ in organization to consumer communication. Visible Language, 46(3): 246–281, September 2012. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V46N3_2012_E.pdf.

Anonymous:2012:BRG

[1645] David Cabianca. Book review: Graphic Design: Now in Production: a Note on Emerging Cultural Relevancy for Graphic Design, Ian Albinson, Rob Giampietro, Andrew Blauvelt, & Ellen Lupton. Minneapolis, MN: Walker Art Center, 2011. ISBN 978-0-935640-98-4. Paperback, 224 pages, illustrated, full color, $40.00. Visible Language, 46(3):284–287, December 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V46N3_2012_E.pdf.

Poggenpohl:2012:BRCa

[1646] Sharon Helmer Poggenpohl. Book review: Chasing the White Whale, The Moby-Dick Marathon; or, What Melville Means Today, David Dowling. Iowa City, IA: University of Iowa Press, 2010. ISBN 978-1-58729-906-3. Paperback, 242 pages, a few illustrations, black and white, $24.95. Visible Language, 46(3):288–289, December 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V46N3_2012_E.pdf.

Poggenpohl:2012:BRCb

[1647] Sharon Helmer Poggenpohl. Book review: China’s Design Revolution, Lorraine Justice. Cambridge, MA: MIT Press, 2012. 978-0-262-01742-8. Hardcover, 144 pages, illustrated black and white, $21.95. Visible Language, 46(3):290–292, December 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V46N3_2012_E.pdf.

Poggenpohl:2012:BRD

[1648] Sharon Helmer Poggenpohl. Book review: Design & Designing, a Critical Introduction, Steve Garner and Chris Evans, editors. London, UK: Berg, 2012. ISBN 978-1-84788-576-0. Paperback, 483 pages, illustrated, some color, $49.95. Visible Language, 46(3):293–295, December 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V46N3_2012_E.pdf.

Poggenpohl:2012:BRR

[1649] Sharon Helmer Poggenpohl. Book review: Reading Letters, Designing for Legibility, Sofie Beier. Amsterdam, NL: BIS Publishers, 2012. ISBN 978-90-6369-271-1. Hardbound, 181 pages, illustrated, full color, $40.00. Visible Language, 46(3):296–297, December 2012. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V46N3_2012_E.pdf.

Anonymous:2012:IV

[1650] Anonymous. Index to Volume 45. Visible Language, 46(3):298– 300, December 2012. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V46N3_2012_E.pdf.

Anonymous:2012:JIc

[1651] Anonymous. Journal information. Visible Language, 46(3):301– 303, December 2012. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V46N3_2012_E.pdf.

Anonymous:2012:CTM

[1652] Anonymous. Colophon: Typefaces: Miller Test, Miller Display, Super Grotesk. Visible Language, 46(3):304, December 2012. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V46N3_2012_E.pdf.

Zender:2013:ABa

[1653] Mike Zender. Advisory Board. Visible Language, 47(1):2, January 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V47N1_2013_E.pdf.

Poggenpohl:2013:VLT

[1654] Sharon Poggenpohl and Paul Michael Zender. Visible language in transition. Visible Language, 47(1):4–11, January 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V47N1_2013_E.pdf.

Poggenpohl:2013:RTR

[1655] Sharon Poggenpohl. Reflections on teaching research: a conversation with Meredith Davis, Mary Dyson, Judith Gregory. Visible Language, 47(1):12–37, January 2013. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V47N1_2013_E.pdf.

Baki:2013:BDL

[1656] Randa Abdel Baki. Bilingual design layout systems: Cases from Beirut. Visible Language, 47(1):38–65, January 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V47N1_2013_E.pdf.

Zender:2013:IID

[1657] Mike Zender and Mauricio Mejía. Improving icon design: Through focus on the role of individual symbols in construction of meaning. Visible Language, 47(1):66–89, January 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V47N1_2013_E.pdf.

Anonymous:2013:JIa

[1658] Anonymous. Journal information. Visible Language, 47(1):96, January 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V47N1_2013_E.pdf.

Zender:2013:AB

[1659] Mike Zender. Advisory Board. Visible Language, 47(2):i, May 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V47N2_2013_E.pdf.

Venkatesh:2013:SRD

[1660] Aruna Venkatesh. A study on the revelations of design students’ thinking styles in reflective journals. Visible Language, 47(2):1–36, May 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V47N2_2013_E.pdf.

Black:2013:DQG

[1661] Alison Black, Annette Gibb, Clare Carey, Sarah Barker, Claire Leake, and Luke Solomons. Designing a questionnaire to gather carer input pain assessment for hospitalized people with dementia. Visible Language, 47(2):37–60, May 2013. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V47N2_2013_E.pdf.

Corcoran:2013:MCS

[1662] Heather Corcoran, Matthew Kreuter, and Christina Clarke. Making cancer surveillance data more accessible for the public through Dataspark. Visible Language, 47(2):61–87, May 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V47N2_2013_E.pdf.

Anonymous:2013:CPE

[1663] Anonymous. Call for papers: Environmental communication: a special issue of Visible Language for 2014. Visible Language, 47(2):89–90, May 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V47N2_2013_E.pdf.

Zender:2013:RRR

[1664] Mike Zender. Reliable + relevant research. Visible Language, 47 (2):91–92, May 2013. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V47N2_2013_E.pdf.

Anonymous:2013:JIb

[1665] Anonymous. Journal information. Visible Language, 47(2):93, May 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V47N2_2013_E.pdf.

Zender:2013:ENB

[1666] Mike Zender. Editor’s note: Blunt Conference. Visible Language, 47 (3):5, September 2013. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V47N3_2013_E.pdf.

Griffin:2013:MBJ

[1667] Dori Griffin. Moving beyond ‘just making things‘: Design history in the studio and the survey classroom. Visible Language, 47(3):6– 28, September 2013. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V47N3_2013_E.pdf.

Coorey:2013:CWS

[1668] Jillian Coorey and Gretchen Caldwell Rinnert. Critical writing strategies to improve class critiques. Visible Language, 47(3):30–51, September 2013. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V47N3_2013_E.pdf.

Cooper:2013:LLB

[1669] Alexander Cooper, Rose Gridneff, and Andrew Haslam. Letterpress: Looking backward to look forward. Visible Language, 47(3): 52–71, September 2013. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V47N3_2013_E.pdf.

Beier:2013:ISH

[1670] Sofie Beier and Mary C. Dyson. The influence of serifs on ’h’ and ’i’: Useful knowledge from design-led scientific research. Visible Language, 47 (3):74–95, September 2013. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V47N3_2013_E.pdf.

Moys:2013:IRI

[1671] Jeanne-Louise Moys. Investigating readers’ impressions of typographic differentiation using repertory grids. Visible Language, 47(3):96– 123, September 2013. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V47N3_2013_E.pdf.

Anonymous:2013:ABJ

[1672] Anonymous. [author biography: Jorge Frascara]. Visible Language, 47 (3):124, September 2013. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V47N3_2013_E.pdf.

Anonymous:2013:ABS

[1673] Anonymous. [author biography: Stan Ruecker]. Visible Language, 47 (3):125, September 2013. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V47N3_2013_E.pdf.

Zender:2014:ABa

[1674] Mike Zender. Advisory Board. Visible Language, 48(1):2, May 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N1_2014_E.pdf.

Mollerup:2014:SPS

[1675] Per Mollerup. Slide presentations, seriously. Visible Language, 48 (1):4–21, May 2014. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V48N1_2014_E.pdf.

Peterson:2014:ITI

[1676] Matthew O. Peterson. The integration of text and image in media and its impact on reader interest. Visible Language, 48(1):22–39, May 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N1_2014_E.pdf.

Moys:2014:TLF

[1677] Jeanne-Louise Moys. Typographic layout and first impressions — testing how changes in text layout influence reader’s judgments of documents. Visible Language, 48(1):40–67, May 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V48N1_2014_E.pdf.

Zender:2014:MUI

[1678] Mike Zender and Amy Cassedy. (mis)understanding: icon comprehension in different cultural contexts. Visible Language, 48(1):68–95, May 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N1_2014_E.pdf.

Anonymous:2014:JIa

[1679] Anonymous. Journal information. Visible Language, 48(1):96, May 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N1_2014_E.pdf.

Zender:2014:ABb

[1680] Mike Zender. Advisory Board. Visible Language, 48(2):2, August 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N2_2014_E.pdf.

Fernandez:2014:GEN

[1681] Oscar Fernández. Guest Editor notes. Visible Language, 48(2):4, August 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N2_2014_E.pdf.

Cheng:2014:MDC

[1682] Karen Cheng and Sarah Pérez-Kriz. Map design for complex architecture: a user study of maps & wayfinding. Visible Language, 48 (2):6–33, August 2014. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V48N2_2014_E.pdf.

Walton:2014:ESH

[1683] Ashley Walton. The environment is (still) not in the head: Harry Heft & contemporary methodological approaches to navigation and wayfinding. Visible Language, 48(2):34–47, August 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V48N2_2014_E.pdf.

Fontaine:2014:LDT

[1684] Lisa Fontaine. Learning design thinking by designing learning experiences: a case study in the development of strategic thinking skills through the design of interactive museum exhibitions. Visible Language, 48(2):48–69, August 2014. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V48N2_2014_E.pdf.

Cue:2014:WDA

[1685] Patricia Cué. On the wall: Designers as agents for change in environmental communication. Visible Language, 48(2):70–83, August 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N2_2014_E.pdf.

Schwanbeck:2014:RPU

[1686] Andrew T. Schwanbeck. Rebuilding perceptions: Using experiential graphic design to reconnect neighborhoods to the greater city population. Visible Language, 48(2):84–107, August 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V48N2_2014_E.pdf.

Anonymous:2014:CPV

[1687] Anonymous. Call for papers: Vl special edition: Critical making: Design and the digital humanities. Visible Language, 48(2):108–109, August 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N2_2014_E.pdf.

Frascara:2014:BRDa

[1688] Jorge Frascara. Book review: Design for Information: An Introduction to the Histories, Theories, & Best Practices Behind Effective Information Visualizations, Isabel Meirelles. Beverly, MA: Rockport Publishers, 2013. Visible Language, 48(2):110, August 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V48N2_2014_E.pdf.

Anonymous:2014:JIb

[1689] Anonymous. Journal information. Visible Language, 48(2):112, August 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N2_2014_E.pdf.

Zender:2014:ABc

[1690] Mike Zender. Advisory Board. Visible Language, 48(3):2, November 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N3_2014_E.pdf.

Borkent:2014:VIC

[1691] Mike Borkent. Visual improvisation: Cognition, materiality, and postlinguistic visual poetry. Visible Language, 48(3):4–27, November 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N3_2014_E.pdf.

Anonymous:2014:TFT

[1692] Anonymous. Typographic features of text: Outcomes from research and practice. Visible Language, 48(3):28–67, November 2014. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N3_2014_E.pdf.

Anonymous:2014:SAV

[1693] Anonymous. A statistical approach for visualizing the quality of multi-hospital data. Visible Language, 48(3):68–85, November 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N3_2014_E.pdf.

Anonymous:2014:LDP

[1694] Anonymous. Linking design principles with educational research theories to teach sound to symbol reading correspondence with multisensory type. Visible Language, 48(3):86–108, November 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V48N3_2014_E.pdf.

Frascara:2014:BRDb

[1695] Jorge Frascara. Book review: Design for information, an introduction to the histories, theories, and best practices behind effective information visualizations, by Isabel Meirelles, Rockport Publishers, 2013. Visible Language, 48(3):109–110, November 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V48N3_2014_E.pdf.

Anonymous:2014:BRI

[1696] Anonymous. Book review: Isotype: Design and contexts 1925–1971. Visible Language, 48(3), November 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V48N3_2014_E.pdf.

Mollerup:2014:BRI

[1697] Per Mollerup. Book review: Isotype: Design and contexts 1925– 1971, Burke, Christopher; Kindel, Eric; Walker, Sue (Eds.). Hyphen Press, London, 2013. Visible Language, 48(3):111–121, November 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V48N3_2014_E.pdf.

Zender:2014:BRM

[1698] Mike Zender. Book reviews: The most important design books most designers have never read: The Case for Mental Imagery, Stephen M. Kosslyn, William L. Thompson, and Giorgio Ganis. Visible Language, 48(3): 122–127, November 2014. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V48N3_2014_E.pdf.

Zender:2014:ABW

[1699] Mike Zender. Advisory Board welcome: Keith Crutcher. Visible Language, 48(3):127, November 2014. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V48N3_2014_E.pdf.

Anonymous:2014:JIc

[1700] Anonymous. Journal information. Visible Language, 48(3):128– 130, November 2014. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V48N3_2014_E.pdf.

Hanke:2015:TCR

[1701] Regina Hanke. Targeted communication to reduce antibiotic prescription. Visible Language, 49(1–2):12–23, April 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL

http://visiblelanguage.herokuapp.com/issue/161/article/891;
 https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/V49N12_
2015_E.pdf.

Bennett:2015:CLH

[1702] Audrey Bennett. Connotative localization of an HIV prevention image to promote safer sex practices in Ghana. Visible Language, 49(1–2):24–39, April 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (elec-

tronic). URL http://visiblelanguage.herokuapp.com/issue/161/
article/901;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V49N12_2015_E.pdf.

Zender:2015:DEH

[1703] Mike Zender and David K. Plate. Designing and evaluating a health program in Africa: Hygiene matters. Visible Language, 49(1–2):40–61, April 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/161/article/
911;   https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V49N12_2015_E.pdf.

Noel:2015:DVT

[1704] Guillermina Noël. Designing a visual tool to interview people with communication disabilities: a user-centered approach. Visible Language, 49 (1–2):62–79, April 2015. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL http://visiblelanguage.herokuapp.
com/issue/161/article/921;   https://s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V49N12_2015_E.pdf.

Patton:2015:UIO

[1705] Amina Patton, Morgan Griffin, Ana Tellez, Mary Ann Petti, and Xanthi Scrimgeour. Using icons to overcome communication barriers during emergencies: a case study of the show me interactive tools. Visible Language, 49(1–2):80–95, April 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL

http://visiblelanguage.herokuapp.com/issue/161/article/931;
 https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/V49N12_
2015_E.pdf.

Taylor:2015:DDB

[1706] Andrea Taylor, Tara French, Jeni Lennox, and Jeremy Keen. Developing a design brief for a virtual hospice using design tools and methods: a preliminary exploration. Visible Language, 49(1–2):96–111, April 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL

http://visiblelanguage.herokuapp.com/issue/161/article/941;
 https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/V49N12_
2015_E.pdf.

Lowe:2015:TDP

[1707] Sarah Lowe, Tami H. Wyatt, Xueping Li, and Susan Fancher. Transdisciplinary partnerships in it health software development: the benefits to learning. Visible Language, 49(1–2):112–127, April 2015. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL

http://visiblelanguage.herokuapp.com/issue/161/article/951;
 https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/V49N12_
2015_E.pdf.

Napier:2015:CDH

[1708] Pamela Napier and Terri Wada. Co-designing for healthcare: visual designers as researchers and facilitators. Visible Language, 49(1–2):128–143, April 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (elec-

tronic). URL http://visiblelanguage.herokuapp.com/issue/161/
article/961;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V49N12_2015_E.pdf.

Paulovich:2015:DIH

[1709] Belinda Paulovich. Design to improve the health education experience: using participatory design methods in hospitals with clinicians and patients. Visible Language, 49(1–2):144–159, April 2015. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL

http://visiblelanguage.herokuapp.com/issue/161/article/971;
 https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/V49N12_
2015_E.pdf.

Sanematsu:2015:DLI

[1710] Helen Sanematsu, Brenda Hudson, Amanda Nyhuis, Paul Dexter, and Siu Hui. Design and language impact on study volunteerism in medical research: learnings from a controlled study of recruitment letters. Visible Language, 49(1–2):160–171, April 2015. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL

http://visiblelanguage.herokuapp.com/issue/161/article/981;
 https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/V49N12_
2015_E.pdf.

Thiessen:2015:BTL

[1711] Myra Thiessen, Mark Kohler, Owen Churches, Scott Coussens, and Hannah Keage. Brainy type: a look at how the brain processes typographic information. Visible Language, 49(1–2):174–189, April 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL

http://visiblelanguage.herokuapp.com/issue/161/article/991;
 https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/V49N12_
2015_E.pdf.

Zender:2015:ABc

[1712] Mike Zender. Advisory Board. Visible Language, 49(3):2, December 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V49N3_2015_E.pdf.

Barness:2015:CME

[1713] Jessica Barness and Amy Papaelias. Critical making at the edges. Visible Language, 49(3):4–11, December 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V49N3_2015_E.pdf.

Burdick:2015:MMM

[1714] Anne Burdick. Meta! Meta! Meta! A speculative design brief for the digital humanities. Visible Language, 49(3):12–33, December 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V49N3_2015_E.pdf.

Ricci:2015:CAU

[1715] Donato Ricci, Robin de Mourat, Christophe Leclercq, and Bruno Latour. Clues. Anomalies. Understanding. Detecting underlying assumptions and expected practices in the Digital Humanities through the AIME project. Visible Language, 49(3):34–61, December 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V49N3_2015_E.pdf.

Willis:2015:WIC

[1716] Holly Willis. Writing images and the cinematic humanities. Visible Language, 49(3):62–77, December 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V49N3_2015_E.pdf.

Allen:2015:BMU

[1717] Tania Allen and Sara Queen. Beyond the map: Unpacking critical cartography in the digital humanities. Visible Language, 49(3):78–99, December 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V49N3_2015_E.pdf.

BoydDavis:2015:IIH

[1718] Stephen Boyd Davis and Florian Kräutli. The idea and image of historical time: Interactions between design and digital humanities. Visible Language, 49(3):100–119, December 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V49N3_2015_E.pdf.

Anderson:2015:CID

[1719] Steve Anderson. Critical interfaces and digital making. Visible Language, 49(3):120–139, December 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V49N3_2015_E.pdf.

Murray:2015:MCL

[1720] Padmini Ray Murray and Chris Hand. Making culture: Locating the digital humanities in India. Visible Language, 49(3):140–155, December 2015. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V49N3_2015_E.pdf.

Sayers:2015:PP

[1721] Jentery Sayers. Prototyping the past. Visible Language, 49(3):156– 177, December 2015. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V49N3_2015_E.pdf.

McCarthy:2015:BAC

[1722] Steven McCarthy. Book art: a critical remix of The Electric Information Age Book. Visible Language, 49(3):178–203, December 2015.

CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V49N3_2015_E.pdf.

Zender:2016:ABa

[1723] Mike Zender. Advisory Board. Visible Language, 50(1):2, April 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V50N1_2016_E.pdf.

Davis:2016:NSC

[1724] Meredith Davis. “Normal science” and the changing practices of design and design education. Visible Language, 50(1):6–23, April 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/182/article/
1252;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N1_2016_E.pdf.

Poggenpohl:2016:DJC

[1725] Sharon Helmer Poggenpohl. Design journals: Context, serendipity, and value. Visible Language, 50(1):24–47, April 2016. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/182/article/
1282;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N1_2016_E.pdf.

Zender:2016:DRP

[1726] Mike Zender. Design research pioneer Josef Albers: a case for design research. Visible Language, 50(1):48–77, April 2016. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/182/article/
1272;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N1_2016_E.pdf.

dosSantosLonsdale:2016:TFT

[1727] Maria dos Santos Lonsdale. Typographic features of text and their contribution to the legibility of academic reading materials: an empirical study. Visible Language, 50(1):79–111, April 2016. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/182/article/
1262;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N1_2016_E.pdf.

Pena:2016:CLL

[1728] Ernesto Peña. Calculating line length: an arithmetic approach. Visible Language, 50(1):112–125, April 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/182/article/1292;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N1_2016_E.pdf.

DelRe:2016:PCT

[1729] Louis Del Re, Régis Vaillancourt, Gilda Villarreal, and Annie Pouliot. Pictograms: Can they help patients recall medication safety instructions? Visible Language, 50(1):126–151, April 2016. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/182/article/
1312;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N1_2016_E.pdf.

Kelly:2016:RAR

[1730] Meghan Kelly and Russell Kennedy. Recognizing appropriate representation of indigenous knowledge in design practice. Visible Language, 50(1):152–173, April 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/182/article/1302;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N1_2016_E.pdf.

Zender:2016:BRD

[1731] Mike Zender. Book review: Data Design by Per Mollerup. Visible Language, 50(1):175, April 2016. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V50N1_2016_E.pdf.

Anonymous:2016:JIa

[1732] Anonymous. Journal information. Visible Language, 50(1):176, April 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V50N1_2016_E.pdf.

Baudelaire:2016:XAF

[1733] Patrick Baudelaire. The Xerox Alto font design system. Visible Language, 50(2):12–25, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/202/article/1342;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N2_2016_E.pdf.

Carter:2016:DT

[1734] Matthew Carter. The digital typefoundry. Visible Language, 50 (2):26–37, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/202/article/1352;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N2_2016_E.pdf.

Schmoller:2016:TL

[1735] Hans Schmoller and ?. Matteson. Two letters: 1968 & 2016. Visible Language, 50(2):38–39, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/202/article/1352;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N2_2016_E.pdf.

Beeton:2016:CMT

[1736] Barbara Beeton and Richard Palais. Communication of mathematics with TEX. Visible Language, 50(2):40–51, August 2016. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/202/article/
1362;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N2_2016_E.pdf.

Mosley:2016:C

[1737] James Mosley. Commercial at @. Visible Language, 50(2):52–63, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/202/article/
1382;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N2_2016_E.pdf.

Beier:2016:LRA

[1738] Sofie Beier. Letterform research: an academic orphan. Visible Language, 50(2):64–79, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/202/article/1372;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N2_2016_E.pdf.

Grainger:2016:OPR

[1739] Jonathan Grainger. Orthographic processing and reading. Visible Language, 50(2):80–101, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/202/article/1392;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N2_2016_E.pdf.

Legge:2016:RDL

[1740] Gordon E. Legge. Reading digital with low vision. Visible Language, 50(2):102–125, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/202/article/1402;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N2_2016_E.pdf.

Lee:2016:ERB

[1741] Ryan Lee and Jeanne-Louise Moys. Exploring the relationship between language and design: a study of Hong Kong newspapers. Visible Language, 50(2):126–149, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/202/article/1412;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N2_2016_E.pdf.

Anonymous:2016:REC

[1742] Anonymous. Remembrances of eminent contributors to Visible Language’s first 50 years …. Visible Language, 50(2):150–173, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/202/article/
1422;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N2_2016_E.pdf.

Anonymous:2016:TTB

[1743] Anonymous. Top 50 typography books of the last 50 years. Visible Language, 50(2):174–177, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/202/article/1432;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N2_2016_E.pdf.

Ojeda:2016:BHS

[1744] Danné Ojeda and Mathieu Lommen. Book history scholarship: creation, transmission of knowledge and archives. Visible Language, 50(2):180–199, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/202/article/1442;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N2_2016_E.pdf.

Seward:2016:TEE

[1745] Reneé Seward and Emily Verba Fischer. Typography education: Ecosystem and excellence. Visible Language, 50(2):200–219, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/202/article/
1452;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N2_2016_E.pdf.

Zender:2016:ABc

[1746] Mike Zender. Advisory Board. Visible Language, 50(3):2, December 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V50N3_2016_E.pdf.

Zender:2016:RYD

[1747] Mike Zender. Reflecting on 50 years of design history + Visible Language. Visible Language, 50(3):5, December 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V50N3_2016_E.pdf.

Griffin:2016:RVL

[1748] Dori Griffin. The role of Visible Language in building and critiquing a canon of graphic design history. Visible Language, 50(3):7–27, December 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/232/article/
1562;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N3_2016_E.pdf.

Wragg:2016:GDS

[1749] Nicole Wragg and Carolyn Barnes. Graphic designers’ sense and knowledge of the user: Is thinking differently the groundwork for acting differently? Visible Language, 50(3):29–62, December 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/232/article/
1572;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V50N3_2016_E.pdf.

Tetlan:2016:HHP

[1750] Lou Tetlan and Douglas Marschalek. How humans process visual information: a focused primer for designing information. Visible Language, 50(3):65–88, December 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/232/article/1582;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V50N3_2016_E.pdf.

Anonymous:2016:VLCa

[1751] Anonymous. Visible Language call for papers: Visible language as social commentary. Visible Language, 50(3):91, December 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V50N3_2016_E.pdf.

Anonymous:2016:VLCb

[1752] Anonymous. Visible Language call for papers: Hermeneutics, communication theory, and visible language. Visible Language, 50(3): 93, December 2016. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V50N3_2016_E.pdf.

Anonymous:2016:JIc

[1753] Anonymous. Journal information. Visible Language, 50(3):96, December 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V50N3_2016_E.pdf.

Zender:2017:ABa

[1754] Mike Zender. Advisory Board. Visible Language, 51(1):2, April 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V51N1_2017_E.pdf.

Pedersen:2017:BIC

[1755] Pia Pedersen. Behind Isotype charts: The design of numberfact pictures. Visible Language, 51(1):12–37, April 2017. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/242/article/
1592;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V51N1_2017_E.pdf.

Hamann:2017:TWP

[1756] Agnieszka Hamann. Tz’ihb’ write/paint’. multimodality in Maya glyphic texts. Visible Language, 51(1):38–57, April 2017. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/242/article/
1632;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V51N1_2017_E.pdf.

Wright-Carr:2017:SRI

[1757] David Charles Wright-Carr. Signs of resistance: Iconography and semasiography in Otomi architectural decoration and manuscripts of the early colonial period. Visible Language, 51(1):58–87, April 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/242/article/
1602;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V51N1_2017_E.pdf.

Kwok:2017:ISP

[1758] Brian Sze-Hang Kwok and Anneke Coppoolse. Illuminography: a survey of the pictorial language of Hong Kong’s neon signs. Visible Language, 51(1):88–111, April 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/242/article/1622;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V51N1_2017_E.pdf.

Beier:2017:LIE

[1759] Sofie Beier, Katrine Sand, and Randi Starrfelt. Legibility implications of embellished display typefaces. Visible Language, 51(1):112–133, April 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/242/article/
1612;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V51N1_2017_E.pdf.

Bigelow:2017:WED

[1760] Chuck Bigelow, Mary Dyson, Maria dos Santos Lonsdale, and Kevin Larson. What exactly is the difference between a text and a display typeface? Visible Language, 51(1):134–142, April 2017. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/242/article/
1652;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V51N1_2017_E.pdf.

Anonymous:2017:VLC

[1761] Anonymous. Visible Language calls for papers: Special student issue and historical evidence. Visible Language, 51(1):143, April 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V51N1_2017_E.pdf.

Anonymous:2017:JIa

[1762] Anonymous. Journal information. Visible Language, 51(1):144, April 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V51N1_2017_E.pdf.

Zender:2017:DCN

[1763] Mike Zender and Alisa Strauss. Design by consensus: a new method for designing effective pictograms. Visible Language, 51(2):6–33, August 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V51N2_2017_E.pdf.

Zender:2017:DUD

[1764] Mike Zender. DrawIt: a user-drawn design research method for symbol design. Visible Language, 51(2):34–61, August 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V51N2_2017_E.pdf.

Ramirez:2017:MMS

[1765] Angélica Baena Ramírez. Metonymic and metaphoric series in the Codex Borgia. plates 33–34. Visible Language, 51(2):62–79, August 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V51N2_2017_E.pdf.

Lama:2017:WAM

[1766] Daniel Salazar Lama and Rogelio Valencia Rivera. The written adornment: the many relations of text and image in Classic Maya visual culture. Visible Language, 51(2):80–105, August 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V51N2_2017_E.pdf.

Ockerse:2017:BWH

[1767] Thomas Ockerse and Per Mollerup. Bespoke wayshowing in hospitals. Visible Language, 51(2):106–125, August 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V51N2_2017_E.pdf.

Renner:2017:PLI

[1768] Michael Renner. Practice-led iconic research: Towards a research methodology for visual communication. Visible Language, 51(3-52.1):

8–33, December 2017. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V51N3521_2017_E.pdf.

Schubbach:2017:PPL

[1769] Arno Schubbach. The practice of practice-led iconic research. Visible Language, 51(3-52.1):34–55, December 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V51N3521_2017_E.pdf.

Ahn:2017:DHV

[1770] Jinsu Ahn. The dynamism of Hangeul’s vertical strokes and the flow of its lines of writing. Visible Language, 51(3-52.1):56–73, December 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V51N3521_2017_E.pdf.

LopezGruninger:2017:IUC

[1771] Paloma López Grüninger. The image as unstable constellation: Rethinking Darwin’s diagram from the perspective of practice-led iconic research. Visible Language, 51(3-52.1):124–147, December 2017. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V51N3521_2017_E.pdf.

Reymond:2017:PIB

[1772] Claire Reymond. Premises for interaction between images. Visible Language, 51(3-52.1):148–173, December 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V51N3521_2017_E.pdf.

Hubner:2017:MTV

[1773] Michael Hübner. Making things visible: Visual strategies for the representation of objects in collections. Visible Language, 51(3-52.1):174– 201, December 2017. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V51N3521_2017_E.pdf.

Gromala:2017:BRF

[1774] Diana J. Gromala and Sharon Poggenpohl. Book review: Fire signs, a semiotic theory for graphic design. Visible Language, 51(3-52.1):204– 207, December 2017. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V51N3521_2017_E.pdf.

Kaser:2017:DIS

[1775] Susanne Käser. Documentary image sequences. Visible Language, 51(3-52.1):96–123, December 2017. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V51N3521_2017_E.pdf.

Aichmaier:2017:IDP

[1776] Helga Aichmaier. Identifying design processes in photography by analyzing photographic strategies in the documentation of public places: ‘It’s hard to be down when you’re up’. Visible Language, 51(3-52.1): 74–85, December 2017. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V51N3521_2017_E.pdf.

Trogu:2018:CLC

[1777] Pino Trogu. Counting but losing count: the legacy of Otto Neurath’s Isotype charts. Visible Language, 52(2):??, August 2018. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V52N2_2018_E.pdf.

Fisher:2018:SDF

[1778] Emma Fisher, Nicolette Lee, and Scott Thompson-Whiteside. Same but different: a framework for understanding conceptions of research in communication design practice and academia. Visible Language, 52(2):??, August 2018. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V52N2_2018_E.pdf.

Ramirez:2018:ROA

[1779] Rodrigo Ramírez. Reviewing open-access icons for emergency: a case study testing meaning performance in guemil. Visible Language, 52 (2):??, August 2018. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V52N2_2018_E.pdf.

Switzer:2018:UDR

[1780] Brian Switzer. Using design research for a better understanding of complex problems: a study of two homes for the elderly. Visible Language, 52(2):??, August 2018. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V52N2_2018_E.pdf.

Bakalka:2018:IPB

[1781] Laura Bakalka and Catarina Lelis. Informing personal branding through self-assessed handwriting analysis: proposal of a supportive online platform. Visible Language, 52(3):??, December 2018. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V52N3_2018_E.pdf.

Klohn:2018:IDP

[1782] Sara C. Klohn and Alison Black. Informative and decorative pictures in health and safety posters for children. Visible Language, 52 (3):??, December 2018. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V52N3_2018_E.pdf.

McIlwraith:2018:TDV

[1783] Andrew McIlwraith. Two-dimensional vs three-dimensional guide maps: which work best for museum visitors? Visible Language, 52(3): ??, December 2018. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V52N3_2018_E.pdf.

Moys:2018:EVP

[1784] Jeanne-Louise Moys, Peter Loveland, and Mary C. Dyson. eInk versus paper: exploring the effects of medium and typographic quality on recall and reading speed. Visible Language, 52(3):??, December 2018. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V52N3_2018_E.pdf.

Moys:2018:EIS

[1785] Jeanne-Louise Moys, Carmen Martínez-Freile, Rachel McCrindle, Lotte Meteyard, Holly Robson, Luke Kendrick, and Maitreyee Wairagkar. Exploring illustration styles for materials used in visual resources for people with aphasia. Visible Language, 52(3):??, December 2018. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V52N3_2018_E.pdf.

Perez:2018:GLD

[1786] Ana Perez and Maria dos Santos Lonsdale. Garment label design and companion information to communicate fashion sustainability issues to young consumers. Visible Language, 52(3):??, December 2018. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V52N3_2018_E.pdf.

John:2018:DNC

[1787] Nicola St John. Designing on Ntaria country: telling stories with new tools. Visible Language, 52(3):??, December 2018. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V52N3_2018_E.pdf.

Traina:2018:HCP

[1788] Rosanna Traina. How can the principles and practices of information design help us produce useful live art documentation? A unique user-centered, experience-design challenge. Visible Language, 52(3): ??, December 2018. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V52N3_2018_E.pdf.

Lysbakken:2018:SSE

[1789] Nina Lysbakken. Spaces that speak: Exploring creative opportunities in the communication design of conversational spaces. Visible Language, 52(3):??, December 2018. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V52N3_2018_E.pdf.

Griffin:2019:HVC

[1790] Dori Griffin. Histories of visual communication design. Visible Language, 53(1), April 2019. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V53N1_2019_E.pdf.

Auji:2019:IMM

[1791] Hala Auji. The implications of media: a material reading of nineteenthcentury Arabic broadsides. Visible Language, 53(1):??, April 2019. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V53N1_2019_E.pdf.

Avni:2019:IDQ

[1792] Shani Avni. Ismar David’s quest for original Hebrew typographic signs. Visible Language, 53(1):??, April 2019. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-2.

amazonaws.com/visiblelanguage/pdf/V53N1_2019_E.pdf.

Witehira:2019:MMS

[1793] Johnson Witehira. Mana matatuhi: a survey of Maori engagement with the written and printed word during the 19th century. Visible Language, 53(1):??, April 2019. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V53N1_2019_E.pdf.

Klevgaard:2019:LCF

[1794] Trond Klevgaard. Lower case in the flatlands: New typography and orthographic reform in a Danish printing calendar. Visible Language, 53(1):??, April 2019. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V53N1_2019_E.pdf.

Kim:2019:VAT

[1795] Min-Soo Kim, Jinsook Kim, and Michael H. Fritsch. For visual attention, are there any tendencies in form interpretation? Visible Language, 53(2):??, August 2019. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V53N2_2019_E.pdf.

Lonsdale:2019:VTT

[1796] Maria dos Santos Lonsdale, David J. Lonsdale, Matthew Baxter, Ryan Graham, Aya Kanafani, Anqi Li, and Chunxinzi Peng. Visualizing the terror threat. The impact of communicating security information to the general public using infographics and motion graphics. Visible Language, 53(2):??, August 2019. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V53N2_2019_E.pdf.

Pedersen:2019:LPP

[1797] Pia Pedersen. Legibility of pharmaceutical pictograms: Towards defining a paradigm. Visible Language, 53(2):??, August 2019. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V53N2_2019_E.pdf.

Martins:2019:DVI

[1798] Tiago Martins, João M. Cunha, João Bicker, and Penousal Machado. Dynamic visual identities: from a survey of the state-of-the-art to a model of features and mechanisms. Visible Language, 53(2):??, August 2019. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V53N2_2019_E.pdf.

Reymond:2019:IDB

[1799] Claire Reymond, Christine Müller, and Indre Grumbinaite. E-inclusion — defining basic image properties for illustrated stimuli in aphasia treatment. Visible Language, 53(3):??, December 2019. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V53N3_2019_E.pdf.

Bessemans:2019:VPS

[1800] Ann Bessemans, Maarten Renckens, Kevin Bormans, Erik Nuyts, and Kevin Larson. Visual prosody supports reading aloud expressively. Visible Language, 53(3):??, December 2019. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V53N3_2019_E.pdf.

Beier:2019:EAF

[1801] Sofie Beier and Chiron Oderkerk. The effect of age and font on reading ability. Visible Language, 53(3):??, December 2019. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-

west-2.amazonaws.com/visiblelanguage/pdf/V53N3_2019_E.pdf.

Zender:2019:LLV

[1802] Mike Zender. Letterform legibility and visual perception: a speculation. Visible Language, 53(3):??, December 2019. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://s3-us-west-

2.amazonaws.com/visiblelanguage/pdf/V53N3_2019_E.pdf.

Moys:2020:REB

[1803] Jeanne-Louise Moys and Laura Marshall. Readers’ experiences of Braille in an evolving technological world. Visible Language, 54(1–2):6–29, April 2020. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL http://visiblelanguage.herokuapp.com/issue/322/article/
2062;   https://s3-us-west-2.amazonaws.com/visiblelanguage/
pdf/V54N12_2020_E.pdf.

Nonaka:2020:LCD

[1804] Angela M. Nonaka, Jean Ann, and Keiko Sagara. Linguistic and cultural design features of the manual syllabary in Japan. Visible Language, 54 (1–2):30–65, April 2020. CODEN VSLGAO. ISSN 0022-2224 (print),

2691-5529 (electronic). URL https://s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V54N12_2020_E.pdf.

Serin:2020:LST

[1805] Efecan Serin, Aprigio Luis Moreira Morgado, and Ricardo Santos. A Latin-script typeface, based on special education teachers’ opinions, to use in literacy education of individuals with autism. Visible Language, 54(1–2):66–97, April 2020. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/322/article/2082;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V54N12_2020_E.pdf.

Peterson:2020:AGB

[1806] Matthew Peterson, Brad Tober, Deborah Littlejohn, and Mac Hill. Anticipating gaze-based HCI applications with the tech receptivity interval: Eye tracking as input. Visible Language, 54(1–2):98–127, April 2020. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V54N12_2020_E.pdf.

dosSantosLonsdale:2020:EBC

[1807] Maria dos Santos Lonsdale, Stephanie Sciberras, Hyejin Ha, and Stephen J. Chapman. Enhancing bowel cancer surgery recovery through information design. Visible Language, 54(1–2):128–169, April 2020. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V54N12_2020_E.pdf.

Kuraityte:2020:IKT

[1808] Milda Kuraitytė, Ann Bessemans, and Erik Nuyts. Impact of kinetic typography on readers’ attention. Visible Language, 54(1–2):170–185, April 2020. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.amazonaws.com/visiblelanguage/pdf/
V54N12_2020_E.pdf.

Parhami:2020:CCW

[1809] Behrooz Parhami. Computers and challenges of writing in Persian: Explorations at the intersection of culture and technology. Visible Language, 54(1–2):182–223, April 2020. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://visiblelanguage.

herokuapp.com/issue/322/article/2122;   https://s3-us-west-
2.amazonaws.com/visiblelanguage/pdf/V54N12_2020_E.pdf.

Krestev:2020:CTD

[1810] Krassen Krestev. The characteristics of text and display sizes in 16th Century Flemish Roman type: comparative analyses of seven types cut by Hendrik Van Den Keere in the period 1570–1580. Visible Language, 54(3):4–31, December 2020. CODEN VSLGAO. ISSN 0022-2224. URL

https://journals.uc.edu/index.php/vl/article/view/4609.

Renckens:2020:CPR

[1811] Maarten Renckens. Consequently positioning the rhythm in type based on the letters’ longest continuous black mass. Visible Language, 54(3):32– 47, December 2020. CODEN VSLGAO. ISSN 0022-2224. URL https:

//journals.uc.edu/index.php/vl/article/view/4610.

Sieghart:2020:IMC

[1812] Sabine Sieghart. The influence of macrotypography on the comprehensibility of texts in easy-to-read language: an empirical study. Visible Language, 54(3):48–93, December 2020. CODEN VSLGAO. ISSN 0022-2224.

URL https://journals.uc.edu/index.php/vl/article/view/4611.

Renckens:2021:VPS

[1813] Maarten Renckens, Leo De Raeve, Erik Nuyts, María Pérez Mena, and Ann Bessemans. Visual prosody supports reading aloud expressively for deaf readers. Visible Language, 55(1):4–47, April 2021. CODEN

VSLGAO. ISSN 0022-2224. URL https://journals.uc.edu/index.
php/vl/article/view/4603.

Renckens:2021:PSE

[1814] Maaten Renckens, Leo De Raeve, Erik Nuyts, María Pérez Mena, and Ann Beesemans. A preliminary study exploring the relation between visual prosody and the prosodic components in sign language. Visible Language, 55(1):48–65, April 2021. CODEN VSLGAO. ISSN 0022-2224.

URL https://journals.uc.edu/index.php/vl/article/view/4604.

Noel:2021:CDI

[1815] Guillermina Noël, Daren Pasay, Denise Campbell-Scherer, and Lynora Saxinger. Co-designing to improve practice in treating urinary tract infections: a case study of reducing inappropriate antibiotic treatment. Visible Language, 55(1):67–96, April 2021. CODEN VSLGAO. ISSN

0022-2224. URL https://journals.uc.edu/index.php/vl/article/
view/4605.

Bakker:2021:RRI

[1816] Wibo Bakker. Railway rules: Image content origins of wayfinding pictograms from the 1960s and 1970s. Visible Language, 55(2):5–26, August 2021. CODEN VSLGAO. ISSN 0022-2224. URL https://journals.

uc.edu/index.php/vl/article/view/4606.

Helmke:2021:TWW

[1817] Christophe Helmke and Jesper Nielsen. Teotihuacan writing: Where are we now? Visible Language, 55(2):28–73, August 2021. CODEN

VSLGAO. ISSN 0022-2224. URL https://journals.uc.edu/index.
php/vl/article/view/4607.

Kazmierczak:2021:FAI

[1818] Elzbieta T. Kazmierczak. Factors affecting interpretation of diagnostic images as a decision process: Ecological psychology, visual heuristics, and design. Visible Language, 55(2):74–111, August 2021. CODEN VSL-

GAO. ISSN 0022-2224. URL https://journals.uc.edu/index.php/
vl/article/view/4608.

Zender:2021:IMM

[1819] Mike Zender. Introduction: The makers and the made. Visible Language, 55(3):??, December 2021. CODEN VSLGAO. ISSN 0022-2224. URL

https://journals.uc.edu/index.php/vl/article/view/4678.

Baron:2021:BJB

[1820] Naomi S. Baron. Beyond judging books by their covers: “Reflections on Interrogating Cultural Anthropology Text Covers”. Visible Language, 55 (3):12–21, December 2021. CODEN VSLGAO. ISSN 0022-2224. URL

https://journals.uc.edu/index.php/vl/article/view/4669.

Bessemans:2021:EBB

[1821] Ann Bessemans and María Pérez Mena. Exploring the boundaries between visual anthropology and visual communication design research. Visible Language, 55(3):??, December 2021. CODEN VSLGAO. ISSN

0022-2224. URL https://journals.uc.edu/index.php/vl/article/
view/4670.

Dyson:2021:CRG

[1822] Mary Dyson. A commentary on: Reappropriation of gendered Irish sign language in one family. Visible Language, 55(3):28–35, December 2021.

CODEN VSLGAO. ISSN 0022-2224. URL https://journals.uc.edu/

index.php/vl/article/view/4671.

Golec:2021:EPV

[1823] Micheal J. Golec. Empathic projections: Visual anthropology design and acknowledgement. Visible Language, 55(3):36–49, December 2021. CODEN VSLGAO. ISSN 0022-2224. URL https://journals.uc.edu/

index.php/vl/article/view/4672.

Griffin:2021:RRB

[1824] Dori Griffin. Re reading the Borderland Imaginary from 2021. Visible Language, 55(3):??, December 2021. CODEN VSLGAO. ISSN 0022-2224.

URL https://journals.uc.edu/index.php/vl/article/view/4673.

Renner:2021:ECM

[1825] Michael Renner. Echoing the call for multimodal representation. Visible Language, 55(3):??, December 2021. CODEN VSLGAO. ISSN 0022-2224.

URL https://journals.uc.edu/index.php/vl/article/view/4674.

Renner:2021:DTV

[1826] Michael Renner. Drawing to tell versus drawing to intrigue? Visible Language, 55(3):??, December 2021. CODEN VSLGAO. ISSN 0022-

2224. URL https://journals.uc.edu/index.php/vl/article/view/
4676.

vanderWaarde:2021:DAV

[1827] Karel van der Waarde. Developing accurate visual conventions? Visible Language, 55(3):??, December 2021. CODEN VSLGAO. ISSN 0022-2224.

URL https://journals.uc.edu/index.php/vl/article/view/4677.

Overmann:2022:EWC

[1828] Karenleigh A. Overmann. Early writing: a cognitive archaeological perspective on literacy and numeracy. Visible Language, 56(1):8–45, ???? 2022. CODEN VSLGAO. ISSN 0022-2224. URL https://journals.

uc.edu/index.php/vl/article/view/4934.

Taffe:2022:BDF

[1829] Simone Taffe, Leon Sterling, and Sonja Pedell. Brainstorm: Do/be/feel and motivational modelling: Applying a new brainstorming process in the design of brandmarks. Visible Language, 56(1):46–71, ???? 2022. CODEN VSLGAO. ISSN 0022-2224. URL https://journals.uc.edu/

index.php/vl/article/view/4935.

Zender:2022:SIS

[1830] Mike Zender. Symbols: Interaction of symbols: Multiple symbols interact to provoke meaning, the foundation for better icon design. Visible Language, 56(1):72–105, ???? 2022. CODEN VSLGAO. ISSN 0022-2224.

URL https://journals.uc.edu/index.php/vl/article/view/4936.

Bessemans:2022:ART

[1831] Ann Bessemans. Academic research into typographic design at the beginning of a new era. Visible Language, 56(2):84–95, August 2022. CODEN

VSLGAO. ISSN 0022-2224. URL https://journals.uc.edu/index.
php/vl/article/view/6057.

Peterson:2022:PFO

[1832] Matthew Peterson. Persistent failure and occasional success: The realities of evidence-based interdisciplinary scholarship by design faculty at research universities. Visible Language, 56(2):8–53, August 2022. CO-DEN VSLGAO. ISSN 0022-2224. URL https://journals.uc.edu/

index.php/vl/article/view/6060.

Schmidt:2022:GDP

[1833] Michael Schmidt, Taghrid Asfar, and Wasim Maziak. Graphic design in public health research: a multiyear pictorial health warning label initiative and recommendation for sustained interdisciplinary collaboration. Visible Language, 56(2):54–83, August 2022. CODEN VSLGAO. ISSN

0022-2224. URL https://journals.uc.edu/index.php/vl/article/
view/6061.

Seward:2022:STC

[1834] Reneé Seward. Studying typography’s capacity to improve reading. Visible Language, 56(2):96–103, August 2022. CODEN VSLGAO. ISSN

0022-2224. URL https://journals.uc.edu/index.php/vl/article/
view/6068.

Thiessen:2022:ETC

[1835] Myra Thiessen, Hannah Keage, Indae Hwang, Jack Astley, and Sofie Beier. Effect of typeface complexity on automatic whole-word reading processes. Visible Language, 56(3):8–31, December 2022. CODEN VSL-

GAO. ISSN 0022-2224. URL https://journals.uc.edu/index.php/
vl/article/view/6393.

Kudo:2022:IDP

[1836] Mao Kudo. The inclusive design of pictograms and easy to understand for people with intellectual disabilities: Focused on graphic elements.

Visible Language, 56(3):??, December 2022. CODEN VSLGAO. ISSN

0022-2224. URL https://journals.uc.edu/index.php/vl/article/
view/5976.

Tjung:2022:DFL

[1837] Caroline Tjung, Simone Taffe, Simon Jackson, and Emily Wright. Design features of learning apps for mobile gamification: Graphic designers use co-design to prompt young children to speak. Visible Language, 56(3):32– 57, December 2022. CODEN VSLGAO. ISSN 0022-2224. URL https:

//journals.uc.edu/index.php/vl/article/view/6394.

Kudo:2022:GDP

[1838] Mao Kudo. Graphic design of pictograms focusing on the comprehension of people with intellectual disabilities — the next step in standardization: Pictogram design and evaluation methods. Visible Language, 56(3):58– 85, December 2022. CODEN VSLGAO. ISSN 0022-2224. URL https:

//journals.uc.edu/index.php/vl/article/view/6395.

Thiessen:2023:IGE

[1839] Myra Thiessen, Daphne Flynn, Leah Heiss, Rowan Page, Nyein Aung, and Indae Hwang. Introduction: Guest Editors. Visible Language, 57(1): 10–13, April 2023. CODEN VSLGAO. ISSN 0022-2224. URL https:/

/journals.uc.edu/index.php/vl/article/view/7322.

Thiessen:2023:FPC

[1840] Myra Thiessen, Leah Heiss, Troy McGee, and Gene Bawden. The future is participatory: Collaborative communication design for global health initiatives. Visible Language, 57(1):14–37, April 2023. CODEN VSL-

GAO. ISSN 0022-2224. URL https://journals.uc.edu/index.php/
vl/article/view/7318.

Ellerton:2023:HMO

[1841] Wendy Ellerton. The human and machine, 2022–23: Open AI, Chat-GPT, Quillbot, Grammarly, Google, Google Docs & humans*. Visible Language, 57(1):38–52, April 2023. CODEN VSLGAO. ISSN 0022-2224.

URL https://journals.uc.edu/index.php/vl/article/view/7319.

Taljaard:2023:SMW

[1842] Darren Taljaard and Myra Thiessen. Show me what you mean: Inclusive augmented typography for students with dyslexia. Visible Language, 57 (1):53–75, April 2023. CODEN VSLGAO. ISSN 0022-2224. URL https:

//journals.uc.edu/index.php/vl/article/view/7320.

McCarthy:2023:DCP

[1843] Steven McCarthy. Dirty concrete poetry and white space: The visual texts of Steve McCaffery and Douglas Kearney. Visible Language, 57(2): 6–35, August 2023. CODEN VSLGAO. ISSN 0022-2224. URL https:/

/journals.uc.edu/index.php/vl/article/view/8183.

Piovesan:2023:PQT

[1844] Andrea Piovesan, Michele Sinico, and Luciano Perondi. The perception of qualities in typeface. Visible Language, 57(2):36–79, August 2023. CODEN VSLGAO. ISSN 0022-2224. URL https://journals.uc.edu/

index.php/vl/article/view/8184.

Halldorsson:2023:TEE

[1845] Halldór Björn Halldórsson. The tacit exhibition: an experimental display of graphic design knowledge. Visible Language, 57(2):80–99, August 2023. CODEN VSLGAO. ISSN 0022-2224. URL https://journals.

uc.edu/index.php/vl/article/view/8185.

Poggenpohl:2023:BTR

[1846] Sharon Helmer Poggenpohl. A book to think with: a review: After the Bauhaus, Before the Internet: a History of Graphic Design Pedagogy, [edited by Geoff Kaplan, no place press, New York, NY, USA (2022). ISBN 1-949484-09-2, ISBN-13 978-1-949484-09-0]. Visible Language, 57 (2):100–112, August 2023. CODEN VSLGAO. ISSN 0022-2224. URL

https://journals.uc.edu/index.php/vl/article/view/8187.